A Complete Bibliography of Publications in the Journal of Theoretical Biology: 2020–2029

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

14 August 2025 Version 1.71

Title word cross-reference

2 [AFTF22, KRSSH21, PD23, RBG^{+}23, SRB^{+}24, Yin21]. 3 [ABB^{+}22, BMB^{+}21, FSC21, GF23, MPTCT^{+}20, MMS^{+}23, YG25]. 4 [CKTW24]. ^{+} [BCK24, LLS^{+}22]. ^{2+} [CBK^{+}20, KK22a, LYE^{+}24, PO20]. _{2} [AG22]. a [PLSS24a, PLSS24b]. β [CSF20, FVB20, GHB22, IMM24, IYL^{+}25, Kuz24a, Kuz24c]. γ [PJVR23]. h [PWC^{+}23]. n [GZ24, KT20a]. Q [YCKK24]. R_{0} [KRSSH21]. T [LLS^{+}22].

-catenin [CSF20]. -cell [IYL^{+}25]. -cells [FVB20]. -Dimensional [AFTF22]. -protein [IMM24]. -serovar [KT20a]. -species [GZ24]. -subtree [CKTW24]. -wave [PLSS24a, PLSS24b].

/IL [PJVR23]. /IL-4 [PJVR23].

01-17 [PLSS24a]. 01-17/111879 [PLSS24a].

1 [BS21, CTL^{+}22, KKK^{+}20, LBF24, NTTI22, PGML20, SF23]. 17/111879 [PLSS24a]. 19 [AGP^{+}21, AGP^{+}23, APF^{+}24, ASS^{+}23, AMR^{+}23, BK21, BGM22, BDdG22, CS21, CSDH^{+}25, DAB^{+}23, DTS21, DPR21, DJ22, ECH^{+}23, FPGOCP^{+}25, GSW^{+}23, GGSTS21, HPGF25, JKH^{+}23, KPA^{+}21, KBW^{+}23, LPLM25, LMS21, LZZ^{+}23, LMW21, MEPS22, NGKB21, PRJ^{+}20, REK20, STCSC^{+}21, SFK^{+}23, SRDC21, SXJ20, Täu20, TNHV21, TWT23, YTY^{+}20, YLT22, ZQHN24, ZZT^{+}21, ZWC^{+}22, ZWW22, dIMM24, dCC^{+}23].

2 [ACF22, BLRV21, BFSA23, CWF23, CLJ^{+}23, CWL20, CPEK23, CDdV23, CWT^{+}23, CSH^{+}25, DNK^{+}24, GFV^{+}23, GFC^{+}22, Hil23, JMWX23, KvLBD25, Kee23, KBD^{+}23, LRST22, NSZ^{+}25, PWC^{+}23, PDD24, Roy21, TAP^{+}23, WUH23, WHRH^{+}23, XCFH23, YY21, dCC^{+}23]. 2020 [Ano20m, Ano20a, Ano20n, Ano20b, Ano20o, Ano20c, Ano20p, Ano20d, Ano20q, Ano20e, Ano20s, Ano20g, Ano20r, Ano20f, Ano20t, Ano20h, Ano20u, Ano20i, Ano20v, Ano20j, Ano20w, Ano20k, Ano20x, Ano20l, YLT22]. 2021 [Ano21m, Ano21a, Ano21n, Ano21b, Ano21o, Ano21c, Ano21p, Ano21d, Ano21q, Ano21e, Ano21s, Ano21g, Ano21r, Ano21f, Ano21t, Ano21h, Ano21u, Ano21i, Ano21v, Ano21j, Ano21w, Ano21k, Ano21x, Ano21l]. 2022 [Ano22l, Ano22a, Ano22m, Ano22b, Ano22n, Ano22c, Ano22o, Ano22d, Ano22p, Ano22e, Ano22r, Ano22g, Ano22q, Ano22f, Ano22s, Ano22h, Ano22t, Ano22i, Ano22u, Ano22j, Ano22v, Ano22w, Ano22k]. 2023 [Ano23j, Ano23k, Ano23b, Ano23l, Ano23c, Ano23m, Ano23d, Ano23o, Ano23f, Ano23n, Ano23e, Ano23p, Ano23g, Ano23q, Ano23h, Ano23r, Ano23s, Ano23t, Ano23i]. 2024 [Ano24i, Ano24a, Ano24j, Ano24b, Ano24k, Ano24l, Ano24c, Ano24m, Ano24d, Ano24o, Ano24f, Ano24n, Ano24e, Ano24p, Ano24g, Ano24q, Ano24h, Ano24r, Ano24s, Ano24t]. 2025 [Ano25h, Ano25i, Ano25j, Ano25b, Ano25k, Ano25c, Ano25m, Ano25e, Ano25l, Ano25d, Ano25n, Ano25f, Ano25o, Ano25g, Ano25p, Ano25q]. 21 [Ano20a, Ano20b, Ano20c, Ano20d, Ano20e, Ano20g, Ano20f, Ano20h, Ano20i, Ano20j, Ano20k, Ano20l, Ano21a, Ano21b, Ano21c, Ano21d, Ano21e, Ano21g, Ano21f, Ano21h, Ano21i, Ano21j, Ano21k, Ano21l, Ano22a, Ano22b, Ano22c, Ano22d, Ano22e, Ano22g, Ano22f, Ano22h, Ano22i, Ano22j, Ano22k, Ano23b, Ano23c, Ano23d, Ano23f, Ano23e, Ano23g, Ano23h, Ano23i, Ano24a, Ano24b, Ano24c, Ano24d, Ano24f, Ano24e, Ano24g, Ano24h, Ano25b, Ano25c, Ano25e, Ano25d, Ano25f, Ano25g]. 27 [LBF24]. 289 [JCLS^{+}21].

3a [AAL^{+}21].

4 [KK23, PJVR23]. 40 [MP22]. 417 [Sch23]. 439 [SSKS20]. 447 [Lio24]. 454 [Sch21a]. 456 [GHS20]. 483 [WLW20]. 486 [D’A20a, RB22]. 492 [Lei24]. 498 [PP24]. 4T1 [GYY^{+}23].

526 [Cor22]. 527 [HPSA23]. 530 [Xie22]. 537 [Akh23]. 545 [CBH^{+}23]. 592 [PLSS24a].

605 [LK25a].

7 [Ano20m, Ano20n, Ano20o, Ano20p, Ano20q, Ano20s, Ano20r, Ano20t, Ano20u, Ano20v, Ano20w, Ano20x, Ano21m, Ano21n, Ano21o, Ano21p, Ano21q, Ano21s, Ano21r, Ano21t, Ano21u, Ano21v, Ano21w, Ano21x, Ano22l, Ano22m, Ano22n, Ano22o, Ano22p, Ano22r, Ano22q, Ano22s, Ano22t, Ano22u, Ano22v, Ano22w, Ano23j, Ano23k, Ano23l, Ano23m, Ano23o, Ano23n, Ano23p, Ano23q, Ano23r, Ano23s, Ano23t, Ano24i, Ano24j, Ano24k, Ano24l, Ano24m, Ano24o, Ano24n, Ano24p, Ano24q, Ano24r, Ano24s, Ano24t, Ano25h, Ano25i, Ano25j, Ano25k, Ano25m, Ano25l, Ano25n, Ano25o, Ano25p, Ano25q].

9 [SMS20].

abandoning [LYYN20]. abdominal [HBWLLPA21]. abductive [PAM24]. ablation [FFGR25]. abnormal [HRD^{+}24]. abruptness [STB20]. absence [HGZM25, LYE^{+}24]. absenteeism [AMR^{+}23]. absolute [WCL^{+}25]. Absorbing [DHC25]. abstraction [ASK20]. abundance [KL21b, SS21, TBNS23, vBW20]. abundance-dependent [SS21]. abundances [AWC25]. abuse [UV23]. acceleration [KP21]. accident [LLJ25]. accompanying [FFGR25]. according [LKP^{+}20]. account [BKP23, KK23, RLL25]. Accounting [MCM20, WHH22]. accumulation [JCLS^{+}11, JCLS^{+}21, PS22b, TS23]. accuracy [GFV^{+}23]. Accurate [YFF25]. accurately [LHG^{+}22]. ACE2 [PRJ^{+}20]. acetaminophen [GBRM21]. acid [CWT^{+}23, MBF21]. acids [IYL^{+}25]. acquires [APF^{+}24]. acquisition [BSL24]. across [AMR^{+}23, BYHH20, JJR20, KPH^{+}24, PJB21, RWHK21, Wor20, XML22, ZZT^{+}21]. actin [RBG^{+}23, YL25, dLFO21]. acting [Sak20, Sak22a, Sak22b]. action [RS24, Vol24]. activated [VTDM21]. activating [SHY21]. activation [AAL^{+}21, BCK24, CPMF^{+}22, DB22a, KB20, MP22, SBC^{+}20, SSKP24]. Active [MI20, AJR^{+}20, JJR20]. activities [SSX^{+}22, SY25, ZSL23]. activity [AKHM23, IYL^{+}25, MCP20, PR21, YLZ^{+}21]. acute [BLP20, PBB20]. adapt [Haj20]. adaptation [RKCvB25]. adaptations [SL23]. adapted [SC23b]. Adapting [MMI22]. Adaptive [CC22, LZZ^{+}23, RID23, AJR^{+}20, BP23, Bau20, BBDB22, GF23, GGG21, JKSD23, MH22b, NVS20, Osb24, PAV21, PLR24, Täu20]. adaptively [For24]. adenocarcinoma [WCZL25]. ADHD [AGTJ23]. adherence [DHC25, YHT24]. adhesion [IHO^{+}20, MDS25, PJB21]. adhesive [SK23]. adipocyte [GMR^{+}24]. adiponectin [SNK20]. Adipose [DWNW24, HBWLLPA21]. adjacent [KMA^{+}21]. adjustments [SW22]. administration [BLP20]. ADRC [SSLMGC25]. adrenergic [MCL21]. ADT [LF25]. adult [AWC25, LSZ21, WMHV24]. advanced [HKL25, KT20b]. advantage [Gav25, KPP20, MGCH21, ZWC^{+}22]. advantageous [LK25a, LK25b]. adversarial [RMSS24]. Aedes [APR^{+}22a]. aegypti [APR^{+}22a]. aerodynamic [PCR21]. affect [AKHM23, BTS24, HPSA21, HPSA23, HDM20, PD23, WDHB20, ZWC^{+}22]. affected [BEBA20]. affecting [Xie23]. affects [AJR^{+}20, KIM^{+}22, PS22a, PBB20, RGBD25, SYH20, TIS25]. Africa [LPP^{+}25]. African [ZRL^{+}21]. after [AOB21, BPP^{+}20, CNK24, CPMF^{+}22, HHI24, WP21, YS25]. against [BPP^{+}20, BS25, KT20a, KK22b, MMI22, OKF^{+}21, PRJ^{+}20, PGC^{+}22]. Age [APR^{+}22b, DRK23, KHHI25, AWC25, BK21, DDDMT22, DKSP24, HK21, HPGF25, LRST22, PLSS24a, PLSS24b, SSBG20, WSY21, ZF21]. Age-dependence [KHHI25]. age-related [PLSS24a, PLSS24b]. age-structure [HPGF25]. Age-structured [APR^{+}22b, DDDMT22, HK21, LRST22, SSBG20]. aged [WMHV24]. ageing [BC20]. Agelessness [RR24]. Agent [YTG24, Ala20, GTGB21, GGG21, IH21, Pro24, SLML20, SSK^{+}25, TT20]. Agent-based [YTG24, Ala20, GTGB21, GGG21, IH21, SSK^{+}25]. agent-model [SLML20]. aggregate [GWZ^{+}22]. aggregates [Kuz24a]. aggregation [BERR20, FSKG20, ND22, OYT25, RB22]. aggregations [AKHM23]. aggregative [CRDD22]. Aging [BF20, JSCW21]. Aging-induced [JSCW21]. agricultural [BB21b]. AIDS [ZXLH23]. aimed [AHW21]. air [HHT25, OTCE24, SHH20, ZZ22]. airborne [ABOR20]. Airway [Don23, NM23]. airways [DWNW24, DEM^{+}23]. Akt [GC20, KLLL22]. alcohol [GBRM21]. algae [OTI^{+}20, TMTK22]. algae-phosphorus [TMTK22]. algal [SGJ^{+}24]. algebraic [MPTCT^{+}20, MZLJ20]. algorithm [AFTF22, CLW^{+}24, YFF25]. Algorithmic [Pat22]. algorithms [FPGOCP^{+}25]. alignment [KF20]. alike [HGD^{+}25]. Allee [AHEB^{+}22, Leh21, LG23, MH22a, SSB20, YM23, ZM20]. Allele [SC23b, Sak22b]. allergy [KHHI25]. allocation [BCBP25, CD22, FHTK20, IDMG24, KHST21, MB21]. allografts [GYY^{+}23]. allows [KK22b]. alone [Avv21]. along [HI21, LKP^{+}20]. alpha [CWT^{+}23, MCP20, RE25]. alpha-cell [MCP20]. alphabet [MBF21]. alter [SCH24]. alternate [EHA^{+}21]. alternative [KNY25, PBP23]. altruism [GCM23, Gar24]. altruistic [FFAV21, Pag25]. alveolar [ACF22]. Alzheimer [IMM24, LLZ^{+}22]. American [LZ23]. amino [CWT^{+}23, MBF21]. among [ADLM24, CBPG23, Kan21, MH22b, Sak22a, TP24, WS24, XSH^{+}24, YJBF20]. among-parent [Sak22a]. amplitude [MG21]. amyloid [Kuz24c]. amyloid- [Kuz24c]. analog [KK22a]. Analogy [YTW21]. analyse [TCMA21]. analyses [WS24]. Analysis [GFV^{+}23, KT21, PCR21, SY25, WBKR22, Ala23, APF^{+}24, BJJM25, BCLJ24, BSC20a, BMB^{+}21, BPM^{+}22, CSHH20, CPHL^{+}20, CF20, CGP^{+}20, Cor21, Cor22, DSdP22, DMR21, DMAC21, ES24, FNGJ21, FLR^{+}21, GGS^{+}24, GSMG22, HJKS20, HXZ^{+}22, HLW24, HGF23, Ish23, KT20a, KG22, KvLBD25, KPR20, KRG24, LPP^{+}25, LMB22, LB20, LS25a, LML24, OGM^{+}25, PBP23, Pat22, PPGB20, dMQHRGA^{+}22, RNP^{+}25, RRAO21, RKCvB25, Sak22b, Sal24, SYI20, SCHH20, SSKS18, SSKS20, SL23, TKRA20, WLJ^{+}20, WRB22, YZSS20, ZB22, ZSL23, ZWW22]. analytical [KK22a, NSZ^{+}25]. analyze [MMS^{+}23]. anatomical [LKP^{+}20]. anatomy [WS22]. Ancestral [DHR25, MGVS21]. ancestry [Di 20]. androdioecious [YI21]. Angiogenesis [GDd^{+}21, DMAC21, FLR23, HAP22, HYM^{+}22, NNOBS23]. angiogenic [HIA^{+}22]. angiotensin [TDRC23]. angles [BBR^{+}21]. angular [SF24a]. Animal [CCSF21, ABBP21, BLRV21, BERR20, HIA^{+}22, YWWF25]. animals [HL24, KM20, SKNvRL22, TP24]. anisogamy [JWKA21]. announcement [FHM20]. annual [Tay23]. answer [KI21]. ant [HDM20]. antagonism [YJBF20]. antagonistic [Bau20]. antennal [PR21, TKR^{+}21]. anthesis [WS24]. anthrax [KM22, MK21]. anti [Ala20, BH20, DMAC21, GYY^{+}23, HIA^{+}22, HL21, KKL20, LDF23, LLS^{+}22, LBF24, SF24b]. anti-angiogenesis [DMAC21]. anti-angiogenic [HIA^{+}22]. anti-cancer [BH20, KKL20, LBF24]. anti-conformity [LDF23]. anti-diabetic [SF24b]. anti-metastatic [GYY^{+}23]. anti-microbial [Ala20]. anti-PD-1 [LBF24]. anti-poaching [HL21]. anti-tumour [LLS^{+}22]. antibiotic [MBM^{+}25, SP20]. antibodies [BS21, GFV^{+}23, Sch17, Sch23, TEWN21, dCC^{+}23]. Antibody [CWF23, BKP23, KT20a]. antibody-dependent [KT20a]. Anticipating [AJR^{+}20]. Antigenic [BRS24]. antimicrobial [GHF22]. Antisocial [LDL^{+}24]. antitumor [KK23]. antiviral [KKK^{+}20, LZZ20, OKF^{+}21]. antral [Avv21]. ants [GGG21, IH21]. APOBEC3 [KKK^{+}20]. apoptosis [LMA21, QYY^{+}23]. apoptotic [MF24]. Application [GSW^{+}23, GYY^{+}23, KHAK21, MEPS22, NCL20, Roy21, STCSC^{+}21, YLT22, ABBP21, CGK23, CSDH^{+}25, DTS21, IB21, MSL23, SSLMGC25, SSK^{+}25]. applications [RK22b, WHH22, YG25]. applied [AG22, BGM22, RRAO21, VMRSL23]. Approach [GT20, Avv23, BDdG22, CCH^{+}24, CHMM23, CMRMC^{+}21, CTL^{+}22, CDdV23, Gar24, GBHNHVL23, GtHH20, GEHL^{+}22, Haj20, HAGADDM24, HLS^{+}25, JAW25, KT20a, KH20, KTP^{+}20, MCL20, NGKB21, OSAR21, Sal25, SRB^{+}24, SJS^{+}25, SLML20, TKB^{+}20, TCMA21, VKS^{+}20, WNHK21, Wu24, vBBB^{+}25]. approaches [CCA20, EPF20, KV23, MY20, PBP23]. approximate [LZW23, SD20, SF24a, SBB^{+}22]. Approximating [HTD22]. approximation [ESW24, dJHP20]. approximations [ACMT24, BS22a, GTGB21, LK20]. April [Ano20m, Ano20a, Ano21m, Ano21a, Ano22l, Ano22a, Ano23j, Ano24i, Ano24a, Ano25h, Ano23a]. aptamers [EPF20]. aquatic [ACP20, CG21, PTK^{+}24, QWZK24]. Arabidopsis [ADP21]. arbitrary [Wad22]. arboviral [SUA^{+}20]. arbuscular [MGZ^{+}21]. arcades [RKCvB25]. architecture [SAK21, TS23, TIS25]. area [GSC23, MMS^{+}23, SKNvRL22]. argument [Bar21]. arising [RG21]. arithmetic [GCM23]. arsenal [Pat22]. Arterial [RKCvB25, SVHM^{+}21, JJR20, TDRC23, YTW21]. arteries [ZL20]. arterioles [FFR22]. arthropod [PCR21]. arthropods [PCR23]. article [PMCLC^{+}23, RB22]. articular [MFGACST23]. Artificial [FBD23, KT21]. Artiodactyl [SCD^{+}23]. Ascertainment [CPEK23]. Asgardarchaeota [Di 20]. Asia [TWT23]. aspect [KT20b]. aspects [Ste21a]. aspirin [AAL^{+}21]. assay [FSC21]. assembly [GGS^{+}24, MDF21, SCH24, YRMWT20]. assess [AY22, BDdG22, GW20].

Assessing [HCR20, IHG21, Ngo22, VCS24, VGTY21, JKH^{+}23, KT20b, KHB^{+}22]. assessment [HPD20, KRG24, PGML20, RLW^{+}21, TEWN21, ZXLH23]. assist [MAHMH20, ZMKF22]. associate [APR^{+}22b]. associated [BNPL20, HHT25, Nah20, SHH20, WFL^{+}24, YFZ^{+}20]. Association [NR20, MZLJ20]. associations [WBKR22]. assortment [IIH25, ML24]. assuming [WS24]. assumptions [GHF22]. asthma [DLN20]. astrocyte [HS20, KK22a, RLL25]. astrocytes [SS22]. asymmetric [KC22, YWQX23, ZYF25]. asymmetry [CSVB25, SCD^{+}23]. asymptomatic [GSW^{+}23, GGSTS21, MEPS22, WUH23]. asymptotic [Pat22]. atherogenesis [YE21]. Atheroprotective [YE21]. atherosclerosis [SJNRS20, EE23]. atherosclerotic [JJR20]. atomistic [FSC21]. atomistic-continuum [FSC21]. atopic [PJVR23]. atrial [DRK23]. atrophy [AHS23]. Attack [Sin21, NAR^{+}20]. attainment [Hon20]. attenuation [BEBA20, SC21]. attractants [TML20]. atypical [HGZM25]. August [Ano20n, Ano20b, Ano21n, Ano21b, Ano22m, Ano22b, Ano23k, Ano23b, Ano24j, Ano24b, Ano25i, Ano25a]. aureus [BJB^{+}20, SHH20]. auris [KRG24]. Austria [KPA^{+}21]. auto [YST21]. auto-regulated [YST21]. autocatalysts [SR24]. Autocatalytic [SSS21, Kuz24a, Rav21]. Autocatalytic-protection [SSS21]. Autoimmune [HI20, vdVPW^{+}22]. automata [DEM^{+}23, LPSMCP23, LSZ21, HCR20]. Automated [RW20]. automaton [LSWVB24]. autonomous [BH23, OKT^{+}21, WSEK23]. autoregression [MCMdAM24]. autoregulation [ADP21]. availability [FHTK20, MSM23]. avascular [FBSD22, MPA23]. average [KL21b]. averaging [vBBB^{+}25]. avian [XYK^{+}20]. avidity [KKL20]. avium [WPEP22]. avoidance [OIK^{+}24, QPL21]. avoiding [GQG23]. aware [BK20a]. awareness [LWXM20]. Axin2 [CSF20]. axis [LKP^{+}20]. axisymmetric [YTW21]. axonal [KK22b]. axons [CCH^{+}24, CHMM23].

B [CMRMC^{+}21, KPH^{+}24, LGSR21, PD23, RY22, SD25]. B/L [PD23]. B4 [NSZ^{+}25]. background [BCC^{+}21, Kau23]. background-induced [BCC^{+}21]. bacteria [ATP22, GT22, Kan21, KHAK21, MK24, TTBL21]. bacteria-particle [MK24]. Bacterial [Sch17, Sch23, WPEP22, BJB^{+}20, ECBR20, MBM^{+}25, MDF21, SHH20, TML20]. bacteriophage [Bol22]. badger [LZ23]. balance [AKHM23, AA22, KNK21, LB20, ZB22]. balanced [VG25]. ball [BFG21, YZB20]. ball-rolling [YZB20]. bambara [DSM^{+}23]. band [RE25]. banded [JMD20]. banks [PB22]. basal [HXZ^{+}22]. based [Akh22, Akh23, Ala20, Ala23, ACD20, AGBLGA^{+}22, AR21, BRN^{+}22, BC21, BGB^{+}21, BEBA20, BMB^{+}21, BS22b, CGS^{+}25, ECH^{+}23, EBK21, FHR22, FLR23, GT22, GTGB21, GCC20, GSMG22, GGG21, HJKS20, HHT25, HBWLLPA21, Hoe24, HXS21, IH21, JMD20, JFCK20, JCC20, KR22, KCW^{+}25, KK22a, KIO^{+}25, LPP^{+}25, LPBMA21, LMS21, LFZ^{+}24, LMC^{+}20, LB20, LCYD23, MPTCT^{+}20, MZLJ20, MI23, NR20, OGM^{+}25, PASO21, PHL^{+}23, PZG20, RT24a, RKCvB25, SSLMGC25, SJS^{+}25, SRDC21, SC21, SHK20, SSK^{+}25, SVHM^{+}21, TCMA21, VVD24, WCM^{+}21, WSL25, YTG24, YCKK24, ZPP23, dLFO21]. basic [DF22, GHF22, GHS18, GHS20, GBHNHVL23]. BASiCS [OV25]. basis [MI20]. bat [GtHH20]. Batesian [QHS23, TT24]. Battle [CK20b]. Bayes [CGS^{+}25]. Bayesian [CRG^{+}23, CGS^{+}25, HGF23, LG23, LZW23, YOIS24]. be [BC20, LBF24, MC20b, SSR21]. beam [LR21]. bearded [CKF^{+}24]. bearing [EHA^{+}21, MUMS24]. bee [CFH^{+}25, SWK^{+}20]. bees [HZS^{+}21]. beetle [BFG21, CGY^{+}23, KCW^{+}25]. beetles [YZB20]. behavior [DMPIG24, IH21, Leh22, Lei20, Lei24, NSMH25, PAV21, SLML20, SMS20, VKS^{+}20, YZB20]. Behavioral [HDM20, BLP20, MSM23, SBSB23, SZR^{+}25]. behaviors [ETTT24, LZZ^{+}23]. behaviour [ABC^{+}21, BEBA20, FFAV21, IHO^{+}20, KHW^{+}22, MW20, NSD22, SOKZ22, Wil23]. behavioural [BDdG22]. behaviours [Rav21, RSPB^{+}22]. behind [Avv23]. Belgium [APF^{+}24]. believing [PZ20]. belowground [KHST21]. bending [HH21]. beneficial [Kau23]. beneficiaries [MCL20]. benefit [PPS^{+}21]. benevolent [FPX20]. benthic [IYY22]. Bernoulli [CGR22, ILGI20]. Best [FBA21, BD21, SH21, TATT21]. best-response [SH21]. bet [ZX22, HKBK25]. bet-hedging [ZX22]. beta [CWT^{+}23, RE25]. beta-band [RE25]. better [EBA21]. between [APR^{+}22a, AJR^{+}20, AAL^{+}21, BHW^{+}20, BRAHV25, CLJ^{+}23, FG21, FG22, GFN23, GMS20, HVC^{+}21, HGD^{+}25, JME24, KM22, Kuz24a, KMA^{+}21, MCL20, NNNHA24, NO24, Pfe22, PJPH25, QHS23, SA25, SUA^{+}20, SH21, SZR^{+}25, TAP^{+}23, TK25, Xie23, YLZ^{+}21, YJBF20]. between-host [BRAHV25, FG22, SA25, SZR^{+}25]. Beverton [FS22]. Beyond [DB22b, RCM^{+}25, PK23, Whi20]. bi [GK21]. bi-stable [GK21]. bias [DPR21, SL22]. Biased [JTB25]. biases [HPSA21, HPSA23, Par22]. Bidirectional [KK22b, HXZ^{+}22]. biennial [MUMS24]. biennial-bearing [MUMS24]. Bifurcation [FLR^{+}21, DBS23, HXZ^{+}22, KG22, LS25a, LS25b, WA22]. bifurcations [GK21]. big [KI21]. binarization [PZG20]. binding [CGG^{+}25, Sch21b, TML20, WDT^{+}20]. binomial [CGR22, LMS21]. bio [MPTCT^{+}20]. bio-macro-molecular [MPTCT^{+}20]. Biochemical [Sal24]. biocontrol [AY22]. bioelectric [Car23]. bioelectrical [CMLM23]. Bioengineering [GT20]. biofilm [GJDS20, WWR^{+}24]. Biofilms [WPEP22]. biogenic [IHS25]. Biol [Akh23, CBH^{+}23, Cor22, HPSA23, JCLS^{+}21, Lei24, Lio24, LK25a, PLSS24a, PP24, Sch21a, Sch23, SSKS20, Xie22]. biological [BH24, BEO^{+}22, CBBD22, FFNG^{+}21, GWZ^{+}22, GSMG22, HCSAB21, IKYW21, LSW20, LMC^{+}20, MMS^{+}23, SJS^{+}25, SCH24, Tsu21, vdBMB^{+}23]. Biology [D’A20a, Don23, GADDVG^{+}24, GHS20, WLW20, Wil23, Dal22, KTP^{+}20, Pru25, RLB25, Sal24, SMM24, Tha24, WBS20]. biomarkers [LPSMCP23, LLZ^{+}22]. Biomathematical [MMK^{+}24]. biomechanical [BC21]. Biomechanics [PTK^{+}24]. biomedical [NR20]. Biomimetic [LMLP20]. biomolecular [GCC20, Xie20]. bioprocess [SSLMGC25]. bipedal [HXS21]. biphasic [SGM^{+}25]. birds [ELM25, HBGV^{+}21]. birth [LT23, PLM24, SHK20]. birth-death [LT23, PLM24, SHK20]. Bistability [JCLS^{+}11, JCLS^{+}21, Koc20, SOKZ22, GBF20]. bistable [BW25, DB22a, FK21, LS25a]. bitter [BSS^{+}20]. bitter-off [BSS^{+}20]. blankets [PRP^{+}20]. bleaching [DCP^{+}22]. blebbing [NNOBS23]. blind [CLM^{+}25]. bliss [CK20b, SW22]. blobs [Ost23]. blockade [SF23, TKRA20]. blockers [DCL23]. blocking [BS21]. blood [AG22, CHT22, Gre25, NM23, SVHM^{+}21, TCH21, TH25]. blue [MPH^{+}22]. Board [Ano20y, Ano20z, Ano20-27, Ano20-28, Ano20-29, Ano20-30, Ano20-31, Ano20-32, Ano20-33, Ano20-34, Ano20-35, Ano20-36, Ano20-37, Ano20-38, Ano20-39, Ano20-40, Ano20-41, Ano20-42, Ano20-43, Ano20-44, Ano20-45, Ano20-46, Ano20-47, Ano20-48, Ano21y, Ano21z, Ano21-27, Ano21-28, Ano21-29, Ano21-30, Ano21-31, Ano21-32, Ano21-33, Ano21-34, Ano21-35, Ano21-36, Ano21-37, Ano21-38, Ano21-39, Ano21-40, Ano21-41, Ano21-42, Ano21-43, Ano21-44, Ano21-45, Ano21-46, Ano21-47, Ano21-48, Ano22x, Ano22y, Ano22z, Ano22-27, Ano22-28, Ano22-29, Ano22-30, Ano22-31, Ano22-32, Ano22-33, Ano22-34, Ano22-35, Ano22-36, Ano22-37, Ano22-38, Ano22-39, Ano22-40, Ano22-41, Ano22-42, Ano22-43, Ano22-44, Ano22-45, Ano22-46, Ano23u, Ano23v, Ano23w, Ano23x, Ano23y, Ano23z, Ano23-27, Ano23-28, Ano23-29, Ano23-30, Ano23-31, Ano23-32, Ano23-33, Ano23-34, Ano23-35, Ano23-36, Ano23-37]. Board [Ano23-38, Ano23-39, Ano23-40, Ano24u, Ano24v, Ano24w, Ano24x, Ano24y, Ano24z, Ano24-27, Ano24-28, Ano24-29, Ano24-30, Ano24-31, Ano24-32, Ano24-33, Ano24-34, Ano24-35, Ano24-36, Ano24-37, Ano24-38, Ano24-39, Ano24-40, Ano25r, Ano25s, Ano25t, Ano25u, Ano25v, Ano25w, Ano25x, Ano25y, Ano25z, Ano25-27, Ano25-28, Ano25-29, Ano25-30, Ano25-31, Ano25-32, Ano25-33, Ano25-34]. body [CKF^{+}24, HYM^{+}22, Kuz24b]. Bohr’s [BGGM23]. BOLD [BSC20a, SVHM^{+}21]. bolster [TRKB25]. bone [GADDVG^{+}24, LZ23, SCD^{+}23]. Boolean [AZYF^{+}23, Tha24, ZMKF22]. booster [SFK^{+}23]. boosting [CCF^{+}20]. born [YST21]. borne [HBGW^{+}25, TNW21]. both [vBW20]. bottleneck [NBBR23]. bound [AN24]. boundaries [WA22]. boundary [KMA^{+}21, dGSNK22]. bounded [CLF^{+}22, CCSF21, CN20, FZE22]. Bovine [SYI20, HCJT23, BGB^{+}21]. Brain [GT20, AHS23, AGTJ23, Bar21, FBJR20, FSKG20, GEHL^{+}22, ISC23, dGSNK22, MPA23, MDCK23, RB22, SBSB23, TCH21, ZLZ^{+}25]. branch [BBR^{+}21]. Branching [TIS25, IS20, LMS21, LZW23, RBG^{+}23, SS23, ZWT24]. break [AZYF^{+}23]. breaking [WH21, Wu24]. breakpoint [HVTA20]. breaks [BH24, LR21]. breast [BT22, GHB22, GYY^{+}23, PHL^{+}23, SC23a]. bribery [FPX20]. Bridging [AAF^{+}25]. broad [LYE^{+}24]. broadcast [BYHH20]. broader [MBF21]. Bronchial [DLN20]. broods [KMA^{+}21]. brucellosis [YWWF25]. Bt [BB21b]. bubbles [Hil23]. budding [AMSC20]. Budget [EHA^{+}21, MUMS24]. building [HPG^{+}24]. bulk [HAGADDM24, PLXD^{+}20]. bulk-surface [HAGADDM24]. burden [FGC24, KBW^{+}23, MBM^{+}25]. burdens [SXL^{+}20]. bursting [FVB20]. butterfly [LCMW21]. buzzards [ABC^{+}21]. bystander [ND24].

Ca [CBK^{+}20, KK22a, LYE^{+}24, PO20]. CA3 [MP22]. cabbage [KCW^{+}25]. caffeine [BLP20]. Calcaneus [SCD^{+}23]. calcification [WDA^{+}23]. Calcium [DYS^{+}21, RLL25, SL21, SL23]. Calcium-dependent [DYS^{+}21]. calculate [KHAK21]. calibration [LH20, SSK^{+}25]. Can [SSK^{+}25, BNP^{+}21, BD21, BW25, Dal22, DCP^{+}22, GBF20, GSC23, HSAH20, IIHH21, LS21, LBF24, MK24, MC20b, NNNHA24, PB22, RCM^{+}25, SPLS^{+}23, SMS20, SSR21, TK25, Wad22, ZMKF22, MGCH21]. Canada [AMR^{+}23]. Cancer [GBF20, SF23, AMSC20, ATAJS20, Ala23, AABG20, BBDB22, BS21, BT22, BKPT25, BAG^{+}24, BH20, CCH^{+}24, CHMM23, CWL20, CPM21, DMR21, DCS20, EWK20, EB21, FLR23, FBC23, FSC21, GDLF20, GFL23, GYY^{+}23, HPK^{+}20, HKL25, KV23, KKL20, LF25, LBF24, LG23, LS25b, MHI23, Pat22, PHL^{+}23, PAM22, PGC^{+}22, PJB21, PJPH25, RNP^{+}21, SK21, SOKZ22, SXL^{+}20, SC23a, ST20, SAG21, TMC^{+}23, VVD24, WvdDL^{+}21, YJZ^{+}24, YXW21]. Cancer-induced [GBF20]. cancer-stem-cell-driven [FBC23]. cancers [BE22, WSL25]. Candida [KRG24]. capacitive [YMR24]. capacity [BJB^{+}20, CD22]. capillary [AG22, CHT22, TCH21, Ter23]. capture [KKS23]. captures [KHW^{+}22]. Capturing [ACMT24]. CAR [CF20]. carbachol [MP22]. carbachol-induced [MP22]. carbon [MGZ^{+}21]. carcinogenesis [SBW^{+}22b]. carcinoma [GHB22]. cardiac [DCL23, MCL21]. care [KRG24]. Cargo [YSK23]. cargos [KK22b]. Carlo [LGT21]. cartilage [MFGACST23]. Cas [SSS21]. cascade [Han23]. Case [CKI20, MC20b, WWR^{+}24, CRDD22, FHM20, GMMR22, GDC^{+}25, GW20, LRST22, NAR^{+}20, RBG^{+}23, TRKB25, AGP^{+}23]. cases [KBW^{+}23, LMW21, MEPS22]. caspase [MF24]. caspase-mediated [MF24]. castration [CPM21, VVD24]. castration-resistant [CPM21, VVD24]. catastrophes [Sch18, Sch21a]. catastrophic [Leh21]. catenin [CSF20]. Cathepsin [PD23]. Causal [KP21]. cause [GSC23]. caused [ABvH^{+}20, BCS^{+}22, TWT23, Xie25]. causing [ATP22]. cautionary [TRKB25]. cavity [CKF^{+}24]. CD8 [BCK24, LLS^{+}22, WMHV24]. Cell [GLLP23, KTU22, KIO^{+}25, MI23, New20, WZP^{+}23, ASK20, AVSTS23, AMSC20, ABB^{+}22, BCK24, BE22, BKZ^{+}22, BKPT25, BC20, BM20, Car23, CBH^{+}22, CBH^{+}23, CF20, CLD21, CLSB23, FLO25, FH23, FBC23, GHB22, GO21, GFL23, HGCC22, HPK^{+}20, HCSAB21, HRD^{+}24, HAGADDM24, HLW24, HSL24, IJROTE22, IHO^{+}20, IYL^{+}25, JS23, JFCK20, JHFMP^{+}21, KPH^{+}24, KJM20, KF20, LGT21, LSWVB24, LSZ21, Lei20, Lei24, LLS^{+}22, LGSR21, LG23, LH20, MCO21, MPC^{+}24, MCP20, PD23, PPFB24, PCLL21, PJVR23, PASO21, PK23, PGC^{+}22, SPLS^{+}23, SGJ^{+}24, ST20, SWVW25, SSKS18, SSKS20, SBW^{+}22b, TCMA21, VG25, VGC22, WH21, WCL^{+}25, WA22, WvdDL^{+}21, WMHV24, WSEK23, WHH22, XZW^{+}24, XZZ^{+}23, YWQX23, YL25, YFZ^{+}20]. cell-based [JFCK20]. cell-cell [GHB22, PJVR23]. Cell-center-based [MI23]. cell-cycle [CBH^{+}22, CBH^{+}23]. cell-fate [YFZ^{+}20]. cell-matrix [IHO^{+}20]. cell-niche [PASO21]. cell-nutrient [ABB^{+}22]. cell-to-cell [CF20]. cell-type [HSL24]. Cells [Haj20, AABG20, BLRV21, CWL20, CGG^{+}25, DLSA24, EWK20, FMO23, FVB20, GHB22, HYM^{+}22, IHO21, KV23, KHHI25, KIO^{+}25, KKL20, MF24, MDS25, NSZ^{+}25, PWC^{+}23, PJPH25, RS24, RY22, SK21, SC23a, UIY22, XSH^{+}24]. Cellular [LSZ21, XSH^{+}24, BS25, Di 20, DEM^{+}23, FB23, HFD23, HCR20, LSWVB24, LPSMCP23, Osb24, Pru25, TCMA21]. cellulose [CLD21]. center [MI23]. central [BB21a, CRGGSC20, MSM23]. central-place [MSM23]. cerebral [CHT22, TCH21, Ter23, TH25]. Cerebrospinal [YTW21]. cerevisiae [San21]. cervical [CWL20]. Chad [EHP^{+}21]. chain [ACP20, DHC25, FPGC21, GCC20, SSS21, SLCC22]. challenges [HH20]. Change [KPP20, DMPIG24, MGFS^{+}20, RT24b]. changes [APR^{+}22b, CD22, CRG^{+}23, DWNW24, Ter23, Wor20]. changing [AWC25, Kau23, MY20]. Channel [EE25, DCL23, JHFMP^{+}21, RLL25]. channels [LYE^{+}24]. Chaos [FNGJ21, Kan21, YG25]. chaperones [GCNZ^{+}21]. characterises [KGP^{+}22]. characterising [OGD^{+}20]. characteristics [HPSA21, HPSA23]. characterization [MDCK23, SVVPPR^{+}24]. Chargaff [For24]. chattering [BAU^{+}22]. cheap [CS22]. Check [GHF22, ZMKF22]. checkpoint [SF23, TKRA20, XZZ^{+}23, vdVPW^{+}22]. chemical [KJV20, PPGB20]. chemicals [OYT25]. chemo [WCZL25]. chemo-switch [WCZL25]. chemoimmunotherapy [KV23]. chemoinformatic [OSAR21]. chemomechanical [Xie20]. chemoresistance [WvdDL^{+}21]. chemostat [CG21]. chemotactic [VKS^{+}20]. Chemotaxis [Pro24, FH23, GDd^{+}21, TML20]. chemotherapeutic [MPA23]. chemotherapies [HPK^{+}20]. chemotherapy [AVSTS23, LF25, PV24, WCZL25]. chickens [XYK^{+}20]. childhood [ASL^{+}24]. children [AGTJ23, Pag25]. China [APR^{+}22b, BGM22, CZN^{+}21, LMW21, RLW^{+}21, SXL^{+}20, XDT^{+}20, ZZ22, ZXLH23, ZWW22]. Chinese [YSP^{+}23]. chip [BAG^{+}24]. Chlamydia [AMA20]. chloride [UIY22]. chlorophyll [SC21]. Choice [RGBD25, WMO23, AZYF^{+}23, BCBP25, EBA21, EWS23, NGKB21, San21]. cholera [GDC^{+}25, TTBL21]. chondrocyte [CNK24]. choose [ZMKF22]. chorus [AKHM23]. Christmas [Hil23]. chromatin [FZE22, MB22]. chromatin-mediated [MB22]. chronic [CWR21, DOF25, DEM^{+}23, XML22]. cicada [KMA^{+}21]. cilia [PTK^{+}24]. circadian [ACD20, ADP21, BCK24, DB22b, GK20, GMS20, HLS^{+}25, KMI23, KTU22, ZG21]. circle [TH21]. circuit [KK22a]. circular [ACMT24, GW21]. Circulating [TDRC23, CGG^{+}25]. circulation [BAJ^{+}21, Han23]. citrus [ILGI20]. class [Lio18, Lio24, San21]. class-structured [Lio18, Lio24]. classification [BKP23, EE25, HXK^{+}20, YWWF25]. clearance [BJB^{+}20, Swe20]. climate [JAW25, WXS25]. climate-sensitive [JAW25]. climates [AWC25]. climatic [APR^{+}22a]. climbing [Sac22]. clinical [DHC25, Han23, PGC^{+}22]. clock [ACD20, ADP21, GMS20, HLS^{+}25, RK22b]. clonal [BV24, PASO21]. clones [BC20]. closed [IHG21, YWTC25]. closed-loop [YWTC25]. Closing [FG21]. clot [Gre25]. club [VB21]. cluster [VGC22]. clustered [dDA21]. clustering [FH23, LCYD23, SL24, YG25]. clusters [EP20]. Co [BW25, KHBS20, MGZ^{+}21, PAV21, GDd^{+}21, KT20b, Ost23]. Co-adaptive [PAV21]. Co-evolution [BW25, KHBS20]. co-evolutionary [KT20b]. Co-inoculation [MGZ^{+}21]. co-option [GDd^{+}21]. coagulation [IB21, SSKP24]. coalescent [CLF^{+}22, KLC24, RK22b]. coconut [CGY^{+}23]. code [Di 21, LLJ25, RP21b]. codes [GW21, PMCLC^{+}23]. coding [FSNS20, GW21]. codon [San21]. coefficients [SC23b]. Coevolution [BG21, AL23]. Coevolutionary [ABvH^{+}20, KRSG24, MWO21]. coevolving [BA24]. Coexistence [YST21, BP20, Cai22, DTLS24, ES20, GZ24, MMG21, PMD21]. cognitive [Par22, VA22, Wal20]. cognitive-evolutionary [VA22]. coherent [Wad22]. cohort [ATH^{+}24, KBD^{+}23]. coincidence [BFD25]. coinfection [LMG22, ND24]. coli [BF20, VKS^{+}20, YRMWT20]. collagen [BC21]. collagen-based [BC21]. collapse [BB25, ZM20]. collaterals [RKCvB25]. collect [HZS^{+}21]. collective [BEBA20, FLO25, KF20, RKP20, Vol24]. Colombia [VMRSL23]. colonies [GCA^{+}20, HDM20, MHI23]. Colonization [WPEP22, Cai22, LNT23]. colony [GTGB21, HDM20, KGP^{+}22, LSWVB24, ZM20]. colony-level [HDM20]. Color [RG21]. colorectal [HKL25, TMC^{+}23]. colored [YFZ^{+}20, YM23]. columns [Ost23]. combination [ATH^{+}24, BCGG21, CWL20, LBF24, SBM^{+}20]. combinatorial [SHK20]. combinatorics [GPDW20]. Combined [HIA^{+}22, BFG21, FPS24, GCNZ^{+}21, RMS^{+}25, SLML20, ZZGG21]. Combining [BB21b, LF25]. commensal [ATP22]. commensalisms [SLCC22]. Comment [RB22, For25]. commitment [LPD23, Qui20]. Common [Di 20, RP21b, FBA21, SdAM24, Sin21]. common-pool [FBA21]. commons [KHST21]. communication [GGG21, XSH^{+}24]. communities [BSC20b, CCA20, Cor21, Cor22, FGES24, FBD23, LW20, NV22, RK22a, Sin21]. community [CPEK23, MDF21, SYH20, SCH24, Wor20]. Comparative [LML24, PZ20]. compared [FBA21, SWK^{+}20]. Comparing [GTGB21]. Comparison [Mit20, SH21, AJR^{+}20, BPM^{+}22, HD22, Jac25, KHB^{+}22, LDaH22, ZBT22]. Compartment [FBMG25, AZdV^{+}22]. compartmentalization [BNPL20]. compartments [GGSTS21]. compensate [MGCH21]. compensatory [KHW^{+}22]. Competition [BCGG21, KM22, SS21, BHW^{+}20, CWR21, HZM^{+}21, JME24, KHST21, Kuz24a, MMG21, MY20, NNG20, PPFB24, WDH^{+}20, WK25, Cai22, CLJ^{+}23]. Competitive [CCA20, ECBR20, WSC^{+}22, ASF22, Gav25, RK22a]. competitors [GQG23]. completing [ZYLW21]. complex [ASF22, BD22, CLM^{+}25, DEM^{+}23, MLG21, MSB24, TMTK22, YDS24, ZF21]. complexities [BRN^{+}22]. complexity [GPDW20, RR24]. compliance [ABOR20, For24]. component [KHBS20]. components [LPP^{+}25, ORW^{+}20]. composed [BKZ^{+}22]. composition [OLF^{+}24, UIY22, Wor20]. compositional [ZX22]. compositions [KJV20, YJZ^{+}24]. compounds [IHS25]. comprehensive [BJJM25, KR22]. compressible [LGBG23]. compressive [DH22]. Computation [DMAC21, Akh22, Akh23, CSHH20, LZW23]. Computational [BSC20a, BAG^{+}24, EPF20, KLLL22, MFGACST23, ASK20, Avv23, BSS^{+}20, Car23, CWT^{+}23, EP20, GBF20, HS20, HGZM25, HLS^{+}25, KKS23, NMC^{+}21, OKT^{+}21, RKCvB25, SKS21, WBS20]. compute [Wad22]. concentration [KK22b]. concept [SCD^{+}23]. conceptual [ZWC^{+}22]. concern [CWT^{+}23, CSDH^{+}25]. concerns [SUA^{+}20]. Conclusions [DN22]. condition [BSV^{+}24, SLCC22]. conditional [Qui20]. conditioned [MGVS21]. Conditioning [CLF^{+}22]. Conditions [MMI21, NO25, SSR21, AZdV^{+}22, BW24, CLSB23, DMAC21, Kuz24c, dGSNK22, LL21]. cones [ABK24]. configuration [SWVW25, YdKdJ^{+}20]. conflict [CT22, GCX^{+}20]. conformation [ZLYL20]. Conformity [LDF23]. conjugation [AZdV^{+}22]. Connectance [Bau20]. connected [Car23, LPP^{+}25]. Connectivity [NNNHA24, RNP^{+}25]. Connectome [GT20, FSKG20, RB22]. connexins [XSH^{+}24]. consciousness [RBW20, RSPB^{+}22, Ste21a]. consensus [PP22]. consequences [GGSTS21, Leh21, NM23, YJBF20]. Conservation [Wor20, ACMT24, Tsu21]. conserved [GHA^{+}20]. consideration [AN24, NVS20]. considering [BGA21, SGJ^{+}24]. Consistency [BK20a]. consolidation [ASL^{+}24, MB24]. constant [AN24, BBR^{+}21]. constrained [Ros21]. constraint [CP23, LB20]. constraint-based [LB20]. constraints [BK20b, FSNS20]. constricted [Don23]. constructing [PZG20]. construction [KPR20]. consumer [MRK24, Wor20]. consumer-resource [MRK24, Wor20]. consumption [GBRM21]. Contact [AHEB^{+}22, KvLBD25, AH25, BGM22, GHB22, YSK23]. contact-network [AH25]. contacts [ZWT24]. contagious [GJCLMQC20]. containment [BK21]. contaminated [CZN^{+}21]. contamination [KHAK21]. contemporary [Hon20]. content [ACP20, BRBF20]. contest [MGCH21]. context [CCPD22, MBF21]. continuous [CCH^{+}24, DSB^{+}21, JCC20, LMB22, OTI^{+}20, SKNvRL23]. continuously [AG22]. continuum [ACF22, BCC^{+}21, FSC21, GTGB21, GF23, MY20, SBB^{+}22]. contraction [Avv23, SdB23, SL21]. contrastive [WSL25]. contributes [PZJPZ23]. contribution [HPGF25, PL20]. Control [PTZ21, YZSS20, AHW21, ATH^{+}24, AJR^{+}20, APL21, BAU^{+}22, BGB^{+}21, BCS^{+}22, CGY^{+}23, CGW22, CS21, CR20, FPGC21, FSK21, HCSAB21, IMM24, JTB25, KHW^{+}22, KYI21, KRG24, Leh22, LZL^{+}25, LTB25, MB21, MSL23, MCM20, MB22, NVS20, Ngo22, NDB20, PBP23, RSPB^{+}22, Sal25, SMJ20, SJS^{+}25, SCD^{+}23, SUA^{+}20, TTBL21, TNW21, VMRSL23, ZRL^{+}21, ZBT22]. controlled [TAP^{+}23]. controller [GFN23]. controlling [CPMF^{+}22, ZZB^{+}21]. controls [AGP^{+}21, ACD20, ASMN23, SBM^{+}20, TDRC23, YCKK24]. Convergence [MPA24]. conversion [PGP23]. convolutional [WDT^{+}20]. Cooperation [SSBG20, BRS24, CKB^{+}21, CSVB25, CS22, EBA21, KL21a, KCO25, Kro25, Kur21, LL21, ML24, PHN^{+}23, SdAM24, YTG24]. Cooperative [BSC20b, Gre25, NCL20, RP21a, FZE22]. cooperativity [NWVN20]. coordinate [DAB^{+}23]. coordinated [HBG21, KCO25]. Coordination [WCD21, FLO25]. coral [DCP^{+}22, HPG^{+}24, KGP^{+}22, RCM^{+}25, ZYF25]. coral-dwelling [DCP^{+}22]. Corallium [KGP^{+}22]. corals [WDA^{+}23]. cord [LSZ21]. core [BAG^{+}24, KBS^{+}19, KBS^{+}20, KB22, YST21]. core-shell [KBS^{+}19, KBS^{+}20, KB22]. corneal [HCJT23]. coronavirus [YY21, Yin21, OIK^{+}24]. correction [DPR21]. correlated [BS22b, YM23]. correlation [LCYD23, TK25]. correlations [PR21, vBBB^{+}25]. Corrigendum [CBH^{+}23, Cor22, D’A20a, GHS20, HPSA23, JCLS^{+}21, KBS^{+}20, Lei24, Lio24, LK25a, PP24, Sch21a, Sch23, WLW20, Xie22]. cortex [HGZM25, Ost23]. Cortical [LRAP22, CGK23, CRGGSC20, SCD^{+}23]. Cost [CPHL^{+}20, FHM20, KT20a, ASF22, KRG24, LB20, OGM^{+}25, PPS^{+}21, WM21b]. cost-effectiveness [KRG24, OGM^{+}25]. Cost-efficiency [KT20a]. Costly [Mur25]. costs [CK20b, GFAP^{+}25, KC22, KC24]. could [BC20]. count [GYY^{+}23, SMM24]. Counterintuitive [NKE23]. countermeasure [SSS21]. Counting [SS23]. countries [KRSSH21, LZZ^{+}23]. counts [ABBP21, HRD^{+}24]. County [REK20]. Coupled [KB22, MG21, SZR^{+}25, ABB^{+}22, BGGM23, IH23, JAB22, MCMdAM24, MB24, PZXC22, PLXD^{+}20, Sak22a, YMR24]. Coupling [TTBL21, BSC20a, CGK23, PZXC22, SL21, SHY21, SA25, DO21, YI21]. couplings [Xie20]. courtship [CK20b]. CoV [ACF22, BLRV21, BFSA23, CWF23, CLJ^{+}23, CPEK23, CWT^{+}23, CSH^{+}25, GFV^{+}23, GFC^{+}22, Hil23, JMWX23, KvLBD25, KRSSH21, Kee23, KBD^{+}23, LRST22, NSZ^{+}25, PD23, PWC^{+}23, PDD24, Roy21, TAP^{+}23, WUH23, WHRH^{+}23, XCFH23, YY21, Yin21, dCC^{+}23]. covariation [HGD^{+}25]. COVID [AGP^{+}21, AGP^{+}23, APF^{+}24, ASS^{+}23, AMR^{+}23, BW24, BK21, BGM22, BDdG22, CS21, CSDH^{+}25, DAB^{+}23, DTS21, DPR21, DJ22, ECH^{+}23, FPGOCP^{+}25, GSW^{+}23, GGSTS21, HPGF25, JKH^{+}23, KPA^{+}21, KBW^{+}23, LPLM25, LMS21, LZZ^{+}23, LMW21, MEPS22, NGKB21, PRJ^{+}20, REK20, STCSC^{+}21, SFK^{+}23, SRDC21, SXJ20, Täu20, TNHV21, TWT23, YTY^{+}20, YLT22, ZQHN24, ZZT^{+}21, ZWC^{+}22, ZWW22, dIMM24, dCC^{+}23]. COVID-19 [AGP^{+}21, AGP^{+}23, APF^{+}24, ASS^{+}23, AMR^{+}23, BK21, BGM22, BDdG22, CS21, CSDH^{+}25, DAB^{+}23, DTS21, DPR21, DJ22, ECH^{+}23, FPGOCP^{+}25, GSW^{+}23, GGSTS21, HPGF25, JKH^{+}23, KPA^{+}21, KBW^{+}23, LPLM25, LMS21, LZZ^{+}23, LMW21, MEPS22, NGKB21, PRJ^{+}20, REK20, STCSC^{+}21, SFK^{+}23, SRDC21, SXJ20, Täu20, TNHV21, TWT23, YTY^{+}20, YLT22, ZQHN24, ZZT^{+}21, ZWC^{+}22, ZWW22, dIMM24, dCC^{+}23]. cows [HRD^{+}24]. Crabtree [SKS21]. CRAC [LYE^{+}24]. CRAC-mediated [LYE^{+}24]. cracking [HH21]. creole [NO25]. Crick [LLJ25]. CRISPR [SSS21]. CRISPR-Cas [SSS21]. criteria [TAP^{+}23, VB21]. Critical [GSR23, HLW24, IYL^{+}25]. crops [BSJ20, BB21b]. Cross [HL20, FG23, SHY21, Wu24, YAN21, YSP^{+}23]. cross-immunity [FG23]. Cross-inhibition [HL20]. cross-propagation [YSP^{+}23]. cross-sectional [YAN21]. crossing [RKP20]. Crosstalk [GMS20, SMS20]. crowders [Xie24]. crown [dAPdBMTS20]. Crozier [Sco24]. CRPC [VVD24]. cruise [MBHG^{+}24]. CSF [SF23]. CSF-1 [SF23]. CT [BAJ^{+}21]. cube [Kuz24b]. cues [BH23, MLG21]. Cultural [Whi20]. culture [KYHN23, OTI^{+}20]. curbing [LZZ^{+}23]. cure [MSK^{+}22]. current [AJR^{+}20, KP21]. currents [JHFMP^{+}21, SHY21]. curvature [HCSAB21, LZ23]. curve [DRO23, KYHN23]. curved [HH21]. Cusp [DBS23]. cyanobacteria [KTU22]. cycle [ASSL22, ASK20, AGBLGA^{+}22, CBH^{+}22, CBH^{+}23, DRO23, FHR22, IYY22, KCW^{+}25, PBD^{+}20, SBW^{+}22b, WSEK23, YYI21]. cycles [HSAH20, MK21]. cyclic [BP20, BEO^{+}22, CBH^{+}22, CBH^{+}23, CBPG23, FET^{+}24, NKE23]. cyclical [ZF21]. cycling [YST21]. cylindrical [YSK23]. cyst [EBK21]. cytokine [PJVR23]. cytoplasmic [LHG^{+}22]. cytoskeleton [MDS25]. cytosolic [LYE^{+}24]. cytotoxic [HIA^{+}22].

D [Gar23, ABB^{+}22, BMB^{+}21, FSC21, GF23, MPTCT^{+}20, MMS^{+}23, RBG^{+}23, SRB^{+}24, YG25]. daily [DF22]. dairy [HRD^{+}24]. Daisyworld [MC22]. damage [FK21]. damicornis [HPG^{+}24]. dark [FFNG^{+}21, MK24]. Darwin [CGS^{+}25]. Darwinian [BH20, NBBR23]. Data [AGP^{+}23, ASL^{+}24, BLP20, CF20, PHL^{+}23, Sim25, BNP^{+}21, BMB^{+}21, DF22, DO21, GWZ^{+}22, GMR^{+}24, GYY^{+}23, GMV^{+}20, HKL25, HRD^{+}24, KvLBD25, LGT21, LPP^{+}25, LTB25, MGVS21, MEPS22, RZRS^{+}22, SKNvRL23, SMM24, SXJ20, YAN21]. Data-based [PHL^{+}23]. Data-driven [AGP^{+}23, ASL^{+}24, BLP20, CF20, GWZ^{+}22, HKL25, LTB25, RZRS^{+}22]. Data-informed [Sim25]. datasets [ZQHN24]. date [CLF^{+}22]. dawn [MB21]. dead [SKNvRL23]. death [LT23, LSZ21, PLM24, SHK20]. decay [DHC25, ECBR20, SGM^{+}25]. deceased [HL21]. December [Ano20o, Ano20c, Ano21o, Ano21c, Ano22n, Ano22c, Ano24k]. decentralized [Pro24]. deception [QHS23]. Deciphering [CGG^{+}25, MLK21]. decision [Bai21, BBDB22, HLW24, JCC20, PP22, VKS^{+}20]. decision-making [JCC20, PP22]. decisions [AJR^{+}20, MSM23, OIK^{+}24]. decline [BCS^{+}22]. Decoding [LS25b, RG21]. Decomposing [ZQHN24]. decomposition [dAPdBMTS20]. decreased [KHHI25, PD23]. dedifferentiation [FB23]. Deep [MDCK23, LFZ^{+}24, WDT^{+}20, YCKK24]. deeper [Pie21]. deer [XML22]. defection [Mur25]. defects [CNK24]. defense [IHS25, PPS^{+}21]. deformable [ZLZ^{+}25]. deformation [MI23, Tsu20]. degeneration [PLSS24a, PLSS24b]. degraders [JME24]. degree [WDA^{+}23]. Delay [CSF20, BJB^{+}20, FPGOCP^{+}25, HKL25, Tak24]. Delay-driven [CSF20]. Delayed [ES24, YST21]. delays [BLRV21, BCGG21, FBMG25, GC20, MSLW23]. deleterious [AL23, Sak22b]. deletion [HD22]. delineating [YJZ^{+}24]. delivery [ASP23, Ala23, BKZ^{+}22, FFGR25, GGS^{+}24]. delta [CWT^{+}23]. dementia [HD24]. demes [Kro25]. demographic [APR^{+}22b, BJR^{+}22, BPS25]. demography [Hen20]. dendritic [EWS23, Ros22, XZZ^{+}23]. dengue [CAARG21, GW20, RY22, VCS24]. Denise [CSW24]. Density [EHA^{+}21, vBW20, EP20, GJDS20, GSC23]. density-area [GSC23]. density-dependent [EP20]. dependence [BKP23, GB23, IKYW21, KHHI25, RWHK21, vBW20]. dependency [Kuz24b]. dependent [ABK24, BH24, CWL20, DCL23, DYS^{+}21, EP20, EHA^{+}21, FBMG25, GJDS20, KT20a, LRAP22, LR21, LWM23, NCL20, OLF^{+}24, RS24, SD20, SBC^{+}20, SS21, WSW21, WXS25, WSY21]. depends [IHO21]. depolarization [CGK23, Car23, WM21b]. depolymerization [YL25]. deposited [BH23]. deposition [CLD21, HMS21, Kuz24a]. depression [CRGGSC20, Sak20, Sak22a]. deprivation [PBB20]. depth [LRAP22, RLB25]. depth-dependent [LRAP22]. deregulated [JS23]. derived [POS21]. deriving [Ros21]. dermatitis [PJVR23]. described [PLAC22]. describing [AABG20, BRBF20, DSM^{+}23, PP20, PP24]. description [CK20a, Han23]. descriptions [SBB^{+}22]. descriptors [MPTCT^{+}20]. Design [PRJ^{+}20, SGM^{+}25, Ala23, BCLJ24, BH20, LTS^{+}21, MBT22, TRKB25]. Designing [CLSB23, SBM^{+}20]. Destabilization [PS22a]. destabilizes [CSVB25]. destabilizing [BYHH20]. destination [SKNvRL22]. destruction [DRK23]. detachment [SK23]. detailed [DN22]. detectability [Mur25]. detecting [CRG^{+}23]. detection [BH24, MBM^{+}25, PCR23, PTK^{+}24, RMSS24, Täu20, Wad22, vdBMB^{+}23, BDB25]. determinant [AM22, CWW22, MMI21, MMI22]. Determinants [HPWvdB20]. determine [MBT22]. determined [FMO23, Kuz24a]. determines [LZZ20, PMD21]. determining [BB21a, dAPdBMTS20]. Deterministic [JHFMP^{+}21]. deterministically [MY20]. develop [MGCH21]. Developing [GDC^{+}25, KV23]. Development [BBB^{+}25, KKS23, LR21, AML^{+}22, CSP21, DSM^{+}23, DNK^{+}24, GJCLMQC20, IYY22, JA23, Kuz24c, Osb24, Ost23, PV24, Pru25, SS22]. Developmental [BPM^{+}22, WCD21]. device [BAG^{+}24]. diabetes [DNK^{+}24, SF24b, Swe20]. diabetic [SF24b]. diagnostic [VGTY21]. diameter [RKCvB25]. diameters [Ter23]. diamond [ASMN23]. diamond-shaped [ASMN23]. Diarrhea [BGB^{+}21]. dictates [RMG^{+}21]. did [KI21]. diel [BEBA20]. diet [Swe20, VTDM21]. diet-linked [VTDM21]. difference [HHT25]. differences [BLRV21, LB25, SL23, WHH22]. different [Car23, CMLM23, CSH^{+}25, GCM23, HXS21, KSSM20, TG22]. Differential [BGM22, CSP21, DSB^{+}21, HKL25, HH20, JCLS^{+}11, JCLS^{+}21, LCYD23, MMS^{+}23, RMS^{+}25, SMM24, TP24, VGC22]. differentiating [BC20, FBC23]. differentiation [CN21, Haj20, New20, SAG21, SS22]. differing [LZ23, MGZ^{+}21, SPLS^{+}23]. diffuse [PGC^{+}22]. diffuse-interface [PGC^{+}22]. Diffusion [BS22a, BNP^{+}21, CGW22, EIK^{+}21, GWZ^{+}22, GJDS20, ISC23, JMD20, KG22, KGP^{+}22, KBS^{+}19, KBS^{+}20, KB22, MTB24, MDCK23, PLXD^{+}20, PLAC22, Sch18, Sch21a]. diffusive [GDd^{+}21]. diffusivity [Kuz24c]. digestion [Avv23]. digital [HIA^{+}22]. dilemma [FS24, KT20b, KL21a, MMI22, WLW19, WLW20]. dilemmas [CSVB25, CS22, FPX20, ML24]. dimension [LPSMCP23]. Dimensional [AFTF22, ABBP21, BRN^{+}22, BC21, EP20, FET^{+}24, MI23, WHH22, YL25]. dimensionality [PMD21, WK25]. dimensions [HAGADDM24]. dimerisation [WRB22]. dimorphism [YYI21]. dinosaurs [KI21]. dip [HZS^{+}21]. Dipole [ZPP23]. Dipole- [ZPP23]. direct [Goy21, IYY22, KHB^{+}22, KP21]. directed [WA22]. directional [JMD20, ORW^{+}20]. disadvantage [MGCH21]. discharge [SCHH20]. Disclaimer [KCS20]. discontinuation [HHI24]. discount [TATT21]. discourages [MW20]. Discovering [WXS25]. discovery [LMC^{+}20]. Discrete [KPH^{+}24, ACF22, LHG^{+}22, LTS^{+}21, MC20b]. discrete-continuum [ACF22]. Discrete-state [KPH^{+}24]. discretisation [RGBD25, WMO23]. discrimination [RG21]. Disease [dDA21, AHEB^{+}22, AH25, CPHL^{+}20, CRG^{+}23, DHR25, DEM^{+}23, EHP^{+}21, FGC24, GSR23, GCA^{+}20, HPSA21, HPSA23, HHT25, IMM24, JKSD23, KHW^{+}22, LX21, LWM23, LLZ^{+}22, MBHG^{+}24, OGM^{+}25, OIK^{+}24, PGP23, RNP^{+}25, RCM^{+}25, RGBD25, RLW^{+}21, SYI20, SXL^{+}20, SL22, SA25, SY24, SOGK^{+}23, SYH20, SCH24, TT20, TATT21, WXS25, XML22, YDS24, ZWT24]. diseases [AN22a, GBHNHVL23, GJCLMQC20, HI20, KT20a, KPH^{+}24, KHB^{+}22, MD22, MN23, TNW21, YFF25]. Disentangling [KYHN23, LMG22]. disorder [DB22b, HBK^{+}20]. disorders [YTH^{+}21]. disparate [SK21]. Dispersal [RT24b, AR20, BERR20, FPS24, POW23, ZYF25]. dispersal-induced [AR20]. dispersed [D’A20a, D’A20b]. dispersing [ZZB^{+}21]. dispersion [NO24]. displacement [LMLP20, TH21]. displays [LPD23, SBJR21]. disposable [RR24]. disruptive [ORW^{+}20]. dissemination [LX21]. dissipative [LFY^{+}24, WFL^{+}24]. dissociation [AN24]. distance [COB23]. distances [PZXC22]. distancing [BD21, CS20, CS21, CC22, GGSTS21, LPLM25]. distinct [DB22a, Mit20]. Distinguishing [SSM^{+}22]. distorted [IS20]. distorting [JTB25]. distributed [AG22, ISC23, SD22, YDS24]. distribution [APR^{+}22a, Coh20, D’A20a, D’A20b, DHC25, GMV^{+}20, HI21, KJV20, MGVS21, MK24, MMS^{+}23, RK22a, KC24]. Distributions [CKTW24, ABC^{+}21, FVC20, GMR^{+}24, HTD22, KG23, PZ20, PLM24, PTZ21, SSR21]. disturbance [HDM20, PRHS20]. disturbance-prone [PRHS20]. disturbances [Bau20, BGA21, GSMG22]. divergence [BBR^{+}21, HGF23, Pie21, PLR24]. diverse [RWHK21]. diversification [Nak25, QVE25, RID23]. diversifies [PJB21]. diversity [ACMT24, FOY21, Kau23, MSB24, PB22, SYH20, TGC22, TIS25, Whi20]. division [KTU22, LSZ21, XZW^{+}24, Yam24, YWQX23]. dMMR [HKL25]. DNA [CBH^{+}23, ABC25, BH24, CBH^{+}22, CK20a, DKSP24, FK21, Goy21, IB21, LAP24, RP21b, Van20, YG25]. DNA-structured [CBH^{+}23, CBH^{+}22].

Do [BBR^{+}21, AS21, EBA21, HPSA21, HPSA23, LSW20, MTB24, SKNvRL22]. Does [Jam21, HD25, LDL^{+}24]. DOM [OTI^{+}20]. domain [BSV^{+}24, WHH22]. domains [Di 20, EP20]. domestication [ASF22]. Dominance [HPD20, BP20, CCA20, CBPG23, OU22, Ost23, PASO21]. dominant [GC20]. dopamine [ZSL23]. dormancy [BKPT25, MSLW23]. dose [BCBP25, CLLL20, CWL20, CKM^{+}20, Wal20, XCFH23]. dose-dependent [CWL20]. dose-response [CKM^{+}20, Wal20, XCFH23]. double [AZYF^{+}23, BH24, YMR24]. double-strand [AZYF^{+}23]. double-stranded [BH24]. Doubling [AN22a]. Dove [HPD20, KC22]. down [SMJ20]. downshift [LS21]. Downstream [SLCC22]. Drafting [ETTT24]. drift [LSZ21]. drive [DWNW24, LMG22]. driven [AGP^{+}23, ASL^{+}24, BGC^{+}20, BLP20, CSF20, CF20, FBC23, GWZ^{+}22, GZM^{+}20, HKL25, LPA25, LFJ21, LTB25, OYT25, PRFSHV20, RZRS^{+}22, WM21a, YHT24, YTW21, YM23]. drivers [FLR^{+}21, PWC^{+}23, XDZ^{+}21]. drives [FBD23, ON24]. driving [LSZ21]. drop [ETTT24]. Drosophila [JA23, YBF23]. drug [Ala23, BKZ^{+}22, BCGG21, BV24, DHC25, DMR21, GDLF20, HGCC22, HII22, HHI24, MPA23, MS20, PV24, PJPH25, SLML20, SF24b, YHT24]. drug-resistant [PJPH25]. drug-sensitive [PJPH25]. drug-targets [HGCC22]. drugs [RNP^{+}21, UV23]. dsRNA [SBC^{+}20]. Dual [CBK^{+}20, SY25, SWVW25, YTG24]. dual-configuration [SWVW25]. duct [HI21]. ductal [WCZL25]. due [AMR^{+}23, Avv21, BB25, DMR21, FH23, HH21, KHHI25, MMG21, MMK^{+}24, NNG20, SLML20]. Duffing [MUMS24]. dung [BFG21, YZB20]. duodenal [Avv21]. duodenogastric [Avv21]. durability [KT21]. duration [ACD20, PBB20]. during [AGP^{+}23, ASF22, APF^{+}24, AGBLGA^{+}22, BP23, BCGG21, CNK24, FRMK22, GRW^{+}21, IHO^{+}20, JD21, JA23, JKH^{+}23, MSM23, MDG24, Nak25, NNOBS23, ND24, OIK^{+}24, STCSC^{+}21, SdB23, SF24a, SS22, SCH24, SXJ20, TEWN21, TWT23, VGC22, ZWC^{+}22]. dwelling [DCP^{+}22]. Dynamic [AML^{+}22, GT22, SYI20, AW22, BD22, ELRG21, FSKG21, GO21, KV23, LB20, MH22b, NGKB21, ORT24, SHH20, HZS^{+}21]. Dynamical [CMRMC^{+}21, IKYW21, AKHM23, FNGJ21, JWKA21, NWVN20, YHT24, YJBF20, ZWW22]. Dynamics [BM20, D’A20b, GCNZ^{+}21, GCA^{+}20, HSL24, Koc20, MCL21, MHPB23, PPS^{+}21, SFYL20, SHK20, SD25, TKR^{+}21, WSY21, Xie21, Xie22, ZYF25, AVSTS23, AGP^{+}21, AR20, AMA20, ATAJS20, ATH^{+}24, ABB^{+}22, APF^{+}24, AKUN20, ATP22, AM22, AR21, AHEB^{+}22, BCC^{+}21, BW24, BE22, BS21, BYHH20, BRAHV25, BK21, Bol22, BGGM23, BGA21, CD22, CAARG21, CBZK21, CLM^{+}25, CWW22, CFH^{+}25, CKB^{+}21, DDDMT22, DKSP24, DBW22, EBK21, FG22, FBA21, FBMG25, FFNG^{+}21, FSKG20, FPS24, FRMK22, GJDS20, GRPG24, GDC^{+}25, GBHNHVL23, GK21, Gre25, GMS20, GDQR23, HPK^{+}20, HC24, HGD^{+}25, HYM^{+}22, HBGV^{+}21, HBGW^{+}25, HL21, IH23, KPA^{+}21, KHW^{+}22, KBD^{+}23, KM22, LS21, LMG22, LNT23, LHG^{+}22, LK20, LZZ20, LX21, LMC21, LWM23, LFT^{+}23, LFY^{+}24, LS25b, MP22, MLK21, MUMS24, MSB24, MDG24, MCM20, MSK^{+}22, MTLZ22, MM22, MG21, MEBNZ24]. dynamics [MC20b, MWO21, NCL20, NWP^{+}24, PR21, POS21, PABS23, PO20, PK23, QWZK24, RB22, Rev25, RDDSA25, RBG^{+}23, RLB25, RLL25, SPLS^{+}23, SK21, STB20, SF24a, SSX^{+}22, ST20, Sin21, SA25, SS21, SH21, SZR^{+}25, SBW^{+}22b, TMTK22, TEWN21, TGC22, TT24, TK25, UV23, VGTY21, VP22, Wal20, WFL^{+}24, WRB22, WDH^{+}20, WK25, Xie24, XKV20, YTG24, YSP^{+}23, YM23, ZWWW23, dDA21, D’A20a].

E2F [LMA21, LS25b]. E2F/Myc [LS25b]. Early [BDB25, APF^{+}24, ASL^{+}24, CSP21, GTE20, GZGH22, JLK22, KGP^{+}22, NSD22, PB21, RG21, Sak20, Sak22a, Sak22b, SJNRS20, SD22, VGC22, YTY^{+}20, vdBMB^{+}23]. early- [Sak20, Sak22b]. early-acting [Sak22a]. earth [RBD^{+}20]. eating [DMPIG24]. Ebola [BPP^{+}20]. Eccentricity [SCD^{+}23]. eclipse [KBD^{+}23, PWC^{+}23]. Eco [CKB^{+}21, CBPG23, PABS23, CN20, CN21, GRPG24, POS21, TT24, WK25]. Eco-evolutionary [CKB^{+}21, CBPG23, PABS23, CN20, CN21, GRPG24, POS21, TT24, WK25]. Ecological [GSC23, Leh21, BB25, Bol22, EWS23, FVC20, JAB22, NV22, PPGB20, PLR24, SCH24, SH21, Tak24]. ecology [FFNG^{+}21, SAG21, WNHK21]. ecosystem [ACP20, BGA21, MBT22, PAV21, QWZK24]. ecosystems [BvdG20, ES20, FGES24, SMJ20, XDZ^{+}21, ZSS^{+}22]. edge [PLSS24a, PLSS24b, SSR21]. editing [NV22]. editor [PMCLC^{+}23]. Editorial [Ano20y, Ano20z, Ano20-27, Ano20-28, Ano20-29, Ano20-30, Ano20-31, Ano20-32, Ano20-33, Ano20-34, Ano20-35, Ano20-36, Ano20-37, Ano20-38, Ano20-39, Ano20-40, Ano20-41, Ano20-42, Ano20-43, Ano20-44, Ano20-45, Ano20-46, Ano20-47, Ano20-48, Ano21y, Ano21z, Ano21-27, Ano21-28, Ano21-29, Ano21-30, Ano21-31, Ano21-32, Ano21-33, Ano21-34, Ano21-35, Ano21-36, Ano21-37, Ano21-38, Ano21-39, Ano21-40, Ano21-41, Ano21-42, Ano21-43, Ano21-44, Ano21-45, Ano21-46, Ano21-47, Ano21-48, Ano22x, Ano22y, Ano22z, Ano22-27, Ano22-28, Ano22-29, Ano22-30, Ano22-31, Ano22-32, Ano22-33, Ano22-34, Ano22-35, Ano22-36, Ano22-37, Ano22-38, Ano22-39, Ano22-40, Ano22-41, Ano22-42, Ano22-43, Ano22-44, Ano22-45, Ano22-46, Ano23u, Ano23v, Ano23w, Ano23x, Ano23y, Ano23z, Ano23-27, Ano23-28, Ano23-29, Ano23-30, Ano23-31, Ano23-32, Ano23-33, Ano23-34, Ano23-35, Ano23-36, Ano23-37]. Editorial [Ano23-38, Ano23-39, Ano23-40, Ano24u, Ano24v, Ano24w, Ano24x, Ano24y, Ano24z, Ano24-27, Ano24-28, Ano24-29, Ano24-30, Ano24-31, Ano24-32, Ano24-33, Ano24-34, Ano24-35, Ano24-36, Ano24-37, Ano24-38, Ano24-39, Ano24-40, Ano25r, Ano25s, Ano25t, Ano25u, Ano25v, Ano25w, Ano25x, Ano25y, Ano25z, Ano25-27, Ano25-28, Ano25-29, Ano25-30, Ano25-31, Ano25-32, Ano25-33, Ano25-34, CKI20]. educational [Hon20]. EEG [RNP^{+}25]. Effect [AAL^{+}21, Bai21, EE23, HFD23, Kau23, Kur21, KH24, WvdDL^{+}21, Xie24, YS25, ZX22, AHS23, AHEB^{+}22, BP20, BS21, BT22, BCBP25, CHT22, CHSB20, CPM21, EWS23, EE25, FFR23, FPS24, GFAP^{+}25, KL21b, KKK^{+}20, LR21, Leh21, LL21, LG23, MGZ^{+}21, ND22, NBBR23, ND24, PS22a, SKS21, Sak20, STB20, SP20, TGC22, TCH21, YM23, ZM20, ZF21, RMS^{+}25, ZB22]. Effective [EIK^{+}21, AGP^{+}21, CKF^{+}24, HS20, Täu20, TNW21]. effectiveness [ESW24, FHM20, GBF20, JKH^{+}23, KRG24, LDL^{+}24, MBHG^{+}24, OGM^{+}25, ZBT22]. effectors [IHO21]. Effects [BK21, EBK21, FBC23, GCA^{+}20, KCW^{+}25, KB20, Sak22a, SL21, Tak24, YdKdJ^{+}20, ACP20, BvdG20, BBH^{+}20, BH24, BYHH20, BGGM23, CZN^{+}21, Cai22, CF20, CBKS21, DMAC21, DCL23, EWK20, ECBR20, FPGC21, FGC24, HAP22, HKBK25, HS24, HGD^{+}25, Hen20, HD24, LRST22, MP22, MH22a, MSK^{+}22, MWO21, OKF^{+}21, PBD^{+}20, QWZK24, Sak22b, TH25, TG22, UV23, VGTY21, WK25, ZZ22, vBBB^{+}25, vdVPW^{+}22, SC21]. Efficacy [DBTI21, BBK23, BS21, GYY^{+}23, HS24, JTB25, PD23, SC23a, SF24b]. efficiency [AKHM23, AM22, KT20a]. Efficient [SA25, BP23, FSNS20, SBB^{+}22, YCKK24]. eigenmodes [RE25]. Eigenvalue [GSMG22]. eight [KRSSH21, LZZ^{+}23]. elasticity [Wu24]. Electric [CSHH20, SCHH20, ZLYL20]. electrical [FVB20, GEHL^{+}22, MCP20, SY25, YZSS20]. electrocytes [CSHH20, SCHH20]. electrolocation [PCR23]. electron [GCC20]. electroretinogram [PLSS24a, PLSS24b]. electrostatic [PCR21]. element [GEHL^{+}22]. elemental [OLF^{+}24]. elementary [CP23]. elements [MDG24]. elicit [KKL20]. eliminate [LDL^{+}24]. Elimination [BSS^{+}20, HMLO^{+}23, SSX^{+}22]. elliptic [HYM^{+}22]. elliptical [YSK23]. elongation [KF20]. embryogenesis [GZGH22]. embryonic [JA23, NTTI22]. emerge [MK24]. Emergence [DLSA24, LPA25, LYE^{+}24, RBD^{+}20, Ala20, CN20, FBD23, HII22, IDD23, JMWX23, KNY25, MRK24, NNNHA24, PPFB24, PR21, RDDSA25]. Emergent [ADLM24]. emerging [IH23, KHB^{+}22]. emigration [HAP22]. emphasis [SJNRS20]. empirical [FSKG21, ZQHN24]. emptying [Avv21]. enable [GBF20]. enables [ES20]. encode [CKNCG22]. encoding [PR21, Van20]. encounter [MGFS^{+}20]. end [WS22, Xie23, YLT22]. end-plate [WS22]. endemic [ASS^{+}23, MCM20, SLML20]. endgame [ASS^{+}23]. endogenous [NVS20]. endometrial [AGBLGA^{+}22]. endoplasmic [DLSA24]. endoscopic [SXL^{+}20]. endothelial [HYM^{+}22, SJNRS20, VGC22]. endothelium [HCJT23, JJR20]. energetic [YWTC25]. energetics [OKT^{+}21]. energy [AKHM23, AA22, BP21, CKM^{+}22, GCC20, GZM^{+}20, HXS21, ISC23, RZRS^{+}22, TCMA21, WM21b, Wu24]. energy-based [GCC20, TCMA21]. energy-driven [GZM^{+}20]. energy-optimization [Wu24]. enforcement [HL21]. engineered [CWL20, PRJ^{+}20]. England [Hil23, HPGF25, Kee23]. enhance [JTB25]. Enhanced [KMI23, MMG21, FPGOCP^{+}25, YSK23]. enhancement [BCC^{+}21, KT20a]. Enhancing [JFCK20, HS24]. Enlightening [CLM^{+}25]. enough [DWNW24]. enrichment [MSB24, YBF23]. ensemble [CBKS21]. entire [Nak25]. entrainment [DB22b, HLS^{+}25]. Entropy [LL20, LPSMCP23, Tsu21]. entry [PD23, TAP^{+}23]. environment [FOY21, FBSD22, KF20, MY20, RT24b, RBZP24, SS21, UIY22, EBA21]. environmental [Bar23, Cai22, GCA^{+}20, KRG24, PS22a, PBD^{+}20, SUA^{+}20, TG22, PMD21]. environments [CZN^{+}21, HCC^{+}21, Kau23, KHB^{+}22, KL21a, MLG21, MSLW23, NKE23, Pro24, WW24]. enzymatic [TH21]. enzyme [CTL^{+}22, CDdV23, LB20]. enzyme-cost [LB20]. Epidemic [BDB25, LWM23, MM22, AHW21, BP23, BD21, CD22, CGW22, DTS21, HGD^{+}25, HPGF25, HTD22, KT20b, KPR20, LWXM20, LZL^{+}25, LK25a, LK25b, MEPS22, dVRA20, RMG^{+}21, SXJ20, TK25, WCM^{+}21, WHRH^{+}23, XCH^{+}20, YTY^{+}20, YLT22, ZXLH23]. epidemics [CLLL20, CSDH^{+}25, EB24, LMW21, MZG^{+}20, TWT23, VMRSL23, ZWC^{+}22, dIMM24]. epidemiologic [CS20]. epidemiological [AJR^{+}20, AH25, BJJM25, FGC24, GGSTS21, GBHNHVL23, Hil23, IH23, MSK^{+}22, PB21, RDDSA25, TTBL21, WHRH^{+}23, YDS24]. epidemiology [Gav25, RT24a]. epigenetic [HSL24]. epilepsy [PZXC22, YZSS20]. epithelia [MCO21]. epithelial [GHB22, LSWVB24, LS25a, RP21a, YOIS24]. epizootic [PGP23]. equation [BNP^{+}21, IB21, JCLS^{+}11, JCLS^{+}21, Lio18, Lio24, SMM24, VGC22]. equations [DSB^{+}21, HKL25, RMS^{+}25]. equilibria [GZ24]. equilibrium [AMSC20, AABG20, Lio18, Lio24, LW20]. equity [FBA21]. equivalent [HGZM25]. eradicate [TNW21]. eradication [SLML20]. Erratum [Akh23, PLSS24a, SSKS20]. error [BNPL20, LB25, Tak24, WHRH^{+}23]. error-proneness [LB25]. errors [MMI21]. erythrocyte [DN22]. escape [IHO21]. Escaping [HII22, MGCH21]. Escherichia [BF20, VKS^{+}20, YRMWT20]. essential [ES24, Kan21, LMC^{+}20]. essentiality [AFTF22]. Establishment [YWWF25, BPP^{+}20, NO25]. estimate [AKUN20, DHC25, GFV^{+}23, LGT21]. estimated [KvLBD25]. estimates [KBD^{+}23, SL22]. Estimating [DF22, HPGF25, KRSSH21, LRST22, LZW23, SKNvRL23, SOGK^{+}23, YAN21]. Estimation [MS20, SKNvRL22, XDT^{+}20, AN24, BKP23, DCS20, DO21, FPGOCP^{+}25, GFH20, GMR^{+}24, IMM24, SBB^{+}22]. Ethnic [KIM^{+}22]. Etosha [KM22]. eukaryogenesis [KI20]. eukaryotes [Di 20]. European [KRSSH21]. eutrophic [SMJ20]. eutrophication [SH21, TMTK22]. Evading [DBD23]. evaluate [GFV^{+}23, GYY^{+}23, RY22, TNHV21]. Evaluating [MN23, MBHG^{+}24]. Evaluation [GF23, LPSMCP23, GSMG22, KIO^{+}25]. evasive [GtHH20]. evenness [Pie21]. Event [SJS^{+}25]. Event-based [SJS^{+}25]. events [BS25, BFSA23, GCA^{+}20, JLK22, SSR21, YY21]. Evidence [CRGGSC20, ASF22, CPMF^{+}22, HGZM25]. Evolution [CTL^{+}22, CDdV23, FG23, IY22, Kur23, ML24, POW23, Pis25, Qui20, Sak22a, SdAM24, WSW21, YYI21, Yam24, AFTF22, ASF22, ASS^{+}23, APL21, BJR^{+}22, BB21a, BTS24, BCGG21, BSL24, BRBF20, BW25, CG21, CLG20, CWT^{+}23, EBA21, FG21, Gar24, GG23, HK21, HH20, HXS21, Jam21, KHBS20, KJV20, KL21a, KCO25, Kur21, Leh22, LZZ20, LL21, LT22, MPH^{+}22, ML21, NSZ^{+}25, NBBR23, OYT25, OU22, OSAR21, Par22, PHN^{+}23, PS22b, Sme21, SP20, TG22, VA22, WS24, WDHB20, WSC^{+}22, Wil23, YTG24, YJZ^{+}24, YY21, HD25]. Evolutionarily [MS25, MC20b]. Evolutionary [Cai22, CWW22, CGS^{+}25, IJROTE22, IS20, IYY22, MRK24, MTLZ22, POS21, YI21, ZWWW23, Bol22, BH20, BA24, CKB^{+}21, CBPG23, CN20, CN21, CGR22, Dal22, FPGOCP^{+}25, GRPG24, IDD23, KT20b, LYYN20, Leh21, LFT^{+}23, MGCH21, NO25, NKE23, NWP^{+}24, PV24, PABS23, QVE25, QHS23, Sch17, Sch23, Ste21b, SSR21, TT24, TRKB25, VA22, WNHK21, WK25, YTG24, Yin21]. evolving [HCSAB21, WLW19, WLW20]. exacerbate [DCP^{+}22]. Exaggerated [LPD23]. examination [GBF20]. example [GFAP^{+}25, YYI21]. exaptation [CGS^{+}25]. excess [ACP20]. exchange [Xie25]. exchanges [BGGM23]. excitability [OGD^{+}20]. Excitable [PK23]. Excitation [CKM^{+}20, SL21]. excitation-contraction [SL21]. Excitatory [AKHM23]. excited [Tsu20]. exclusion [Gav25, GW20, LDL^{+}24, SR24, TAP^{+}23]. exercise [BSV^{+}24]. Exergy [DRK23, DRK23]. exhibit [IH21]. exhibited [AFTF22]. existence [MMI21]. existing [HH20]. exit [Kur21]. exocytosis [SWVW25]. exogenous [NVS20]. expansion [IB21, KF20, MMS^{+}23, RBZP24, TS23]. expenditure [HXS21]. expensive [CS22]. experienced [Haj20]. experiment [KJV20, dGSNK22]. experimental [CFH^{+}25, CLSB23, DO21, GLLP23, NSZ^{+}25, ZZGG21, ZZB^{+}21]. experiments [KKS23, WZKM21]. explain [BNP^{+}21, Dal22, Wad22]. explained [RBW20]. explaining [GRW^{+}21, HH20, KHBS20, WXS25]. explains [HL20, SdB23, SBC^{+}20]. explanation [JME24]. explicit [BYHH20, FG22, GGSTS21, PK23, RW20]. exploitation [MC20a, MW20]. Exploration [SCD^{+}23, SD25, FHR22, KKS23]. explored [LLJ25]. Exploring [BFSA23, NSD22, PDD24, SMJ20, TWT23, SUA^{+}20, VB21]. Exponential [DKSP24, SR24]. exponentially [YDS24]. exposed [JKH^{+}23, RT24b]. exposure [CWF23, FPGC21, GFC^{+}22, SHH20]. expression [BS21, BS25, GWZ^{+}22, PL20, RMSS24]. extended [FSKG20, MWO21, PLM24, RB22, SSB20]. Extending [Pat22]. extent [Goy21]. external [CT22, KF20, LR21, LZ23]. extinction [CSDH^{+}25, EWS23, FVC20, HII22, Kro25, Leh21, ON24, SSB20, Sch18, Sch21a, SD22, ZWT24]. Extinctions [MTB25]. extracellular [PGML20, SL21, ST20]. extraction [NR20]. extreme [PWC^{+}23]. extremes [AH25]. eyespots [LCMW21].

Fåhræus [FFR23]. face [NGKB21, Ngo22]. facilitate [Jam21]. facilitated [ADP21]. facilitates [FLO25, HZS^{+}21, HGF23]. facilitation [MCL20]. facilities [KRG24]. factor [KB20, RLW^{+}21, Rob22, SSKP24, ZZGG21]. factors [Bau20, Cor21, Cor22, TWT23, ZZ22]. facultative [GZ24, KM22]. failure [APR^{+}22b, BC21, BB25, HH21]. fairness [WFT23]. falciparum [DDDMT22, TEWN21]. farm [SHH20]. farmers [LYYN20]. farming [BPS25, CBBD22]. farms [ZRL^{+}21]. fascia [LMLP20]. fascicle [CDC20]. fastidiosa [BCS^{+}22, WWR^{+}24]. fat [VTDM21]. fatality [NAR^{+}20]. fate [AVSTS23, GO21, HLW24, WA22, YFZ^{+}20]. fatigability [BSV^{+}24]. fatigue [MMK^{+}24, Tay23]. Fatty [IYL^{+}25]. favor [LL21]. fear [BSL24, MSB24, PPS^{+}21]. fear-acquisition [BSL24]. fear-induced [MSB24, PPS^{+}21]. feasibility [AY22, MTB25]. feather [ELM25]. feature [LFZ^{+}24]. features [KIO^{+}25]. February [Ano20p, Ano20d, Ano21p, Ano21d, Ano22o, Ano22d, Ano23l, Ano23c, Ano24l, Ano24c, Ano25j, Ano25b]. feed [Wad22]. feed-forward [Wad22]. feedback [Ala23, APL21, BBK23, CSF20, CMRMC^{+}21, FG21, LR25, LYE^{+}24, LTB25, NSMH25, Tak24, TDRC23, VB21, WvdDL^{+}21]. feedbacks [ADP21]. feedforward [BSC20a]. Feeding [IYY22, SWK^{+}20]. female [ASSL22, IY22, SL23]. females [Dal22, IIHH21]. femoral [SCD^{+}23]. ferroptosis [KTP^{+}20]. Fertilization [DCP^{+}22, NDB20]. fertilizers [CT22]. feto [BAJ^{+}21]. feto-placental [BAJ^{+}21]. fever [GW20, MZG^{+}20, ZRL^{+}21, ZZ22]. few [Dal22]. fiber [JME24]. fibrillar [SK23]. fibrillation [DRK23]. fibrous [ST20]. field [FFGR25, HBK^{+}20, MB24, WBKR22, ZLYL20, dDA21]. fields [LYYN20]. filaments [PEMLD22]. filariasis [OGM^{+}25, ORT24]. Filippov [DTS21]. final [BD21]. findings [GLLP23]. fine [LSW20, TH20]. fine-tuned [LSW20]. fine-tuning [TH20]. fingerprinting [CKM^{+}20]. finite [BD21, GEHL^{+}22, HZM^{+}21, LDF23, LL21, ML24, MTLZ22, PABS23]. finite-time [BD21]. fire [BGA21]. firing [BRN^{+}22, CHT22, ZSL23]. First [BJG21, Ish23, KvLBD25]. Fish [RCM^{+}25, DCP^{+}22, ETTT24, EHP^{+}21, LSZ21, MLK21, MSB24, ZPP23]. Fisher [AAF^{+}25]. fishing [NNNHA24]. fission [HL20]. Fitness [RWHK21, BP20, HMS21, MLG21, PASO21, RID23, Sme21, SS21]. five [ZZ22]. Fixation [WM20, CBBD22, EWK20]. fixed [HBG21]. flatness [Sme21]. flea [KCW^{+}25]. flicker [BCC^{+}21]. flight [GtHH20, MIP^{+}21, Sac22]. floating [CRDD22]. Flock2 [Hoe24]. flocking [Hoe24]. floral [SBJR21]. Flow [SVVPPR^{+}24, AG22, AVP^{+}25, CHT22, GCA^{+}20, JJR20, KB20, MIP^{+}21, MN23, Ros21, SVHM^{+}21, TCH21, TH25, YBF23, YTW21]. flow-induced [KB20]. flowering [ML21, WS24]. flowers [HMS21, WS24]. fluctuating [MSLW23, SS21, UIY22, WM20]. fluctuation [ZX22]. fluctuations [BJR^{+}22, Cai22, FSK21, IDMG24, NSMH25]. Fluid [ELRG21, SAK21, SWK^{+}20, WGRE21, YTW21]. fluid-feeding [SWK^{+}20]. fluorescence [SC21]. fluorescence-based [SC21]. flux [CP23, LB20, ZB22]. focal [MDS25]. focus [ACF22]. folded [WA22]. folding [San21]. follicle [DBW22, ELM25, FHR22]. following [dCC^{+}23]. food [ASMN23, ACP20, BACS24, BD22, KHHI25, LPA25, LLZS25, QHS23, QPL21, WDHB20]. food-web [WDHB20]. foot [HXS21, RLW^{+}21]. foot-and-mouth [RLW^{+}21]. forager [MSM23]. foragers [GCX^{+}20, MH22b]. foraging [GQG23, Ios21, MBT22, QPL21]. Force [Coh20, CDC20, KPP20, LMLP20, SL21]. force-displacement [LMLP20]. forecast [JAW25]. forecasting [GDC^{+}25]. foreign [AKUN20]. forest [ZYLW21]. form [BTS24, DRO23, IHG21]. formation [AZdV^{+}22, BH24, BAG^{+}24, BPM^{+}22, CLM^{+}25, DEM^{+}23, EP20, HYM^{+}22, HGF23, JMD20, KG22, KGP^{+}22, KMA^{+}21, LFY^{+}24, MDS25, NNOBS23, NO24, PLXD^{+}20, Sal25, VGC22, WH21, WFL^{+}24, YXW21, ZSS^{+}22, ZYLW21]. formations [CGW22, Nak25]. formed [FS24]. forms [Bar23]. Formula [GRW^{+}21, IHG21]. forward [CCSF21, Wad22]. four [Ros22, Wad22]. four-node [Wad22]. Fourier [DH22]. fractal [LPSMCP23]. fractional [CLLL20, DSB^{+}21, GFN23, ANGJ25]. fractional-dose [CLLL20]. fractional-order [GFN23, ANGJ25]. fractionated [LR21]. fractionation [KK23]. fractions [WvdDL^{+}21]. fracture [SCD^{+}23]. fragility [DN22, JSCW21]. fragmentation [FSKG20, RB22]. fragments [CMLM23]. framework [ASF22, BSV^{+}24, BEO^{+}22, HL24, Lei20, Lei24, LB20, MAHMH20, OKT^{+}21, PPFB24, PPGB20, SUA^{+}20, VVD24, WXS25, YJZ^{+}24]. frameworks [GHF22]. France [LMW21]. Free [KC24, TATT21, BP21, BAU^{+}22, CG21, CCPD22, D’A20a, D’A20b, Kuz24a, WM21b]. free-swimming [CG21]. French [MEPS22]. frequencies [GEHL^{+}22, LMG22, MPH^{+}22, SC23b]. frequency [CWR21, SL22, SHY21]. frog [PO20]. frogs [AKHM23]. Frozen [LLJ25]. frustrated [dLFO21]. FtsZ [YRMWT20]. function [BTS24, BAJ^{+}21, FNGJ21, For24, PRFSHV20, PZG20, San21, ZMKF22]. functional [AGTJ23, Bar23, DWNW24, KSSM20, LAP24, RNP^{+}25, SY25, Wor20]. functions [BRS24, FOY21, PAV21]. fundament [DRO23]. funds [HL21]. fungal [SVVPPR^{+}24, dJHP20]. fungi [MGZ^{+}21]. Further [GHF22]. fusion [LFZ^{+}24]. future [AJR^{+}20].

Gaia [AN22b]. gain [BAU^{+}22]. gain-scheduled [BAU^{+}22]. gains [SSLMGC25]. gaits [SF24a]. game [AR21, CGP^{+}20, CS20, CK20b, FFAV21, GCX^{+}20, HPD20, IYY22, JCC20, KT20a, KT20b, Kan21, KHST21, KC22, Kur23, LYYN20, MH22b, MTLZ22, NO25, NGKB21, ORT24, PAV21, QHS23, SSBG20, YI21, YTG24, YG25, Whi21]. game-theoretic [CS20]. game-theoretical [CGP^{+}20]. games [BK20b, CWW22, HC24, KCO25, KL21b, MMI21, MMI22, MS25, RP21a, Ste21b, WSW21, WM20]. gametocytes [DDDMT22]. gamma [CWT^{+}23, LMS21, MP22]. ganglia [HXZ^{+}22]. gap [ACG24]. gas [BGGM23]. gastric [Avv21, SXL^{+}20, YXW21]. gastrointestinal [ATP22]. gated [RLL25]. gathering [Pro24]. GC [BRBF20]. geese [YdKdJ^{+}20]. Gene [ON24, AFTF22, ACG24, BS25, DMAC21, FLRR23, GT22, GWZ^{+}22, GTE20, HJKS20, MB22, NCL20, PL20, RMSS24, RW20, Tha24, WW22, YST21]. gene-regulatory [FLRR23, Tha24]. genealogy [CLF^{+}22]. genera [HXS21]. general [Lei20, Lei24, QHS23, WH21]. Generalist [XKV20, BB25]. generality [BNPL20]. generalization [AM22]. Generalized [BACS24, GFH20, LR21, MPTCT^{+}20, SSBG20]. generalized-growth [GFH20]. generate [SHY21]. Generating [XWS24]. generation [AN24, CSHH20, GBF20, RMSS24, SL21]. generations [ZZT^{+}21]. generative [RMSS24]. generic [SUA^{+}20]. Generosity [MC20a]. genes [Kan21]. Genetic [CKNCG22, RP21b, ZF21, AFTF22, ABC25, CPMF^{+}22, Di 21, IY22, JTB25, Kau23, LLJ25, MWO21, TIS25, TG22, WH21, Whi20, YST21]. genetical [Wil23]. genetics [BS22a]. genome [FFNG^{+}21, GPDW20, HJKS20, LAP24, YY21, Yin21]. Genomic [For24, BRBF20, WBKR22]. genotype [AN24, WBKR22]. genotype-phenotype [AN24]. genotypic [Sme21]. geographic [APR^{+}22a, WDHB20]. geometric [Gar24]. geometrical [Haj20]. Geometrically [VG25]. geometries [Car23]. geometry [FLO25, GSC23, KBS^{+}19, KB22, MTB25, Pie21, WLJ^{+}20, YSK23, KBS^{+}20]. germ [TG22]. germ-soma [TG22]. Germany [BGC^{+}20, LMW21]. germination [KP21]. get [HBG21]. gets [ABOR20]. Getting [DAB^{+}23]. Gibbs [HBK^{+}20]. GIFT [WBKR22, IY22]. given [MSLW23]. gleaned [TKRA20]. Glia [GT20]. gliding/climbing [Sac22]. glioblastoma [ATH^{+}24, HBK^{+}20, LS21]. glioblastoma-immune [ATH^{+}24]. Glioma [CGGG20, ZLZ^{+}25]. Global [DMR21, GZ24, RRAO21, dIMM24, DSdP22, LNT23, RBG^{+}23, YXW21]. globally [VCS24]. glucagon [GZM^{+}20]. glucose [DSB^{+}21, HBWLLPA21, LR25, LPBMA21, MCP20, SK21]. glucose-induced [MCP20]. glycolysis [JS23]. glycolysis-OXPHOS [JS23]. go [SKNvRL22]. goals [SSX^{+}22]. going [Hon20]. golden [Gar23]. gonorrhoeae [TKB^{+}20]. goods [CPS20, FPX20, FFAV21, JCC20, RP21a]. Governance [CPS20]. governing [VB21]. gradient [KK22b, LKP^{+}20, Lio18, Lio24, MTB24, SLCC22]. gradients [DBTI21]. gradual [APL21]. graduated [CPS20]. Graph [HJKS20, LLZ^{+}22, POS21]. graphs [BK20b, BPM^{+}22, CGR22]. grazing [CBBD22, MTB24]. greatly [TK25]. greedy [MC20a]. Green [WSY21]. grounded [ABOR20]. groundnut [DSM^{+}23]. group [HXK^{+}20, Kur21, MH22b, MCMdAM24, ND22, ORW^{+}20, PPS^{+}21, TBNS23, Yam24]. group-living [ND22, TBNS23]. group-structured [ORW^{+}20]. grouping [TBNS23]. groups [FS24, KCO25, MHPB23, XML22]. grow [KI21]. growing [GJDS20, PHN^{+}23]. grows [HD25]. growth [AMSC20, AW22, AML^{+}22, AGBLGA^{+}22, BCLJ24, BMB^{+}21, CBZK21, CLD21, DCP^{+}22, DSM^{+}23, FBSD22, FHR22, FB23, FB24, FLR23, GTGB21, GDd^{+}21, GFH20, IHO21, KHAK21, KPR20, KYHN23, Kuz24a, Kuz24b, LSWVB24, LHG^{+}22, LPSMCP23, LSZ21, LGBG23, MGZ^{+}21, MMS^{+}23, NNG20, PBP23, PP20, PP24, PTZ21, dMQHRGA^{+}22, SRB^{+}24, Sch18, Sch21a, SBW^{+}22a, SWVW25, Tay23, Tsu20, VG25, WWR^{+}24, WZP^{+}23, Wu24, ZLZ^{+}25, ZYLW21, dJHP20]. growth-elasticity [Wu24]. GTPase [KJM20]. Guam [CGY^{+}23]. guidance [HCSAB21]. guide [SOGK^{+}23, WBS20]. guided [HGF23, MBT22]. guidelines [NV22, PRHS20]. Guinea [EHP^{+}21]. gut [AZdV^{+}22, HI21, JME24, TMC^{+}23, Wu24].

h [DB22b]. H/dMMR [HKL25]. H1N1 [WSC^{+}22]. H3N2 [WSC^{+}22]. H9N2 [XYK^{+}20]. habitability [MC22]. habitat [PAV21]. habitats [BP20, BYHH20, CRDD22, PHN^{+}23]. Habitual [PBB20]. haemorrhagic [WP21]. hair [DBW22]. hairs [HZS^{+}21, PCR21]. Haldane [BGGM23]. Haldane’s [GCM23]. Hamilton [APL21, FS24, Gar23, HZM^{+}21, Leh22]. handling [SL21]. haptotactic [BH23]. harmonic [KPR20, SHY21]. harmonics [RE25]. harms [CS22]. Harnessing [Kan21]. harvested [FS22]. harvester [HDM20]. harvesting [SSB20]. Hawk [HPD20, KC22]. HBV [Goy21]. HCV [LZZ20]. HDL [YE21]. head [ASSL22, CMLM23, YBF23]. healing [HPG^{+}24, HCJT23, LS25a]. health [BW24, Swe20]. healthcare [BP23, CPEK23, FG21, FG22, FG23, STCSC^{+}21]. Heaps’ [LAP24]. heart [DO21, DRO23]. Heat [FFGR25, ASP23, CKF^{+}24, ZZGG21]. hedging [ZX22]. height [JCLS^{+}11, JCLS^{+}21, PZG20]. helminths [HVTA20]. help [PB22]. helper [Mit20, Pag25]. helps [CS22]. hematopoietic [CSH^{+}25, PASO21]. hemodynamic [LRAP22]. hemodynamics [RKCvB25]. hemodynamics-based [RKCvB25]. Hepatitis [SD25]. hepatocytes [CBK^{+}20, DYS^{+}21]. herbivore [MTB24]. herd [Bai21, GFC^{+}22, KPA^{+}21, TNHV21]. herring [SSB20]. Hes [Pfe22]. Hes1 [SMS20]. hesitancy [BDdG22, JMWX23]. heteroclinic [YST21]. Heterogeneity [RMG^{+}21, TK25, Bai21, BE22, BS21, BYHH20, BV24, CBK23, CF20, Don23, EB24, GFAP^{+}25, GCA^{+}20, LGT21, LLS^{+}22, MCM20, POW23, PBD^{+}20, SKS21, SMJ20, TMC^{+}23, TCH21, WZP^{+}23, YTG24, ZWT24]. Heterogeneous [MCP20, DJ22, LSWVB24, Lei20, Lei24, LH20, MF24, MG21, ORW^{+}20, Pro24, dVRA20, RBZP24, SLTD22, SS21, Ter23, WGRE21, WFT23, dKCA22]. heterologous [RY22]. heterotypic [ST20]. heuristic [MS20]. HFMD [XDT^{+}20]. hierarchical [LPA25, LH20, NSMH25, PRP^{+}20, TS23, VKS^{+}20, YG25]. hierarchy [Ros22]. high [Ala23, ASS^{+}23, KP21, KKL20, Sak20, VTDM21, WDA^{+}23]. high-avidity [KKL20]. high-vaccinated [ASS^{+}23]. higher [Cor21, Cor22, PR21, Wad22]. higher-order [Wad22]. highlights [ISC23]. highly [dVRA20]. hindering [Xie20]. Hindmarsh [ANGJ25]. hippocampus [AOB21, HS20, MP22]. history [APR^{+}22a, AL23, KI21, KYI21]. HIV [GQK^{+}21, GDQR23, HKBK25, HCR20, KKK^{+}20, LPP^{+}25, Mar24, NTTI22, PZJPZ23, SdB23, UV23, YHT24, YCKK24, ZXLH23]. HIV-1 [KKK^{+}20, NTTI22]. HIV/AIDS [ZXLH23]. hive [ZM20]. Hodgkin [EE25]. Holt [FS22]. home [XML22]. Homeostasis [Koc20, KTU22, SL23]. hominid [HXS21]. Homo [Koc20]. Homo-Oligomerisation [Koc20]. homogeneous [IHG21, Ter23]. Homogenization [BNP^{+}21]. homology [Ste21b]. homophily [KIM^{+}22]. honey [CFH^{+}25, HZS^{+}21, LZ23, SWK^{+}20]. Hopf [HXZ^{+}22, YST21]. Hopf-born [YST21]. hopping [ELRG21]. Hormonal [FHR22, RNP^{+}21]. hormone [COB23, YTH^{+}21]. hormones [ASSL22]. Horton [Ish23]. hospital [KBW^{+}23]. hospitalization [AGP^{+}23]. hospitalized [KBD^{+}23]. host [AMA20, BB21a, BLRV21, BRAHV25, Bol22, BW25, BRS24, CHSB20, Cor21, Cor22, FG22, GW20, HII22, JLK22, Kis24, KBD^{+}23, KRSG24, KH24, LMC21, MBM^{+}25, MW20, OKF^{+}21, Sin21, SA25, SYH20, SZR^{+}25, TGC22, TNW21, TK25, UV23, VB21, VP22, XYK^{+}20, XCFH23, dCC^{+}23, WDH^{+}20]. host-parasitoid [KH24]. host-symbiont [KRSG24, VB21]. host-virus [CHSB20]. hosts [EHP^{+}21, MGCH21, Roy21, SPLS^{+}23]. Hot [YSP^{+}23]. Hot-topics [YSP^{+}23]. household [Hil23, HTD22]. HPTN [LPP^{+}25]. HPV [SBW^{+}22b]. HPV-related [SBW^{+}22b]. HPVsim [SSK^{+}25]. hub [TT20]. Hubei [SXJ20]. Human [NGKB21, OIK^{+}24, APR^{+}22a, BBH^{+}20, BLRV21, BC20, BGGM23, CDC20, CPM21, CSH^{+}25, Dal22, GMMR22, GTE20, HVTA20, HVC^{+}21, HL24, LAP24, MN23, OKF^{+}21, Par22, Sal25, SC23a, SCD^{+}23, SHH20, TEWN21, ZWT24]. human-antibodies [TEWN21]. human-to-human [ZWT24]. humanized [GDQR23, KKK^{+}20]. humans [FHR22, HD24]. Humeri [LZ23]. hunt [LTB25]. Huxley [EE25]. Hybrid [ZLZ^{+}25, ASK20, AKHM23, ACF22, CSP21, Hen20, JAW25, MS20, ML21, PCLL21]. hybridising [BHW^{+}20]. hydraulic [LKP^{+}20, Ros22]. hydrofoil [ETTT24]. hydrophobicity [KB20]. hyperparasites [NWP^{+}24]. hypha [dJHP20]. Hypnozoite [MSK^{+}22]. hypotheses [HH20, WDA^{+}23]. hypothesis [HGF23, KKS23, Kuz24b, Rob22, SLCC22]. hypoxia [CBZK21, CBH^{+}22, CBH^{+}23]. hypoxic [SLTD22]. Hz [MP22].

ICP [DRO23, DRO23]. ideal [D’A20a, D’A20b, KC24, RP21b]. idealized [DCL23, NM23]. ideas [Di 21]. Identifiability [KLC24, LT23, BK20a, KBD^{+}23, Mar24, SBW^{+}22a, GMR^{+}24]. identifiable [Sim25]. Identification [CR20, TKB^{+}20, AH25, KH20, XDT^{+}20]. identifiers [SS23]. identifies [JLK22, PWC^{+}23]. identify [CP23, Cor21, Cor22, HRD^{+}24, KPH^{+}24, Roy21]. Identifying [HLW24, OGD^{+}20, LLZ^{+}22]. identities [FS22]. idiosyncrasy [ND22]. IFN [PJVR23]. II’ [Pru25, Wil23, GADDVG^{+}24, Sal24, Swe20]. III [Sal24]. IIT [MSL23]. IL-1 [PGML20]. IL-2 [CWL20]. IL-27 [LBF24]. IL-4 [PJVR23]. illness [SSM^{+}22]. illusion [RBW20]. image [ZLZ^{+}25]. images [LG23]. Imaging [Bar21, AVP^{+}25, MDCK23]. imitation [FBA21]. immediately [HHI24]. immigration [FVC20, SS23]. immobile [CG21]. Immune [WPEP22, ATH^{+}24, ACF22, AABG20, BBB^{+}25, BPP^{+}20, BT22, BRAHV25, CAARG21, EB21, FGA^{+}24, GF23, IHO21, JSCW21, JLK22, LLS^{+}22, LT22, LS25a, MHI23, MAHMH20, OKF^{+}21, OGD^{+}20, RS24, SPLS^{+}23, SFK^{+}23, SBC^{+}20, SMEG22, SF23, TKRA20, WMHV24, XZZ^{+}23, vdVPW^{+}22]. immunisation [RDDSA25]. immunity [Bai21, CCF^{+}20, EB24, FG23, GFC^{+}22, JKSD23, KT21, KPA^{+}21, SFK^{+}23, TNHV21]. immunization [SSX^{+}22]. immuno [HPK^{+}20, YDS24]. immuno- [HPK^{+}20]. immuno-epidemiological [YDS24]. immunodeficiency [OKF^{+}21]. immunology [Pat22]. immunomodulatory [MF22]. immunosuppression [GBF20]. immunotherapies [CPM21]. immunotherapy [ATH^{+}24, CWL20, DCS20, GBF20, HS24, LFZ^{+}24, RMS^{+}25]. Impact [ASSL22, BW24, CBBD22, NAR^{+}20, OU22, PEMLD22, QVE25, SRDC21, YTY^{+}20, AH25, BKPT25, BDdG22, CBK23, CLJ^{+}23, FSK21, GFH20, GW20, HCR20, KvLBD25, LLS^{+}22, MN23, Ngo22, OGD^{+}20, RDDSA25, San21, SXL^{+}20, SSX^{+}22, TEWN21, TK25, WCZL25, WUH23]. impacting [SCH24]. impacts [APR^{+}22a, BGM22, CFH^{+}25, DTS21, PAV21, ZWT24]. imperfect [Kis24]. imperfectness [KT21]. implantation [CNK24, MFGACST23]. Implementation [YL25, BH20]. implementing [HMLO^{+}23]. implicates [KLLL22]. implication [WS22, YTH^{+}21]. Implications [DH22, GBRM21, HXS21, MEBNZ24, SBW^{+}22b, CWT^{+}23, Han23, Hil23, KRSSH21, LZ23, MCL21, SSX^{+}22, YL25, Yin21, AWC25, BH24, TMTK22]. imply [LSW20]. importance [Bol22, GG23, KV23, PRHS20, SY24, TBNS23]. important [AH25, KHBS20]. improve [FBA21, GFV^{+}23, SSK^{+}25]. Improved [dGSNK22, LLZ^{+}22, RLB25]. Improving [LK20, LCYD23]. in-depth [RLB25]. in-host [KBD^{+}23]. in-silico} [SBSB23, TKB^{+}20]. in-stent [LFY^{+}24]. in-vitro [AMSC20, NSZ^{+}25]. inactive [DN22]. inattention [Whi21]. inbreeding [Sak20, Sak22a]. incidence [LZL^{+}25, PP20, PP24]. inclusion [TAP^{+}23]. inclusion/exclusion [TAP^{+}23]. Incoherent [GK21]. incompatible [MSL23]. inconsistent [AA22]. Incorporating [AWC25, ASK20, ABC25, SHH20]. Increased [HMS21, PD23]. index [YWWF25, DRK23]. India [SRDC21]. indication [VKS^{+}20]. Indirect [PAM24, LMB22, MCL20, MPA24, Mur25, PJVR23]. indirectly [TML20]. Individual [AM22, BGB^{+}21, GFC^{+}22, ABBP21, BEBA20, CA21, EB24, EP20, FLR23, For25, HGD^{+}25, HRD^{+}24, KT21, LB25, MCM20, OU22, TMC^{+}23, TIS25, WS24]. Individual-based [BGB^{+}21, BEBA20]. individuals [GG23, MHPB23, PZJPZ23]. induced [AR20, ASP23, AHEB^{+}22, ACF22, BCC^{+}21, DBW22, ETTT24, EE25, GBF20, GBRM21, HLS^{+}25, JAB22, JSCW21, KB20, MP22, MSB24, MFGACST23, MHV^{+}21, MCP20, OKF^{+}21, PPS^{+}21, PLSS24a, PLSS24b, SOKZ22, SF24b, SZR^{+}25, WRB22, YFZ^{+}20, YM23]. induces [KF20, MGZ^{+}21]. induction [DMR21, YFZ^{+}20]. inductive [YFZ^{+}20]. inertia [MMK^{+}24]. infarction [MF22]. infected [BLRV21, HVTA20, LK20, MEPS22, PWC^{+}23]. Infection [SD25, WPEP22, AGP^{+}23, APF^{+}24, AKUN20, ACF22, BLRV21, CCF^{+}20, CAARG21, CWF23, Cor21, Cor22, CSH^{+}25, GMMR22, Goy21, GQK^{+}21, GDQR23, HI20, HVC^{+}21, HBGV^{+}21, HSAH20, IH23, Jac25, JLK22, Kee23, KBD^{+}23, KM22, KKK^{+}20, LMC21, NM23, PWC^{+}23, RY22, SPLS^{+}23, Sal25, SdB23, SMEG22, WMHV24, WMO23, XYK^{+}20, YdKdJ^{+}20]. Infections [HGD^{+}25, CPEK23, HPSA21, HPSA23, HBGW^{+}25, JD21, JAW25, KPH^{+}24, XDT^{+}20]. infectious [AN22a, AH25, DHR25, GBHNHVL23, GJCLMQC20, JKH^{+}23, KHB^{+}22, MD22, MN23, MBHG^{+}24, OGM^{+}25, SGM^{+}25, SA25, SOGK^{+}23, YFF25, ZWT24]. infectious-contagious [GJCLMQC20]. infectiousness [CR20, LWM23]. infective [LPLM25]. infectivity [NAR^{+}20]. infer [CLSB23]. Inference [GFH20, WZKM21, WW22, CGS^{+}25, GFL23, HHT25, LT23, LG23, MMS^{+}23, OV25, RW20, SGM^{+}25, Sim25, TMC^{+}23, VTDM21, WBS20, YOIS24]. inferred [RGBD25, WHRH^{+}23]. Inferring [MCMdAM24, RBG^{+}23, SWVW25, WCL^{+}25]. infertility [Pag25]. infestation [BPS25]. infinite [GPDW20]. inflammation [ATP22, MHV^{+}21, NM23, QYY^{+}23]. inflammation-causing [ATP22]. inflammatory [EE23]. Inflected [MIP^{+}21]. inflight [YBF23]. inflorescence [WS24]. Influence [WFL^{+}24, BCK24, GHB22, LS21, LWXM20, MSB24, MSM23, SMM^{+}23, TCMA21]. influenced [CWT^{+}23, MPC^{+}24]. influences [WMO23]. influenza [AMR^{+}23, BNP^{+}21, LMC21, NAR^{+}20, SPLS^{+}23, SD20, SMEG22, SSM^{+}22, WMHV24, WDH^{+}20, WSC^{+}22, XYK^{+}20, XCH^{+}20, YAN21]. influenza-like [SSM^{+}22]. information [CKM^{+}22, CRRRP^{+}20, FSNS20, GCA^{+}20, Kan21, LL20, LX21, LMC^{+}20, MB22, TBNS23]. informational [WBKR22]. informative [MLG21]. informed [PGC^{+}22, RRAO21, Sim25]. inhaled [BBB^{+}25]. inhibition [BCLJ24, BSC20a, CPMF^{+}22, DB22a, HL20, MCP20, Wal22]. inhibitor [SF23]. inhibitors [DOF25, HKBK25, PD23, TAP^{+}23, XZZ^{+}23, vdVPW^{+}22]. inhibitory [AKHM23, NSMH25]. inhomogeneities [SP20]. initial [GB23, MBM^{+}25]. initiated [HI20]. initiation [EE23, Gre25]. initiatives [DAB^{+}23]. injury [AHS23, GBRM21, WFL^{+}24]. Innate [WPEP22, ACF22, BBB^{+}25, CAARG21, EB21, GF23]. inner [QYY^{+}23]. innocent [GRPG24]. innocent-spreader [GRPG24]. innovation [SHK20]. inoculation [MGZ^{+}21]. inosine [YS25]. input [SHY21, Wad22]. input-output [SHY21]. inputs [ACD20]. insect [BB21b, CCPD22, IIHH21, MSL23, PR21]. insecticide [HPWvdB20, Ngo22, ORT24]. insecticide-treated [Ngo22]. insects [GCA^{+}20, OYT25, SAK21]. insight [LX21, SXJ20, ZBT22]. Insights [GDC^{+}25, OGM^{+}25, WRB22, FPS24, TKRA20, DCL23, ES24, GGG21, Sco24]. Insightsl [DDDMT22]. inspired [BFG21]. instability [AW22, MTB25, XDZ^{+}21]. insulin [DYS^{+}21, GC20, GZM^{+}20, IYL^{+}25, LML24, Swe20]. integrated [JA23, KKS23]. integrating [BvdG20, BJJM25]. integration [CFH^{+}25, TKR^{+}21]. integrations [Goy21]. integrative [BSV^{+}24]. integro [CSP21, HKL25]. integro-differential [CSP21, HKL25]. Intelligent [RMSS24, TRKB25]. intensity [LPLM25]. intensive [KRG24]. inter [FMO23]. inter-determined [FMO23]. interact [MBT22]. interacting [FLR23, PB22, PAV21, vBBB^{+}25]. interaction [AFTF22, AAL^{+}21, CLW^{+}24, CLSB23, CA21, EPF20, FBSD22, KHBS20, MTLZ22, Nah20, PJVR23, Pru25, Rev25, TMC^{+}23, TTBL21, Xie23, ZSS^{+}22]. Interactions [MCL20, YJBF20, AKHM23, AABG20, ADP21, BT22, BW25, BS22b, DEM^{+}23, FBJR20, FGA^{+}24, FBD23, GF23, GCA^{+}20, HYM^{+}22, HLW24, LTB25, LLZS25, MMG21, MK21, MCMdAM24, MK24, MWO21, OGD^{+}20, RCM^{+}25, WCWL21]. interactive [QWZK24]. interbreeding [RT24b]. intercellular [DEM^{+}23]. interface [BM20, PGC^{+}22]. interferes [Mit20]. interferon [ACF22, SD20, XYK^{+}20]. interferon-dependent [SD20]. interferons [GCNZ^{+}21]. intergalactic [KJV20]. intermediate [EHP^{+}21, PEMLD22, Roy21, TATT21]. intermittency [VP22]. intermittent [CCSF21]. Internal [TMTK22]. Interplay [AVSTS23, CGGG20, PJPH25]. interpretation [BAJ^{+}21]. Interrelating [BFD25]. interruption [DMPIG24]. interruptions [HCR20]. interspecific [vBBB^{+}25]. interstellar [KJV20]. interstitial [WGRE21]. interval [AOB21, WS24]. intervals [ZZT^{+}21]. intervention [IHG21, PRHS20, SRDC21, LPLM25]. interventions [BSS^{+}20, CC22, DJ22, FG21, FG22, FG23, FGC24, HD24, JKH^{+}23, LRST22, PAM22, ZWC^{+}22]. intestinal [Avv23, HI21]. intra [BV24, BRS24, LLS^{+}22, SY24, WS24]. intra-clonal [BV24]. intra-host [BRS24]. intra-inflorescence [WS24]. intra-modular [SY24]. intra-tumour [LLS^{+}22]. Intracellular [SD25, CAARG21, PLXD^{+}20, PEMLD22]. intracranial [DRO23]. intraguild [SYH20]. intraspecific [vBBB^{+}25]. intratumoral [WGRE21]. intrinsic [LMA21, SSKP24]. introduction [BB25]. invader [MS25]. invaders [vdBMB^{+}23]. invagination [JA23]. invasion [AZdV^{+}22, BH23, CGY^{+}23, CGGG20, DTLS24, FSC21, KMA^{+}21, MH22a, MCM20, PJB21, WHH22]. invasions [SCH24]. invasive [DP22, JTB25, MGCH21, NSMH25, PZ20]. inverse [NNOBS23]. invertebrates [CG21, IYY22]. Inverted [YY21, ABOR20, Wal20]. inverted-U [Wal20]. inverted-U/U [Wal20]. investigate [BS21, FPGC21]. Investigating [HS24, LYYN20]. Investigation [AGTJ23, AHS23, BE22, EB21, GB23, HIA^{+}22, HS20, LR21, MPA23, PP22, YS25]. investigations [WGRE21]. involved [BBK23, MHI23, PJVR23]. involving [HHI24]. iodate [PLSS24a, PLSS24b]. iodate-induced [PLSS24a, PLSS24b]. ion [MI20]. ionizing [DBW22, For25]. irregular [Wu24]. irreversible [CG21]. irruptions [VP22]. ischaemia [AVP^{+}25]. ischaemic [WP21]. ischemia [NMC^{+}21]. ischemia-reperfusion [NMC^{+}21]. Ising [LLJ25]. isokinetic [BSV^{+}24]. Isola [GK21]. isolate [NGKB21]. isolation [GGSTS21, PLR24]. Isotropic [KF20]. Italy [LMW21]. iterated [HPD20]. IV [KSSM20].

J [Akh23, CBH^{+}23, Cor22, HPSA23, JCLS^{+}21, Lei24, Lio24, LK25a, PLSS24a, PP24, Sch21a, Sch23, SSKS20, Xie22]. jackal [KM22, MK21]. January [Ano20q, Ano20e, Ano21q, Ano21e, Ano22p, Ano22e, Ano23m, Ano23d, Ano24m, Ano24d, Ano25k, Ano25c]. Japan [AKUN20, OIK^{+}24, SFK^{+}23]. jet [DB22b]. jetting [KM20]. Jiangsu [APR^{+}22b, CZN^{+}21]. Joint [DTS21, Ost23]. Josephson [YMR24]. Journal [D’A20a, GHS20, WLW20]. JTB [CKI20]. July [Ano20r, Ano20f, Ano21r, Ano21f, Ano22q, Ano22f, Ano23n, Ano23e, Ano24n, Ano24e, Ano25l, Ano25d]. jump [GWZ^{+}22]. junction [WS22, YMR24]. junction-coupled [YMR24]. June [Ano20s, Ano20g, Ano21s, Ano21g, Ano22r, Ano22g, Ano23o, Ano23f, Ano24o, Ano24f, Ano25m, Ano25e]. juvenile [MGCH21].

keratocytes [MLK21]. kernels [EIK^{+}21]. key [BC20, CRRRP^{+}20, FLR^{+}21, GRW^{+}21, LSZ21, TWT23, ZYLW21]. kidneys [LPBMA21]. Kill [RS24]. Killer [Whi20, HPK^{+}20]. kills [Pis25]. kin [HZM^{+}21, Sco24, SBJR21, WFT23]. kin-selection [WFT23]. kinase [NTTI22]. kinesin [Xie21, Xie22, Xie23, Xie24, Xie25]. Kinetic [GGS^{+}24, KHAK21, LML24, PTZ21]. kinetically [PRFSHV20]. Kinetics [Jac25, SSKP24, SD20, SGJ^{+}24, WZP^{+}23, dCC^{+}23]. Kingdom [LMW21]. Kir4.1 [RLL25]. Kirschner [CSW24]. kleptoparasitism [GCX^{+}20]. known [LKP^{+}20]. Kolmogorov [XKV20]. Korea [LMW21]. Kullback [HGF23].

L [PD23]. L. [CGY^{+}23]. L1 [BS21]. labor [Yam24]. laboratory [WW24]. Lactate [SBSB23]. lactation [MSM23]. lactylation [LR25]. lag [DB22b]. lake [SH21]. Lamarckian [DMR21]. lamellipodium [MLK21]. laminar [Ost23]. land [CBBD22, PGP23]. landscape [MBT22, RGBD25, SFK^{+}23]. landscapes [BB21b, RID23, dKCA22]. lane [FLO25]. languages [NO25]. Laplacian [LLZ^{+}22]. lapse [LG23]. Large [LSW20, LPP^{+}25, FFR23, Gav25, MZG^{+}20, Täu20, ZRL^{+}21]. large-scale [MZG^{+}20, Täu20, ZRL^{+}21]. larvae [IYY22]. Lassa [MZG^{+}20]. Lasso [HXK^{+}20]. late [GRW^{+}21, KT20b, Sak20, Sak22b]. late-acting [Sak20, Sak22b]. latency [BK21]. Latent [Yin21, HHI24, PZJPZ23]. lateral [CSP21]. lattice [Ala20, Xie23]. lattices [Xie25]. law [CLM^{+}25, CLG20, Ish23, LAP24, MK24, Ros21, Wad22]. layer [AG22]. lead [FSNS20, HSAH20, NNNHA24]. leaders [FPX20]. leadership [PP22]. Leading [PLSS24a, PLSS24b]. leads [CWR21, Nak25, ZM20]. leaf [LKP^{+}20]. leakage [SLCC22]. learn [GWZ^{+}22]. learned [New20]. Learning [Ala23, LTB25, AJR^{+}20, BJJM25, CA21, LFZ^{+}24, MDCK23, NR20, OU22, Par22, WXS25, YCKK24]. Learning-based [Ala23]. learning-explaining [WXS25]. learnt [DAB^{+}23]. leaves [Nak25, Ros22]. leg [ABOR20, KPP20]. legal [HL21]. LEGO [MPTCT^{+}20]. LEGO-based [MPTCT^{+}20]. Leibler [HGF23]. length [CLW^{+}24, HL24, LKP^{+}20, SBC^{+}20, SSR21]. length-dependent [SBC^{+}20]. lengths [GW21]. Lepidopteran [SAK21]. leprosy [GSR23]. leptin [Swe20]. lesions [AOB21]. lessons [DAB^{+}23, LPLM25]. lethal [BS25]. letter [PMCLC^{+}23]. leucine [NTTI22]. Levant [MUMS24]. level [ASK20, GG23, HDM20, KHAK21, SKS21, WHRH^{+}23]. levels [BKP23, CTL^{+}22, CDdV23, CN20, HKBK25, Ros22, TCH21]. Lewis [Ano21-49]. Lewy [Kuz24b]. Li [KK22a]. lies [ZYF25]. Life [AL23, APR^{+}22a, BBDB22, Di 20, HD25, HL24, IYY22, KI21, KCW^{+}25, KYI21, PBD^{+}20, TH21, YYI21]. life-cycle [IYY22, PBD^{+}20]. lift [SXJ20]. ligand [CKM^{+}20, WRB22]. ligand-induced [WRB22]. light [BEBA20, MIP^{+}21, PP20, PP24, SC21]. like [FSKG20, FGA^{+}24, FET^{+}24, HD25, RB22, RBD^{+}20, SSM^{+}22]. likelihood [GFH20]. limit [BM20, Gav25, MBM^{+}25, SBB^{+}22]. limitations [BJB^{+}20, ESW24]. limited [BF20, CD22, CLLL20, Kuz24c, LGBG23, LAP24, SPLS^{+}23]. limits [DTLS24, DB22b, Sch18, Sch21a]. line [CLSB23, CSP21, KP21, LG23]. lineage [NSMH25, SSKS18, SSKS20]. linear [KL21b, MPTCT^{+}20, SF24a]. Link [CLW^{+}24, SY24, ZYLW21]. linked [KNK21, Swe20, VTDM21]. Linking [BRAHV25, DHLS20, LX21, AN24, GHS18, GHS20]. lipid [Swe20]. Listeria [Jac25]. literature [NR20]. little [BF20]. lived [TS23, ZF21]. liver [GBRM21, HBWLLPA21]. livestock [BGB^{+}21, CBBD22, KHW^{+}22, SHH20]. livestock-associated [SHH20]. living [Bar21, ND22, TBNS23]. LMI [SJS^{+}25]. load [BNP^{+}21, LZ23, PWC^{+}23, SdB23, SHH20]. loading [IHO^{+}20]. loads [Xie20, Xie21, Xie22]. lobe [PR21, TKR^{+}21]. local [BP21, BYHH20, ETTT24, FOY21, FPS24, HCC^{+}21, LNT23, LRST22, NSMH25, RBG^{+}23, SMEG22, Ter23, WCD21]. localities [ENMW21]. Locally [BGC^{+}20, HKL25, SC23b]. location [PCR23, ZLYL20]. loci [KNK21]. lockdown [KvLBD25, SXJ20]. lockdowns [BGM22]. Lockhart [CLD21]. locomotion [CCSF21, ELRG21]. locus [ABC25, Sak22b, SSS21]. lodging [BSJ20]. logarithm [Wad22]. logical [ZMKF22]. logistic [Sch18, Sch21a]. Long [COB23, FGC24, Rav21, BW24, BBH^{+}20, HD24, MPH^{+}22, MGFS^{+}20, MWO21, PL20, SAK21, SXL^{+}20, SWK^{+}20, TS23, WCD21, ZF21]. Long-distance [COB23]. long-lived [TS23, ZF21]. long-range [MGFS^{+}20]. Long-term [FGC24, Rav21, MWO21, PL20, SXL^{+}20]. longitudinal [DNK^{+}24, LKP^{+}20, Xie21, Xie22, ZLZ^{+}25]. look [HD25]. loop [BH24, DB22a, FG21, Wad22, YWTC25]. loops [For24]. losers [PPFB24]. loss [BD22, CBKS21, ES24]. lost [CK20b, KC22]. Lotka [CLSB23, MMG21]. Low [ABOR20, ASS^{+}23, CWR21, KHAK21, Mur25]. low- [ASS^{+}23]. lower [Cor21, Cor22]. lowers [GFC^{+}22]. LRBmat [TMC^{+}23]. lumen [NNOBS23]. luminal [Avv23]. lung [ACF22, BE22, BGGM23, Don23, EB21, MHV^{+}21, PBD^{+}20]. Luxembourg [KPA^{+}21]. Lyme [CPHL^{+}20]. lymph [PZJPZ23]. lymphatic [OGM^{+}25, ORT24, RRZDF20]. lymphopoiesis [CMRMC^{+}21]. lysogen [CWR21].

MacArthur [Bar23]. machine [BJJM25, HXK^{+}20]. machinery [GHS18, GHS20]. machines [TH20]. macro [MPTCT^{+}20]. macroevolution [DHLS20]. macromolecular [OLF^{+}24]. Macrophage [FLRR23, BE22, FLR^{+}21, LMC21]. macrophages [BJB^{+}20, EB21]. Macroscale [MK24]. macroscopic [GHS18, GHS20]. macular [PLSS24a, PLSS24b]. magnetic [ASP23]. mainland [RLW^{+}21]. maintain [PB22]. maintained [Sak22b]. Maintaining [MCO21]. maintains [vBW20]. Maintenance [Sak20, BPP^{+}20, Osb24]. major [TWT23]. makes [MIP^{+}21, RR24]. making [Bai21, JCC20, PP22]. malaria [AY22, Hen20, MSL23, MSK^{+}22, Ngo22, SLML20]. Malawi [GDC^{+}25]. male [AKHM23, IY22, KNY25, YI21]. males [IIHH21, Jam21]. Malpractice [CKI20]. mammal [Dal22]. Mammalian [GT20, ASK20, WSEK23, YL25]. Management [PRHS20, AJR^{+}20, BB21b, CBBD22, FG21, HPWvdB20, MH22a, OKT^{+}21, SYI20, Tak24, TMTK22]. mandatory [FHM20]. maneuver [RRAO21]. manifest [YY21]. manipulate [SMS20]. manure [CBBD22]. Many [BTS24, KCO25, Kro25]. Many-strategy [KCO25]. Many-to-many [BTS24]. map [BRN^{+}22, SMM^{+}23]. map-based [BRN^{+}22]. Mapping [GT20, BTS24]. March [Ano20t, Ano20h, Ano21t, Ano21h, Ano22s, Ano22h, Ano23p, Ano23g, Ano24p, Ano24g, Ano25n, Ano25f]. MARCKS [FET^{+}24]. marginal [WBS20]. marine [BvdG20, Dal22, IYY22, TP24, XDZ^{+}21]. marital [CK20b, SW22]. mark [SKNvRL23]. mark-dead-recovery [SKNvRL23]. Markov [CP23, DHC25, FPGC21, LZW23, PRP^{+}20]. Markovian [BBDB22, VCS24]. marsh [ZSS^{+}22]. mass [IYL^{+}25, KJV20, RS24]. massive [AN24]. master [GFN23]. master-slave [GFN23]. mastitis [SYI20]. matching [LPA25]. Maternal [NTTI22, SL23, ZF21]. Mathematical [AGP^{+}21, AN24, AMA20, AY22, BE22, BLRV21, BT22, BSV^{+}24, BGA21, EB21, ELM25, GMR^{+}24, HI20, LHG^{+}22, LCMW21, MPH^{+}22, MHV^{+}21, PASO21, Rob22, SL23, Swe20, TEWN21, TCH21, Vol24, WP21, YE21, vdVPW^{+}22, ABB^{+}22, AML^{+}22, ASL^{+}24, AMR^{+}23, APR^{+}22b, BBK23, BBB^{+}25, BCK24, BJB^{+}20, BGC^{+}20, CGY^{+}23, CNK24, CRGGSC20, CBH^{+}22, CBH^{+}23, CHMM23, CFH^{+}25, CMRMC^{+}21, CGG^{+}25, DAB^{+}23, DOF25, DSdP22, DSM^{+}23, DYS^{+}21, DRO23, EHP^{+}21, FH23, FLR23, GDd^{+}21, GBF20, GO21, GBRM21, GSR23, GZGH22, GLLP23, Han23, HPG^{+}24, HCJT23, JME24, KR22, KV23, KKS23, KK23, LR25, Lei20, Lei24, LLS^{+}22, LGBG23, MF21, MF22, MDF21, NWVN20, OGD^{+}20, PPFB24, PAM22, PGC^{+}22, RT24a, RLB25, RKCvB25, Roy21, RY22, SXL^{+}20, SSX^{+}22, SF23, SF24b, SNK20, SOGK^{+}23, SBJR21, TKRA20, WWR^{+}24, WRB22]. mathematical [WCWL21, WSC^{+}22, WSEK23, XYK^{+}20, XZZ^{+}23, YS25, YTH^{+}21, dCC^{+}23, LS25a, Pru25]. maths [DAB^{+}23]. mating [CT22, GCM23, RWHK21, Sak22a]. Matrix [PJB21, ABC^{+}21, BH23, DOF25, GHA^{+}20, IHO^{+}20, ST20]. matrix-deposited [BH23]. matter [FFNG^{+}21]. maximizes [Sme21]. maximizing [HMS21]. maximum [GFH20, GRW^{+}21, NNNHA24]. May [Ano20u, Ano20i, Ano21u, Ano21i, Ano22t, Ano22i, Ano23q, Ano23h, Ano24q, Ano24h, Ano25o, Ano25g, FVB20, For24, LPD23, PD23, ZWC^{+}22]. MCF7 [GHB22]. mChialvo [SMM^{+}23]. mean [GFAP^{+}25, TDRC23, dDA21]. mean-arterial [TDRC23]. meaningful [TT20]. means [LYYN20]. measles [FHM20, SSX^{+}22]. measure [AGTJ23, CN21]. measured [CK20b, KC22]. measures [BERR20, KHW^{+}22, LL20, NKE23, Pie21]. Mechanical [LZ23, Tay23, GHS18, GHS20, KIO^{+}25, TKR^{+}21]. mechanics [Coh20, GADDVG^{+}24, GRW^{+}21, SS22, YOIS24]. Mechanism [HH21, MP22, ACD20, FZE22, GT22, HL20, NNOBS23, NO24, OY21, SAK21, SVHM^{+}21, WXS25, Xie23, YXW21, ZLYL20]. Mechanisms [SK23, WS22, CLD21, CBK^{+}20, GGS^{+}24, HD22, HGZM25, Jac25, JKSD23, LSZ21, Nah20, SWK^{+}20, SBW^{+}22b, TH21, ZQHN24]. Mechanistic [DRO23, MZG^{+}20, SBC^{+}20, XCH^{+}20, BS21, CF20, KHW^{+}22, TKRA20, VVD24]. Mechanistic-mathematical [DRO23]. mechanobiochemical [HAGADDM24]. mechanochemical [KJM20]. mechanoreceptor [PCR21]. mechanosensory [PTK^{+}24]. media [DTS21, YTY^{+}20, YSP^{+}23]. mediated [BJG21, CNK24, DBD23, FK21, HSL24, KJM20, LYE^{+}24, MF24, MB22, PJVR23, SK21, SBC^{+}20, TDRC23, PL20]. medicine [NDB20]. Mediterranean [KGP^{+}22]. medium [FFR23]. medium-large [FFR23]. meet [GCX^{+}20]. melanoma [TKRA20, XZZ^{+}23]. MELK [NTTI22]. mellitus [DNK^{+}24]. membership [VB21]. membrane [Car23, NNOBS23, PLXD^{+}20, QYY^{+}23, SHY21, WM21a, XSH^{+}24]. membrane-bulk [PLXD^{+}20]. membranes [SSKP24, YMR24]. Memory [LFJ21, MBT22, FNGJ21, IHO21, KR22, LGSR21, RY22, ZSN22]. Memory-driven [LFJ21]. Memory-guided [MBT22]. memory-TFT [KR22]. memristive [ANGJ25, KR22]. memristors [SY25]. menopausal [Dal22]. menstrual [ASSL22, AGBLGA^{+}22]. Menten [CLM^{+}25]. mesenchymal [GHB22]. mesh [LMLP20]. mesoderm [JA23]. meta [MS20]. meta-heuristic [MS20]. metabolic [CP23, CKM^{+}20, CTL^{+}22, CDdV23]. Metabolism [CRGGSC20, CKM^{+}22, ISC23, JS23, LR25, LPBMA21, Swe20]. metabolites [SL24]. metacommunities [MG21, QPL21, RH23]. metacommunity [BACS24, Cai22, VB21]. metalloproteinases [DOF25]. metapopulation [ASKG25, BYHH20, HAP22, LNT23, MH22a, MM22, dKCA22]. metapopulations [ON24]. metastatic [DBS23, GYY^{+}23, Han23, LF25, PV24]. meteorological [ZZ22]. methicillin [SHH20]. methicillin-resistant [SHH20]. method [CRG^{+}23, DSdP22, GWZ^{+}22, HXK^{+}20, KK22a, LGT21, LFZ^{+}24, LLZ^{+}22, MMS^{+}23, Osb24, PZG20, RGBD25, TMC^{+}23, YAN21, YCKK24, YLT22]. methods [BKP23, KHB^{+}22, MS20, PTZ21, SOGK^{+}23, TH20, WBS20]. methylation [ABC25, DKSP24, GK20, LS21]. metrics [PB21, ZMKF22]. metronomic [WCZL25]. Mexico [FPGOCP^{+}25]. MGMT [LS21]. MI [REK20]. mice [GDQR23, WMHV24]. Michaelis [CLM^{+}25]. micro [DHLS20, IDMG24, BAJ^{+}21]. micro-CT [BAJ^{+}21]. micro-organisms [IDMG24]. microalgae [PP20, PP24, MB21]. microalgal [SSLMGC25]. microalgal-based [SSLMGC25]. microbe [MC22]. microbes [HI21]. microbial [Ala20, BRBF20, CLG20, KHBS20, KYHN23, TMC^{+}23]. microbicidal [BJB^{+}20]. microblog [YSP^{+}23]. microenvironment [GF23, HAP22, HS24, JS23]. microenvironmental [FSK21]. microfibril [CLD21]. microfluidic [CGG^{+}25]. microglia [LR25]. microplastics [WW24]. microRNA [PL20, SOKZ22]. microRNA-mediated [PL20]. microscale [MK24]. microscopic [GHS18, GHS20]. microscopy [LG23]. microtubule [AW22, JMD20, Xie23, Xie25, YSK23]. microtubule-based [JMD20]. microtubules [RKP20]. microvascular [KG23]. microvasculature [HBK^{+}20]. microvesicle [DMR21]. microvessels [FFR23]. midbrain [ZSL23]. migrating [SKNvRL22]. migration [BEBA20, DHLS20, FLO25, HVTA20, HH20, HAGADDM24, MPC^{+}24, OKT^{+}21, PABS23, YL25, YdKdJ^{+}20]. migrations [KH24, LFJ21]. Miller [KJV20]. mimicry [QHS23, TT24]. mineral [RBD^{+}20]. mineral-rich [RBD^{+}20]. mineralization [CPMF^{+}22]. Minimal [LSWVB24, CA21, KG22, MAHMH20]. minimising [FB24]. minimization [Ros21, SL22]. minimize [Ros22]. minimum [CLF^{+}22]. miR [SMS20]. miR-9 [SMS20]. miR34 [NWVN20]. miRNA [BJG21, LS25b]. miRNA-mediated [BJG21]. misuse [TRKB25]. mitigated [Tay23]. mitigation [MBHG^{+}24]. Mitochondrial [QYY^{+}23, FMO23, GCC20]. mixed [Ala23, CBBD22, CBKS21, EB21, KRSG24, RWHK21, SKS21, Sak22a, PK23]. mixed-mode [KRSG24, PK23]. mixing [Avv23, DJ22]. mixotrophs [ADLM24]. mixture [FGA^{+}24]. mixture-like [FGA^{+}24]. MMM [ZSN22]. mobility [MM22, MDF21, RH23]. modal [LFZ^{+}24]. modalities [HAP22]. mode [BAU^{+}22, CP23, HFD23, KRSG24, PK23]. Model [ASK20, BKZ^{+}22, BMB^{+}21, FHR22, PLSS24a, SD25, AGP^{+}21, AG22, AKHM23, Ala20, Ala23, AZdV^{+}22, AHW21, ATH^{+}24, ABK24, ABOR20, ABB^{+}22, AML^{+}22, AGBLGA^{+}22, ACF22, ACP20, AABG20, AMR^{+}23, AZYF^{+}23, ABC25, BCC^{+}21, BRN^{+}22, BvdG20, BBB^{+}25, BPP^{+}20, Bar23, BC21, BGB^{+}21, BBDB22, BH24, BJB^{+}20, BGC^{+}20, BKPT25, BRBF20, Bol22, BM20, BAG^{+}24, BV24, BGGM23, BGA21, BGM22, BPS25, CGY^{+}23, CNK24, CGK23, CRGGSC20, CLF^{+}22, Car23, CBH^{+}22, CBH^{+}23, CMLM23, CF20, CGW22, CDC20, CLSB23, CS20, CCSF21, CGG^{+}25, CBKS21, CSP21, CGGG20, CA21, DHC25, DMR21, DTS21, DN22, DCS20, DDDMT22, DSM^{+}23, DYS^{+}21, DO21, DJ22, ETTT24, ES20, EWS23, EHP^{+}21, ECH^{+}23, EBK21, FG22, FPX20, FFR23, FH23, FPGC21, FBSD22, FBMG25, FPGOCP^{+}25, FS24, FGA^{+}24, FET^{+}24, FPS24, FLR^{+}21]. model [FLRR23, FSC21, DNK^{+}24, GTGB21, GDd^{+}21, GFH20, GMMR22, GSW^{+}23, GCX^{+}20, GCC20, GO21, GGSTS21, GJDS20, GFN23, GBRM21, GRPG24, GBHNHVL23, GW20, GF23, GGG21, GPDW20, Gre25, GHF22, GHS18, GHS20, GYY^{+}23, HIA^{+}22, HBK^{+}20, HAP22, HPG^{+}24, HYM^{+}22, HHT25, Hen20, HCJT23, HBWLLPA21, Hoe24, HXZ^{+}22, HGZM25, ISC23, IH21, IY22, JS23, JME24, JFCK20, JA23, JWKA21, JCLS^{+}11, JCLS^{+}21, JLK22, JKH^{+}23, KT20b, KG22, KR22, KCW^{+}25, KKS23, KK22a, KHW^{+}22, KSSM20, KJM20, KHAK21, Kro25, KRG24, KKK^{+}20, LSWVB24, LR25, LS21, LPP^{+}25, LF25, LR21, LMB22, LYE^{+}24, LPBMA21, LLS^{+}22, LMS21, LX21, LGBG23, LZL^{+}25, LFJ21, LTB25, LMA21, LCYD23, MP22, MLK21, MUMS24, MSB24, MD22, MBMA25, MFGACST23, Mar24, MEPS22, MH22b, MDS25, MI23, MPC^{+}24, MEBNZ24, MF21, MF22, ML21, MDF21]. model [MB24, NCL20, NNOBS23, NWVN20, OLF^{+}24, OGD^{+}20, Ost23, PPS^{+}21, PLXD^{+}20, PLSS24b, PK23, PGC^{+}22, dAPdBMTS20, Pro24, dMQHRGA^{+}22, dVRA20, RZRS^{+}22, RRAO21, Rev25, RLB25, Roy21, RY22, RBW20, SPLS^{+}23, Sak22b, Sal25, SSB20, SLTD22, SKNvRL22, SKNvRL23, SD20, Sch17, Sch23, SMEG22, SF24a, SRDC21, SL21, SLML20, SWK^{+}20, SF23, SF24b, SBW^{+}22a, Sim25, SNK20, SBSB23, SFYL20, SHH20, SWVW25, SD22, SSK^{+}25, SUA^{+}20, SVHM^{+}21, SXJ20, SH21, SBJR21, SZR^{+}25, Swe20, TMTK22, Tay23, TTBL21, TH20, Tsu20, TKRA20, VKS^{+}20, VG25, VA22, VGC22, Vol24, WWR^{+}24, WSY21, WDH^{+}20, WCWL21, WFT23, WDA^{+}23, WSEK23, XYK^{+}20, XZW^{+}24, Xie25, YWQX23, YHT24, YTH^{+}21, YDS24, YL25, YM23, ZSN22, ZG21, ZWC^{+}22, ZWT24, ZYF25, ZLZ^{+}25, dLFO21, dCC^{+}23, EHA^{+}21]. Model-based [BMB^{+}21, FHR22, ECH^{+}23, LPP^{+}25]. modeled [EE23, IJROTE22, ZSS^{+}22]. modelers [ZMKF22]. Modeling [ATAJS20, AW22, BW24, BBH^{+}20, CCF^{+}20, CAARG21, CHT22, CFH^{+}25, CLJ^{+}23, CPM21, DSB^{+}21, DBW22, FFR22, GHB22, Goy21, GQK^{+}21, GDQR23, HBGV^{+}21, JKSD23, KG23, KBS^{+}19, KNY25, MK21, MF24, MDG24, ND24, PGP23, PWC^{+}23, RNP^{+}21, RSPB^{+}22, SJNRS20, Ter23, TH25, UV23, Wal22, WMHV24, ZRL^{+}21, ZZ22, ASL^{+}24, Avv21, BCK24, BS21, BYHH20, BT22, BSV^{+}24, BLP20, CCH^{+}24, CC22, CSH^{+}25, DRO23, ELM25, FLR23, GADDVG^{+}24, GMR^{+}24, GTE20, GFV^{+}23, GEHL^{+}22, HI20, HKBK25, HS20, JHFMP^{+}21, KV23, KLLL22, KTP^{+}20, KM20, KK23, Kuz24c, LRAP22, LHG^{+}22, LSZ21, LCMW21, LB20, LS25a, MMK^{+}24, MHV^{+}21, MAHMH20, MCP20, NDB20, PCLL21, PRFSHV20, Pat22, PHL^{+}23, PBB20, PAM22, RRZDF20, SVVPPR^{+}24, SBC^{+}20, SL23, TCH21, TTBL21, Tha24, VTDM21, WLJ^{+}20, WSC^{+}22, XWS24, XZZ^{+}23, YSP^{+}23]. modeling [YE21, ZQHN24, CCA20, FGES24, KBS^{+}20, LZZ^{+}23]. modelled [CG21]. Modelling [ASSL22, AMSC20, BHW^{+}20, BDB25, CZN^{+}21, CSHH20, CLD21, CLLL20, FG22, GTGB21, GZM^{+}20, HPK^{+}20, HCSAB21, Hil23, Hon20, KPA^{+}21, LMC21, LFY^{+}24, NNG20, PD23, PJVR23, PBD^{+}20, STCSC^{+}21, SMM24, SCHH20, Ste21a, TML20, TGC22, TS23, dJHP20, ASK20, ABBP21, AMA20, AY22, APR^{+}22b, BLRV21, BEBA20, BEO^{+}22, BDdG22, BCBP25, CBK23, CMRMC^{+}21, CN21, DAB^{+}23, DOF25, DLN20, DEM^{+}23, GCC20, GZGH22, GLLP23, GtHH20, HAGADDM24, HMLO^{+}23, MZG^{+}20, NMC^{+}21, NGKB21, PCR23, PASO21, RKP20, SXL^{+}20, SSX^{+}22, WP21, WRB22, XCH^{+}20, ZZGG21, ZBT22, vdVPW^{+}22, SC23a]. Models [BERR20, YRMWT20, AGP^{+}23, ABBP21, AAF^{+}25, AWC25, ASMN23, AR21, ADP21, ASKG25, BJJM25, BCLJ24, BK20a, CK20a, CN20, DSdP22, DCL23, FS22, FSKG20, FVC20, FGC24, GZ24, HTD22, KPH^{+}24, LT23, LK20, LLJ25, LH20, MMG21, MPH^{+}22, MCO21, OGM^{+}25, PLAC22, Pru25, RB22, RT24a, RGBD25, Rob22, RMG^{+}21, Sco24, SBW^{+}22a, SBB^{+}22, SMM24, Sim25, TNHV21, VCS24, WCM^{+}21, WW24, WCL^{+}25, WHRH^{+}23, Wor20, XKV20, XCFH23, YWWF25, YFF25, ZPP23, dDA21]. modes [PJB21, Tsu20, XZW^{+}24]. modification [BSS^{+}20]. modifications [HSL24]. Modified [JKH^{+}23, BAU^{+}22, NAR^{+}20, PDD24]. modular [BEO^{+}22, SY24, TS23, YLZ^{+}21]. modularity [PMCLC^{+}23, SY24]. modularity-disease [SY24]. modulate [KPP20]. modulates [CCPD22, CBZK21, WCD21]. modulating [HS24]. modulation [GK21, HKBK25]. modules [RW20, YLZ^{+}21]. moiety [GHA^{+}20]. Molecular [CRRRP^{+}20, Sch21b, SMS20, Xie23, ACD20, CLM^{+}25, MPTCT^{+}20, Nah20, OSAR21, RK22b, RBG^{+}23, SS23, TH20, Xie24, ZSN22]. molecularly [PK23]. molecularly-explicit [PK23]. molecule [CNK24, KJV20]. molecule-mediated [CNK24]. molecules [AN24, KJV20]. moment [HH21]. moments [Sch18, Sch21a]. monitoring [DSB^{+}21, PLAC22]. monocytogenes [Jac25]. monogamous [HH20]. monolayer [MI23]. monomers [Kuz24c]. Monte [LGT21]. Moon [RBW20]. Moran [AAF^{+}25]. Morphogen [WA22]. Morphogen-directed [WA22]. morphogenesis [MMS^{+}23]. morphogenetic [ELM25]. morphology [BFD25, HXS21, YL25]. mortality [MGCH21, NAR^{+}20, SFYL20, ZQHN24]. mosquito [AWC25, TEWN21]. most [DAB^{+}23]. moth [TKR^{+}21]. moths [GtHH20]. Motif [GW21, Wad22]. motile [MDS25]. motility [Avv21, BM20, PK23, TCMA21]. motion [BFG21, PP20, PP24]. motions [YBF23]. motor [RKP20, Sch21b, Xie21, Xie22, Xie25]. motors [Xie20, Xie23, Xie24]. mound [FOY21]. mouse [CPHL^{+}20, KKK^{+}20, dGSNK22]. mouth [RLW^{+}21]. mouthpart [SAK21]. movement [ABBP21, EP20, FH23, HVC^{+}21, HL24, LFJ21, MI20, OIK^{+}24]. movements [CCPD22]. moving [CCSF21]. mpox [GFAP^{+}25]. MRIs [ZLZ^{+}25]. mRNA [ECBR20]. MSI [HKL25]. MSI-H [HKL25]. MSI-H/dMMR [HKL25]. mtDNA [HD22]. mTORC1 [GMS20]. mu [RE25]. much [DWNW24]. mucin [JME24]. mucosal [AMA20]. mucus [JME24]. muddy [RBD^{+}20]. Muller [BS22a]. Muller’s [KNK21]. Multi [DSdP22, NNOBS23, ZSL23, ASK20, BCGG21, Coh20, Cor21, Cor22, CSDH^{+}25, GSW^{+}23, HSAH20, LPA25, LX21, LFZ^{+}24, LMC^{+}20, LCYD23, MEPS22, MCM20, Osb24, Pro24, SS23, Sin21, TGC22, ZB22, dIMM24]. multi-agent [Pro24]. multi-cellular [Osb24]. multi-drug [BCGG21]. multi-host [Cor21, Cor22, Sin21]. multi-level [ASK20]. Multi-method [DSdP22]. multi-modal [LFZ^{+}24]. multi-objective [ZB22]. multi-omics [LCYD23]. multi-period [dIMM24]. multi-period/multi-phasic [dIMM24]. multi-phasic [dIMM24]. Multi-scale [NNOBS23, Coh20, LX21]. multi-source [LMC^{+}20]. multi-strain [GSW^{+}23, MEPS22, MCM20, TGC22]. Multi-timescale [ZSL23]. multi-trophic [LPA25]. multi-type [CSDH^{+}25, SS23]. multi-year [HSAH20]. multicellular [KIO^{+}25, MCO21, NSMH25, SMEG22, WMO23]. multidrug [GSR23]. multifarious [KKS23]. multifractality [HD25]. multihost [SYH20]. multimers [KB20]. multinomial [SKNvRL22, SKNvRL23]. multiplayer [CWW22]. Multiple [KBD^{+}23, MHI23, dIMM24, AHS23, BvdG20, BV24, CDC20, FRMK22, GQK^{+}21, LMG22, MF21, SBM^{+}20, SXJ20, WSL25, XCH^{+}20, Yam24, vBW20]. multiplied [CLW^{+}24]. Multiscale [PP20, PP24, ST20, BCC^{+}21, CGGG20, DSM^{+}23, SMEG22, SZR^{+}25, YJZ^{+}24, YFF25]. Multisolitons [FET^{+}24]. Multisolitons-like [FET^{+}24]. multispecies [KLC24, RK22b]. Multistability [Pfe22, FLR^{+}21]. multistable [GT22]. multivariate [PLM24]. mumps [APR^{+}22b]. murine [ATH^{+}24]. Murray’s [Ros21]. Muscle [GRW^{+}21, AG22, CDC20, Coh20, GHS18, GHS20, IHO^{+}20, PCLL21, PO20, SL21, XSH^{+}24]. Mushroom [GK21]. mussels [ZSS^{+}22]. mutagenic [SSS21]. Mutant [RDDSA25, EWK20, HII22, HD22]. mutated [ZWC^{+}22]. mutation [AL23, KNK21, KL21b, LZW23, TS23]. mutation-selection [KNK21]. mutations [CWT^{+}23, HBG21, Sak20, Sak22b]. mutual [DB22a, LL20, MHI23]. mutualism [GZ24]. mutualisms [MRK24]. mutualistic [BSC20b, DBD23, KYI21]. mutualists [YJBF20]. mutually [GQG23, YST21]. mutually-repressed [YST21]. Myc [LS25b]. Mycobacteria [FK21]. Mycobacterium [AKUN20, AML^{+}22, JLK22, WPEP22]. mycorrhizal [MGZ^{+}21]. myeloma [BV24]. myocardial [ABB^{+}22, MF22, NMC^{+}21]. myocarditis [vdVPW^{+}22]. myocytes [MCL21].

n [Kur23]. n-player [Kur23]. nACh [WM21a, WM21b]. naïve [KHHI25]. nanoparticle [ASP23]. nasal [CKF^{+}24, PWC^{+}23]. National [KM22]. natural [Akh22, Akh23, Hon20, HL24, Jam21, KT21, WW24]. naturally [CN20, HL21]. naturally-bounded [CN20]. nature [DB22a, GRW^{+}21]. navigation [HCC^{+}21, OTCE24]. NCHB [PZG20]. near [BM20]. nearest [Ios21]. necessary [SLCC22]. neck [SCD^{+}23]. necrotic [BAG^{+}24]. nectar [HZS^{+}21, Rev25]. needed [GFV^{+}23]. negative [BBK23, IIH25, LYE^{+}24, LMS21, SVHM^{+}21]. neighbor [Ios21]. neighboring [ENMW21]. Neisseria [TKB^{+}20]. neither [HHI24]. nematode [EBK21]. neoadjuvant [HKL25]. neocortical [GFN23]. Nepal [AGP^{+}21, AGP^{+}23]. nervous [CGGG20]. nested [NBBR23]. nestedness [Bau20]. net [RP21b]. nets [ASK20, Ngo22, ORT24]. Network [CSH^{+}25, PB21, RT24a, TNHV21, VTDM21, ACG24, AH25, AZYF^{+}23, BSC20a, CLW^{+}24, CPMF^{+}22, CC22, DBS23, DLSA24, EWS23, FLRR23, GFN23, GSMG22, HHT25, LK20, LK25a, LK25b, Nah20, PZXC22, PABS23, dAPdBMTS20, dVRA20, RBG^{+}23, SOKZ22, SMM^{+}23, SCH24, WDT^{+}20, WLW19, WLW20, YZSS20, YST21, YdKdJ^{+}20]. Network-based [RT24a]. network-structured [PABS23]. networks [AGTJ23, ANGJ25, BGB^{+}21, Bau20, BFD25, CP23, CKM^{+}20, CKNCG22, DBD23, EIK^{+}21, EWS23, GT22, IKYW21, KG23, LPP^{+}25, LL20, LWXM20, LS25b, MM22, PJVR23, PPGB20, PB21, PZG20, RMSS24, RW20, RBG^{+}23, Ter23, Tha24, WSL25, WW22, XWS24, YDS24, YLZ^{+}21, dDA21]. Neural [GtHH20, HHT25, LL20, MB24, WDT^{+}20]. Neuro [GT20]. Neuro-Glia [GT20]. neurobiology [Bar21]. neurodegenerative [FSKG20, RB22]. neuromechanical [YWTC25]. neuromodulation [YWTC25]. neuromuscular [WS22]. neuron [BSC20a, SY25, YMR24]. neuron-system [BSC20a]. neuronal [BBK23, BFD25, CHT22, YLZ^{+}21, ZSL23]. neurons [ANGJ25, BRN^{+}22, EE25, GFN23, SHY21, TKR^{+}21]. neurovascular [BSC20a, CGK23]. Neutral [MY20, BJR^{+}22, FVC20, PLM24]. neutralization [PS22a]. neutralizing [dCC^{+}23]. next [AN24]. next-generation [AN24]. niche [MCO21, PASO21, RID23]. Nigeria [MZG^{+}20]. Nile [BGC^{+}20]. NIRS [BSC20a]. nitric [GJDS20]. nitrogen [CBBD22]. NMDA [MP22]. No [Bar23, Goy21, LMA21]. node [ADP21, PZJPZ23, Wad22]. node2vec [MM22]. nodulation [KYI21]. Noise [HLS^{+}25, YM23, BS25, EE25, HGCC22, PEMLD22, SOKZ22, Tsu20, WS22, WNHK21, YFZ^{+}20]. Noise-induced [HLS^{+}25, YM23, SOKZ22]. noises [YM23]. noisy [DF22]. Non [DB22b, FZE22, Xie20, AMSC20, BE22, BB21b, For24, FGC24, GTGB21, HL24, IHG21, JKH^{+}23, Kan21, LRST22, MDS25, MTLZ22, MWO21, Sim25, Täu20, TBNS23, TH21, VCS24, XKV20, YDS24, ZWC^{+}22]. Non-24-h [DB22b]. non-adaptive [Täu20]. non-adaptively [For24]. Non-cooperative [FZE22]. non-enzymatic [TH21]. non-equilibrium [AMSC20]. non-essential [Kan21]. non-exponentially [YDS24]. non-grouping [TBNS23]. non-homogeneous [IHG21]. non-human [HL24]. non-identifiable [Sim25]. non-Kolmogorov [XKV20]. non-Markovian [VCS24]. non-motile [MDS25]. non-pharmaceutical [FGC24, JKH^{+}23, LRST22, ZWC^{+}22]. non-random [MWO21]. non-recessive [BB21b]. non-small [BE22]. Non-tight [Xie20]. non-uniform [GTGB21, MTLZ22]. nonfeeding [IYY22]. Nonlinear [FBJR20, NVS20, Tsu21, ZB22, AHW21, AGTJ23, FNGJ21, GSC23, LPA25, SVHM^{+}21, vBBB^{+}25]. nonlocal [EIK^{+}21, MMG21, VGC22]. nonparametric [CRG^{+}23]. nor [HHI24]. normalization [HS24]. normoxia [CBZK21]. norms [LB25]. norovirus [JAW25]. northwest [XDT^{+}20, ZZ22]. Notch [Pfe22, SMS20, WSL25]. note [ASKG25, TRKB25]. Nothing [EBA21]. Novel [GT20, IB21, ASF22, FSC21, GGSTS21, HXK^{+}20, KRG24, LMC^{+}20, MZLJ20, MS20, PRJ^{+}20, RP21b, Roy21, SLTD22, TMC^{+}23, VGC22, Wu24, YJZ^{+}24]. November [Ano20v, Ano20j, Ano21v, Ano21j, Ano22u, Ano22j, Ano23r, Ano24r]. nuclear [LHG^{+}22]. nucleation [FSKG20, IMM24, RB22]. nucleotide [BEO^{+}22]. nucleus [HGZM25]. null [CCSF21]. number [BTS24, DF22, HMS21, KRSSH21, LMW21, XDT^{+}20]. numbers [WCL^{+}25]. Numerical [KG22, Kuz24c, MPA23, AKHM23, NSZ^{+}25, Osb24, PP20, PP24]. nuptial [IY22]. nurse [MCL20]. nut [ZLYL20]. nutlin [AAL^{+}21]. nutlin-3a [AAL^{+}21]. Nutrient [MSB24, ABB^{+}22, AML^{+}22, ACP20, IDMG24, LGBG23, OTI^{+}20, SFYL20]. nutrient- [AML^{+}22]. nutrient-limited [LGBG23]. nutrient-phytoplankton [SFYL20]. nutrition [HBGW^{+}25]. nutritional [FHTK20].

O [AG22]. obesity [SF24b, VTDM21]. obesity-induced [SF24b]. objective [ZB22]. objects [CRDD22]. obligate [KM22, VB21]. observability [Mar24]. observation [GSR23, MMI21]. Observations [JFCK20]. observer [BAU^{+}22]. obstacles [Pro24]. obstructive [DEM^{+}23]. occur [LS21]. occurrence [GMV^{+}20, HHI24, MGVS21]. ocean [Ios21]. oceans [MK24]. October [Ano20w, Ano20k, Ano21w, Ano21k, Ano22v, Ano23s, Ano24s, Ano25p]. ocular [Ost23]. ODEs [ILGI20]. off [APR^{+}22a, BSS^{+}20, Cai22, MEA21, VP22, WK25]. offs [ADLM24, SUA^{+}20]. Ohio [KBW^{+}23]. oilseed [KCW^{+}25]. old [SL24]. Olfactory [HCC^{+}21, PR21, Bar21, OTCE24, TKR^{+}21]. Oligomerisation [Koc20]. olive [BCS^{+}22, WWR^{+}24]. olives [MUMS24]. Omicron [PD23]. omics [LCYD23]. omnivores [BD22]. on-going [Hon20]. Oncolytic [SC23a, JFCK20]. oncotically [JJR20]. One [GHS18, GHS20, BRN^{+}22, BC21, FET^{+}24]. one-dimensional [BRN^{+}22, BC21, FET^{+}24]. ones [Dal22]. Ongoing [PZJPZ23]. only [Dal22]. onset [CR20, ML21, NSMH25]. opinion [YSP^{+}23]. opinion-transfer [YSP^{+}23]. Optic [YBF23]. optically [SC21]. Optimal [ASP23, ATH^{+}24, CGW22, CS20, CS21, GQG23, HMS21, IDMG24, IMM24, IHS25, Kis24, KYI21, LPLM25, LZL^{+}25, MSLW23, MH22a, PLAC22, UIY22, VMRSL23, WSL25, YCKK24, APL21, BKP23, CD22, DBD23, GCX^{+}20, HPWvdB20, KRG24, LK25a, LK25b, MB21, MSL23, MC20a, NDB20, PHL^{+}23, RG21, Sal25, SYI20, SBM^{+}20, TATT21, TTBL21, ZZB^{+}21]. optimality [BF20, MBF21]. optimally [NTTI22]. Optimisation [BS25, HKL25, KKL20]. optimising [FG22]. Optimization [KK23, DCS20, SGM^{+}25, Wu24]. optimizing [SLTD22]. opting [Kur23, LL21]. opting-out [Kur23, LL21]. option [GDd^{+}21, Kur21, Kur23]. OQDS [BCS^{+}22]. orbit [MC20b]. orbits [LW20]. orchestrates [AVSTS23, GK21]. orchid [SWK^{+}20]. order [GFN23, LMB22, LPD23, PR21, RMG^{+}21, SSR21, Wad22, ANGJ25]. Ordovician [ELRG21]. organ [CKF^{+}24]. organic [IHS25]. organisation [ES20, PRP^{+}20]. organised [AMSC20]. organisms [HL21, IDMG24, ND22, PTK^{+}24]. organization [HJKS20]. organize [STCSC^{+}21]. organized [HL20]. organophosphates [GJCLMQC20]. organotypic [FSC21]. orientation [GLLP23, Hoe24]. orientation-based [Hoe24]. origin [Di 21, JWKA21, KI20, SK21, TH21, Wal20]. originate [AS21]. originated [For24]. Oryctes [CGY^{+}23]. Oscillation [MBMA25, HXZ^{+}22]. Oscillations [SGJ^{+}24, CSF20, CBK^{+}20, LYE^{+}24, MP22, PLXD^{+}20, PK23]. oscillator [MUMS24]. oscillatory [IHO^{+}20, MP22]. OSIRIS [BvdG20]. osmoregulation [UIY22]. osmotic [DN22]. osmotically [DN22]. osteoblast [CPMF^{+}22]. osteochondral [CNK24]. osteogenic [DBTI21]. otter [LZ23]. outbreak [BGM22, CR20, GDC^{+}25, LRST22, MBHG^{+}24, SOGK^{+}23, YLT22, ZRL^{+}21]. outbreaks [APR^{+}22b, BCS^{+}22, BFSA23, DHR25, ENMW21]. outcome [BCBP25, Goy21]. outcomes [BW25, DTLS24, FBA21, JLK22]. outcrossing [YI21]. outer [BCC^{+}21, QYY^{+}23]. output [KMI23, SHY21]. outward [SHY21]. ovarian [FHR22]. overdispersion [GFH20]. overlapping [XML22]. overproduction [Sak22a]. ovules [Sak22a]. ownership [Cai22]. oxide [GJDS20]. OXPHOS [JS23]. oxygen [NM23, TCH21]. oxygenation [SLTD22]. oyster [JCLS^{+}11, JCLS^{+}21].

P [HKBK25]. P-TEFb [HKBK25]. p38 [KLLL22]. p53 [AVSTS23, AAL^{+}21, BAU^{+}22, NWVN20]. packing [RID23]. paddy [LYYN20]. paediatric [SdB23]. PafBC [FK21]. PafBC-mediated [FK21]. pair [AZdV^{+}22, CT22, Sal25]. pair-formation [AZdV^{+}22]. pairwise [LK20]. pancreatic [CCH^{+}24, CHMM23, FVB20, PAM22, WCZL25]. Pandemic [HMLO^{+}23, AGP^{+}23, APF^{+}24, ASS^{+}23, DAB^{+}23, JKH^{+}23, LPLM25, NGKB21, XCH^{+}20, dIMM24]. pandemics [STCSC^{+}21]. paper [WM20]. papillomavirus [GMMR22, Sal25, GTE20]. paradigm [AA22, BH20]. paradox [Sco24]. Parameter [DCS20, FPGOCP^{+}25, KBD^{+}23, SBW^{+}22a, DO21, FHR22, GMR^{+}24, IMM24, KH20, SBB^{+}22, YOIS24]. parameterized [CGG^{+}25]. parameters [GB23, GFV^{+}23, SC21, WHRH^{+}23]. Parasite [HK21]. parasites [CBPG23, TEWN21]. parasitic [MCL20]. parasitoid [KH24, Sin21, VP22]. parasitoids [MGCH21]. parent [Sak22a]. parents [Pag25]. parity [For24]. Park [KM22]. Parkinson [RNP^{+}25]. Parkinson’s [HXZ^{+}22]. parsimonious [FGC24]. Part [GADDVG^{+}24, DAB^{+}23, CTL^{+}22, CDdV23]. partial [SMM24, VGC22]. particle [BLRV21, LFY^{+}24, MK24, WFL^{+}24]. particles [HYM^{+}22, SGM^{+}25]. partner [EBA21, ZWWW23]. passage [BJG21, WA22]. Passive [PCR23, AJR^{+}20, BC21]. past [CKNCG22, Pro24]. patch [FPS24, GSC23, SXJ20, ZZB^{+}21, ZYF25]. paternal [HMS21]. path [LKP^{+}20]. pathogen [BW25, HI20, LT22, MW20, TGC22]. pathogen-host [BW25]. pathogens [FG23, GG23]. pathologies [GADDVG^{+}24]. paths [CLW^{+}24, MMI22, Swe20, YZB20]. pathway [AZYF^{+}23, CSF20, CKM^{+}22, FSNS20, GMS20, KPH^{+}24, LMA21, VTDM21]. Pathways [IYL^{+}25, BBK23, CTL^{+}22, CDdV23, SL24, SMM^{+}23, XYK^{+}20]. Patient [DO21, BG21, DLN20]. Patient-specific [DO21, DLN20]. patients [DEM^{+}23, PAM22, TKRA20, YCKK24]. Pattern [HYM^{+}22, PLXD^{+}20, ZSS^{+}22, BPM^{+}22, CGW22, DEM^{+}23, HS20, HII22, HH20, JMD20, KG22, Kan21, NO24, PP20, PP24, RK22b, WH21, ZWW22]. patterning [LTS^{+}21, WCD21]. Patterns [Kee23, AR20, ABvH^{+}20, CKTW24, CWT^{+}23, EWS23, FET^{+}24, HRD^{+}24, HL20, KH20, LKP^{+}20, MTB24, Pfe22, Pis25, SMM^{+}23, TWT23, TKR^{+}21, WS24, WDHB20]. pause [AW22]. payoffs [KL21b]. PD [BS21, LBF24, SSLMGC25]. PD-1 [BS21]. PD-based [SSLMGC25]. PD-L1 [BS21]. peak [CSDH^{+}25]. peaks [RE25]. peer [LDL^{+}24]. PEF [ZLYL20]. pellets [SVVPPR^{+}24]. pembrolizumab [HKL25]. pendulum [ABOR20]. peptide [KKL20, WDT^{+}20, ZLYL20]. Perceived [SZR^{+}25]. perception [MGFS^{+}20]. performance [CBKS21, FGES24, WM21b]. period/multi [dIMM24]. Periodic [MC20b, ES24, IDMG24, LFJ21, LW20, NKE23, SF24a]. periodical [KMA^{+}21]. periodicity [XDT^{+}20, Yin21]. periodicity-2 [Yin21]. Peripheral [Nak25, PS22b]. Perivascular [dGSNK22, YTW21]. permeability [EE23, SJNRS20]. permeabilization [QYY^{+}23]. permits [ABOR20]. peroxisome [VTDM21]. persistence [BS25, CHSB20, JAB22]. Persistent [Ste21b, SCH24, AABG20]. Personalized [NDB20, YTH^{+}21]. perspective [BCK24, GADDVG^{+}24, MS25, NO25, PV24, RSPB^{+}22, SKS21, Ste21a]. perspectives [ZZB^{+}21]. perturbation [HLW24]. perturbations [QPL21]. pest [BB21b, BPS25, SJS^{+}25]. pesticide [CFH^{+}25]. pesticides [GJCLMQC20]. pests [IIHH21]. petri [ASK20]. pH [AML^{+}22]. pH-stress [AML^{+}22]. phage [Mit20, TTBL21]. phage-bacteria [TTBL21]. phagocytosis [dLFO21]. Phanerozoic [TP24]. Phantom [FVB20]. pharmaceutical [FGC24, JKH^{+}23, LRST22, QWZK24, ZWC^{+}22]. pharmacodynamical [AAL^{+}21]. pharmacodynamics [MPA23]. pharmacokinetics [MPA23]. pharmacological [GCNZ^{+}21, MCL21]. pharmacology [BS21]. Phase [RNP^{+}25, APF^{+}24, AGTJ23, CKM^{+}20, GT22, KBD^{+}23, MG21, PWC^{+}23, SVHM^{+}21]. phase-amplitude [MG21]. phasic [dIMM24]. Phenomenological [HBWLLPA21, AGBLGA^{+}22, LPBMA21]. Phenomenological-based [HBWLLPA21, AGBLGA^{+}22, LPBMA21]. phenomenon [NNNHA24]. Phenotype [BV24, DLN20, AN24, FBSD22, FLRR23, MEA21, RMSS24, WZP^{+}23, WBKR22]. Phenotype- [DLN20]. phenotype-detection [RMSS24]. phenotype-genotype [WBKR22]. Phenotype-structured [BV24, FBSD22, MEA21]. Phenotypes [WPEP22, EB21]. Phenotypic [CBZK21, LS21, CBK23, EWK20, GDLF20, GFL23, MSLW23, PCLL21]. phenotypically [APL21]. Philip [Whi21]. Philippines [ECH^{+}23]. phloem [COB23]. phosphate [SGJ^{+}24]. phosphoinositide [dLFO21]. phosphoinositide-based [dLFO21]. phospholipid [SSKP24]. phosphorus [MDF21, TMTK22]. photoacclimation [MB21]. photobioreactor [PP20, PP24]. Photoinhibition [SC21]. photoreceptor [WCWL21]. phyllotaxis [OY21, Wal22]. phylogenetic [BK20a, CKTW24, GMV^{+}20, Ish23, LT23, MGVS21, PZG20, SSR21]. phylogeny [BPM^{+}22]. physical [IYL^{+}25, LPBMA21, YXW21]. Physically [GCC20]. Physically-plausible [GCC20]. physics [Avv23]. Physiological [FSNS20, FGES24]. physiology [ASSL22]. phytoflagellates [BEBA20]. phytophagous [OYT25]. phytoplankton [MSB24, OLF^{+}24, PB22, SFYL20]. PID [GFN23]. pig [SHH20, ZRL^{+}21]. pigmentosa [Rob22]. Pillars [Don23, Pru25, Sal24, Tha24, Wil23]. PIN7 [Nah20]. pine [ZLYL20]. pinnate [Ros22]. pipeline [ASK20]. pirate [Mit20]. PKC [LYE^{+}24]. place [MSM23]. placental [BAJ^{+}21]. Placoparia [ELRG21]. planaria [CMLM23]. planarian [HL20]. plankton [MSB24]. plankton-fish [MSB24]. planktonic [SMJ20]. planning [MAHMH20]. Plant [MC22, Rev25, ASF22, CCA20, CLD21, DP22, DLSA24, LKP^{+}20, MRK24, MTB24, MGZ^{+}21, NNG20, OYT25, RLB25, Wal22, WWR^{+}24, WLJ^{+}20]. plant-herbivore [MTB24]. Plant-microbe [MC22]. Plant-pollinator [Rev25]. plants [KHST21, MDG24, MCL20, MI20, QHS23, ZG21]. plaque [Kuz24c, MF21, PCLL21]. plaques [Kuz24a]. plasmid [AZdV^{+}22]. Plasmodium [MSK^{+}22, DDDMT22, TEWN21]. plastic [APL21]. plasticity [ADLM24, CHSB20, EWK20, KKS23, WCL^{+}25]. plate [WS22]. platelets [BEO^{+}22]. platform [BS21, RRZDF20]. plausible [GCC20]. player [KL21b, Kur23]. plays [CRGGSC20]. PLC [LYE^{+}24]. pleiotropic [Nah20]. Poacher [HL21]. Poacher-population [HL21]. poaching [HL21]. Pocillopora [HPG^{+}24]. point [HBK^{+}20, LMA21]. points [CSH^{+}25, DBD23, LS25b]. polarization [BBK23, Car23, FLR^{+}21, KJM20, PLXD^{+}20]. policies [WUH23]. policing [CKB^{+}21]. policy [FHM20, TT20, TATT21]. poliomyelitis [CGP^{+}20]. pollen [HMS21, Rev25]. pollinator [MRK24, QHS23, Rev25]. pollinators [HMS21]. pollutants [ZZ22]. pollution [HHT25]. Pólya [PLM24]. Pólya-splitting [PLM24]. polymerization [IMM24, YL25]. Polymorphic [RBZP24]. polymorphism [BK20a, ZF21]. polymorphism-aware [BK20a]. polymorphisms [MRK24]. polyploidization [MDG24]. polyploids [PS22b]. pool [FBA21, SdAM24]. pools [RBD^{+}20]. Poor [ZM20, SdB23]. poorly [BC20]. PopART [LPP^{+}25]. popular [Mar24]. Population [FSKG21, HC24, SKS21, APR^{+}22a, AR20, AWC25, AKUN20, Bai21, BKPT25, BS22a, D’A20a, D’A20b, EWS23, EP20, EBK21, FPS24, GB23, GFAP^{+}25, GG23, HK21, HZM^{+}21, HSAH20, HL21, KHAK21, LGT21, LDF23, LL21, LH20, MGVS21, MTLZ22, MCP20, MB22, MS25, NKE23, PP20, PP24, RDDSA25, RBZP24, RH23, SSB20, SXL^{+}20, ST20, SBW^{+}22a, Sin21, SD22, TBNS23, WZP^{+}23, WSY21, WHRH^{+}23, YI21, ZZB^{+}21]. Population-level [SKS21, GG23, WHRH^{+}23]. populations [AS21, ASS^{+}23, BJR^{+}22, BB21a, BKZ^{+}22, BRS24, CN21, IHG21, LDL^{+}24, LG23, Lio18, Lio24, LH20, LRST22, ML24, MCMdAM24, MEA21, NBBR23, ORW^{+}20, PLAC22, PABS23, PRHS20, PAV21, dVRA20, Sak20, Sak22b, SC23b, SS21, WH21, WM20, WFT23, vBBB^{+}25]. portrait [GZGH22]. position [KIO^{+}25, YBF23]. position-based [KIO^{+}25]. Positive [IIH25, LR25]. Possible [HGZM25, BBK23, JME24, KI21, RR24]. post [PWC^{+}23]. posterior [CSP21, HGF23]. potassium [SL21]. potato [EBK21]. potential [CSHH20, KPA^{+}21, LMA21, MBHG^{+}24, NSD22, PAM22, RMS^{+}25, TKB^{+}20, Wal22, WS22, WUH23, XSH^{+}24]. Potts [TCMA21]. Power [CSVB25, FMO23, MK24, Ros22, WM21a]. power-law [MK24]. Powered [Sac22]. Powered-gliding [Sac22]. Powered-gliding/climbing [Sac22]. practical [SOGK^{+}23, WBS20]. Pre [JJR20, FPGC21, PGC^{+}22]. Pre-atherosclerotic [JJR20]. pre-clinical [PGC^{+}22]. pre-exposure [FPGC21]. prebiotic [CRRRP^{+}20, PPGB20, RBD^{+}20]. precancerous [WHH22]. precision [KMI23]. predation [AM22, BSC20b, BSL24, GtHH20, IDD23, RCM^{+}25, SYH20]. Predator [ASMN23, QPL21, AM22, BB25, BG21, FSKG21, GRPG24, KHBS20, KSSM20, Leh21, LTB25, PPS^{+}21, STB20, XKV20]. predator-prey [AM22, GRPG24, KSSM20, Leh21]. predators [CBPG23, Pis25]. predict [CKM^{+}20, dMQHRGA^{+}22]. predictability [PL20, YLZ^{+}21]. Predicting [LMW21, REK20, SL24, WDT^{+}20, AWC25, LFZ^{+}24, Tsu21]. prediction [BJJM25, CLW^{+}24, DTLS24, Sim25, KYI21]. predictions [CCA20, GFV^{+}23]. Predictive [YRMWT20, AHW21, EPF20, JLK22, PGC^{+}22]. Predictor [BDB25]. predicts [CF20, LSWVB24, LHG^{+}22, Swe20]. preference [OYT25]. preferences [ZWWW23]. preferentially [KKL20]. pregnancy [ASSL22]. PrEP [DHC25]. preparedness [LPLM25]. presence [Ala23, CKB^{+}21, DP22, DHLS20, SFK^{+}23, WZP^{+}23]. present [Kur21]. Preservation [LYYN20]. preserved [AOB21]. preserves [RWHK21]. pressure [DRO23, ETTT24, KM20, TDRC23, WGRE21]. prestressed [ZL20]. presuming [Avv21]. Prevalence [BKP23, AKUN20, Cor21, Cor22, YdKdJ^{+}20]. prevent [LYYN20]. Prevention [BCS^{+}22, IYL^{+}25, ORT24]. prey [ABvH^{+}20, AM22, BSC20b, BG21, CBPG23, FSKG21, GRPG24, KHBS20, KSSM20, Leh21, LTB25, MSM23, PPS^{+}21, PTK^{+}24, Pis25, STB20, TBNS23]. Price [Lio18, Lio24]. primary [Ost23]. principal [PWC^{+}23]. principle [Gav25, WH21]. principles [AN22b, LTS^{+}21]. printed [BMB^{+}21]. prion [FSKG20, GCNZ^{+}21, RB22]. prion-like [FSKG20, RB22]. prioritization [KIM^{+}22]. priority [TT20]. prisoner [FS24, MMI22]. pro [LBF24]. pro-cancer [LBF24]. Probabilistic [BCLJ24, DEM^{+}23, JCC20, GBHNHVL23]. probabilities [RK22b]. Probability [CSDH^{+}25, FVC20, CCSF21, DHC25, EWK20, FSK21, GMV^{+}20, MGVS21, MCM20, YLT22, ZWT24]. probe [PJPH25]. problem [Ios21, MB21]. problems [FBA21]. proboscid [SWK^{+}20]. proboscises [SAK21]. procedure [NDB20]. process [AMSC20, Bai21, BH20, DHR25, EE23, EBK21, GWZ^{+}22, GADDVG^{+}24, HS20, IB21, KCW^{+}25, LMS21, LZW23, NNOBS23, NKE23, NTTI22, SS23, SRB^{+}24, San21, SHK20, ZWT24]. process-based [EBK21, KCW^{+}25]. processes [Bol22, CNK24, GC20, PLM24, Wad22]. processing [DB22a, MB22, SLCC22]. produce [BW25]. produces [ZM20]. producing [IHS25]. production [BLRV21, DDDMT22, Kuz24a, OTI^{+}20, TH25, Tsu21, YI21]. productivity [MTB24, WDHB20]. Professor [Ano21-49, CSW24]. profile [KBD^{+}23, RMSS24]. profiles [DRO23, SWVW25]. progenitor [VGC22]. programmatic [LT22]. programmes [SXL^{+}20]. programming [SYI20]. progression [CBH^{+}22, CBH^{+}23, PCLL21, WMO23, YXW21]. project [YLT22]. Projecting [KBW^{+}23, SFK^{+}23]. projection [HGZM25]. projective [RBW20]. proliferation [Car23, CGR22, GCNZ^{+}21, GFL23, HD22, SK21]. proliferative [MCO21]. proliferator [VTDM21]. proliferator-activated [VTDM21]. promise [MB21]. promote [PHN^{+}23, SW22]. promoter [GTE20, LS21]. promotes [CK20b, DCP^{+}22, JMWX23, LK25a, LK25b, TMTK22]. promoting [Cor21, Cor22, Kro25]. prone [PRHS20]. proneness [LB25]. Propagation [DP22, NM23, LMS21, YSP^{+}23]. properties [BJG21, CK20a, CMRMC^{+}21, GBHNHVL23, HDM20, Mar24, MS20, NKE23, OSAR21, dAPdBMTS20, RNP^{+}21, RG21, WM20, YSK23]. property [AOB21]. Prophylactic [MW20]. proportion [WCL^{+}25]. proposing [KV23]. prosocial [LDL^{+}24]. prospect [HGCC22]. prostate [CPM21, LF25, RNP^{+}21, VVD24]. protection [CWF23, KT21, SSS21]. protein [AVSTS23, BJG21, CLW^{+}24, CWT^{+}23, CSH^{+}25, FSKG20, FET^{+}24, IMM24, KLLL22, LDaH22, LMC^{+}20, Nah20, RKP20, RB22, San21, WDT^{+}20]. protein-peptide [WDT^{+}20]. protein-protein [Nah20]. protein-spreading [FSKG20, RB22]. proteins [BP21, Nah20, Xie21, Xie22]. protest [Vol24]. prothrombinase [GGS^{+}24]. protoxylem [JMD20]. provide [Wad22]. provides [JME24, ZYLW21]. Province [APR^{+}22b, SXJ20]. provinces [ZZ22]. provisions [KT20b]. pruning [SSK^{+}25]. pseudo [WBS20]. pseudo-marginal [WBS20]. public [CPS20, FPX20, FFAV21, JCC20, RP21a]. Puerto [VMRSL23]. Pulmonary [LS25a, DEM^{+}23]. PULSAR [RMS^{+}25]. Pulsatile [DB22a]. pulsations [YTW21]. pulse [Bau20, CKM^{+}20, DB22a]. pulsed [FFGR25, ZLYL20]. pumping [dGSNK22]. punishment [CPS20, FFAV21, JCC20, Mur25]. purine [RP21b]. pyloric [Avv21]. pyrimidine [RP21b].

QAP14 [GYY^{+}23]. QSARs [MPTCT^{+}20]. qualitative [DMPIG24, GDC^{+}25]. quality [BBDB22, BYHH20, EBA21, HL24]. Quantification [YRMWT20, DLN20, OGM^{+}25, RT24a, YHT24]. quantifies [HBK^{+}20]. quantify [YG25]. Quantifying [ACG24, KKK^{+}20, OKF^{+}21, OTI^{+}20, SD22, WCZL25, vBBB^{+}25]. Quantitative [PRFSHV20, PGML20, ABC25, DH22, IY22, YLT22]. quantities [AH25]. quarantine [AHW21]. Quasi [BJR^{+}22, BFG21, MDF21, Sme21, ESW24, Lio18, Lio24]. quasi-equilibrium [Lio18, Lio24]. Quasi-neutral [BJR^{+}22].

Quasi-random [BFG21]. Quasi-species [Sme21]. Quasi-steady [MDF21, ESW24]. quick [BCS^{+}22]. quiescence [Avv21]. quota [SGJ^{+}24].

R [BH24, KK23]. R-loop [BH24]. R0 [IHG21]. rabies [MK21]. radiation [DBW22, For25, MAHMH20, RMS^{+}25, YS25]. radiations [PLR24]. radical [MSK^{+}22]. radiobiology [KK23]. radiotherapy [CBZK21, CBK23, KK23, MEBNZ24, MAHMH20, SLTD22, WCZL25]. radiovirotherapy [ATAJS20]. radius [Kuz24b]. random [ABBP21, ABC^{+}21, BFG21, CKTW24, FS24, FSK21, LWXM20, MWO21, Sch18, Sch21a, SFYL20, YHT24]. randomized [TAP^{+}23]. randomly [UIY22]. Range [LAP24, MGFS^{+}20, MC22, Wad22, WCD21]. Range-limited [LAP24]. ranges [BRN^{+}22, XML22]. rank [Pie21]. Rankings [MBM^{+}25]. rape [KCW^{+}25]. Rapid [Täu20, CRG^{+}23]. rare [MZLJ20]. rat [GMR^{+}24, GEHL^{+}22]. ratchet [BS22a, KNK21]. Rate [DCL23, AN24, AL23, BS22a, CLM^{+}25, CPEK23, CT22, DHC25, GFAP^{+}25, IHO21, KHAK21, Kuz24a, LZL^{+}25, MMS^{+}23, NAR^{+}20, PZJPZ23, RGBD25, Tsu21]. Rate-dependent [DCL23]. rates [AM22, BRN^{+}22, BLRV21, FVC20, GFL23, LZW23, MGFS^{+}20, MTLZ22, RS24, ZYLW21]. Ratio [RS24, Gar23, HPSA21, HPSA23, JTB25, NAR^{+}20, WSY21]. Ratio-dependent [RS24, WSY21]. ratios [JTB25, SSR21]. rats [SL23]. RdCVFL [WCWL21]. Reacting [ENMW21]. reaction [BNP^{+}21, CGW22, EIK^{+}21, GJDS20, KG22, KBS^{+}19, KBS^{+}20, KB22, PLXD^{+}20, PLAC22, PPGB20]. reaction-diffusion [KB22]. reactions [SSS21]. reactivation [BPP^{+}20, NVS20]. Reactivity [LW20, LLZS25, RE25]. real [EE25, HCC^{+}21, KK22a, Wal22]. real-time [KK22a]. realizations [LML24]. really [TT20]. rearrangement [GPDW20]. reasoning [PAM24]. rebound [GDQR23, HHI24]. receptor [VTDM21, WM21a, WM21b]. receptors [CCVB^{+}21]. recessive [BB21b]. recipient [HMS21]. reciprocal [CSVB25]. reciprocity [LMB22, MPA24, Mur25, PAM24]. reckless [BG21]. recognition [Sco24, ZWW22]. recombination [RWHK21]. recommended [GSR23]. reconcile [dGSNK22]. reconstructed [GMV^{+}20, MGVS21]. Reconstructing [AKUN20]. recorded [HRD^{+}24]. recovered [JKH^{+}23]. recovery [BAU^{+}22, PBB20, SKNvRL23]. recurrence [AGTJ23, YS25]. recycling [TMTK22]. red [HDM20]. reduce [BD21]. reduced [JS23, RH23]. Reducing [CK20b, SF24b, SCD^{+}23]. reduction [ASK20, RRAO21, Sim25]. reductionism [Bar21]. Redundancy [MEA21]. Redundancy-selection [MEA21]. reef [HPG^{+}24, JCLS^{+}11, JCLS^{+}21]. reef-building [HPG^{+}24]. reefs [ZYF25]. reencounter [SKNvRL22]. reflection [GFH20, SKS21]. reflux [Avv21]. refuges [BB21b]. refusal [PB21]. regeneration [ABB^{+}22, CNK24, CMLM23, HFD23, Lei20, Lei24, MFGACST23]. regenerative [DBW22]. regime [SF24a]. regimen [WCZL25]. regimens [MBM^{+}25]. region [ACF22]. regions [HMLO^{+}23]. registration [ZLZ^{+}25]. registry [HRD^{+}24]. regress [RKCvB25]. regular [BK20b]. regularization [LLZ^{+}22]. regularly [SMM^{+}23]. regulated [YST21]. regulates [NTTI22, QWZK24]. regulating [QYY^{+}23]. regulation [AG22, CCH^{+}24, CHMM23, DOF25, FHR22, FHTK20, GTE20, GK20, KLLL22, KTU22, LMA21, MB22, PGML20, PEMLD22, PL20, SOKZ22, SJNRS20, SSKP24, WvdDL^{+}21, XZW^{+}24, YBF23, YST21, YTH^{+}21]. regulator [LMA21]. regulators [FSKG21]. Regulatory [LB20, CPMF^{+}22, DBS23, FLRR23, GT22, HL20, HBG21, HLW24, HGZM25, KV23, RW20, Tha24, WW22]. reinfection [BRAHV25, CLJ^{+}23, Kee23, NVS20]. relapses [FRMK22]. relapsing [FRMK22]. relapsing-remitting [FRMK22]. related [MGCH21, NNNHA24, PLSS24a, PLSS24b, SBW^{+}22b]. relatedness [QVE25]. relating [HL24]. relations [Wal20]. Relationship [TAP^{+}23, JS23, SW22, SYH20]. relationships [BAJ^{+}21, GSC23]. relative [CR20, GFAP^{+}25, TBNS23, WCL^{+}25]. relatives [KCO25]. relaxation [CLM^{+}25]. relaxed [RK22b]. release [GSC23]. releases [BB21b]. Reliable [SBB^{+}22]. remating [IY22]. Remembrance [Ano21-49]. remitting [FRMK22]. remodeling [FZE22, GADDVG^{+}24, PJB21, SRB^{+}24]. removal [LCYD23, WHH22, Xie25]. renin [TDRC23]. renin-angiotensin [TDRC23]. reopening [ASS^{+}23]. reorientation [CLD21, YZB20]. repair [AZYF^{+}23]. repeat [IB21, WUH23]. repeated [CWW22, MMI21, MMI22]. repeats [YY21]. reperfusion [NMC^{+}21]. replace [LMLP20]. replication [BJB^{+}20, LZZ20, PZJPZ23, TH21]. Replicator [AR21, FBMG25, LFT^{+}23, MC20b, SH21]. repopulation [LR21]. Report [CKI20]. reported [Kee23, LMW21]. reports [YTY^{+}20]. representation [GtHH20, Kan21, LDaH22, NR20, YG25]. Representing [CKM^{+}22]. repressed [YST21]. repressive [ADP21, Vol24]. reproduce [FBSD22, SPLS^{+}23]. reproduction [BB21a, DF22, HSAH20, XDT^{+}20, Yam24]. reproductive [AM22, KRSSH21, MSM23, PLR24, Sme21]. reputations [MPA24]. research [DAB^{+}23]. reservoir [HHI24]. reservoirs [HVC^{+}21]. residence [Xie23]. residency [MGFS^{+}20]. residents [AKUN20]. resistance [Ala20, Avv21, BCGG21, BV24, BB21b, DMR21, DYS^{+}21, EBK21, FB24, GHF22, GDLF20, HDM20, HII22, HHI24, HPWvdB20, LKP^{+}20, Ngo22, PV24, Ros21, Sch17, Sch23, SNK20, SP20]. resistant [CPM21, PJPH25, SHH20, VVD24]. resolution [AVP^{+}25, KPH^{+}24]. Resolving [CT22]. Resource [KHST21, MUMS24, FHTK20, IDMG24, MB21, MRK24, MC20a, OKT^{+}21, RK22a, SdAM24, SLCC22, VVD24, WK25, Wor20, YTG24, EHA^{+}21]. resource-based [VVD24]. resources [BP23, FBA21]. respiratory [CWF23, GJCLMQC20]. Response [LB25, TKRA20, WPEP22, AVSTS23, ACD20, ACF22, BBB^{+}25, BPP^{+}20, Bau20, BT22, BLP20, CAARG21, CBZK21, CLLL20, CKM^{+}20, DAB^{+}23, DCP^{+}22, FG23, FBA21, FK21, For25, FHTK20, KHW^{+}22, KSSM20, LPSMCP23, LLS^{+}22, LFZ^{+}24, LT22, MF24, MEBNZ24, OKF^{+}21, QPL21, SBC^{+}20, SMEG22, SVHM^{+}21, SH21, Ter23, TKR^{+}21, Vol24, Wal20, WMHV24, XYK^{+}20, XCFH23, ZZGG21]. responses [ACF22, EB21, FGES24, HDM20, LRAP22, SPLS^{+}23]. responsible [SSM^{+}22]. resting [AG22]. restrict [BJB^{+}20]. restriction [JMD20]. result [LPD23]. resulting [ZXLH23]. results [GDQR23]. retaining [CKF^{+}24]. reticulum [DLSA24, PO20]. retina [ABK24, BCC^{+}21, FSNS20, YBF23]. retinal [SS22]. retinitis [Rob22]. Retrospective [KvLBD25]. return [LMA21]. reveal [DB22a, FLR^{+}21, LWM23]. Revealing [RMS^{+}25]. reveals [ASF22, AABG20, BLRV21, CLD21, GGSTS21, Goy21]. reversal [Leh21]. review [GLLP23]. revisit [JS23]. revisited [Bol22, CK20b]. Revisiting [BPS25, Gav25, Mar24, MY20, SY24]. reward [JCC20]. rhinoceros [CGY^{+}23, CGY^{+}23]. rho [KJM20]. rhythm [KMI23]. rhythms [BCK24, GK20, RE25, ZG21]. ribosomal [NCL20]. riboswitches [PRFSHV20]. ribozyme [RBD^{+}20]. rice [LYYN20]. rich [RBD^{+}20, TMTK22]. richness [FVC20, PS22a]. Rico [VMRSL23]. RIG [SBC^{+}20]. RIG-I [SBC^{+}20]. ring [YRMWT20]. Rinzel [KK22a]. risk [AGP^{+}23, BRAHV25, GtHH20, LWM23, MHPB23, NVS20, RLW^{+}21, SSB20, SCD^{+}23, SZR^{+}25, VCS24, YWWF25, ZXLH23]. risks [LRST22, SOGK^{+}23]. risky [CPS20]. rituximab [BT22]. RNA [Akh23, Akh22, GK20, RP21b, TH21]. RNA-based [Akh23, Akh22]. RNAP [NCL20]. road [DP22]. Robust [BAU^{+}22, JMD20, LH20, CKF^{+}24]. robustness [ACG24, ADP21, DMAC21, IKYW21, WDA^{+}23]. rock [BFG21, WM20]. rods [ABK24]. Role [Avv23, PCLL21, PZXC22, WPEP22, AMA20, AM22, BE22, BB21a, BFSA23, CRGGSC20, CLM^{+}25, EB21, FOY21, GDd^{+}21, GBRM21, Goy21, GDLF20, HS20, HBGW^{+}25, ISC23, JD21, KT20b, Kro25, LPP^{+}25, Leh21, LPBMA21, LWM23, LT22, LMA21, LS25b, MMS^{+}23, NSMH25, NWVN20, NSD22, PGP23, PR21, PP22, PJPH25, RZRS^{+}22, RY22, RSPB^{+}22, RH23, SMJ20, SCD^{+}23, WA22, WCWL21, XSH^{+}24, YTG24, YE21]. roll [ECH^{+}23]. roll-out [ECH^{+}23]. Rolling [TH21, YZB20]. Rolling-circle [TH21]. root [CLF^{+}22, FS22, KYI21, Kuz24b, LKP^{+}20, NNG20]. root-to-leaf [LKP^{+}20]. rooted [PZG20]. roots [WLJ^{+}20]. ROP [JMD20]. Rose [ANGJ25]. Rosenzweig [Bar23]. rotation [YST21]. route [Lio18, Lio24]. routes [SD22]. routinely [HRD^{+}24]. RSC [LCYD23]. RSC-based [LCYD23]. rubrum [KGP^{+}22]. rugged [RID23]. Rule [WCM^{+}21, APL21, For24, HZM^{+}21, JCC20, Leh22]. Rule-based [WCM^{+}21]. Rules [GRW^{+}21, Leh22, ZMKF22]. rumor [GRPG24]. Run [GRW^{+}21]. Runaway [LPD23]. running [ABOR20, GRW^{+}21, SF24a]. ruthless [Bar21].

Saccharomyces [San21]. sacoglossan [YYI21]. saddles [WA22]. Saharan [LPP^{+}25]. same [SW22]. sample [LSW20, TAP^{+}23]. samples [SC21]. Sampling [SL22, CGS^{+}25, DPR21]. sampling-based [CGS^{+}25]. sanctuary [PZJPZ23]. sarcoplasmic [PO20]. SARS [ACF22, BLRV21, BFSA23, CWF23, CLJ^{+}23, CPEK23, CWT^{+}23, CSH^{+}25, GFV^{+}23, GFC^{+}22, Hil23, JMWX23, KvLBD25, KRSSH21, Kee23, KBD^{+}23, LRST22, NSZ^{+}25, PD23, PWC^{+}23, PDD24, Roy21, TAP^{+}23, WUH23, WHRH^{+}23, XCFH23, YY21, Yin21, dCC^{+}23]. SARS-CoV- [KRSSH21, PD23, Yin21]. SARS-CoV-2 [ACF22, BLRV21, BFSA23, CWF23, CLJ^{+}23, CPEK23, CWT^{+}23, CSH^{+}25, GFV^{+}23, GFC^{+}22, Hil23, JMWX23, KvLBD25, Kee23, KBD^{+}23, LRST22, NSZ^{+}25, PWC^{+}23, PDD24, Roy21, TAP^{+}23, WUH23, WHRH^{+}23, XCFH23, YY21, dCC^{+}23]. sauropod [KI21]. savanna [ES20, BGA21]. scaffold [MFGACST23]. scaffolding [KLLL22]. scaffolds [BMB^{+}21]. Scalable [OV25]. scalar [AOB21]. scale [CCPD22, Coh20, FGES24, LX21, MBM^{+}25, MZG^{+}20, NNOBS23, dMQHRGA^{+}22, Täu20, ZRL^{+}21]. scale-free [CCPD22]. scales [GHS18, GHS20, HVC^{+}21, PJB21, RCM^{+}25, vBW20]. scaling [LHG^{+}22]. scarlet [ZZ22]. scavengers [KM22]. scenario [KI20]. schedule [IHS25]. scheduled [BAU^{+}22]. scheme [LMC^{+}20]. schistosomiasis [VGTY21]. Schooling [CRDD22]. scientific [TRKB25]. scissors [WM20]. scleractinian [WDA^{+}23]. sclerosis [AHS23, FRMK22, MF21]. scrambled [HJKS20]. screening [SXL^{+}20]. scrutiny [GHF22]. Sea [WSY21, OKT^{+}21, YYI21]. seal [CKF^{+}24]. search [CGS^{+}25, HCC^{+}21, IH21]. searching [Kan21]. seasonal [IHS25, ML21, STB20]. seasonality [APR^{+}22b]. seasonally [KHB^{+}22]. seasons [HSAH20]. Second [AZdV^{+}22, For24, LMB22, LPD23, REK20]. second-order [LMB22, LPD23]. secondary [RY22, San21]. secretion [GZM^{+}20]. secrets [CKNCG22]. sectional [YAN21]. sections [Wu24]. sediment [JCLS^{+}11, JCLS^{+}21, ZSS^{+}22]. Seed [PB22, KP21]. seeing [PZ20]. segment [HJKS20]. segmental [Avv23]. segregation [dKCA22]. seizures [HGZM25, RLL25]. Selection [AN22b, TP24, AN24, ASF22, ABvH^{+}20, AR21, BC20, CS22, FBD23, GO21, Hon20, Jam21, Kau23, KNK21, LS21, LL21, LFT^{+}23, LGSR21, Lio18, Lio24, MEA21, MEBNZ24, ORW^{+}20, PAV21, RRAO21, RWHK21, SC23b, SBW^{+}22a, SBJR21, WS24, WFT23, ZF21]. selective [IIH25, Kro25]. Self [dKCA22, AMSC20, CPMF^{+}22, ES20, HL20, NGKB21, PRP^{+}20]. self-inhibition [CPMF^{+}22]. self-isolate [NGKB21]. self-organisation [ES20, PRP^{+}20]. self-organised [AMSC20]. self-organized [HL20]. Self-segregation [dKCA22]. selfing [Sak20, Sak22b]. Selfish [Pag25, MHPB23]. selfishness [MC20a]. Semi [BH23, LPBMA21]. Semi-autonomous [BH23]. semi-physical [LPBMA21]. senescence [BF20]. senile [Kuz24a, Kuz24c]. senolytic [LF25]. sensing [DH22, PCR21]. sensitive [JAW25, MBM^{+}25, PJPH25, Wad22]. sensitivities [FHR22]. Sensitivity [KRG24, Ala23, APF^{+}24, Bar23, BD22, Cor21, Cor22, DSdP22, FLR^{+}21, GSMG22, IYL^{+}25, KBD^{+}23, RRAO21, Swe20, VGTY21, DMAC21]. sensitivity-based [GSMG22]. sensitizes [FK21]. sensory [TKR^{+}21, Wad22]. Seoul [JKH^{+}23]. Separate [HSAH20, Rev25]. Separation [BA24, HS20, dMQHRGA^{+}22, SVHM^{+}21]. sepsis [RZRS^{+}22]. September [Ano20x, Ano20l, Ano21x, Ano21l, Ano22w, Ano22k, Ano23t, Ano23i, Ano24t, Ano25q]. sequence [NCL20, YG25]. sequence-dependent [NCL20]. sequences [CK20a, GW21, HBG21, LDaH22]. sequencing [AN24, SS23, WCZL25]. serial [YAN21, ZZT^{+}21]. series [LDaH22, Wad22]. seroepidemiological [YAN21]. serological [GFV^{+}23]. serovar [KT20a]. services [MBT22, QHS23]. sessile [CG21]. set [MPTCT^{+}20]. Sets [Rav21]. setting [AJR^{+}20]. settings [MCM20]. seven [Nah20]. severe [BSV^{+}24]. sex [ASSL22, GMMR22, Gar23, HPSA21, HPSA23, JTB25, SW22, SL23]. sex-ratio [JTB25]. Sexes [CK20b]. sexual [LPP^{+}25, ZXLH23]. sexually [HPSA21, HPSA23]. SGL [HXK^{+}20]. SGL-SVM [HXK^{+}20]. Shanghai [ZWW22]. shape [CRRRP^{+}20, LKP^{+}20, MTB24, Nak25, QVE25, QPL21, Ste21b, Tsu20, WK25, dIMM24]. shaped [ASMN23]. shapes [APR^{+}22a, Haj20, KSSM20, SWVW25]. sharing [HPD20, MC20a, MHPB23]. Sharp [KMA^{+}21, BM20]. shear [DBTI21]. shedding [AGBLGA^{+}22]. sheet [YTW21, ZPP23, dJHP20]. sheet-based [ZPP23]. shell [KBS^{+}19, KBS^{+}20, KB22]. shift [DB22b, KR22]. ships [MBHG^{+}24]. shock [ZZGG21]. shocks [DSB^{+}21]. shorebird [HH20]. short [For25, PL20, SWK^{+}20]. short- [PL20]. shortening [AW22]. shows [BJB^{+}20, IH21, WDA^{+}23]. Shrinkage [ZZT^{+}21]. sib [GCM23]. sideways [Xie21, Xie22]. Sidney [Whi21]. sigmoid [SBW^{+}22a]. Signal [Koc20, Wal20, BSC20a, Tsu21, Wad22, YLZ^{+}21]. signaling [ACD20, BEO^{+}22, CF20, DB22a, GHB22, LPD23, LML24, MF24, MCL21, MB22, Pfe22, WSL25, WCD21, XWS24]. signalling [CNK24, CSF20, CMRMC^{+}21, CKM^{+}20, GZGH22, HGCC22]. signals [LPD23, NSD22, PB21, SD22]. significantly [PAV21]. silico [BAJ^{+}21, FSC21, WGRE21, XWS24, SBSB23, TKB^{+}20, AVP^{+}25, HIA^{+}22, BH20]. simian [OKF^{+}21]. simian/human [OKF^{+}21]. similarity [BFD25, CLW^{+}24, LDaH22, YG25]. Simple [OTCE24, dVRA20, BRN^{+}22, BRBF20, DPR21, HCC^{+}21]. Simplified [XCFH23, DNK^{+}24, GCC20, HXZ^{+}22, Sco24]. Simulating [MEBNZ24, Wu24, YZB20, MI23]. Simulation [GEHL^{+}22, WS24, BTS24, ELRG21, dGSNK22, PO20, SCHH20, YFF25, ZG21]. Simulations [PJPH25, AKHM23, FSC21, LZZ^{+}23, Osb24]. Sina [YSP^{+}23]. Sina-microblog [YSP^{+}23]. Singaporean [FPX20]. single [ABC25, DRO23, FVB20, HAGADDM24, JHFMP^{+}21, KPH^{+}24, SSKS18, SSKS20, TCMA21, WM21a, WM21b, YM23]. single-cell [KPH^{+}24]. single-channel [JHFMP^{+}21]. single-locus [ABC25]. sink [MGZ^{+}21]. SIR [LK25a, CGW22, FPGOCP^{+}25, Gre25, LK25b, Whi21]. SIRS [EB24, MBMA25]. SIS [LMG22]. Sisyphus [BFG21]. Site [RK22b, Sch21b]. sites [CRDD22, GPDW20, NNNHA24, PZJPZ23, WDT^{+}20]. size [AABG20, BD21, BGM22, FMO23, GB23, GMR^{+}24, HC24, KTU22, Kur21, KB20, MGVS21, MGCH21, NBBR23, RH23, TAP^{+}23, XDZ^{+}21, ZYLW21]. size-related [MGCH21]. size-spectrum [XDZ^{+}21]. sizes [LSWVB24]. skeletal [AG22, GHS18, GHS20, PO20, SL21]. slave [GFN23]. sleep [ASL^{+}24, BBH^{+}20, CBKS21, DB22b, FNGJ21, GEHL^{+}22, MMK^{+}24, PBB20]. sliding [BAU^{+}22]. SLIP [SF24a]. slow [PZJPZ23, Sch21b, WA22]. Slowly [SHY21]. slugs [YYI21]. small [Avv23, BJR^{+}22, BE22, DWNW24, Xie24]. smegmatis [AML^{+}22]. smooth [Coh20, IHO^{+}20, PCLL21, XSH^{+}24]. snapshot [LGT21]. Social [CCPD22, SW22, BD21, CS20, CS21, CSVB25, CC22, CA21, DB22b, GGSTS21, GCA^{+}20, Hoe24, IH23, JAB22, KT20b, LPLM25, LB25, ML24, OU22, PB21, SH21, TBNS23, WLW19, WLW20, Wil23, YSP^{+}23]. social-ecological [JAB22, SH21]. social-epidemiological [PB21]. socially [TATT21]. societies [Hon20]. socio [DBD23]. socio-mutualistic [DBD23]. sodium [DCL23, PLSS24a, PLSS24b]. soft [BC21, Tsu20]. soil [HVTA20, MDF21, NNG20]. soil-phosphorus [MDF21]. soil-transmitted [HVTA20]. solid [CBZK21, HIA^{+}22, dMQHRGA^{+}22]. solute [JJR20]. solution [SD20, SF24a]. solutions [DMAC21, PLM24]. Solving [FS24]. soma [RR24, TG22]. Somatic [BC20, HRD^{+}24, TS23]. song [MPH^{+}22]. source [CRRRP^{+}20, LMC^{+}20, SXJ20]. South [LMW21]. space [AABG20, CCSF21, HAGADDM24, PJPH25, SKNvRL23]. spaces [IS20, LSW20, YTW21]. spacing [ABC^{+}21, HZS^{+}21]. sparse [HXK^{+}20]. Spatial [ATP22, ES20, FHTK20, HVC^{+}21, HI21, LNT23, XDZ^{+}21, CS22, EP20, GCA^{+}20, ILGI20, MBM^{+}25, OYT25, PLAC22, PZ20, PJB21, RH23, SMJ20, SSBG20, WMO23, ZLYL20, ZWT24, vBW20]. spatial-temporal [ZWT24]. Spatially [ABK24, BYHH20, FSKG20, LZZ20, RB22, BB21a, D’A20a, D’A20b, ISC23, KMA^{+}21, SD22]. Spatially-dependent [ABK24]. Spatially-extended [FSKG20, RB22]. Spatio [CBK23, LF25, WDH^{+}20, BE22, MD22, POW23, Tsu20, WDA^{+}23]. spatio-stochastic [MD22]. Spatio-temporal [CBK23, LF25, WDH^{+}20, BE22, POW23, Tsu20, WDA^{+}23]. Spatiotemporal [ZWW22, Pfe22, SMEG22, SP20]. spawners [BYHH20]. specialization [AM22, TG22, Whi20]. speciation [KLC24]. species [BP20, BA24, DP22, ES20, FVC20, GZ24, GRPG24, JTB25, MMG21, PS22a, PMD21, PZ20, PB22, QVE25, RT24b, RP21b, SdAM24, Sme21, UV23, YM23, ZF21, ZX22, RH23]. specific [CWT^{+}23, DCL23, DO21, DLN20, GT20, KPH^{+}24]. spectra [ES24]. spectral [RE25]. spectrum [WM21a, XDZ^{+}21]. speed [GRW^{+}21, KPP20, XML22, BDB25]. speeds [ABOR20]. Spike [YLZ^{+}21, CWT^{+}23, CSH^{+}25]. spiking [BRN^{+}22, LL20, PR21]. spinal [LSZ21]. spiral [OY21]. spite [BS22b, Kur23]. Split [ACMT24]. splitting [GHA^{+}20, PLM24]. spontaneous [Gre25, WH21]. sporogony [TEWN21]. spot [CLM^{+}25]. spots [Tsu20]. spread [AHW21, BGB^{+}21, BGC^{+}20, HPSA21, HPSA23, HFD23, ILGI20, LK25a, LK25b, MD22, MN23, PGP23, RGBD25, RLW^{+}21, SRDC21, SHH20, WWR^{+}24, YTY^{+}20]. spreader [GRPG24]. Spreading [SS22, XML22, CGK23, CRGGSC20, FSKG20, GCA^{+}20, LWXM20, RB22, TT20, TATT21]. sprouting [NNOBS23]. Square [FS22]. Stability [LLZS25, ASMN23, Don23, DLSA24, ES24, GZ24, GSMG22, KRSG24, LMB22, LFT^{+}23, MTB25, SF24a]. stabilizes [Sin21]. stabilizing [Bau20, BYHH20]. stabilizing/destabilizing [BYHH20]. stable [GK21, MIP^{+}21, MC20b, MS25]. staff [WUH23]. stage [EBK21, Han23, LWM23, PPS^{+}21, Sak20, SGJ^{+}24, SJNRS20, YDS24]. stage-dependent [LWM23]. stage-structured [EBK21, PPS^{+}21]. stages [MPA23, RG21, VGC22]. stance [GRW^{+}21, SF24a]. standard [LZL^{+}25, MB24]. Staphylococcus [BJB^{+}20, SHH20]. state [AR21, CG21, DCL23, ESW24, HTD22, KPH^{+}24, WSW21]. state-dependent [WSW21]. state-specific [DCL23]. states [ACD20, AABG20, CKM^{+}20, KPH^{+}24, SOKZ22, ZL20, GFV^{+}23, KRSSH21, WSC^{+}22]. Stationary [Sch18, Sch21a, PLM24]. Statistical [GFL23, HHT25, KH20, WLJ^{+}20, CK20a, SSKS18, SSKS20, TH20]. statistics [MZLJ20]. status [LS21, RDDSA25]. Staying [VB21]. steady [ESW24, HTD22, MDF21, SOKZ22, SVHM^{+}21]. steady-flow [SVHM^{+}21]. stealing [SVHM^{+}21]. stem [ABB^{+}22, BC20, EWK20, For24, FBC23, KCW^{+}25, LSZ21, Lei20, Lei24, PASO21, SK21, WvdDL^{+}21, XZW^{+}24, YWQX23]. stem-cell [LSZ21]. stem-loops [For24]. stems [HH21]. stent [LFY^{+}24, WFL^{+}24]. step [OY21]. Sterile [IIHH21]. stevioside [BSS^{+}20]. STI [YCKK24]. stickiness [HMS21]. stiffness [Haj20]. stimulation [GEHL^{+}22, YZSS20]. stimuli [TKR^{+}21]. stimulus [CKM^{+}22, DBTI21, YFZ^{+}20]. Stochastic [CK20a, GTE20, HKBK25, KGP^{+}22, RKP20, ZG21, BKPT25, BRBF20, DCS20, EE25, FLRR23, HII22, HHT25, HCR20, KHAK21, LS21, LFT^{+}23, LTB25, MH22a, MD22, MCMdAM24, MPC^{+}24, NKE23, OGM^{+}25, SSB20, SBB^{+}22, WSW21, WZP^{+}23, WCL^{+}25, WHH22, YWWF25, YFF25, BDB25]. stochastic-based [OGM^{+}25]. Stochasticity [JAB22, Bar23, PS22a]. Stochasticity-induced [JAB22]. Stoichiometric [WW24, GHA^{+}20, SMJ20]. stomatitis [PDD24]. storage [PS22a]. straightness [Nak25]. strain [CLJ^{+}23, DBTI21, GSW^{+}23, LMG22, MEPS22, MCM20, TGC22]. strains [TWT23]. strand [AZYF^{+}23, TH21]. strand-displacement [TH21]. stranded [BH24]. Strategic [Whi21, GRPG24]. strategies [ADLM24, BSL24, CWW22, CS20, CS21, GMMR22, GW20, GQG23, HPWvdB20, HCC^{+}21, IHG21, KT21, KIM^{+}22, Kis24, KNY25, KRG24, MSLW23, MMI21, MC20a, Mit20, MMI22, MS25, RT24b, RWHK21, Sal25, STCSC^{+}21, TNHV21, TNW21, WSW21, ZBT22]. strategy [AHW21, BP23, BK20b, BPM^{+}22, DBD23, ECH^{+}23, FBMG25, HMLO^{+}23, KPP20, KCO25, LL21, MSL23, SKS21, ZZB^{+}21]. strategy-dependent [FBMG25]. streaming [DLSA24]. strength [ASMN23, MGZ^{+}21, STB20]. stress [AML^{+}22, HH21, SLCC22, ZL20]. stress-gradient [SLCC22]. stress/uniform [ZL20]. stresses [MIP^{+}21]. stressors [BvdG20]. stretch [GLLP23, ZL20]. stroke [WP21]. strong [Bar23, MIP^{+}21]. Structural [CWT^{+}23, GHA^{+}20, CK20a, IKYW21, PB21, SVVPPR^{+}24]. structural-statistical [CK20a]. Structure [BAJ^{+}21, Van20, AH25, BSS^{+}20, BK21, DBD23, EPF20, GHS18, GHS20, HPGF25, LPA25, RBG^{+}23, San21, SY24, Tak24, WW22, WSY21, Wor20, ZF21, ZYLW21].

Structure-function [BAJ^{+}21]. structured [AABG20, APR^{+}22b, BB21a, BV24, CBH^{+}22, CBH^{+}23, DDDMT22, EBK21, FPGC21, FBSD22, FBD23, HK21, HTD22, LL20, LDL^{+}24, Lio18, Lio24, LRST22, MCMdAM24, MEA21, ORW^{+}20, PPS^{+}21, PABS23, SSBG20, SMM^{+}23]. structures [BAJ^{+}21, FOY21, NKE23, PS22b, RK22a, RBD^{+}20]. strut [WFL^{+}24]. studies [BEO^{+}22, GW20, PGC^{+}22, XWS24]. Study [CWL20, AGP^{+}23, AVP^{+}25, Avv23, APR^{+}22b, BJJM25, BBK23, BTS24, BCBP25, CWT^{+}23, EP20, FNGJ21, GMMR22, GJDS20, GSR23, GDC^{+}25, GK20, GBHNHVL23, HS20, JAW25, KBD^{+}23, LLS^{+}22, LX21, LRST22, MC20b, NMC^{+}21, NKE23, NSZ^{+}25, PCLL21, PZ20, PBB20, PO20, RBG^{+}23, RLB25, Sch17, Sch23, SRDC21, SXL^{+}20, SSX^{+}22, WWR^{+}24, WFL^{+}24, WDH^{+}20, WSC^{+}22, XZZ^{+}23, YSP^{+}23, ZZGG21, ZQHN24, ZBT22]. Studying [ILGI20, KV23, HKBK25, OKT^{+}21]. sub [SR24, SHY21, LPP^{+}25]. sub-exponential [SR24]. sub-harmonic [SHY21]. Sub-Saharan [LPP^{+}25]. subclonal [YJZ^{+}24]. subcutaneous [HBWLLPA21]. subfield [MP22]. subgenomes [PS22b]. subject [DSB^{+}21]. subjective [RSPB^{+}22]. subordinated [LMS21]. subpopulation [HPK^{+}20]. subsidizing [TT20]. subsidy [TATT21]. Subsistence [GCM23]. subspecies [BHW^{+}20]. substrate [GGS^{+}24]. subthalamic [HGZM25]. subtree [CKTW24]. Success [BB25, BKPT25, MSM23, RP21a]. successful [AS21]. succession [PLR24]. suction [NNG20]. suggests [ELRG21, LSZ21, SAK21]. SUMOylation [ZZGG21]. sunfish [Ios21]. Super [AVP^{+}25, BB25]. super-predator [BB25]. Super-resolution [AVP^{+}25]. superinfection [GW20, Hen20]. superoxide [TH25]. superspreading [BFSA23, DHR25]. supplementary [SSX^{+}22]. supply [CLLL20, KHHI25, RK22a]. support [BCLJ24, DAB^{+}23, HXK^{+}20, YBF23]. suppress [IIHH21, TT20, TATT21]. Suppression [Tsu20, MHI23]. surae [CDC20]. surface [HAGADDM24, NSD22, TCMA21, YSK23]. surgical [LMLP20]. surrounding [FBSD22]. surveillance [OGM^{+}25, vdBMB^{+}23]. surveys [GFV^{+}23]. survival [RT24b, SKNvRL23, TP24, ZX22]. survive [SKNvRL22]. Survivor [KL21a]. Susceptibility [GJCLMQC20, GFC^{+}22, HGD^{+}25, Kis24, RMG^{+}21, TK25]. susceptible [GGSTS21, JKH^{+}23, LK20]. susceptible-exposed-infectious-recovered [JKH^{+}23]. susceptible-infected-susceptible [LK20]. sustainability [FBA21]. Sustainable [NNNHA24]. sustains [Mur25]. SVM [HXK^{+}20]. Sweden [KPA^{+}21]. sweeps [Kau23]. swimming [CG21, YTW21, ZPP23]. swine [ZRL^{+}21]. switch [FRMK22, GZM^{+}20, SBSB23, WCZL25]. switches [DB22a]. Switching [IHO^{+}20, ASMN23, GDLF20, GFL23, MSLW23, PCLL21]. symbiont [KRSG24, VB21]. symbionts [BRBF20, VB21]. symbiosis [MC22]. Symbiotic [FFAV21]. symmetric [KL21b, LSZ21, SF24a, YWQX23]. Symmetry [RP21b, WH21, Wu24]. symptom [CR20]. synapse [KR22]. synaptic [KKS23, SMM^{+}23]. Synchronization [ANGJ25, AR20, AGTJ23, BACS24, GFN23, RNP^{+}25, SMM^{+}23]. synchronized [MUMS24]. synchrony [KH24, QPL21, YdKdJ^{+}20]. syncytia [JD21]. syndrome [BCS^{+}22]. synergies [RMS^{+}25]. synergistic [DMAC21, HS24, MGZ^{+}21, RNP^{+}21, SCD^{+}23]. synergistically [IIHH21]. synergy [MAHMH20, PD23]. synthesis [ZLZ^{+}25]. synthetic [RBG^{+}23, XWS24]. system [BSC20a, BB25, BT22, CGG^{+}25, FLR23, FGA^{+}24, FSKG21, GF23, JSCW21, KMI23, KYI21, KH24, LS25a, MTB24, MHI23, MAHMH20, PASO21, RR24, RRZDF20, SdAM24, YWTC25, ZZB^{+}21]. Systematic [HGF23, HLW24, LL20]. systemic [Jac25]. Systems [KTP^{+}20, ACMT24, BP20, BS21, BEO^{+}22, CBBD22, CHSB20, GCM23, GCC20, JAB22, KIO^{+}25, KRSG24, LSW20, LTS^{+}21, MMS^{+}23, OKT^{+}21, RG21, Sal24, SK23, TDRC23, TH20, Wal20, WBS20].

T [BCK24, CF20, CWL20, HPK^{+}20, KV23, KHHI25, KKL20, PGC^{+}22, RY22, WMHV24]. T-cell [HPK^{+}20]. T2D [IYL^{+}25]. table [RP21b]. Tag [BS22b]. Tag-based [BS22b]. tagging [KKS23]. Tale [CCVB^{+}21]. tandem [GGG21]. target [DMAC21, SPLS^{+}23]. target-cell [SPLS^{+}23]. targeted [IIHH21, WSL25]. targeting [ZLYL20]. targets [AN24, HGCC22, TKB^{+}20]. task [EE25]. tasks [BTS24, Yam24]. taste [BSS^{+}20]. tau [RE25]. Taylor [YTW21]. TBD [TNW21]. TCP [LR21]. teaching [VA22]. technique [IIHH21, MSL23]. techniques [JTB25]. TEFb [HKBK25]. teleost [LSZ21]. telescoping [KPP20]. Temperate [CWR21]. Temperature [ADP21, QWZK24, BGC^{+}20, EBK21, FGES24, KCW^{+}25]. temperature-driven [BGC^{+}20]. Temporal [AR20, HII22, MB22, SWK^{+}20, BCC^{+}21, BE22, CBK23, ILGI20, LF25, PS22a, POW23, PR21, Tsu20, WDH^{+}20, WDA^{+}23, ZWT24]. temporarily [FS24]. tenase [SSKP24]. tension [LKP^{+}20]. term [BBH^{+}20, FGC24, HD24, MPH^{+}22, MWO21, PL20, Rav21, SXL^{+}20]. termite [FOY21]. terrain [BFG21]. terrestrial [PCR23]. Territorial [ABC^{+}21]. tertiary [JTB25]. test [MZLJ20]. Testing [WDA^{+}23, BP23, CD22, CS21, CPEK23, Täu20, WUH23]. testosterone [Jam21]. tethering [KB20]. TFT [KR22]. TGF [GHB22]. TGF- [GHB22]. thalamocortical [FNGJ21, PZXC22, YZSS20]. thanks [CGG^{+}25]. their [ABBP21, BH24, CKNCG22, CN21, Haj20, KK22b, LPP^{+}25, MWO21, RE25, SKNvRL22, Xie23]. Theor [Akh23, CBH^{+}23, Cor22, HPSA23, JCLS^{+}21, Lio24, LK25a, PLSS24a, Sch21a, Sch23, SSKS20, Xie22]. Theoret [Lei24, PP24]. theoretic [AR21, CS20, NO25]. Theoretical [Avv21, CPMF^{+}22, D’A20a, Don23, GB23, GK20, GHS20, NV22, PCR23, WLW20, CCA20, CGP^{+}20, CTL^{+}22, CDdV23, FFR23, GCX^{+}20, MDS25, MCL20, Pru25, dMQHRGA^{+}22, Sal24, Tay23, ZZB^{+}21]. Theory [KJV20, MPTCT^{+}20, RK22b, Akh22, Akh23, AA22, DMPIG24, KI21, Kan21, KYI21, KP21, dGSNK22, LYYN20, Leh22, LLJ25, MB24, POS21, PLM24, RR24, TRKB25, WBKR22, ZYLW21]. therapeutic [TKB^{+}20]. therapeutics [PRJ^{+}20]. therapies [SLML20, SBM^{+}20]. therapy [Ala23, ABB^{+}22, BCGG21, BKPT25, BH20, CGGG20, DMAC21, FBC23, GSR23, HKL25, HD24, LR21, PHL^{+}23, PGC^{+}22, RMS^{+}25, SC23a, SF23, SAG21, WSL25, YS25]. thereby [Wad22]. Thermal [FGES24]. thermodynamics [Tsu21]. thermoperiods [HLS^{+}25]. thermoplasty [DLN20]. thermoregulation [BBH^{+}20, ZM20]. Thermotogales [KI20]. thick [SC21]. thickness [SCD^{+}23]. thin [BMB^{+}21]. thinning [JME24]. Thornley [RLB25]. threat [KHB^{+}22]. Three [ABBP21, Di 20, Ros22, ADP21, BSC20a, CLW^{+}24, DAB^{+}23, MI23, TWT23, XZW^{+}24, YST21]. Three-dimensional [ABBP21, MI23]. three-gene [YST21]. three-node [ADP21]. threshold [CS22, GFC^{+}22, KR22]. thresholds [BNPL20, Jac25]. thrombus [LFY^{+}24, WFL^{+}24]. throughout [OSAR21]. throughput [OTI^{+}20]. thyroid [YTH^{+}21]. tick [CPHL^{+}20, TNW21]. tick-borne [TNW21]. tick-mouse [CPHL^{+}20]. ticket [TATT21]. tight [Xie20]. Time [FPGOCP^{+}25, AN22a, BKP23, BJG21, BD21, BRBF20, BK20b, CK20b, CSDH^{+}25, EWS23, FBMG25, HBG21, HD25, KG23, KK22a, KC22, Kuz24b, LR21, LDaH22, LWM23, LG23, OLF^{+}24, PRHS20, SSLMGC25, TH21, YST21, GC20]. Time-delay [FPGOCP^{+}25]. time-delayed [YST21]. time-dependent [LR21, OLF^{+}24]. time-lapse [LG23]. time-varying [LWM23, SSLMGC25]. timeline [IJROTE22]. times [FVC20, KLC24, Sch18, Sch21a, Xie23]. timescale [ZSL23]. timescales [BA24]. timing [AOB21, CR20, GGSTS21, KNY25, RDDSA25]. tip [RCM^{+}25, SWVW25, dJHP20]. tipping [DBD23]. tissue [BC21, BMB^{+}21, Car23, CGGG20, DOF25, DCL23, DWNW24, GO21, HBWLLPA21, KF20, LGBG23, MI23, Osb24, SLTD22, TCH21, WZKM21]. tissues [CBK23, Coh20, HCSAB21, WvdDL^{+}21, WMO23]. TMPRSS2 [PD23]. toe [ASSL22]. tolerance [JKSD23]. tomato [HLS^{+}25]. tongue [HZS^{+}21]. top [BD22, SMJ20]. top-down [SMJ20]. topics [YSP^{+}23]. topographically [MPC^{+}24]. topography [RE25, TCMA21]. Topological [Ish23, SD22, BP21, Ste21a]. topologies [EWS23]. topology [dAPdBMTS20, RH23, SCH24, WLJ^{+}20]. tornadoes [BSJ20]. total [ESW24, FGC24, LKP^{+}20, PZJPZ23, RH23]. toxicants [BBB^{+}25]. toxicity [MSB24]. Trabecular [SCD^{+}23]. trachomatis [AMA20]. Tracing [OSAR21, AHEB^{+}22, BGM22]. tracing-induced [AHEB^{+}22]. track [FLO25]. tract [ATP22, CWF23]. Trade [APR^{+}22a, ADLM24, BGB^{+}21, Cai22, HL21, MEA21, SUA^{+}20, WK25]. Trade-off [APR^{+}22a, Cai22, MEA21, WK25]. trade-offs [ADLM24, SUA^{+}20]. Tradeoffs [YWTC25]. trading [KHW^{+}22]. traffic [Sch21b]. tragedy [KHST21]. training [HHT25]. trait [CN21, DH22, IS20, LPA25, vBBB^{+}25]. trait-matching [LPA25]. traits [ASF22, APL21, BTS24, CN20, CN21, LMG22, LT22, MWO21, vBW20]. transactivation [HKBK25]. Transcription [ACD20, BH24, NCL20]. Transcription-based [ACD20]. transcription-dependent [BH24]. transcriptional [LMA21, OV25, PRFSHV20]. Transduction [Koc20, PTK^{+}24, Tsu21, Wal20]. transfer [CBBD22, DMR21, NM23, NR20, YSP^{+}23]. transform [DH22]. transformation [WP21, BRS24]. transgenic [BB21b]. Transient [KM20, BNPL20, SOKZ22]. transit [KG23]. Transition [SLML20, ANGJ25, CG21, GHB22, HLW24, HSL24, RLL25]. Transitions [YFZ^{+}20, FLRR23, NO24, Pfe22]. translation [BJG21, San21]. transmissibility [HGD^{+}25, TK25, YAN21]. Transmission [AGP^{+}21, ZXLH23, APF^{+}24, AH25, BRAHV25, BW25, CPHL^{+}20, CRG^{+}23, GSW^{+}23, GFV^{+}23, HFD23, HHT25, KRSG24, KP21, KRG24, LPP^{+}25, LX21, LZZ^{+}23, LWM23, SSX^{+}22, SY24, SCH24, TWT23, WXS25, WUH23, YLZ^{+}21, ZZT^{+}21, ZWC^{+}22]. transmitted [HPSA21, HPSA23, HVTA20]. Transplantation [CMLM23, WZKM21]. transport [AG22, ASP23, Avv23, COB23, EHP^{+}21, GCC20, HBWLLPA21, JJR20, KBS^{+}19, KBS^{+}20, KB22, KK22b, MI20, PEMLD22, RKP20, YSK23]. transported [BFG21]. transporting [KK22b]. transposable [MDG24]. transverse [HH21]. trap [ABBP21, FS24, MGCH21]. travel [KC24]. travelers [LPLM25]. travelling [ES24]. treated [Ngo22, ORT24, PZJPZ23]. Treatment [XZZ^{+}23, BBDB22, CBK23, DTS21, DCS20, FB23, FB24, FRMK22, GCNZ^{+}21, HIA^{+}22, HAP22, HS24, Hen20, HCR20, LF25, LPSMCP23, MCL21, MBM^{+}25, MF22, NMC^{+}21, PZXC22, PDD24, RNP^{+}21, SC23a, WCZL25, YTH^{+}21]. treatments [SLTD22]. tree [GMV^{+}20, HD25, IHS25, MGVS21, NM23, QVE25, TS23, TIS25, ZYLW21]. trees [BBR^{+}21, CKTW24, Ish23, PS22b, SSKS18, SSKS20, SSR21, Tay23, WWR^{+}24]. trial [BCLJ24]. trials [DHC25, ILGI20, TAP^{+}23]. triceps [CDC20]. trilobite [ELRG21]. triple [KT20b]. triple-dilemma [KT20b]. triplets [PZG20]. tristeza [ILGI20]. tritrophic [ACP20, BACS24]. tRNA [RBD^{+}20]. tRNA-like [RBD^{+}20]. trophic [CN20, LPA25, LLZS25, QPL21, Rob22, RH23]. tropical [CRDD22]. True [KPR20]. trust [JCC20]. trust-based [JCC20]. TSTDP [KR22]. tuberculosis [BJJM25, CZN^{+}21, FPGC21, MBM^{+}25, NVS20, PBD^{+}20, AKUN20, JLK22]. tubulin [Xie25]. Tumor [HAP22, SRB^{+}24, AABG20, BCLJ24, BT22, BAG^{+}24, CLSB23, CGG^{+}25, FBSD22, FSK21, FGA^{+}24, GDd^{+}21, GF23, HS24, HXK^{+}20, JS23, LPSMCP23, MPA23, MF24, MS20, MEBNZ24, PTZ21, SLTD22, YS25]. tumor-immune [BT22, FGA^{+}24]. tumor-on-a-chip [BAG^{+}24]. Tumorigenesis [CHMM23, AABG20, CCH^{+}24]. tumors [FBC23, SLTD22, WvdDL^{+}21]. Tumour [PBP23, CBK23, HIA^{+}22, LLS^{+}22, OGD^{+}20, dMQHRGA^{+}22]. tumour-immune [OGD^{+}20]. tumoural [FB23, FB24]. tumours [ASP23, CBZK21, MDCK23, PV24]. tuna [CRDD22]. tunable [LMLP20]. tuned [LSW20]. tuning [TH20]. turbulent [HCC^{+}21]. Turing [HL20, KH20, LTS^{+}21, Pis25]. turtle [OKT^{+}21, WSY21]. twin [HIA^{+}22]. twist [CLD21]. twisting [CLD21]. Two [BK20b, FLO25, KNK21, OY21, AG22, BKZ^{+}22, CCVB^{+}21, Di 20, EP20, FPS24, GMMR22, GRPG24, HYM^{+}22, HAGADDM24, KL21b, MPTCT^{+}20, Mit20, NNNHA24, Sak22b, SdAM24, SGJ^{+}24, TATT21, UV23, YL25, ZZB^{+}21, ZYF25, AZdV^{+}22, Sak20]. two-allele [Sak22b]. two-body [HYM^{+}22]. Two-compartment [AZdV^{+}22]. two-dimensional [EP20, YL25]. Two-lane [FLO25]. two-layer [AG22]. two-locus [Sak22b]. two-patch [FPS24, ZZB^{+}21, ZYF25]. two-player [KL21b]. two-sex [GMMR22]. two-space [HAGADDM24]. two-species [SdAM24]. two-stage [SGJ^{+}24, Sak20]. Two-step [OY21]. Two-strategy [BK20b]. type [CSDH^{+}25, HSL24, IB21, KSSM20, MPH^{+}22, SS23, SF24b, Swe20, Wad22, XYK^{+}20, DNK^{+}24]. type-1 [Wad22]. type-I [XYK^{+}20]. types [BKZ^{+}22, CLSB23, IYY22]. typical [HGZM25].

U [Wal20]. U/U [Wal20]. U87 [HBK^{+}20]. U87-glioblastoma [HBK^{+}20]. UK [CPEK23, DAB^{+}23, KvLBD25]. ultrasensitive [FZE22]. Ultrasensitivity [Koc20]. ultrasound [AVP^{+}25]. uncertainties [Ala23]. Uncertainty [OGM^{+}25, WHRH^{+}23, YHT24, YRMWT20, CLG20, DLN20, RT24a]. uncoating [NTTI22]. unconditional [Qui20]. uncoupling [HD24]. uncover [YTG24]. Uncovering [PAM22, AGP^{+}21, SY24]. uncovers [JKSD23]. underlie [FVB20]. underlying [BW24, Jac25]. Understanding [BEBA20, DOF25, DDDMT22, GDLF20, HGCC22, LGSR21, SBW^{+}22b, Lei20, Lei24, MBF21, Pie21, WvdDL^{+}21]. undesired [SSS21]. unexpected [GGSTS21]. uniaxial [GTGB21, LGBG23]. unidirectional [KK22b]. Unification [RE25]. unified [YTH^{+}21]. uniform [CKTW24, GTGB21, MTLZ22, ZL20, MIP^{+}21]. unifying [BSV^{+}24, LB20]. unimolecular [CP23]. unique [SS23]. uniquely [CP23]. United [GFV^{+}23, KRSSH21, LMW21, WSC^{+}22]. universal [Akh22, Akh23]. universe [KJV20]. unknown [SSS21]. unlabelled [PS22b]. unlike [KK22b]. unlikely [RR24]. unnatural [MGCH21]. Unraveling [SK21]. unreasonable [ESW24]. unreported [LMW21]. unrooted [CKTW24]. until [HBG21]. Unveiling [MMS^{+}23]. upon [BCK24]. uptake [MDF21, PWC^{+}23, SAK21, SGJ^{+}24]. Urey [KJV20]. USA [JAW25]. usage [PD23]. Use [YG25, CLSB23, ORT24, TBNS23]. used [SSR21]. Using [Cor21, Cor22, HRD^{+}24, Pie21, TH20, AKHM23, AGTJ23, BAU^{+}22, BEBA20, CWL20, CN21, DTLS24, FPGOCP^{+}25, GJDS20, GW20, GEHL^{+}22, HKL25, HCR20, IHG21, KT20a, LPSMCP23, LG23, LZW23, MSL23, MS20, NDB20, RW20, SS23, SSLMGC25, SBM^{+}20, SBB^{+}22, SWVW25, VCS24, WBKR22, WDA^{+}23, Wu24, YAN21]. Usutu [HBGV^{+}21]. utilization [BP23]. utilizing [Ish23]. Utkin [BAU^{+}22].

Vacancies [PHN^{+}23]. vaccinated [ASS^{+}23, GG23]. Vaccination [Sal25, WLW19, WLW20, Bai21, BPP^{+}20, BCK24, CCF^{+}20, CPHL^{+}20, CLLL20, CGP^{+}20, CS20, DTS21, FHM20, GFAP^{+}25, GW20, KT20a, KRSSH21, KPA^{+}21, Kis24, LZZ^{+}23, LZL^{+}25, SFK^{+}23, TT20, dCC^{+}23, ASS^{+}23, GMMR22]. Vaccine [APR^{+}22b, JMWX23, AMA20, BDdG22, BCBP25, CLLL20, ECH^{+}23, GFAP^{+}25, KIM^{+}22, NAR^{+}20, PB21, TNHV21]. vaccine-modified [NAR^{+}20]. vaccines [FPGC21, Kis24, KKL20, XZZ^{+}23]. validation [AML^{+}22, MPTCT^{+}20]. Valsalva [RRAO21]. value [GGG21, LYYN20, Sme21, YWTC25, vdBMB^{+}23]. variability [FHR22, KL21b, OV25, PWC^{+}23, QWZK24]. variable [HC24, Kis24]. Variance [ZYLW21]. variant [BC20, CWT^{+}23, LK25a, LK25b, PD23]. variant-specific [CWT^{+}23]. variants [BRS24, CLJ^{+}23, CWT^{+}23, CSDH^{+}25, JMWX23, MZLJ20, ZWC^{+}22]. variation [CBZK21, GT22, GFC^{+}22, Sak22a, vBW20, vBBB^{+}25]. various [HZS^{+}21, PZ20, ZBT22]. varying [KHB^{+}22, KPP20, LWM23, SSLMGC25, SD22]. vascular [BAJ^{+}21, IHO^{+}20, MPA23, PCLL21]. vascularised [ASP23]. vasculature [SLTD22]. vasculogenesis [VGC22]. vasomotion [FFR22]. vector [AY22, GW20, HBGW^{+}25, HXK^{+}20, SUA^{+}20]. vector-borne [HBGW^{+}25]. Vectors [GT20]. vegetation [KG22, ZSS^{+}22]. velocity [RBZP24, YL25]. venation [Ros22]. ventilator [MHV^{+}21]. ventilator-induced [MHV^{+}21]. Vero [NSZ^{+}25]. versatile [Pfe22]. version [SKNvRL23]. versus [ABC^{+}21, Cor21, Cor22, GSMG22, IYY22, XKV20]. vertebrate [BCC^{+}21, GZGH22, Ish23]. vertex [JA23]. vertical [BEBA20, BW25, MK24]. very [PZJPZ23]. vesicular [PDD24]. vessel [GDd^{+}21, Ter23]. vestige [MTB25]. via [BH23, CSF20, COB23, DBD23, GFH20, GF23, HXK^{+}20, KK23, MCMdAM24, NNOBS23, PAM22, RE25, RMS^{+}25, Täu20, WRB22, XSH^{+}24, YBF23, dJHP20]. vicinity [KP21]. view [RLB25]. vigilance [MH22b]. Viral [BGB^{+}21, HHI24, BLRV21, BRS24, CAARG21, CHSB20, CWT^{+}23, GDQR23, HFD23, HBGW^{+}25, JD21, ND24, PWC^{+}23, SD20, SdB23, WMO23]. virotherapy [JFCK20]. virtual [ATH^{+}24, JLK22, SLTD22]. Virulence [FG21, IHO21, BB21a]. Virus [SD25, BNP^{+}21, BPP^{+}20, BGC^{+}20, CHSB20, CWR21, HII22, HBGV^{+}21, HBGW^{+}25, ILGI20, IHO21, LMC21, OKF^{+}21, PDD24, SGM^{+}25, SC23a, VMRSL23, ZRL^{+}21, WDH^{+}20]. viruses [SSM^{+}22, WSC^{+}22]. viscoelasticity [BC21]. viscosities [HZS^{+}21]. viscous [dJHP20]. visual [FSNS20, LRAP22, Ost23, RG21]. vitro [AMSC20, NSZ^{+}25, NDB20, FSC21, AML^{+}22]. vivax [MSK^{+}22]. vivo [LZZ20, MFGACST23]. VLDLTG [Swe20]. volatile [IHS25, KL21a]. voltage [KR22, KP21, RLL25, WM21a, WM21b]. voltage-depolarization [WM21b]. voltage-gated [RLL25]. Volterra [CLSB23, MMG21]. volume [DN22, MEBNZ24]. volumes [LHG^{+}22]. voluntary [KT20a, ORT24]. volunteer [vdBMB^{+}23]. Voronoi [JFCK20]. vortex [ZPP23]. vs [WM21b, dIMM24]. vulture [KM22]. vulture-jackal-anthrax [KM22].

W. [Gar23]. waiting [HBG21]. wake [DB22b]. walk [ABBP21, BFG21]. walking [HXS21, KPP20, Xie25]. walled [SWVW25]. waning [CCF^{+}20, EB24, SFK^{+}23]. Warburg [ZB22]. warming [QWZK24]. warning [NSD22, PB21, SD22]. washing [AN24]. washing/selection [AN24]. Washtenaw [REK20]. wasting [XML22]. water [CKF^{+}24, MI20, OTCE24]. wave [ELM25, FBJR20, PLSS24a, PLSS24b, REK20]. wave-wave [FBJR20]. waveforms [YZSS20]. Waves [IH23, ES24, FPGOCP^{+}25, Vol24, XCH^{+}20, dIMM24, dLFO21]. Weak [CS22, LFT^{+}23]. web [ASMN23, BACS24, WDHB20]. Weber [Wad22]. webs [BD22, LPA25, LLZS25]. weekend [LR21]. weight [SF24b]. weighted [MDCK23]. weights [CP23]. were [DAB^{+}23]. West [BGC^{+}20]. whale [MPH^{+}22]. whales [Whi20]. Where [AS21, SKNvRL22, ABOR20]. which [SSR21, TATT21]. White [ZYF25, WNHK21]. Who [APF^{+}24, GSR23]. whole [JHFMP^{+}21]. whole-cell [JHFMP^{+}21]. whom [APF^{+}24]. wide [BRN^{+}22]. widens [AZdV^{+}22, MC22]. widespread [PB21]. wild [IIHH21, ON24]. wildlife [RGBD25]. Willebrand [KB20]. windkessel [SVHM^{+}21]. windkessel-based [SVHM^{+}21]. wings [LCMW21, MIP^{+}21]. winners [PPFB24]. winter [KCW^{+}25]. WITHDRAWN [PMD21]. within [AMA20, BRAHV25, CT22, CSH^{+}25, FG22, HII22, LMC21, MCMdAM24, PR21, PBD^{+}20, PO20, SW22, ST20, SA25, SZR^{+}25, TEWN21, TGC22, TIS25, TKR^{+}21, UV23, WWR^{+}24, WH21, WMO23, XYK^{+}20, XCFH23, dCC^{+}23]. within- [FG22]. within-and [SA25]. within-host [AMA20, BRAHV25, LMC21, SZR^{+}25, TGC22, UV23, XYK^{+}20, XCFH23, dCC^{+}23]. within-mosquito [TEWN21]. within-pair [CT22]. without [LS21, TTBL21]. wizardry [Leh22]. Wnt [CSF20, GZGH22]. Wnt/ [CSF20]. Wolbachia [AY22, MSL23]. Wolpert [Ano21-49]. woody [dAPdBMTS20]. words [LAP24]. work [DB22b]. workflow [PP20, PP24]. workforce [STCSC^{+}21]. Working [HXS21]. Workplace [AMR^{+}23]. world [HCC^{+}21]. worlds [TATT21]. worm [EHP^{+}21]. wound [BH23, HPG^{+}24, HCJT23, LS25a]. wounds [DOF25]. Wright [AAF^{+}25]. Wuhan [YLT22].

X [SSKP24]. XBB [CLJ^{+}23]. xenograft [HIA^{+}22]. XIAP [AVSTS23]. Xylella [BCS^{+}22, WWR^{+}24]. xylem [WWR^{+}24].

year [HSAH20]. years [New20]. yeast [GTGB21]. yellow [GW20]. Yield [NNNHA24].

zebra [MK21]. zebrafish [BLP20, CSP21]. zero [CWW22, MMI21, MMI22]. zero-determinant [CWW22, MMI21, MMI22]. Zika [VMRSL23]. zipper [NTTI22]. zombie [IH21]. zone [YSK23].

References

Arencibia-Albite:2022:EBT

[AA22] Francisco Arencibia-Albite. The energy balance theory is an inconsistent paradigm. Journal of Theoretical Biology, 550(??):??, October 7, 2022. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519322002351.

Atsou:2020:SSS

[AABG20] Kevin Atsou, Fabienne Anjuère, Véronique M. Braud, and Thierry Goudon. A size and space structured model describing interactions of tumor cells with immune cells reveals cancer persistent equilibrium states in tumorigenesis. Journal of Theoretical Biology, 490(??):??, April 7, 2020. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519320300199.

Alexandre:2025:BWF

[AAF^{+}25] Arthur Alexandre, Alia Abbara, Cecilia Fruet, Claude Loverdo, and Anne-Florence Bitbol. Bridging Wright– Fisher and Moran models. Journal of Theoretical Biology, 599(??):??, February 21, 2025. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519324003151.

Awan:2021:EPI

[AAL^{+}21] Muhammad Suleman Awan, Maria Aslam, Muwahida Liaquat, A. I. Bhatti, and Afrose Liaquat. Effect of pharmacodynamical interaction between nutlin-3a and aspirin in the activation of p53. Journal of Theoretical Biology, 522(??):??, August 7, 2021. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519321001181.

Andreucci:2022:MMC

[ABB^{+}22] D. Andreucci, A. M. Bersani, E. Bersani, F. J. León Trujillo, and S. Marconi. A 3D mathematical model of coupled stem cell-nutrient dynamics in myocardial regeneration therapy. Journal of Theoretical Biology, 537(??):??, March 21, 2022. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519322000212.

Ahmed:2021:TDR

[ABBP21] D. A. Ahmed, S. Benhamou, M. B. Bonsall, and S. V. Petrovskii. Three-dimensional random walk models of individual animal movement and their application to trap counts modelling. Journal of Theoretical Biology, 524(??):??, September 7, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519321001508.

Akemann:2021:TBB

[ABC^{+}21] Gernot Akemann, Michael Baake, Nayden Chakarov, Oliver Krüger, Adam Mielke, Meinolf Ottensmann, and Rebecca Werdehausen. Territorial behaviour of buzzards versus random matrix spacing distributions. Journal of Theoretical Biology, 509(??):??, January 21, 2021. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519320303301.

Ayres:2025:SLQ

[ABC25] L. Ayres, H. Bovenhuis, and M. P. L. Calus. A single-locus quantitative genetic model incorporating DNA methylation. Journal of Theoretical Biology, 607(??):??, June 21, 2025. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519325000761.

Anderson:2024:SDM

[ABK24] Daniel M. Anderson, Danielle C. Brager, and Anthony J. Kearsley. Spatially-dependent model for rods and cones in the retina. Journal of Theoretical Biology, 579(??):??, February 21, 2024. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519323002849.

Andrada:2020:LLC

[ABOR20] Emanuel Andrada, Reinhard Blickhan, Naomichi Ogihara, and Christian Rode. Low leg compliance permits grounded running at speeds where the inverted pendulum model gets airborne. Journal of Theoretical Biology, 494(??):??, June 7, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519320300825.

Araujo:2020:CPC

[ABvH^{+}20] Sabrina B. L. Araujo, Marcelo Eduardo Borges, Francisco W. von Hartenthal, Leonardo R. Jorge, Thomas M. Lewinsohn, Paulo R. Guimarães, Jr., and Minus van Baalen. Coevolutionary patterns caused by prey selection. Journal of Theoretical Biology, 501(??):??, September 21, 2020. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251932030182X.

Almeida:2020:TBC

[ACD20] Sofia Almeida, Madalena Chaves, and Franck Delaunay. Transcription-based circadian mechanism controls the duration of molecular clock states in response to signaling inputs. Journal of Theoretical Biology, 484(??):??, January 7, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519319303856.

Aristotelous:2022:HDC

[ACF22] Andreas C. Aristotelous, Alex Chen, and M. Gregory Forest. A hybrid discrete-continuum model of immune responses to SARS-CoV-2 infection in the lung alveolar region, with a focus on interferon induced innate response. Journal of Theoretical Biology, 555(??):??, December 21, 2022. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322002843.

Andreas:2024:QRG

[ACG24] Elizabeth Andreas, Breschine Cummins, and Tomás Gedeon. Quantifying robustness of the gap gene network. Journal of Theoretical Biology, 580(??):??, March 7, 2024. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519324000018.

Abhari:2024:CDS

[ACMT24] Niloufar Abhari, Caroline Colijn, Arne Mooers, and Paul Tupper. Capturing diversity: Split systems and circular approximations for conservation. Journal of Theoretical Biology, 578(??):??, February 7, 2024. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519323002862.

Asik:2020:EEF

[ACP20] Lale Asik, Ming Chen, and Angela Peace. The effects of excess food nutrient content on a tritrophic food chain model in the aquatic ecosystem. Journal of Theoretical Biology, 491(??):??, April 21, 2020. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519320300394.

Archibald:2024:ETO

[ADLM24] Kevin M. Archibald, Stephanie Dutkiewicz, Charlotte Laufkötter, and Holly V. Moeller. Emergent trade-offs among plasticity strategies in mixotrophs. Journal of Theoretical Biology, 590(??):??, August 7, 2024. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519324001358.

Avello:2021:TRA

[ADP21] Paula Avello, Seth J. Davis, and Jonathan W. Pitchford. Temperature robustness in Arabidopsis circadian clock models is facilitated by repressive interactions, autoregulation, and three-node feedbacks. Journal of Theoretical Biology, 509(??):??, January 21, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519320303507.

Akashi:2022:DGA

[AFTF22] Motohiro Akashi, Ichiro Fujihara, Masaharu Takemura, and Mitsuru Furusawa. 2-dimensional genetic algorithm exhibited an essentiality of gene interaction for evolution. Journal of Theoretical Biology, 538(??):??, April 7, 2022. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251932200042X.

Afas:2022:TLC

[AG22] Keith Christian Afas and Daniel Goldman. A two-layer continuously distributed capillary O_{2} transport model applied to blood flow regulation in resting skeletal muscle. Journal of Theoretical Biology, 539(??):??, April 21, 2022. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251932200056X.

Arbelaez-Gomez:2022:PBM

[AGBLGA^{+}22] Daniela Arbeláez-Gómez, Santiago Benavides-López, Maria Paula Giraldo-Agudelo, Juan Pablo Guzmán-Álvarez, Carolina Ramirez-Mazo, and Lina María Gómez-Echavarría. A phenomenological-based model of the endometrial growth and shedding during the menstrual cycle. Journal of Theoretical Biology, 532(??):??, January 7, 2022. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519321003416.

Adhikari:2021:TDC

[AGP^{+}21] Khagendra Adhikari, Ramesh Gautam, Anjana Pokharel, Kedar Nath Uprety, and Naveen K. Vaidya. Transmission dynamics of COVID-19 in Nepal: Mathematical model uncovering effective controls. Journal of Theoretical Biology, 521(??): ??, July 21, 2021. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519321001028.

Adhikari:2023:DDM

[AGP^{+}23] Khagendra Adhikari, Ramesh Gautam, Anjana Pokharel, Kedar Nath Uprety, and Naveen K. Vaidya. Data-driven models for the risk of infection and hospitalization during a pandemic: Case study on COVID-19 in Nepal. Journal of Theoretical Biology, 574(??):??, October 7, 2023. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519323002199.

Ansarinasab:2023:IPS

[AGTJ23] Sheida Ansarinasab, Farnaz Ghassemi, Zahra Tabanfar, and Sajad Jafari. Investigation of phase synchronization in functional brain networks of children with ADHD using nonlinear recurrence measure. Journal of Theoretical Biology, 560(??): ??, March 7, 2023. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519322003721.

Avraam:2025:ICN

[AH25] Demetris Avraam and Christoforos Hadjichrysanthou. The impact of contact-network structure on important epidemiological quantities of infectious disease transmission and the identification of the extremes. Journal of Theoretical Biology, 599(??):??, February 21, 2025. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519324002959.

Arim:2022:CTI

[AHEB^{+}22] Matías Arim, Daniel Herrera-Esposito, Paola Bermolen, Álvaro Cabana, María Inés Fariello, Mauricio Lima, and Hector Romero. Contact tracing-induced Allee effect in disease dynamics. Journal of Theoretical Biology, 542(??):??, June 7, 2022. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519322001072.

Abdi:2023:IEB

[AHS23] Hamed Abdi, Kamran Hassani, and Shahrokh Shojaei. An investigation of the effect of brain atrophy on brain injury in multiple sclerosis. Journal of Theoretical Biology, 557(??):??, January 21, 2023. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519322003307.

An:2021:NMP

[AHW21] Ran An, Jixin Hu, and Luosheng Wen. A nonlinear model predictive control model aimed at the epidemic spread with quarantine strategy. Journal of Theoretical Biology, 531(??):??, December 21, 2021. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519321003349.

Atkins:2020:AFL

[AJR^{+}20] Benjamin D. Atkins, Chris P. Jewell, Michael C. Runge, Matthew J. Ferrari, Katriona Shea, William J. M. Probert, and Michael J. Tildesley. Anticipating future learning affects current control decisions: a comparison between passive and active adaptive management in an epidemiological setting. Journal of Theoretical Biology, 506(??):??, December 7, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519320302356.

Akhlaghpour:2022:RBT

[Akh22] Hessameddin Akhlaghpour. An RNA-based theory of natural universal computation. Journal of Theoretical Biology, 537(??):??, March 21, 2022. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519321004045. See erratum [Akh23].

Akhlaghpour:2023:ERB

[Akh23] Hessameddin Akhlaghpour. Erratum to “An RNA-based theory of natural universal computation” [J. Theor. Biol. 537 (2022) 110984]. Journal of Theoretical Biology, 566(??):??, June 7, 2023. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519323000772. See [Akh22].

Aihara:2023:EII

[AKHM23] Ikkyu Aihara, Daichi Kominami, Yushi Hosokawa, and Masayuki Murata. Excitatory and inhibitory interactions affect the balance of chorus activity and energy efficiency in the aggregations of male frogs: numerical simulations using a hybrid dynamical model. Journal of Theoretical Biology, 558(??):??, February 7, 2023. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519322003435.

Anzai:2020:RPD

[AKUN20] Asami Anzai, Lisa Kawatsu, Kazuhiro Uchimura, and Hiroshi Nishiura. Reconstructing the population dynamics of foreign residents in Japan to estimate the prevalence of infection with mycobacterium tuberculosis. Journal of Theoretical Biology, 489(??):??, March 21, 2020. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519320300163.

Avila:2023:LHD

[AL23] Piret Avila and Laurent Lehmann. Life history and deleterious mutation rate coevolution. Journal of Theoretical Biology, 573(??):??, September 21, 2023. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519323001959.

Alamino:2020:ABL

[Ala20] R. C. Alamino. An agent-based lattice model for the emergence of anti-microbial resistance. Journal of Theoretical Biology, 486(??):??, February 7, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519319304497.

Alamir:2023:LBS

[Ala23] Mazen Alamir. Learning-based sensitivity analysis and feedback design for drug delivery of mixed therapy of cancer in the presence of high model uncertainties. Journal of Theoretical Biology, 568(??):??, July 7, 2023. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519323001042.

Araujo:2022:ISG

[AM22] Gui Araujo and Rafael Rios Moura. Individual specialization and generalization in predator-prey dynamics: the determinant role of predation efficiency and prey reproductive rates. Journal of Theoretical Biology, 537(??):??, March 21, 2022. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519322000248.

Akinlotan:2020:MMR

[AMA20] Morenikeji Deborah Akinlotan, Dann G. Mallet, and Robyn P. Araujo. Mathematical modelling of the role of mucosal vaccine on the within-host dynamics of Chlamydia trachomatis. Journal of Theoretical Biology, 497(??):??, July 21, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519320301466.

Apiyo:2022:DMM

[AML^{+}22] D. Apiyo, J. M. Mouton, C. Louw, S. L. Sampson, and T. M. Louw. Dynamic mathematical model development and validation of in vitro Mycobacterium smegmatis growth under nutrient- and pH-stress. Journal of Theoretical Biology, 532(??):??, January 7, 2022. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321003404.

Avusuglo:2023:WAD

[AMR^{+}23] W. S. Avusuglo, Rahele Mosleh, Tedi Ramaj, Ao Li, Sileshi Sintayehu Sharbayta, Abdoul Aziz Fall, Srijana Ghimire, Fenglin Shi, Jason K. H. Lee, Edward Thommes, Thomas Shin, and Jianhong Wu. Workplace absenteeism due to COVID-19 and influenza across Canada: a mathematical model. Journal of Theoretical Biology, 572(??):??, September 7, 2023. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251932300156X.

Agosti:2020:MCC

[AMSC20] A. Agosti, S. Marchesi, G. Scita, and P. Ciarletta. Modelling cancer cell budding in-vitro as a self-organised, non-equilibrium growth process. Journal of Theoretical Biology, 492(??):??, May 7, 2020. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519320300588.

Anzai:2022:DTI

[AN22a] Asami Anzai and Hiroshi Nishiura. Doubling time of infectious diseases. Journal of Theoretical Biology, 554(??):??, December 7, 2022. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519322002697.

Arthur:2022:SPG

[AN22b] Rudy Arthur and Arwen Nicholson. Selection principles for Gaia. Journal of Theoretical Biology, 533(??):??, January 21, 2022. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519321003593.

Aita:2024:MCM

[AN24] Takuyo Aita and Naoto Nemoto. Mathematical consideration of massive estimation of dissociation rate constant for genotype-phenotype linking molecules bound to targets through washing/selection and next-generation sequencing. Journal of Theoretical Biology, 595(??):??, December 7, 2024. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0022519324002297.

Ansarinasab:2025:STN

[ANGJ25] Sheida Ansarinasab, Fahimeh Nazarimehr, Farnaz Ghassemi, and Sajad Jafari. Synchronization transition in networks of Fractional-Order memristive Hindmarsh–Rose neurons. Journal of Theoretical Biology, 609(??):??, July 21, 2025. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519325001109.

Anonymous:2020:Ab

[Ano20a] Anonymous. 21 April 2020. Journal of Theoretical Biology, 491(??):??, April 21, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2020:Ad

[Ano20b] Anonymous. 21 August 2020. Journal of Theoretical Biology, 499(??):??, August 21, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2020:Db

[Ano20c] Anonymous. 21 December 2020. Journal of Theoretical Biology, 507(??):??, December 21, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2020:Fb

[Ano20d] Anonymous. 21 February 2020. Journal of Theoretical Biology, 487(??):??, February 21, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2020:Jb

[Ano20e] Anonymous. 21 January 2020. Journal of Theoretical Biology, 485(??):??, January 21, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2020:Jf

[Ano20f] Anonymous. 21 July 2020. Journal of Theoretical Biology, 497(??):??, July 21, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2020:Jd

[Ano20g] Anonymous. 21 June 2020. Journal of Theoretical Biology, 495(??):??, June 21, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2020:Mb

[Ano20h] Anonymous. 21 March 2020. Journal of Theoretical Biology, 489(??):??, March 21, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2020:Md

[Ano20i] Anonymous. 21 May 2020. Journal of Theoretical Biology, 493(??):??, May 21, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2020:Nb

[Ano20j] Anonymous. 21 November 2020. Journal of Theoretical Biology, 505(??):??, November 21, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2020:Ob

[Ano20k] Anonymous. 21 October 2020. Journal of Theoretical Biology, 503(??):??, October 21, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2020:Sb

[Ano20l] Anonymous. 21 September 2020. Journal of Theoretical Biology, 501(??):??, September 21, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2020:Aa

[Ano20m] Anonymous. 7 April 2020. Journal of Theoretical Biology, 490(??):??, April 7, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2020:Ac

[Ano20n] Anonymous. 7 August 2020. Journal of Theoretical Biology, 498(??):??, August 7, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2020:Da

[Ano20o] Anonymous. 7 December 2020. Journal of Theoretical Biology, 506(??):??, December 7, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2020:Fa

[Ano20p] Anonymous. 7 February 2020. Journal of Theoretical Biology, 486(??):??, February 7, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2020:Ja

[Ano20q] Anonymous. 7 January 2020. Journal of Theoretical Biology, 484(??):??, January 7, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2020:Je

[Ano20r] Anonymous. 7 July 2020. Journal of Theoretical Biology, 496(??):??, July 7, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2020:Jc

[Ano20s] Anonymous. 7 June 2020. Journal of Theoretical Biology, 494(??):??, June 7, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2020:Ma

[Ano20t] Anonymous. 7 March 2020. Journal of Theoretical Biology, 488(??):??, March 7, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2020:Mc

[Ano20u] Anonymous. 7 May 2020. Journal of Theoretical Biology, 492(??):??, May 7, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2020:Na

[Ano20v] Anonymous. 7 November 2020. Journal of Theoretical Biology, 504(??):??, November 7, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2020:Oa

[Ano20w] Anonymous. 7 October 2020. Journal of Theoretical Biology, 502(??):??, October 7, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2020:Sa

[Ano20x] Anonymous. 7 September 2020. Journal of Theoretical Biology, 500(??):??, September 7, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2020:EBa

[Ano20y] Anonymous. Editorial Board. Journal of Theoretical Biology, 484(??):??, January 7, 2020. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519319304163.

Anonymous:2020:EBb

[Ano20z] Anonymous. Editorial Board. Journal of Theoretical Biology, 485(??):??, January 21, 2020. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519319304333.

Anonymous:2020:EBc

[Ano20-27] Anonymous. Editorial Board. Journal of Theoretical Biology, 486(??):??, February 7, 2020. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519319304990.

Anonymous:2020:EBd

[Ano20-28] Anonymous. Editorial Board. Journal of Theoretical Biology, 487(??):??, February 21, 2020. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519320300059.

Anonymous:2020:EBe

[Ano20-29] Anonymous. Editorial Board. Journal of Theoretical Biology, 488(??):??, March 7, 2020. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519320300242.

Anonymous:2020:EBf

[Ano20-30] Anonymous. Editorial Board. Journal of Theoretical Biology, 489(??):??, March 21, 2020. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519320300369.

Anonymous:2020:EBg

[Ano20-31] Anonymous. Editorial Board. Journal of Theoretical Biology, 490(??):??, April 7, 2020. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519320300552.

Anonymous:2020:EBh

[Ano20-32] Anonymous. Editorial Board. Journal of Theoretical Biology, 491(??):??, April 21, 2020. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519320300746.

Anonymous:2020:EBi

[Ano20-33] Anonymous. Editorial Board. Journal of Theoretical Biology, 492(??):??, May 7, 2020. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519320300898.

Anonymous:2020:EBj

[Ano20-34] Anonymous. Editorial Board. Journal of Theoretical Biology, 493(??):??, May 21, 2020. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519320301156.

Anonymous:2020:EBk

[Ano20-35] Anonymous. Editorial Board. Journal of Theoretical Biology, 494(??):??, June 7, 2020. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519320301272.

Anonymous:2020:EBl

[Ano20-36] Anonymous. Editorial Board. Journal of Theoretical Biology, 495(??):??, June 21, 2020. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519320301405.

Anonymous:2020:EBm

[Ano20-37] Anonymous. Editorial Board. Journal of Theoretical Biology, 496(??):??, July 7, 2020. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519320301570.

Anonymous:2020:EBn

[Ano20-38] Anonymous. Editorial Board. Journal of Theoretical Biology, 497(??):??, July 21, 2020. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251932030179X.

Anonymous:2020:EBo

[Ano20-39] Anonymous. Editorial Board. Journal of Theoretical Biology, 498(??):??, August 7, 2020. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251932030196X.

Anonymous:2020:EBp

[Ano20-40] Anonymous. Editorial Board. Journal of Theoretical Biology, 499(??):??, August 21, 2020. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519320302113.

Anonymous:2020:EBq

[Ano20-41] Anonymous. Editorial Board. Journal of Theoretical Biology, 500(??):??, September 7, 2020. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519320302204.

Anonymous:2020:EBr

[Ano20-42] Anonymous. Editorial Board. Journal of Theoretical Biology, 501(??):??, September 21, 2020. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519320302502.

Anonymous:2020:EBs

[Ano20-43] Anonymous. Editorial Board. Journal of Theoretical Biology, 502(??):??, October 7, 2020. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519320302654.

Anonymous:2020:EBt

[Ano20-44] Anonymous. Editorial Board. Journal of Theoretical Biology, 503(??):??, October 21, 2020. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251932030285X.

Anonymous:2020:EBu

[Ano20-45] Anonymous. Editorial Board. Journal of Theoretical Biology, 504(??):??, November 7, 2020. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519320302988.

Anonymous:2020:EBv

[Ano20-46] Anonymous. Editorial Board. Journal of Theoretical Biology, 505(??):??, November 21, 2020. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519320303210.

Anonymous:2020:EBw

[Ano20-47] Anonymous. Editorial Board. Journal of Theoretical Biology, 506(??):??, December 7, 2020. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519320303350.

Anonymous:2020:EBx

[Ano20-48] Anonymous. Editorial Board. Journal of Theoretical Biology, 507(??):??, December 21, 2020. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519320303611.

Anonymous:2021:Ab

[Ano21a] Anonymous. 21 April 2021. Journal of Theoretical Biology, 515(??):??, April 21, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2021:Ad

[Ano21b] Anonymous. 21 August 2021. Journal of Theoretical Biology, 523(??):??, August 21, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2021:Db

[Ano21c] Anonymous. 21 December 2021. Journal of Theoretical Biology, 531(??):??, December 21, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2021:Fb

[Ano21d] Anonymous. 21 February 2021. Journal of Theoretical Biology, 511(??):??, February 21, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2021:Jb

[Ano21e] Anonymous. 21 January 2021. Journal of Theoretical Biology, 509(??):??, January 21, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2021:Jf

[Ano21f] Anonymous. 21 July 2021. Journal of Theoretical Biology, 521(??):??, July 21, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2021:Jd

[Ano21g] Anonymous. 21 June 2021. Journal of Theoretical Biology, 519(??):??, June 21, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2021:Mb

[Ano21h] Anonymous. 21 March 2021. Journal of Theoretical Biology, 513(??):??, March 21, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2021:Md

[Ano21i] Anonymous. 21 May 2021. Journal of Theoretical Biology, 517(??):??, May 21, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2021:Nb

[Ano21j] Anonymous. 21 November 2021. Journal of Theoretical Biology, 529(??):??, November 21, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2021:Ob

[Ano21k] Anonymous. 21 October 2021. Journal of Theoretical Biology, 527(??):??, October 21, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2021:Sb

[Ano21l] Anonymous. 21 September 2021. Journal of Theoretical Biology, 525(??):??, September 21, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2021:Aa

[Ano21m] Anonymous. 7 April 2021. Journal of Theoretical Biology, 514(??):??, April 7, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2021:Ac

[Ano21n] Anonymous. 7 August 2021. Journal of Theoretical Biology, 522(??):??, August 7, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2021:Da

[Ano21o] Anonymous. 7 December 2021. Journal of Theoretical Biology, 530(??):??, December 7, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2021:Fa

[Ano21p] Anonymous. 7 February 2021. Journal of Theoretical Biology, 510(??):??, February 7, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2021:Ja

[Ano21q] Anonymous. 7 January 2021. Journal of Theoretical Biology, 508(??):??, January 7, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2021:Je

[Ano21r] Anonymous. 7 July 2021. Journal of Theoretical Biology, 520(??):??, July 7, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2021:Jc

[Ano21s] Anonymous. 7 June 2021. Journal of Theoretical Biology, 518(??):??, June 7, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2021:Ma

[Ano21t] Anonymous. 7 March 2021. Journal of Theoretical Biology, 512(??):??, March 7, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2021:Mc

[Ano21u] Anonymous. 7 May 2021. Journal of Theoretical Biology, 516(??):??, May 7, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2021:Na

[Ano21v] Anonymous. 7 November 2021. Journal of Theoretical Biology, 528(??):??, November 7, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2021:Oa

[Ano21w] Anonymous. 7 October 2021. Journal of Theoretical Biology, 526(??):??, October 7, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2021:Sa

[Ano21x] Anonymous. 7 September 2021. Journal of Theoretical Biology, 524(??):??, September 7, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2021:EBa

[Ano21y] Anonymous. Editorial Board. Journal of Theoretical Biology, 508(??):??, January 7, 2021. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519320303763.

Anonymous:2021:EBb

[Ano21z] Anonymous. Editorial Board. Journal of Theoretical Biology, 509(??):??, January 21, 2021. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519320304045.

Anonymous:2021:EBc

[Ano21-27] Anonymous. Editorial Board. Journal of Theoretical Biology, 510(??):??, February 7, 2021. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251932030429X.

Anonymous:2021:EBd

[Ano21-28] Anonymous. Editorial Board. Journal of Theoretical Biology, 511(??):??, February 21, 2021. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519321000060.

Anonymous:2021:EBe

[Ano21-29] Anonymous. Editorial Board. Journal of Theoretical Biology, 512(??):??, March 7, 2021. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519321000138.

Anonymous:2021:EBf

[Ano21-30] Anonymous. Editorial Board. Journal of Theoretical Biology, 513(??):??, March 21, 2021. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519321000333.

Anonymous:2021:EBg

[Ano21-31] Anonymous. Editorial Board. Journal of Theoretical Biology, 514(??):??, April 7, 2021. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519321000473.

Anonymous:2021:EBh

[Ano21-32] Anonymous. Editorial Board. Journal of Theoretical Biology, 515(??):??, April 21, 2021. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519321000606.

Anonymous:2021:EBi

[Ano21-33] Anonymous. Editorial Board. Journal of Theoretical Biology, 516(??):??, May 7, 2021. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519321000746.

Anonymous:2021:EBj

[Ano21-34] Anonymous. Editorial Board. Journal of Theoretical Biology, 517(??):??, May 21, 2021. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519321000953.

Anonymous:2021:EBk

[Ano21-35] Anonymous. Editorial Board. Journal of Theoretical Biology, 518(??):??, June 7, 2021. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519321001119.

Anonymous:2021:EBl

[Ano21-36] Anonymous. Editorial Board. Journal of Theoretical Biology, 519(??):??, June 21, 2021. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519321001272.

Anonymous:2021:EBm

[Ano21-37] Anonymous. Editorial Board. Journal of Theoretical Biology, 520(??):??, July 7, 2021. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519321001442.

Anonymous:2021:EBn

[Ano21-38] Anonymous. Editorial Board. Journal of Theoretical Biology, 521(??):??, July 21, 2021. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251932100165X.

Anonymous:2021:EBo

[Ano21-39] Anonymous. Editorial Board. Journal of Theoretical Biology, 522(??):??, August 7, 2021. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519321001776.

Anonymous:2021:EBp

[Ano21-40] Anonymous. Editorial Board. Journal of Theoretical Biology, 523(??):??, August 21, 2021. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519321001922.

Anonymous:2021:EBq

[Ano21-41] Anonymous. Editorial Board. Journal of Theoretical Biology, 524(??):??, September 7, 2021. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519321002022.

Anonymous:2021:EBr

[Ano21-42] Anonymous. Editorial Board. Journal of Theoretical Biology, 525(??):??, September 21, 2021. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519321002216.

Anonymous:2021:EBs

[Ano21-43] Anonymous. Editorial Board. Journal of Theoretical Biology, 526(??):??, October 7, 2021. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519321002472.

Anonymous:2021:EBt

[Ano21-44] Anonymous. Editorial Board. Journal of Theoretical Biology, 527(??):??, October 21, 2021. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519321002629.

Anonymous:2021:EBu

[Ano21-45] Anonymous. Editorial Board. Journal of Theoretical Biology, 528(??):??, November 7, 2021. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519321002873.

Anonymous:2021:EBv

[Ano21-46] Anonymous. Editorial Board. Journal of Theoretical Biology, 529(??):??, November 21, 2021. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519321003088.

Anonymous:2021:EBw

[Ano21-47] Anonymous. Editorial Board. Journal of Theoretical Biology, 530(??):??, December 7, 2021. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519321003283.

Anonymous:2021:EBx

[Ano21-48] Anonymous. Editorial Board. Journal of Theoretical Biology, 531(??):??, December 21, 2021. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519321003490.

Anonymous:2021:RPL

[Ano21-49] Anonymous. In remembrance: Professor Lewis Wolpert. Journal of Theoretical Biology, 523(??):??, August 21, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321001156.

Anonymous:2022:Ab

[Ano22a] Anonymous. 21 April 2022. Journal of Theoretical Biology, 539(??):??, April 21, 2022. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2022:Ad

[Ano22b] Anonymous. 21 August 2022. Journal of Theoretical Biology, 547(??):??, August 21, 2022. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2022:Db

[Ano22c] Anonymous. 21 December 2022. Journal of Theoretical Biology, 555(??):??, December 21, 2022. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2022:Fb

[Ano22d] Anonymous. 21 February 2022. Journal of Theoretical Biology, 535(??):??, February 21, 2022. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2022:Jb

[Ano22e] Anonymous. 21 January 2022. Journal of Theoretical Biology, 533(??):??, January 21, 2022. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2022:Jf

[Ano22f] Anonymous. 21 July 2022. Journal of Theoretical Biology, 545(??):??, July 21, 2022. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2022:Jd

[Ano22g] Anonymous. 21 June 2022. Journal of Theoretical Biology, 543(??):??, June 21, 2022. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2022:Mb

[Ano22h] Anonymous. 21 March 2022. Journal of Theoretical Biology, 537(??):??, March 21, 2022. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2022:Md

[Ano22i] Anonymous. 21 May 2022. Journal of Theoretical Biology, 541(??):??, May 21, 2022. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2022:Nb

[Ano22j] Anonymous. 21 November 2022. Journal of Theoretical Biology, 553(??):??, November 21, 2022. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2022:Sb

[Ano22k] Anonymous. 21 September 2022. Journal of Theoretical Biology, 549(??):??, September 21, 2022. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2022:Aa

[Ano22l] Anonymous. 7 April 2022. Journal of Theoretical Biology, 538(??):??, April 7, 2022. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2022:Ac

[Ano22m] Anonymous. 7 August 2022. Journal of Theoretical Biology, 546(??):??, August 7, 2022. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2022:Da

[Ano22n] Anonymous. 7 December 2022. Journal of Theoretical Biology, 554(??):??, December 7, 2022. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2022:Fa

[Ano22o] Anonymous. 7 February 2022. Journal of Theoretical Biology, 534(??):??, February 7, 2022. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2022:Ja

[Ano22p] Anonymous. 7 January 2022. Journal of Theoretical Biology, 532(??):??, January 7, 2022. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2022:Je

[Ano22q] Anonymous. 7 July 2022. Journal of Theoretical Biology, 544(??):??, July 7, 2022. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2022:Jc

[Ano22r] Anonymous. 7 June 2022. Journal of Theoretical Biology, 542(??):??, June 7, 2022. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2022:Ma

[Ano22s] Anonymous. 7 March 2022. Journal of Theoretical Biology, 536(??):??, March 7, 2022. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2022:Mc

[Ano22t] Anonymous. 7 May 2022. Journal of Theoretical Biology, 540(??):??, May 7, 2022. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2022:Na

[Ano22u] Anonymous. 7 November 2022. Journal of Theoretical Biology, 551–552(??):??, ???? 2022. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2022:O

[Ano22v] Anonymous. 7 October 2022. Journal of Theoretical Biology, 550(??):??, October 7, 2022. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2022:Sa

[Ano22w] Anonymous. 7 September 2022. Journal of Theoretical Biology, 548(??):??, September 7, 2022. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2022:EBa

[Ano22x] Anonymous. Editorial Board. Journal of Theoretical Biology, 532(??):??, January 7, 2022. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519321003726.

Anonymous:2022:EBb

[Ano22y] Anonymous. Editorial Board. Journal of Theoretical Biology, 533(??):??, January 21, 2022. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519321003891.

Anonymous:2022:EBc

[Ano22z] Anonymous. Editorial Board. Journal of Theoretical Biology, 534(??):??, February 7, 2022. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519321004112.

Anonymous:2022:EBd

[Ano22-27] Anonymous. Editorial Board. Journal of Theoretical Biology, 535(??):??, February 21, 2022. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519322000091.

Anonymous:2022:EBe

[Ano22-28] Anonymous. Editorial Board. Journal of Theoretical Biology, 536(??):??, March 7, 2022. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519322000340.

Anonymous:2022:EBf

[Ano22-29] Anonymous. Editorial Board. Journal of Theoretical Biology, 537(??):??, March 21, 2022. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519322000479.

Anonymous:2022:EBg

[Ano22-30] Anonymous. Editorial Board. Journal of Theoretical Biology, 538(??):??, April 7, 2022. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519322000698.

Anonymous:2022:EBh

[Ano22-31] Anonymous. Editorial Board. Journal of Theoretical Biology, 539(??):??, April 21, 2022. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519322000807.

Anonymous:2022:EBi

[Ano22-32] Anonymous. Editorial Board. Journal of Theoretical Biology, 540(??):??, May 7, 2022. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519322000972.

Anonymous:2022:EBj

[Ano22-33] Anonymous. Editorial Board. Journal of Theoretical Biology, 541(??):??, May 21, 2022. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519322001126.

Anonymous:2022:EBk

[Ano22-34] Anonymous. Editorial Board. Journal of Theoretical Biology, 542(??):??, June 7, 2022. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519322001291.

Anonymous:2022:EBl

[Ano22-35] Anonymous. Editorial Board. Journal of Theoretical Biology, 543(??):??, June 21, 2022. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519322001400.

Anonymous:2022:EBm

[Ano22-36] Anonymous. Editorial Board. Journal of Theoretical Biology, 544(??):??, July 7, 2022. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519322001540.

Anonymous:2022:EBn

[Ano22-37] Anonymous. Editorial Board. Journal of Theoretical Biology, 545(??):??, July 21, 2022. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519322001667.

Anonymous:2022:EBo

[Ano22-38] Anonymous. Editorial Board. Journal of Theoretical Biology, 546(??):??, August 7, 2022. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519322001783.

Anonymous:2022:EBp

[Ano22-39] Anonymous. Editorial Board. Journal of Theoretical Biology, 547(??):??, August 21, 2022. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519322001886.

Anonymous:2022:EBq

[Ano22-40] Anonymous. Editorial Board. Journal of Theoretical Biology, 548(??):??, September 7, 2022. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519322002144.

Anonymous:2022:EBr

[Ano22-41] Anonymous. Editorial Board. Journal of Theoretical Biology, 549(??):??, September 21, 2022. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519322002260.

Anonymous:2022:EBs

[Ano22-42] Anonymous. Editorial Board. Journal of Theoretical Biology, 550(??):??, October 7, 2022. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519322002454.

Anonymous:2022:EBt

[Ano22-43] Anonymous. Editorial Board. Journal of Theoretical Biology, 551–552(??):??, ???? 2022. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519322002570.

Anonymous:2022:EBu

[Ano22-44] Anonymous. Editorial Board. Journal of Theoretical Biology, 553(??):??, November 21, 2022. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519322002788.

Anonymous:2022:EBv

[Ano22-45] Anonymous. Editorial Board. Journal of Theoretical Biology, 554(??):??, December 7, 2022. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251932200296X.

Anonymous:2022:EBw

[Ano22-46] Anonymous. Editorial Board. Journal of Theoretical Biology, 555(??):??, December 21, 2022. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519322003101.

Anonymous:2023:Ab

[Ano23a] Anonymous. 21 April 2023. Journal of Theoretical Biology, 563(??):??, April 21, 2023. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2023:Ad

[Ano23b] Anonymous. 21 August 2023. Journal of Theoretical Biology, 571(??):??, August 21, 2023. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2023:Fb

[Ano23c] Anonymous. 21 February 2023. Journal of Theoretical Biology, 559(??):??, February 21, 2023. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2023:Jb

[Ano23d] Anonymous. 21 January 2023. Journal of Theoretical Biology, 557(??):??, January 21, 2023. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2023:Jf

[Ano23e] Anonymous. 21 July 2023. Journal of Theoretical Biology, 569(??):??, July 21, 2023. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2023:Jd

[Ano23f] Anonymous. 21 June 2023. Journal of Theoretical Biology, 567(??):??, June 21, 2023. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2023:Mb

[Ano23g] Anonymous. 21 March 2023. Journal of Theoretical Biology, 561(??):??, March 21, 2023. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2023:Md

[Ano23h] Anonymous. 21 May 2023. Journal of Theoretical Biology, 565(??):??, May 21, 2023. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2023:Sb

[Ano23i] Anonymous. 21 September 2023. Journal of Theoretical Biology, 573(??):??, September 21, 2023. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2023:Aa

[Ano23j] Anonymous. 7 April 2023. Journal of Theoretical Biology, 562(??):??, April 7, 2023. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2023:Ac

[Ano23k] Anonymous. 7 August 2023. Journal of Theoretical Biology, 570(??):??, August 7, 2023. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2023:Fa

[Ano23l] Anonymous. 7 February 2023. Journal of Theoretical Biology, 558(??):??, February 7, 2023. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2023:Ja

[Ano23m] Anonymous. 7 January 2023. Journal of Theoretical Biology, 556(??):??, January 7, 2023. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2023:Je

[Ano23n] Anonymous. 7 July 2023. Journal of Theoretical Biology, 568(??):??, July 7, 2023. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2023:Jc

[Ano23o] Anonymous. 7 June 2023. Journal of Theoretical Biology, 566(??):??, June 7, 2023. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2023:Ma

[Ano23p] Anonymous. 7 March 2023. Journal of Theoretical Biology, 560(??):??, March 7, 2023. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2023:Mc

[Ano23q] Anonymous. 7 May 2023. Journal of Theoretical Biology, 564(??):??, May 7, 2023. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2023:N

[Ano23r] Anonymous. 7 November 2023. Journal of Theoretical Biology, 575(??):??, November 7, 2023. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2023:O

[Ano23s] Anonymous. 7 October 2023. Journal of Theoretical Biology, 574(??):??, October 7, 2023. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2023:Sa

[Ano23t] Anonymous. 7 September 2023. Journal of Theoretical Biology, 572(??):??, September 7, 2023. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2023:EBa

[Ano23u] Anonymous. Editorial Board. Journal of Theoretical Biology, 556(??):??, January 7, 2023. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251932200337X.

Anonymous:2023:EBb

[Ano23v] Anonymous. Editorial Board. Journal of Theoretical Biology, 557(??):??, January 21, 2023. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519322003538.

Anonymous:2023:EBc

[Ano23w] Anonymous. Editorial Board. Journal of Theoretical Biology, 558(??):??, February 7, 2023. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519322003800.

Anonymous:2023:EBd

[Ano23x] Anonymous. Editorial Board. Journal of Theoretical Biology, 559(??):??, February 21, 2023. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519322003897.

Anonymous:2023:EBe

[Ano23y] Anonymous. Editorial Board. Journal of Theoretical Biology, 560(??):??, March 7, 2023. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519323000061.

Anonymous:2023:EBf

[Ano23z] Anonymous. Editorial Board. Journal of Theoretical Biology, 561(??):??, March 21, 2023. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519323000218.

Anonymous:2023:EBg

[Ano23-27] Anonymous. Editorial Board. Journal of Theoretical Biology, 562(??):??, April 7, 2023. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519323000371.

Anonymous:2023:EBh

[Ano23-28] Anonymous. Editorial Board. Journal of Theoretical Biology, 563(??):??, April 21, 2023. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519323000516.

Anonymous:2023:EBi

[Ano23-29] Anonymous. Editorial Board. Journal of Theoretical Biology, 564(??):??, May 7, 2023. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251932300070X.

Anonymous:2023:EBj

[Ano23-30] Anonymous. Editorial Board. Journal of Theoretical Biology, 565(??):??, May 21, 2023. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519323000826.

Anonymous:2023:EBk

[Ano23-31] Anonymous. Editorial Board. Journal of Theoretical Biology, 566(??):??, June 7, 2023. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251932300098X.

Anonymous:2023:EBl

[Ano23-32] Anonymous. Editorial Board. Journal of Theoretical Biology, 567(??):??, June 21, 2023. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519323001091.

Anonymous:2023:EBm

[Ano23-33] Anonymous. Editorial Board. Journal of Theoretical Biology, 568(??):??, July 7, 2023. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519323001261.

Anonymous:2023:EBn

[Ano23-34] Anonymous. Editorial Board. Journal of Theoretical Biology, 569(??):??, July 21, 2023. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519323001388.

Anonymous:2023:EBo

[Ano23-35] Anonymous. Editorial Board. Journal of Theoretical Biology, 570(??):??, August 7, 2023. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519323001467.

Anonymous:2023:EBp

[Ano23-36] Anonymous. Editorial Board. Journal of Theoretical Biology, 571(??):??, August 21, 2023. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519323001698.

Anonymous:2023:EBq

[Ano23-37] Anonymous. Editorial Board. Journal of Theoretical Biology, 572(??):??, September 7, 2023. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251932300200X.

Anonymous:2023:EBr

[Ano23-38] Anonymous. Editorial Board. Journal of Theoretical Biology, 573(??):??, September 21, 2023. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251932300214X.

Anonymous:2023:EBs

[Ano23-39] Anonymous. Editorial Board. Journal of Theoretical Biology, 574(??):??, October 7, 2023. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519323002370.

Anonymous:2023:EBt

[Ano23-40] Anonymous. Editorial Board. Journal of Theoretical Biology, 575(??):??, November 7, 2023. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519323002588.

Anonymous:2024:Ab

[Ano24a] Anonymous. 21 April 2024. Journal of Theoretical Biology, 583(??):??, April 21, 2024. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2024:Ad

[Ano24b] Anonymous. 21 August 2024. Journal of Theoretical Biology, 591(??):??, August 21, 2024. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2024:Fb

[Ano24c] Anonymous. 21 February 2024. Journal of Theoretical Biology, 579(??):??, February 21, 2024. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2024:Jb

[Ano24d] Anonymous. 21 January 2024. Journal of Theoretical Biology, 577(??):??, January 21, 2024. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2024:Jf

[Ano24e] Anonymous. 21 July 2024. Journal of Theoretical Biology, 589(??):??, July 21, 2024. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2024:Jd

[Ano24f] Anonymous. 21 June 2024. Journal of Theoretical Biology, 587(??):??, June 21, 2024. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2024:Mb

[Ano24g] Anonymous. 21 March 2024. Journal of Theoretical Biology, 581(??):??, March 21, 2024. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2024:Md

[Ano24h] Anonymous. 21 May 2024. Journal of Theoretical Biology, 585(??):??, May 21, 2024. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2024:Aa

[Ano24i] Anonymous. 7 April 2024. Journal of Theoretical Biology, 582(??):??, April 7, 2024. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2024:Ac

[Ano24j] Anonymous. 7 August 2024. Journal of Theoretical Biology, 590(??):??, August 7, 2024. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2024:D

[Ano24k] Anonymous. 7 December 2024. Journal of Theoretical Biology, 595(??):??, December 7, 2024. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2024:Fa

[Ano24l] Anonymous. 7 February 2024. Journal of Theoretical Biology, 578(??):??, February 7, 2024. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2024:Ja

[Ano24m] Anonymous. 7 January 2024. Journal of Theoretical Biology, 576(??):??, January 7, 2024. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2024:Je

[Ano24n] Anonymous. 7 July 2024. Journal of Theoretical Biology, 588(??):??, July 7, 2024. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2024:Jc

[Ano24o] Anonymous. 7 June 2024. Journal of Theoretical Biology, 586(??):??, June 7, 2024. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2024:Ma

[Ano24p] Anonymous. 7 March 2024. Journal of Theoretical Biology, 580(??):??, March 7, 2024. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2024:Mc

[Ano24q] Anonymous. 7 May 2024. Journal of Theoretical Biology, 584(??):??, May 7, 2024. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2024:N

[Ano24r] Anonymous. 7 November 2024. Journal of Theoretical Biology, 594(??):??, November 7, 2024. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2024:O

[Ano24s] Anonymous. 7 October 2024. Journal of Theoretical Biology, 593(??):??, October 7, 2024. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2024:S

[Ano24t] Anonymous. 7 September 2024. Journal of Theoretical Biology, 592(??):??, September 7, 2024. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2024:EBa

[Ano24u] Anonymous. Editorial Board. Journal of Theoretical Biology, 576(??):??, January 7, 2024. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519323002758.

Anonymous:2024:EBb

[Ano24v] Anonymous. Editorial Board. Journal of Theoretical Biology, 577(??):??, January 21, 2024. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519323002904.

Anonymous:2024:EBc

[Ano24w] Anonymous. Editorial Board. Journal of Theoretical Biology, 578(??):??, February 7, 2024. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519323003053.

Anonymous:2024:EBd

[Ano24x] Anonymous. Editorial Board. Journal of Theoretical Biology, 579(??):??, February 21, 2024. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519324000067.

Anonymous:2024:EBe

[Ano24y] Anonymous. Editorial Board. Journal of Theoretical Biology, 580(??):??, March 7, 2024. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251932400033X.

Anonymous:2024:EBf

[Ano24z] Anonymous. Editorial Board. Journal of Theoretical Biology, 581(??):??, March 21, 2024. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519324000456.

Anonymous:2024:EBg

[Ano24-27] Anonymous. Editorial Board. Journal of Theoretical Biology, 582(??):??, April 7, 2024. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519324000572.

Anonymous:2024:EBh

[Ano24-28] Anonymous. Editorial Board. Journal of Theoretical Biology, 583(??):??, April 21, 2024. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519324000699.

Anonymous:2024:EBi

[Ano24-29] Anonymous. Editorial Board. Journal of Theoretical Biology, 584(??):??, May 7, 2024. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519324000821.

Anonymous:2024:EBj

[Ano24-30] Anonymous. Editorial Board. Journal of Theoretical Biology, 585(??):??, May 21, 2024. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519324000912.

Anonymous:2024:EBk

[Ano24-31] Anonymous. Editorial Board. Journal of Theoretical Biology, 586(??):??, June 7, 2024. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519324001103.

Anonymous:2024:EBl

[Ano24-32] Anonymous. Editorial Board. Journal of Theoretical Biology, 587(??):??, June 21, 2024. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519324001206.

Anonymous:2024:EBm

[Ano24-33] Anonymous. Editorial Board. Journal of Theoretical Biology, 588(??):??, July 7, 2024. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519324001279.

Anonymous:2024:EBn

[Ano24-34] Anonymous. Editorial Board. Journal of Theoretical Biology, 589(??):??, July 21, 2024. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519324001425.

Anonymous:2024:EBo

[Ano24-35] Anonymous. Editorial Board. Journal of Theoretical Biology, 590(??):??, August 7, 2024. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519324001504.

Anonymous:2024:EBp

[Ano24-36] Anonymous. Editorial Board. Journal of Theoretical Biology, 591(??):??, August 21, 2024. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519324001711.

Anonymous:2024:EBq

[Ano24-37] Anonymous. Editorial Board. Journal of Theoretical Biology, 592(??):??, September 7, 2024. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519324001899.

Anonymous:2024:EBr

[Ano24-38] Anonymous. Editorial Board. Journal of Theoretical Biology, 593(??):??, October 7, 2024. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519324002030.

Anonymous:2024:EBs

[Ano24-39] Anonymous. Editorial Board. Journal of Theoretical Biology, 594(??):??, November 7, 2024. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519324002236.

Anonymous:2024:EBt

[Ano24-40] Anonymous. Editorial Board. Journal of Theoretical Biology, 595(??):??, December 7, 2024. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0022519324002662.

Anonymous:2025:Ac

[Ano25a] Anonymous. 21 August 2025. Journal of Theoretical Biology, 611(??):??, August 21, 2025. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2025:Fb

[Ano25b] Anonymous. 21 February 2025. Journal of Theoretical Biology, 599(??):??, February 21, 2025. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2025:Jb

[Ano25c] Anonymous. 21 January 2025. Journal of Theoretical Biology, 597(??):??, January 21, 2025. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2025:Jf

[Ano25d] Anonymous. 21 July 2025. Journal of Theoretical Biology, 609(??):??, July 21, 2025. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2025:Jd

[Ano25e] Anonymous. 21 June 2025. Journal of Theoretical Biology, 607(??):??, June 21, 2025. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2025:Mb

[Ano25f] Anonymous. 21 March 2025. Journal of Theoretical Biology, 601(??):??, March 21, 2025. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2025:Md

[Ano25g] Anonymous. 21 May 2025. Journal of Theoretical Biology, 605(??):??, May 21, 2025. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2025:Aa

[Ano25h] Anonymous. 7 April 2025. Journal of Theoretical Biology, 602–603(??):??, April 7, 2025. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2025:Ab

[Ano25i] Anonymous. 7 August 2025. Journal of Theoretical Biology, 610(??):??, August 7, 2025. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2025:Fa

[Ano25j] Anonymous. 7 February 2025. Journal of Theoretical Biology, 598(??):??, February 7, 2025. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2025:Ja

[Ano25k] Anonymous. 7 January 2025. Journal of Theoretical Biology, 596(??):??, January 7, 2025. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2025:Je

[Ano25l] Anonymous. 7 July 2025. Journal of Theoretical Biology, 608(??):??, July 7, 2025. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2025:Jc

[Ano25m] Anonymous. 7 June 2025. Journal of Theoretical Biology, 606(??):??, June 7, 2025. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2025:Ma

[Ano25n] Anonymous. 7 March 2025. Journal of Theoretical Biology, 600(??):??, March 7, 2025. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2025:Mc

[Ano25o] Anonymous. 7 May 2025. Journal of Theoretical Biology, 604(??):??, May 7, 2025. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2025:Oa

[Ano25p] Anonymous. 7 October 2025. Journal of Theoretical Biology, 613(??):??, October 7, 2025. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2025:S

[Ano25q] Anonymous. 7 September 2025. Journal of Theoretical Biology, 612(??):??, September 7, 2025. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2025:EBa

[Ano25r] Anonymous. Editorial Board. Journal of Theoretical Biology, 596(??):??, January 7, 2025. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519324002893.

Anonymous:2025:EBb

[Ano25s] Anonymous. Editorial Board. Journal of Theoretical Biology, 597(??):??, January 21, 2025. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251932400300X.

Anonymous:2025:EBc

[Ano25t] Anonymous. Editorial Board. Journal of Theoretical Biology, 598(??):??, February 7, 2025. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519324003102.

Anonymous:2025:EBd

[Ano25u] Anonymous. Editorial Board. Journal of Theoretical Biology, 599(??):??, February 21, 2025. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519325000049.

Anonymous:2025:EBe

[Ano25v] Anonymous. Editorial Board. Journal of Theoretical Biology, 600(??):??, March 7, 2025. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0022519325000207.

Anonymous:2025:EBf

[Ano25w] Anonymous. Editorial Board. Journal of Theoretical Biology, 601(??):??, March 21, 2025. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0022519325000323.

Anonymous:2025:EBg

[Ano25x] Anonymous. Editorial Board. Journal of Theoretical Biology, 602–603(??):??, April 7, 2025. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251932500044X.

Anonymous:2025:EBh

[Ano25y] Anonymous. Editorial Board. Journal of Theoretical Biology, 604(??):??, May 7, 2025. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519325000670.

Anonymous:2025:EBi

[Ano25z] Anonymous. Editorial Board. Journal of Theoretical Biology, 605(??):??, May 21, 2025. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519325000815.

Anonymous:2025:EBj

[Ano25-27] Anonymous. Editorial Board. Journal of Theoretical Biology, 606(??):??, June 7, 2025. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519325000918.

Anonymous:2025:EBk

[Ano25-28] Anonymous. Editorial Board. Journal of Theoretical Biology, 607(??):??, June 21, 2025. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519325001298.

Anonymous:2025:EBl

[Ano25-29] Anonymous. Editorial Board. Journal of Theoretical Biology, 608(??):??, July 7, 2025. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519325001183.

Anonymous:2025:EBm

[Ano25-30] Anonymous. Editorial Board. Journal of Theoretical Biology, 609(??):??, July 21, 2025. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519325001390.

Anonymous:2025:EBn

[Ano25-31] Anonymous. Editorial Board. Journal of Theoretical Biology, 610(??):??, August 7, 2025. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251932500147X.

Anonymous:2025:EBo

[Ano25-32] Anonymous. Editorial Board. Journal of Theoretical Biology, 611(??):??, August 21, 2025. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519325001705.

Anonymous:2025:EBp

[Ano25-33] Anonymous. Editorial Board. Journal of Theoretical Biology, 612(??):??, September 7, 2025. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519325001900.

Anonymous:2025:EBq

[Ano25-34] Anonymous. Editorial Board. Journal of Theoretical Biology, 613(??):??, October 7, 2025. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519325002073.

Aft:2021:SPI

[AOB21] Tristan Aft, Sorinel A. Oprisan, and Catalin V. Buhusi. Is the scalar property of interval timing preserved after hippocampus lesions? Journal of Theoretical Biology, 516(??): ??, May 7, 2021. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519321000278.

Angeli:2024:WAI

[APF^{+}24] Leonardo Angeli, Constantino Pereira Caetano, Nicolas Franco, Steven Abrams, Pietro Coletti, Inneke Van Nieuwenhuyse, Sorin Pop, and Niel Hens. Who acquires infection from whom? A sensitivity analysis of transmission dynamics during the early phase of the COVID-19 pandemic in Belgium. Journal of Theoretical Biology, 581(??):??, March 21, 2024. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251932400002X.

Avila:2021:HRG

[APL21] Piret Avila, Tadeas Priklopil, and Laurent Lehmann. Hamilton’s rule, gradual evolution, and the optimal (feedback) control of phenotypically plastic traits. Journal of Theoretical Biology, 526(??):??, October 7, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519321000242.

Abdalgader:2022:TBC

[APR^{+}22a] Tarteel Abdalgader, Michael Pedersen, Dongsheng Ren, Guiquan Sun, and Lai Zhang. Trade-off between climatic and human population impacts on Aedes aegypti life history shapes its geographic distribution. Journal of Theoretical Biology, 535(??):??, February 21, 2022. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321004070.

Azimaqin:2022:VFS

[APR^{+}22b] Nurbek Azimaqin, Zhihang Peng, Xinzhi Ren, Yangjiang Wei, and Xianning Liu. Vaccine failure, seasonality and demographic changes associate with mumps outbreaks in Jiangsu Province, China: Age-structured mathematical modelling study. Journal of Theoretical Biology, 544(??):??, July 7, 2022. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519322001230.

Ahn:2020:TPD

[AR20] Sungwoo Ahn and Leonid L. Rubchinsky. Temporal patterns of dispersal-induced synchronization in population dynamics. Journal of Theoretical Biology, 490(??):??, April 7, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519320300151.

Argasinski:2021:RDG

[AR21] Krzysztof Argasinski and Ryszard Rudnicki. Replicator dynamics for the game theoretic selection models based on state. Journal of Theoretical Biology, 526(??):??, October 7, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519320303957.

Andras:2021:WDS

[AS21] Peter Andras and Adam Stanton. Where do successful populations originate from? Journal of Theoretical Biology, 524(??):??, September 7, 2021. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519321001569.

Allaby:2022:NCF

[ASF22] Robin G. Allaby, Chris J. Stevens, and Dorian Q. Fuller. A novel cost framework reveals evidence for competitive selection in the evolution of complex traits during plant domestication. Journal of Theoretical Biology, 537(??):??, March 21, 2022. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519322000029.

Abroudi:2020:TAC

[ASK20] Ali Abroudi, Sandhya Samarasinghe, and Don Kulasiri. Towards abstraction of computational modelling of mammalian cell cycle: Model reduction pipeline incorporating multi-level hybrid petri nets. Journal of Theoretical Biology, 496(??): ??, July 7, 2020. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519320300679.

Ayabina:2025:NMM

[ASKG25] Diepreye Ayabina, Hasan Sevil, Adam Kleczkowski, and M. Gabriela-M. Gomes. A note on metapopulation models. Journal of Theoretical Biology, 613(??):??, October 7, 2025. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519325001791.

Athanasouli:2024:DDM

[ASL^{+}24] Christina Athanasouli, Shelby R. Stowe, Monique K. LeBourgeois, Victoria Booth, and Cecilia G. Diniz Behn. Data-driven mathematical modeling of sleep consolidation in early childhood. Journal of Theoretical Biology, 593(??):??, October 7, 2024. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519324001760.

Archibald:2023:PSS

[ASMN23] Kevin M. Archibald, Heidi M. Sosik, Holly V. Moeller, and Michael G. Neubert. Predator switching strength controls stability in diamond-shaped food web models. Journal of Theoretical Biology, 570(??):??, August 7, 2023. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519323001327.

AlSariri:2023:OHT

[ASP23] Tahani Al Sariri, Radostin D. Simitev, and Raimondo Penta. Optimal heat transport induced by magnetic nanoparticle delivery in vascularised tumours. Journal of Theoretical Biology, 561(??):??, March 21, 2023. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519322003630.

Are:2023:CEP

[ASS^{+}23] Elisha B. Are, Yexuan Song, Jessica E. Stockdale, Paul Tupper, and Caroline Colijn. COVID-19 endgame: From pandemic to endemic? Vaccination, reopening and evolution in low- and high-vaccinated populations. Journal of Theoretical Biology, 559(??):??, February 21, 2023. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322003599.

Abo:2022:MFP

[ASSL22] Stéphanie Abo, Delaney Smith, Melissa Stadt, and Anita Layton. Modelling female physiology from head to toe: Impact of sex hormones, menstrual cycle, and pregnancy. Journal of Theoretical Biology, 540(??):??, May 7, 2022. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322000728.

Al-Tuwairqi:2020:MDC

[ATAJS20] Salma M. Al-Tuwairqi, Najwa O. Al-Johani, and Eman A. Simbawa. Modeling dynamics of cancer radiovirotherapy. Journal of Theoretical Biology, 506(??):??, December 7, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519320302605.

Anderson:2024:OCC

[ATH^{+}24] Hannah G. Anderson, Gregory P. Takacs, Jeffrey K. Harrison, Libin Rong, and Tracy L. Stepien. Optimal control of combination immunotherapy for a virtual murine cohort in a glioblastoma-immune dynamics model. Journal of Theoretical Biology, 595(??):??, December 7, 2024. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0022519324002364.

Aogo:2022:SDI

[ATP22] Rosemary A. Aogo, Mark M. Tanaka, and Catherine J. Penington. Spatial dynamics of inflammation-causing and commensal bacteria in the gastrointestinal tract. Journal of Theoretical Biology, 548(??):??, September 7, 2022. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322001928.

Arthur:2025:SRU

[AVP^{+}25] Lachlan Arthur, Vasiliki Voulgaridou, Georgios Papageorgiou, Weiping Lu, Steven R. McDougall, and Vassilis Sboros. Super-resolution ultrasound imaging of ischaemia flow: an in silico study. Journal of Theoretical Biology, 599(??):??, February 21, 2025. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519324003035.

Abukwaik:2023:IPX

[AVSTS23] Roba Abukwaik, Elias Vera-Siguenza, Daniel A. Tennant, and Fabian Spill. Interplay of p53 and XIAP protein dynamics orchestrates cell fate in response to chemotherapy. Journal of Theoretical Biology, 572(??):??, September 7, 2023. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519323001595.

Avvari:2021:TMR

[Avv21] Ravi Kant Avvari. Theoretical modeling of the resistance to gastric emptying and duodenogastric reflux due to pyloric motility alone, presuming antral and duodenal quiescence. Journal of Theoretical Biology, 508(??):??, January 7, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519320303155.

Avvari:2023:RSC

[Avv23] Ravi Kant Avvari. Role of segmental contraction in the small intestinal digestion: a computational approach to study the physics behind the luminal mixing and transport. Journal of Theoretical Biology, 561(??):??, March 21, 2023. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519323000140.

Amoah-Darko:2022:MMD

[AW22] Frederick Laud Amoah-Darko, Jr. and Diana White. Modeling microtubule dynamic instability: Microtubule growth, shortening and pause. Journal of Theoretical Biology, 553(??):??, November 21, 2022. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519322002491.

Andrade:2025:IAA

[AWC25] Renato Andrade, Steven M. White, and Christina A. Cobbold. Incorporating adult age into mosquito population models: Implications for predicting abundances in changing climates. Journal of Theoretical Biology, 604(??):??, May 7, 2025. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519325000505.

Andreychuk:2022:MMA

[AY22] Saige Andreychuk and Laith Yakob. Mathematical modelling to assess the feasibility of Wolbachia in malaria vector biocontrol. Journal of Theoretical Biology, 542(??):??, June 7, 2022. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519322001084.

Alderliesten:2022:SCW

[AZdV^{+}22] Jesse B. Alderliesten, Mark P. Zwart, J. Arjan G. M. de Visser, Arjan Stegeman, and Egil A. J. Fischer. Second compartment widens plasmid invasion conditions: Twocompartment pair-formation model of conjugation in the gut. Journal of Theoretical Biology, 533(??):??, January 21, 2022. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321003568.

Ayala-Zambrano:2023:BNM

[AZYF^{+}23] Cecilia Ayala-Zambrano, Mariana Yuste, Sara Frias, Benilde Garcia de Teresa, Luis Mendoza, Eugenio Azpeitia, Alfredo Rodríguez, and Leda Torres. A Boolean network model of the double-strand break repair pathway choice. Journal of Theoretical Biology, 573(??):??, September 21, 2023. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519323002059.

Buckingham:2024:SET

[BA24] Lydia J. Buckingham and Ben Ashby. Separation of evolutionary timescales in coevolving species. Journal of Theoretical Biology, 579(??):??, February 21, 2024. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519323002850.

Bagchi:2024:GST

[BACS24] Dweepabiswa Bagchi, Ramesh Arumugam, V. K. Chandrasekar, and D. V. Senthilkumar. Generalized synchronization in a tritrophic food web metacommunity. Journal of Theoretical Biology, 582(??):??, April 7, 2024. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519324000407.

Bonifacio:2024:CMC

[BAG^{+}24] Elton Diêgo Bonifácio, Cleudmar Amaral Araújo, Marcília Valéria Guimarães, Márcio Peres de Souza, Thiago Parente Lima, Bethânia Alves de Avelar Freitas, and Libardo Andrés González-Torres. Computational model of the cancer necrotic core formation in a tumor-on-a-chip device. Journal of Theoretical Biology, 592(??):??, September 7, 2024. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519324001772.

Bai:2021:EPH

[Bai21] Fan Bai. Effect of population heterogeneity on herd immunity and on vaccination decision making process. Journal of Theoretical Biology, 526(??):??, October 7, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321002149.

Byrne:2021:SFR

[BAJ^{+}21] Monika Byrne, Rosalind Aughwane, Joanna L. James, J. Ciaran Hutchinson, Owen J. Arthurs, Neil J. Sebire, Sebastien Ourselin, Anna L. David, Andrew Melbourne, and Alys R. Clark. Structure-function relationships in the feto-placental circulation from in silico interpretation of micro-CT vascular structures. Journal of Theoretical Biology, 517(??):??, May 21, 2021. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519321000527.

Barwich:2021:ILB

[Bar21] Ann-Sophie Barwich. Imaging the living brain: an argument for ruthless reductionism from olfactory neurobiology.

Journal of Theoretical Biology, 512(??):??, March 7, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251932030415X.

Barraquand:2023:NSF

[Bar23] Frédéric Barraquand. No sensitivity to functional forms in the Rosenzweig–MacArthur model with strong environmental stochasticity. Journal of Theoretical Biology, 572(??):??, September 7, 2023. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519323001637.

Baumgartner:2020:CNS

[Bau20] Matheus T. Baumgartner. Connectance and nestedness as stabilizing factors in response to pulse disturbances in adaptive antagonistic networks. Journal of Theoretical Biology, 486(??):??, February 7, 2020. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519319304424.

Bano:2022:RGP

[BAU^{+}22] Sheher Bano, Muhammad Rizwan Azam, Ali Arshad Uppal, Syed Bilal Javed, and Aamer Iqbal Bhatti. Robust p53 recovery using chattering free sliding mode control and a gain-scheduled modified Utkin observer. Journal of Theoretical Biology, 532(??):??, January 7, 2022. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321003337.

Bartlett:2021:CRH

[BB21a] Lewis J. Bartlett and Mike Boots. The central role of host reproduction in determining the evolution of virulence in spatially structured populations. Journal of Theoretical Biology, 523(??):??, August 21, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321001399.

Brewer:2021:CRT

[BB21b] Tom R. Brewer and Michael B. Bonsall. Combining refuges with transgenic insect releases for the management of an insect pest with non-recessive resistance to Bt crops in agricultural landscapes. Journal of Theoretical Biology, 509(??):??, January 21, 2021. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519320303696.

Basak:2025:SFC

[BB25] Ayanava Basak and Nandadulal Bairagi. Success, failure and collapse of an ecological system due to the introduction of a generalist super-predator. Journal of Theoretical Biology, 612(??):??, September 7, 2025. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251932500164X.

Bakry:2025:DMM

[BBB^{+}25] Aamy Bakry, Emma Brashear, Jacob Brashear, Shannon Z. Jones, and Marcella M. Torres. Development of a mathematical model of the innate immune response to inhaled toxicants. Journal of Theoretical Biology, 611(??):??, August 21, 2025. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519325001250.

Bayer:2022:MDM

[BBDB22] Péter Bayer, Joel S. Brown, Johan Dubbeldam, and Mark Broom. A Markovian decision model of adaptive cancer treatment and quality of life. Journal of Theoretical Biology, 551–552(??):??, ???? 2022. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519322002338.

Banuelos:2020:MLT

[BBH^{+}20] Selenne Bañuelos, Janet Best, Gemma Huguet, Alicia Prieto-Langarica, Pamela B. Pyzza, and Shelby Wilson. Modeling the long term effects of thermoregulation on human sleep. Journal of Theoretical Biology, 493(??):??, May 21, 2020.

CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519320300631.

Bai:2023:MSE

[BBK23] Fan Bai, Richard Bertram, and Bhargav R. Karamched. A mathematical study of the efficacy of possible negative feedback pathways involved in neuronal polarization. Journal of Theoretical Biology, 571(??):??, August 21, 2023. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519323001583.

Beyer:2021:DTC

[BBR^{+}21] Robert M. Beyer, David Basler, Pasi Raumonen, Mikko Kaasalainen, and Hans Pretzsch. Do trees have constant branch divergence angles? Journal of Theoretical Biology, 512(??):??, March 7, 2021. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519320304227.

Bodmer:2020:SSP

[BC20] Walter F. Bodmer and Daniel J. M. Crouch. Somatic selection of poorly differentiating variant stem cell clones could be a key to human ageing. Journal of Theoretical Biology, 489(??):??, March 21, 2020. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519320300096.

Barrett:2021:ODC

[BC21] Jeff M. Barrett and Jack P. Callaghan. A one-dimensional collagen-based biomechanical model of passive soft tissue with viscoelasticity and failure. Journal of Theoretical Biology, 509(??):??, January 21, 2021. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251932030343X.

Burch:2025:EOC

[BCBP25] Eleanor Burch, Hannah Christensen, and Ellen Brooks-Pollock. The effect of outcome choice on vaccine dose allocation: a modelling study. Journal of Theoretical Biology, 610(??):??, August 7, 2025. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519325001225.

Baer:2021:MCM

[BCC^{+}21] Steven M. Baer, Shaojie Chang, Sharon M. Crook, Carl L. Gardner, Jeremiah R. Jones, Christian Ringhofer, and Ralph F. Nelson. A multiscale continuum model of the vertebrate outer retina: the temporal dynamics of backgroundinduced flicker enhancement. Journal of Theoretical Biology, 525(??):??, September 21, 2021. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519321001855.

Berrios-Caro:2021:CDM

[BCGG21] Ernesto Berríos-Caro, Danna R. Gifford, and Tobias Galla. Competition delays multi-drug resistance evolution during combination therapy. Journal of Theoretical Biology, 509(??):??, January 21, 2021. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519320303799.

Balit:2024:ICR

[BCK24] Nasri Balit, Nicolas Cermakian, and Anmar Khadra. The influence of circadian rhythms on CD8^{+} T cell activation upon vaccination: a mathematical modeling perspective. Journal of Theoretical Biology, 590(??):??, August 7, 2024. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519324001334.

Baaz:2024:PAT

[BCLJ24] Marcus Baaz, Tim Cardilin, Torbjörn Lundh, and Mats Jirstrand. Probabilistic analysis of tumor growth inhibition models to support trial design. Journal of Theoretical Biology, 595(??):??, December 7, 2024. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0022519324002546.

Beretta:2022:PCO

[BCS^{+}22] Edoardo Beretta, Vincenzo Capasso, Simone Scacchi, Matteo Brunetti, and Matteo Montagna. Prevention and control of OQDS (olive quick decline syndrome) outbreaks caused by Xylella fastidiosa. Journal of Theoretical Biology, 542(??): ??, June 7, 2022. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519322001163.

Bliman:2021:HBC

[BD21] Pierre-Alexandre Bliman and Michel Duprez. How best can finite-time social distancing reduce epidemic final size? Journal of Theoretical Biology, 511(??):??, February 21, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519320304124.

Baumgartner:2022:SCD

[BD22] Matheus Tenório Baumgartner and Lucas Del Bianco Faria. The sensitivity of complex dynamic food webs to the loss of top omnivores. Journal of Theoretical Biology, 538(??):??, April 7, 2022. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251932200025X.

Bowers:2025:MSS

[BDB25] Kathryn H. Bowers, Daniela De Angelis, and Paul J. Birrell. Modelling with SPEED: a Stochastic Predictor of Early Epidemic Detection. Journal of Theoretical Biology, 607(??):??, June 21, 2025. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519325000864.

Buonomo:2022:BMA

[BDdG22] Bruno Buonomo, Rossella Della Marca, Alberto d’Onofrio, and Maria Groppi. A behavioural modelling approach to assess the impact of COVID-19 vaccine hesitancy. Journal of Theoretical Biology, 534(??):??, February 7, 2022. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321003933.

Bartha:2022:MIR

[BE22] Liza Bartha and Raluca Eftimie. Mathematical investigation into the role of macrophage heterogeneity on the temporal and spatio-temporal dynamics of non-small cell lung cancers. Journal of Theoretical Biology, 549(??):??, September 21, 2022. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519322002053.

Bouderbala:2020:UHC

[BEBA20] Ilhem Bouderbala, Nadjia El Saadi, Alassane Bah, and Pierre Auger. Understanding how the collective behaviour of phytoflagellates is affected by light attenuation and diel vertical migration using individual-based modelling. Journal of Theoretical Biology, 494(??):??, June 7, 2020. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519320300965.

Breitenbach:2022:MSB

[BEO^{+}22] Tim Breitenbach, Nils Englert, Özge Osmanoglu, Natalia Rukoyatkina, Gaby Wangorsch, Katrin Heinze, Andreas Friebe, Elke Butt, Robert Feil, Marcus Dittrich, Stepan Gambaryan, and Thomas Dandekar. A modular systems biological modelling framework studies cyclic nucleotide signaling in platelets. Journal of Theoretical Biology, 550(??):??, October 7, 2022. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251932200220X.

Broom:2020:MMA

[BERR20] Mark Broom, Igor V. Erovenko, Jonathan T. Rowell, and Jan Rychtár. Models and measures of animal aggregation and dispersal. Journal of Theoretical Biology, 484(??):??, January 7, 2020. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519319303716.

Blitvic:2020:ALO

[BF20] Natasha Blitvić and Vicente I. Fernandez. Aging a little: On the optimality of limited senescence in Escherichia coli. Journal of Theoretical Biology, 502(??):??, October 7, 2020.

CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519320301867.

Benatti:2025:INM

[BFD25] Alexandre Benatti, Henrique Ferraz De Arruda, and Luciano Da Fontoura Costa. Interrelating neuronal morphology by coincidence similarity networks. Journal of Theoretical Biology, 606(??):??, June 7, 2025. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519325000700.

Bijma:2021:SHR

[BFG21] Nienke N. Bijma, Alexander E. Filippov, and Stanislav N. Gorb. Sisyphus and his rock: Quasi-random walk inspired by the motion of a ball transported by a dung beetle on combined terrain. Journal of Theoretical Biology, 520(??): ??, July 7, 2021. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519321000813.

Bramble:2023:ERS

[BFSA23] Jordan Bramble, Alexander Fulk, Raul Saenz, and Folashade B. Agusto. Exploring the role of superspreading events in SARS-CoV-2 outbreaks. Journal of Theoretical Biology, 558(??):??, February 7, 2023. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519322003447.

Berardo:2021:CRP

[BG21] Cecilia Berardo and Stefan Geritz. Coevolution of the reckless prey and the patient predator. Journal of Theoretical Biology, 530(??):??, December 7, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321002927.

Brhane:2021:MMD

[BGA21] Kewani Welay Brhane, Michael Gidey Gebru, and Abdulaziz Garba Ahmad. Mathematical model for the dynamics of Savanna ecosystem considering fire disturbances. Journal of Theoretical Biology, 509(??):??, January 21, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519320303702.

Bassett:2021:IBM

[BGB^{+}21] Jason Bassett, Jörn Gethmann, Pascal Blunk, Franz J. Conraths, and Philipp Hövel. Individual-based model for the control of Bovine Viral Diarrhea spread in livestock trade networks. Journal of Theoretical Biology, 527(??):??, October 21, 2021. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519321002393.

Bhowmick:2020:LTD

[BGC^{+}20] Suman Bhowmick, Jörn Gethmann, Franz J. Conraths, Igor M. Sokolov, and Hartmut H. K. Lentz. Locally temperature-driven mathematical model of West Nile virus spread in Germany. Journal of Theoretical Biology, 488(??): ??, March 7, 2020. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519319304862.

Boudin:2023:CMD

[BGGM23] Laurent Boudin, Céline Grandmont, Bérénice Grec, and Sébastien Martin. A coupled model for the dynamics of gas exchanges in the human lung with Haldane and Bohr’s effects. Journal of Theoretical Biology, 573(??):??, September 21, 2023. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251932300187X.

Browne:2022:DIC

[BGM22] Cameron J. Browne, Hayriye Gulbudak, and Joshua C. Macdonald. Differential impacts of contact tracing and lockdowns on outbreak size in COVID-19 model applied to China. Journal of Theoretical Biology, 532(??):??, January 7, 2022. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321003386.

Brutovsky:2020:SIE

[BH20] B. Brutovsky and D. Horvath. In Silico implementation of evolutionary paradigm in therapy design: Towards anti-cancer therapy as Darwinian process. Journal of Theoretical Biology, 485(??):??, January 21, 2020. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319304072.

Baldwin:2023:SAW

[BH23] Scott A. Baldwin and Jason M. Haugh. Semi-autonomous wound invasion via matrix-deposited, haptotactic cues. Journal of Theoretical Biology, 568(??):??, July 7, 2023. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519323001029.

Belotserkovskii:2024:MTD

[BH24] Boris P. Belotserkovskii and Philip C. Hanawalt. A model for transcription-dependent R-loop formation at double-stranded DNA breaks: Implications for their detection and biological effects. Journal of Theoretical Biology, 595(??):??, December 7, 2024. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0022519324002479.

Beeton:2020:MCB

[BHW^{+}20] Nicholas J. Beeton, Geoffrey R. Hosack, Andrew Wilkins, Lawrence K. Forbes, Adrien Ickowicz, and Keith R. Hayes. Modelling competition between hybridising subspecies. Journal of Theoretical Biology, 486(??):??, February 7, 2020. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319304412.

Best:2020:MMS

[BJB^{+}20] Alex Best, Jamil Jubrail, Mike Boots, David Dockrell, and Helen Marriott. A mathematical model shows macrophages delay Staphylococcus aureus replication, but limitations in microbicidal capacity restrict bacterial clearance. Journal of Theoretical Biology, 497(??):??, July 21, 2020. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519320301119.

Biswas:2021:FPT

[BJG21] Kuheli Biswas, Mohit Kumar Jolly, and Anandamohan Ghosh. First passage time properties of miRNAmediated protein translation. Journal of Theoretical Biology, 529(??):??, November 21, 2021. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519321002824.

B:2025:CST

[BJJM25] Hamna Mariyam K. B., Sayooj Aby Jose, Anuwat Jirawattanapanit, and Karuna Mathew. A comprehensive study on tuberculosis prediction models: Integrating machine learning into epidemiological analysis. Journal of Theoretical Biology, 597(??):??, January 21, 2025. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251932400273X.

Balasekaran:2022:QNE

[BJR^{+}22] Madhumitha Balasekaran, Michal Johanis, Jan Rychtár, Dewey Taylor, and Jackie Zhu. Quasi-neutral evolution in populations under small demographic fluctuations. Journal of Theoretical Biology, 538(??):??, April 7, 2022. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322000388.

Borges:2020:CIP

[BK20a] Rui Borges and Carolin Kosiol. Consistency and identifiability of the polymorphism-aware phylogenetic models. Journal of Theoretical Biology, 486(??):??, February 7, 2020. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319304436.

Broom:2020:TSG

[BK20b] Mark Broom and Vlastimil Krivan. Two-strategy games with time constraints on regular graphs. Journal of Theoretical Biology, 506(??):??, December 7, 2020. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519320302812.

Blyuss:2021:ELA

[BK21] K. B. Blyuss and Y. N. Kyrychko. Effects of latency and age structure on the dynamics and containment of COVID-19. Journal of Theoretical Biology, 513(??):??, March 21, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519321000096.

Bedekar:2023:PEO

[BKP23] Prajakta Bedekar, Anthony J. Kearsley, and Paul N. Patrone. Prevalence estimation and optimal classification methods to account for time dependence in antibody levels. Journal of Theoretical Biology, 559(??):??, February 21, 2023. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322003666.

Blath:2025:SPM

[BKPT25] Jochen Blath, Anna Kraut, Tobias Paul, and András Tóbiás. A stochastic population model for the impact of cancer cell dormancy on therapy success. Journal of Theoretical Biology, 597(??):??, January 21, 2025. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519324002807.

Becker:2022:MDD

[BKZ^{+}22] Sid Becker, Andrey V. Kuznetsov, Dan Zhao, Filippo de Monte, and Giuseppe Pontrelli. Model of drug delivery to populations composed of two cell types. Journal of Theoretical Biology, 534(??):??, February 7, 2022. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321003660.

Burbano-L:2020:DDM

[BLP20] Daniel A. Burbano-L. and Maurizio Porfiri. Data-driven modeling of zebrafish behavioral response to acute caffeine administration. Journal of Theoretical Biology, 485(??):??, January 21, 2020. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519319304230.

Bernhauerova:2021:MMS

[BLRV21] Veronika Bernhauerová, Bartek Lisowski, Veronica V. Rezelj, and Marco Vignuzzi. Mathematical modelling of SARS-CoV-2 infection of human and animal host cells reveals differences in the infection rates and delays in viral particle production by infected cells. Journal of Theoretical Biology, 531(??):??, December 21, 2021. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519321003143.

Bolle:2020:DCM

[BM20] Nicolas Bolle and Matthew S. Mizuhara. Dynamics of a cell motility model near the sharp interface limit. Journal of Theoretical Biology, 505(??):??, November 21, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519320302757.

Browning:2021:MBD

[BMB^{+}21] Alexander P. Browning, Oliver J. Maclaren, Pascal R. Buenzli, Matthew Lanaro, Mark C. Allenby, Maria A. Woodruff, and Matthew J. Simpson. Model-based data analysis of tissue growth in thin 3D printed scaffolds. Journal of Theoretical Biology, 528(??):??, November 7, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251932100271X.

Baabdulla:2021:HRD

[BNP^{+}21] Arwa Abdulla Baabdulla, Hesung Now, Ju An Park, Woo-Jong Kim, Sungjune Jung, Joo-Yeon Yoo, and Thomas Hillen. Homogenization of a reaction diffusion equation can explain influenza A virus load data. Journal of Theoretical Biology, 527(??):??, October 21, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321002356.

Blokhuis:2020:GTC

[BNPL20] Alex Blokhuis, Philippe Nghe, Luca Peliti, and David Lacoste. The generality of transient compartmentalization and its associated error thresholds. Journal of Theoretical Biology, 487(??):??, February 21, 2020. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319304795.

Boldin:2022:IED

[Bol22] Barbara Boldin. The importance of ecological dynamics in evolutionary processes: a host–bacteriophage model revisited. Journal of Theoretical Biology, 539(??):??, April 21, 2022. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519322000558.

Baker:2020:EHF

[BP20] Ryan Baker and Michel Pleimling. The effect of habitats and fitness on species coexistence in systems with cyclic dominance. Journal of Theoretical Biology, 486(??):??, February 7, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519319304539.

Baldwin:2021:LTF

[BP21] Quenisha Baldwin and Eleni Panagiotou. The local topological free energy of proteins. Journal of Theoretical Biology, 529(??):??, November 21, 2021. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519321002733.

Balakrishnan:2023:ATS

[BP23] Sreenath Balakrishnan and Safvan Palathingal. An adaptive testing strategy for efficient utilization of healthcare resources during an epidemic. Journal of Theoretical Biology, 571(??):??, August 21, 2023. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519323001522.

Butuzova:2022:DGC

[BPM^{+}22] Oksana Butuzova, Nikolay Pakudin, Andrey Minarsky, Nikolay Bessonov, and Nadya Morozova. Developmental graphs comparison strategy for analysis of pattern formation and phylogeny. Journal of Theoretical Biology, 532(??):??, January 7, 2022. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519321003441.

Balelli:2020:MEM

[BPP^{+}20] Irene Balelli, Chloé Pasin, Mélanie Prague, Fabien Crauste, Thierry Van Effelterre, Viki Bockstal, Laura Solforosi, and Rodolphe Thiébaut. A model for establishment, maintenance and reactivation of the immune response after vaccination against Ebola virus. Journal of Theoretical Biology, 495(??): ??, June 21, 2020. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519320301090.

Buffoni:2025:RDM

[BPS25] Giuseppe Buffoni, Sara Pasquali, and Cinzia Soresina. Revisiting a demographic model of pest infestation in farming. Journal of Theoretical Biology, 611(??):??, August 21, 2025. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519325001572.

Blanco-Rodriguez:2025:LWH

[BRAHV25] Rodolfo Blanco-Rodriguez, Alejandro Anderson, and Esteban Hernandez-Vargas. Linking within-host immune dynamics to between-host transmission and reinfection risk. Journal of Theoretical Biology, 613(??):??, October 7, 2025. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519325001766.

Bohlin:2020:SSM

[BRBF20] Jon Bohlin, Brittany Rose, Ola Brynildsrud, and Birgitte Freiesleben De Blasio. A simple stochastic model describing genomic evolution over time of GC content in microbial symbionts. Journal of Theoretical Biology, 503(??):??, October 21, 2020. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519320302447.

Bahramian:2022:SOD

[BRN^{+}22] Alireza Bahramian, Janarthanan Ramadoss, Fahimeh Nazarimehr, Karthikeyan Rajagopal, Sajad Jafari, and Iqtadar Hussain. A simple one-dimensional map-based model of spiking neurons with wide ranges of firing rates and complexities. Journal of Theoretical Biology, 539(??):??, April 21, 2022. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519322000601.

Bunimovich:2024:ACV

[BRS24] Leonid Bunimovich, Athulya Ram, and Pavel Skums. Antigenic cooperation in viral populations: Transformation of functions of intra-host viral variants. Journal of Theoretical Biology, 580(??):??, March 7, 2024. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519323003168.

Bazzazi:2021:MSP

[BS21] Hojjat Bazzazi and Azar Shahraz. A mechanistic systems pharmacology modeling platform to investigate the effect of PD-L1 expression heterogeneity and dynamics on the efficacy of PD-1 and PD-L1 blocking antibodies in cancer. Journal of Theoretical Biology, 522(??):??, August 7, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321001193.

Brautigam:2022:DAP

[BS22a] Camila Bräutigam and Matteo Smerlak. Diffusion approximations in population genetics and the rate of Muller’s ratchet. Journal of Theoretical Biology, 550(??):??, October 7, 2022. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519322002314.

Bruner:2022:TBS

[BS22b] Justin P. Bruner and Rory Smead. Tag-based spite with correlated interactions. Journal of Theoretical Biology, 540(??):

??, May 7, 2022. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519322000509.

Bokes:2025:OGE

[BS25] Pavol Bokes and Abhyudai Singh. Optimisation of gene expression noise for cellular persistence against lethal events. Journal of Theoretical Biology, 598(??):??, February 7, 2025. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519324002819.

Bandyopadhyay:2020:CAN

[BSC20a] Anirban Bandyopadhyay, Gaurav Sharma, and Shubhajit Roy Chowdhury. Computational analysis of NIRS and BOLD signal from neurovascular coupling with three neuron-system feedforward inhibition network. Journal of Theoretical Biology, 498(??):??, August 7, 2020. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519320301521.

Banerjee:2020:CPM

[BSC20b] Swarnendu Banerjee, Amar Sha, and Joydev Chattopadhyay. Cooperative predation on mutualistic prey communities. Journal of Theoretical Biology, 490(??):??, April 7, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519320300126.

Baker:2020:LCT

[BSJ20] Chris Baker, Mark Sterling, and Mike Jesson. The lodging of crops by tornadoes. Journal of Theoretical Biology, 500(??):??, September 7, 2020. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519320301648.

Bishop:2024:EFA

[BSL24] Megan E. Bishop, Maria R. Servedio, and Brian A. Lerch. The evolution of fear-acquisition strategies under predation. Journal of Theoretical Biology, 595(??):??, December 7, 2024.

CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0022519324002340.

Bhardwaj:2020:EBT

[BSS^{+}20] Vijay Bhardwaj, Rahul Singh, Pooja Singh, Rituraj Purohit, and Sanjay Kumar. Elimination of bitter-off taste of stevioside through structure modification and computational interventions. Journal of Theoretical Biology, 486(??):??, February 7, 2020. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519319304631.

Bowen:2024:MME

[BSV^{+}24] M. Bowen, P. Samozino, M. Vonderscher, D. Dutykh, and B. Morel. Mathematical modeling of exercise fatigability in the severe domain: a unifying integrative framework in isokinetic condition. Journal of Theoretical Biology, 578(??):??, February 7, 2024. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251932300293X.

Bitsouni:2022:MMT

[BT22] Vasiliki Bitsouni and Vasilis Tsilidis. Mathematical modeling of tumor-immune system interactions: the effect of rituximab on breast cancer immune response. Journal of Theoretical Biology, 539(??):??, April 21, 2022. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321004215.

Bergmann:2024:MMM

[BTS24] Philip J. Bergmann and Isabel Tonelli-Sippel. Many-to-many mapping: a simulation study of how the number of traits and tasks affect the evolution of form and function. Journal of Theoretical Biology, 581(??):??, March 21, 2024. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519324000250.

Bouchnita:2024:PSM

[BV24] Anass Bouchnita and Vitaly Volpert. Phenotype-structured model of intra-clonal heterogeneity and drug resistance in multiple myeloma. Journal of Theoretical Biology, 576(??):??, January 7, 2024. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519323002497.

Bailey:2020:OMI

[BvdG20] Richard M. Bailey and Jesse M. A. van der Grient. OSIRIS: a model for integrating the effects of multiple stressors on marine ecosystems. Journal of Theoretical Biology, 493(??): ??, May 21, 2020. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519320300667.

Bai:2024:MLC

[BW24] Jie Bai and Jin Wang. Modeling long COVID dynamics: Impact of underlying health conditions. Journal of Theoretical Biology, 576(??):??, January 7, 2024. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519323002667.

Brotman:2025:CEP

[BW25] Samantha Brotman and Geoff Wild. Co-evolution of pathogen-host interactions with vertical transmission can produce bistable outcomes. Journal of Theoretical Biology, 604(??):??, May 7, 2025. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0022519325000396.

Bessho:2020:SEM

[BYHH20] Kazuhiro Bessho, Kenta Yashima, Toyomitsu Horii, and Masakazu Hori. Spatially explicit modeling of metapopulation dynamics of broadcast spawners and stabilizing/ destabilizing effects of heterogeneity of quality across local habitats. Journal of Theoretical Biology, 492(??):??, May 7, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519320300138.

Cox:2021:MMI

[CA21] Kingsley J. A. Cox and Paul R. Adams. A minimal model of the interaction of social and individual learning. Journal of Theoretical Biology, 527(??):??, October 21, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251932100134X.

Castano-Arcila:2021:MID

[CAARG21] Mauricio Castaño-Arcila, Luis U. Aguilera, and Jesús Rodríguez-González. Modeling the intracellular dynamics of the dengue viral infection and the innate immune response. Journal of Theoretical Biology, 509(??):??, January 21, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519320303842.

Cai:2022:ECM

[Cai22] Yuhua Cai. Evolutionary coexistence in a metacommunity: Competition–colonization trade-off, ownership effects, environmental fluctuations. Journal of Theoretical Biology, 533(??):??, January 21, 2022. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519321003635.

Carvalho:2023:CMC

[Car23] Joao Carvalho. A computational model of cell membrane bioelectric polarization and depolarization, connected with cell proliferation, in different tissue geometries. Journal of Theoretical Biology, 557(??):??, January 21, 2023. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322003290.

Casenave:2022:IBN

[CBBD22] Céline Casenave, Anne Bisson, Simon Boudsocq, and Tanguy Daufresne. Impact of biological nitrogen fixation and livestock management on the manure transfer from grazing land in mixed farming systems. Journal of Theoretical Biology, 545(??):??, July 21, 2022. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519322001345.

Celora:2022:DSM

[CBH^{+}22] Giulia L. Celora, Samuel B. Bader, Ester M. Hammond, Philip K. Maini, Joe M. Pitt-Francis, and Helen M. Byrne. A DNA-structured mathematical model of cell-cycle progression in cyclic hypoxia. Journal of Theoretical Biology, 545(??):??, July 21, 2022. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519322001023. See corri-

gendum [CBH^{+}23].

Celora:2023:CDS

[CBH^{+}23] Giulia L. Celora, Samuel B. Bader, Ester M. Hammond, Philip K. Maini, Joe M. Pitt-Francis, and Helen M. Byrne. Corrigendum to “A DNA-structured mathematical model of cell-cycle progression in cyclic hypoxia” [J. Theor. Biol. 545 (2022) 111104]. Journal of Theoretical Biology, 571(??):??, August 21, 2023. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519323001030. See [CBH^{+}22].

Cloete:2020:DMC

[CBK^{+}20] Ielyaas Cloete, Paula J. Bartlett, Vivien Kirk, Andrew P. Thomas, and James Sneyd. Dual mechanisms of Ca 2+ oscillations in hepatocytes. Journal of Theoretical Biology, 503(??):??, October 21, 2020. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519320302459.

Celora:2023:STM

[CBK23] Giulia L. Celora, Helen M. Byrne, and P. G. Kevrekidis. Spatio-temporal modelling of phenotypic heterogeneity in tumour tissues and its impact on radiotherapy treatment. Journal of Theoretical Biology, 556(??):??, January 7, 2023. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322002417.

Cochrane:2021:EME

[CBKS21] Courtney Cochrane, Demba Ba, Elizabeth B. Klerman, and Melissa A. St. Hilaire. An ensemble mixed effects model of sleep loss and performance. Journal of Theoretical Biology, 509(??):??, January 21, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519320303520.

Chowdhury:2023:EEC

[CBPG23] Sayantan Nag Chowdhury, Jeet Banerjee, Matjaz Perc, and Dibakar Ghosh. Eco-evolutionary cyclic dominance among predators, prey, and parasites. Journal of Theoretical Biology, 564(??):??, May 7, 2023. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519323000425.

Celora:2021:PVM

[CBZK21] Giulia L. Celora, Helen M. Byrne, Christos E. Zois, and P. G. Kevrekidis. Phenotypic variation modulates the growth dynamics and response to radiotherapy of solid tumours under normoxia and hypoxia. Journal of Theoretical Biology, 527(??):??, October 21, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519321002113.

Corcoran:2022:ANM

[CC22] Carl Corcoran and John Michael Clark. Adaptive network modeling of social distancing interventions. Journal of Theoretical Biology, 546(??):??, August 7, 2022. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322001497.

Capitan:2020:CDP

[CCA20] José A. Capitán, Sara Cuenda, and David Alonso. Competitive dominance in plant communities: Modeling approaches and theoretical predictions. Journal of Theoretical Biology, 502(??):??, October 7, 2020. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519320302046.

Carlsson:2020:MWB

[CCF^{+}20] Rose-Marie Carlsson, Lauren M. Childs, Zhilan Feng, John W. Glasser, Jane M. Heffernan, Jing Li, and Gergely Röst. Modeling the waning and boosting of immunity from infection or vaccination. Journal of Theoretical Biology, 497(??):??, July 21, 2020. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251932030120X.

Chaaya:2024:CAM

[CCH^{+}24] Marie-Jose Chaaya, Sophie Chauvet, Florence Hubert, Fanny Mann, Mathieu Mezache, and Pierre Pudlo. A continuous approach of modeling tumorigenesis and axons regulation for the pancreatic cancer. Journal of Theoretical Biology, 595(??):??, December 7, 2024. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0022519324002522.

Castiblanco:2022:SCM

[CCPD22] Julieth Castiblanco, Paulo F. Cristaldo, Leticia R. Paiva, and Og DeSouza. Social context modulates scale-free movements in a social insect. Journal of Theoretical Biology, 542(??): ??, June 7, 2022. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519322001047.

Christensen:2021:AIL

[CCSF21] Kim Christensen, Luca Cocconi, and Ana B. Sendova-Franks. Animal intermittent locomotion: a null model for the probability of moving forward in bounded space. Journal of Theoretical Biology, 510(??):??, February 7, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251932030388X.

Cloete:2021:TTR

[CCVB^{+}21] Ielyaas Cloete, Juliana C. Corrêa-Velloso, Paula J. Bartlett, Vivien Kirk, Andrew P. Thomas, and James Sneyd. A tale of two receptors. Journal of Theoretical Biology, 518(??):??, June 7, 2021. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519321000515.

Calabrese:2022:HOA

[CD22] Justin M. Calabrese and Jeffery Demers. How optimal allocation of limited testing capacity changes epidemic dynamics. Journal of Theoretical Biology, 538(??):??, April 7, 2022. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322000157.

Chiu:2020:MFM

[CDC20] Loren Z. F. Chiu, Torstein E. Dæhlin, and Jason P. Carey. A multiple fascicle muscle force model of the human triceps surae. Journal of Theoretical Biology, 495(??):??, June 21, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519320301065.

Coton:2023:EEL

[CDdV23] Charlotte Coton, Christine Dillmann, and Dominique de Vienne. Evolution of enzyme levels in metabolic pathways: a theoretical approach. Part 2. Journal of Theoretical Biology, 558(??):??, February 7, 2023. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519322003459.

Cess:2020:DDA

[CF20] Colin G. Cess and Stacey D. Finley. Data-driven analysis of a mechanistic model of CAR T cell signaling predicts effects of cell-to-cell heterogeneity. Journal of Theoretical Biology, 489(??):??, March 21, 2020. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319304941.

Chen:2025:MPI

[CFH^{+}25] Jun Chen, Adrian Fisher, Jon F. Harrison, Gloria DeGrandi-Hoffman, Jennifer H. Fewell, and Yun Kang. Modeling pesticide impacts on honey bee dynamics from mathematical and experimental integration. Journal of Theoretical Biology, 611(??):??, August 21, 2025. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251932500133X.

Cai:2021:EIT

[CG21] Yuhua Cai and Stefan A. H. Geritz. The evolution of the irreversible transition from a free-swimming state to an immobile sessile state in aquatic invertebrates modelled in a chemostat. Journal of Theoretical Biology, 522(??):??, August 7, 2021. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251932100103X.

Ciavolella:2025:DCT

[CGG^{+}25] Giorgia Ciavolella, Julien Granet, Jacky G. Goetz, Naël Osmani, Christèle Etchegaray, and Annabelle Collin. Deciphering circulating tumor cells binding in a microfluidic system thanks to a parameterized mathematical model. Journal of Theoretical Biology, 600(??):??, March 7, 2025. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S002251932400314X.

Conte:2020:GII

[CGGG20] Martina Conte, Luca Gerardo-Giorda, and Maria Groppi. Glioma invasion and its interplay with nervous tissue and therapy: a multiscale model. Journal of Theoretical Biology, 486(??):??, February 7, 2020. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519319304576.

Cao:2023:MNC

[CGK23] Jiaming Cao, Pulkit Grover, and Jana M. Kainerstorfer. A model of neurovascular coupling and its application to cortical spreading depolarization. Journal of Theoretical Biology, 572(??):??, September 7, 2023. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519323001777.

Cheng:2020:GTA

[CGP^{+}20] Emily Cheng, Neeha Gambhirrao, Rohani Patel, Aufia Zhowandai, Jan Rychtář, and Dewey Taylor. A gametheoretical analysis of poliomyelitis vaccination. Journal of Theoretical Biology, 499(??):??, August 21, 2020. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519320301533.

Cuesta:2022:BBP

[CGR22] Fernando Alcalde Cuesta, Gustavo Guerberoff, and Álvaro Lozano Rojo. Bernoulli and binomial proliferation on evolutionary graphs. Journal of Theoretical Biology, 534(??):??, February 7, 2022. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519321003611.

Csillag:2025:BDE

[CGS^{+}25] Márton Csillag, Hamza Giaffar, Eörs Szathmáry, Mauro Santos, and Dániel Czégel. From Bayes to Darwin: Evolutionary search as an exaptation from sampling-based Bayesian inference. Journal of Theoretical Biology, 599(??):??, February 21, 2025. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519324003175.

Chang:2022:OCP

[CGW22] Lili Chang, Shupeng Gao, and Zhen Wang. Optimal control of pattern formations for an SIR reaction–diffusion epidemic model. Journal of Theoretical Biology, 536(??):??, March 7, 2022. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519322000017.

Caasi:2023:MMI

[CGY^{+}23] Jovic Aaron S. Caasi, Alex Leon Guerrero, Kangsan Yoon, Leslie J. Camacho Aquino, Aubrey Moore, Hyunju Oh, Jan Rychtář, and Dewey Taylor. A mathematical model of invasion and control of coconut rhinoceros beetle Oryctes rhinoceros (L.) in Guam. Journal of Theoretical Biology, 570(??):??, August 7, 2023. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519323001212.

Chauvet:2023:TAR

[CHMM23] Sophie Chauvet, Florence Hubert, Fanny Mann, and Mathieu Mezache. Tumorigenesis and axons regulation for the pancreatic cancer: a mathematical approach. Journal of Theoretical Biology, 556(??):??, January 7, 2023. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322002922.

Choua:2020:EVP

[CHSB20] Melinda Choua, Michael R. Heath, Douglas C. Speirs, and Juan A. Bonachela. The effect of viral plasticity on the persistence of host-virus systems. Journal of Theoretical Biology, 498(??):??, August 7, 2020. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519320301181.

Chen:2022:MEC

[CHT22] Liam Chen, Yousef Hannawi, and David Terman. Modeling the effect of cerebral capillary blood flow on neuronal firing. Journal of Theoretical Biology, 537(??):??, March 21, 2022. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519322000169.

Chaley:2020:SMD

[CK20a] Maria Chaley and Vladimir Kutyrkin. Stochastic models for description of structural-statistical properties in DNA sequences. Journal of Theoretical Biology, 496(??):??, July 7, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519319304953.

Cressman:2020:RCT

[CK20b] Ross Cressman and Vlastimil Krivan. Reducing courtship time promotes marital bliss: the Battle of the Sexes game revisited with costs measured as time lost. Journal of Theoretical Biology, 503(??):??, October 21, 2020. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251932030237X.

Chowdhury:2021:EED

[CKB^{+}21] Sayantan Nag Chowdhury, Srilena Kundu, Jeet Banerjee, Matjaz Perc, and Dibakar Ghosh. Eco-evolutionary dynamics of cooperation in the presence of policing. Journal of Theoretical Biology, 518(??):??, June 7, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251932100028X.

Cheon:2024:NCB

[CKF^{+}24] Hyejeong Lee Cheon, Nataliya Kizilova, Eirik G. Flekkøy, Matthew J. Mason, Lars P. Folkow, and Signe Kjelstrup. The nasal cavity of the bearded seal: an effective and robust organ for retaining body heat and water. Journal of Theoretical Biology, 595(??):??, December 7, 2024. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0022519324002182.

Chaplain:2020:JEM

[CKI20] Mark Chaplain, Denise Kirschner, and Yoh Iwasa. JTB Editorial: Malpractice: a case report. Journal of Theoretical Biology, 488(??):??, March 7, 2020. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519320300278.

Coggan:2020:ESM

[CKM^{+}20] Jay S. Coggan, Daniel Keller, Henry Markram, Felix Schürmann, and Pierre J. Magistretti. Excitation states of metabolic networks predict dose-response fingerprinting and ligand pulse phase signalling. Journal of Theoretical Biology, 487(??):??, February 21, 2020. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319304928.

Coggan:2022:RSI

[CKM^{+}22] Jay S. Coggan, Daniel Keller, Henry Markram, Felix Schürmann, and Pierre J. Magistretti. Representing stimulus information in an energy metabolism pathway. Journal of Theoretical Biology, 540(??):??, May 7, 2022. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322000881.

Crawford-Kahrl:2022:GNE

[CKNCG22] Peter Crawford-Kahrl, Robert R. Nerem, Bree Cummins, and Tomas Gedeon. Genetic networks encode secrets of their past. Journal of Theoretical Biology, 541(??):??, May 21, 2022. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251932200090X.

Choi:2024:DSP

[CKTW24] Kwok Pui Choi, Gursharn Kaur, Ariadne Thompson, and Taoyang Wu. Distributions of 4-subtree patterns for uniform random unrooted phylogenetic trees. Journal of Theoretical Biology, 584(??):??, May 7, 2024. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519324000754.

Chakraborty:2021:LTM

[CLD21] Jeevanjyoti Chakraborty, Jingxi Luo, and Rosemary J. Dyson. Lockhart with a twist: Modelling cellulose microfibril deposition and reorientation reveals twisting plant cell growth mechanisms. Journal of Theoretical Biology, 525(??):??, September 21, 2021. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519321001582.

Carson:2022:BCM

[CLF^{+}22] Jake Carson, Alice Ledda, Luca Ferretti, Matt Keeling, and Xavier Didelot. The bounded coalescent model: Conditioning a genealogy on a minimum root date. Journal of Theoretical Biology, 548(??):??, September 7, 2022. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322001849.

Crona:2020:ULM

[CLG20] Kristina Crona, Mengming Luo, and Devin Greene. An uncertainty law for microbial evolution. Journal of Theoretical Biology, 489(??):??, March 21, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519320300114.

Cheng:2023:MXS

[CLJ^{+}23] Ziqiang Cheng, Yinglei Lai, Kui Jin, Mengping Zhang, and Jin Wang. Modeling the XBB strain of SARS-CoV-2: Competition between variants and impact of reinfection. Journal of Theoretical Biology, 574(??):??, October 7, 2023. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519323002084.

Chen:2020:MEF

[CLLL20] Zhimin Chen, Kaihui Liu, Xiuxiang Liu, and Yijun Lou. Modelling epidemics with fractional-dose vaccination in response to limited vaccine supply. Journal of Theoretical Biology, 486(??):??, February 7, 2020. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519319304540.

Chae:2025:EBS

[CLM^{+}25] Junghun Chae, Roktaek Lim, Thomas L. P. Martin, Cheol-Min Ghim, and Pan-Jun Kim. Enlightening the blind spot of the Michaelis–Menten rate law: the role of relaxation dynamics in molecular complex formation. Journal of Theoretical Biology, 597(??):??, January 21, 2025. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519324002741.

Cho:2023:DEC

[CLSB23] Heyrim Cho, Allison L. Lewis, Kathleen M. Storey, and Helen M. Byrne. Designing experimental conditions to use the Lotka–Volterra model to infer tumor cell line interaction types. Journal of Theoretical Biology, 559(??):??, February 21, 2023. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251932200368X.

Cai:2024:LPP

[CLW^{+}24] Wangmin Cai, Peiqiang Liu, Zunfang Wang, Hong Jiang, Chang Liu, Zhaojie Fei, and Zhuang Yang. Link prediction in protein–protein interaction network: a similarity multiplied similarity algorithm with paths of length three. Journal of Theoretical Biology, 589(??):??, July 21, 2024. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519324001310.

Cervera:2023:TFD

[CMLM23] Javier Cervera, José A. Manzanares, Michael Levin, and Salvador Mafe. Transplantation of fragments from different planaria: a bioelectrical model for head regeneration. Journal of Theoretical Biology, 558(??):??, February 7, 2023. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322003472.

Chulian:2021:DPF

[CMRMC^{+}21] Salvador Chulián, Álvaro Martínez-Rubio, Anna Marciniak-Czochra, Thomas Stiehl, Cristina Blázquez Goñi, Juan Francisco Rodríguez Gutiérrez, Manuel Ramírez Orellana, Ana Castillo Robleda, Víctor M. Pérez-García, and María Rosa. Dynamical properties of feedback signalling in B lymphopoiesis: a mathematical modelling approach. Journal of Theoretical Biology, 522(??):??, August 7, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321001077.

Cropp:2020:ENT

[CN20] Roger Cropp and John Norbury. The emergence of new trophic levels in eco-evolutionary models with naturallybounded traits. Journal of Theoretical Biology, 496(??):??, July 7, 2020. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519320301193.

Cropp:2021:EEM

[CN21] Roger Cropp and John Norbury. The eco-evolutionary modelling of populations and their traits using a measure of trait differentiation. Journal of Theoretical Biology, 531(??):??, December 21, 2021. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251932100312X.

Campbell:2024:MMS

[CNK24] Kelly Campbell, Shailesh Naire, and Jan Herman Kuiper. A mathematical model of signalling molecule-mediated processes during regeneration of osteochondral defects after chondrocyte implantation. Journal of Theoretical Biology, 592(??):??, September 7, 2024. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519324001589.

Collis:2023:LDH

[COB23] Heather L. Collis, Markus R. Owen, and Leah R. Band. Long-distance hormone transport via the phloem. Journal of Theoretical Biology, 562(??):??, April 7, 2023. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519323000115.

Cohen:2020:FDM

[Coh20]  Noy Cohen.  Force distribution and multi-scale me-

chanics in smooth muscle tissues. Journal of Theoretical Biology, 491(??):??, April 21, 2020. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519320300448.

Cortez:2021:USA

[Cor21] Michael H. Cortez. Using sensitivity analysis to identify factors promoting higher versus lower infection prevalence in multi-host communities. Journal of Theoretical Biology, 526(??):??, October 7, 2021. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S0022519321001880. See corrigendum [Cor22].

Cortez:2022:CUS

[Cor22] Michael H. Cortez. Corrigendum to “Using sensitivity analysis to identify factors promoting higher versus lower infection prevalence in multi-host communities” [J. Theor. Biol. 526 (2021) 110766]. Journal of Theoretical Biology, 534(??):??, February 7, 2022. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519321003799. See [Cor21].

Chitpin:2023:MCU

[CP23] Justin G. Chitpin and Theodore J. Perkins. A Markov constraint to uniquely identify elementary flux mode weights in unimolecular metabolic networks. Journal of Theoretical Biology, 575(??):??, November 7, 2023. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519323002291.

Colman:2023:ARS

[CPEK23] Ewan Colman, Gavrila A. Puspitarani, Jessica Enright, and Rowland R. Kao. Ascertainment rate of SARS-CoV-2 infections from healthcare and community testing in the UK. Journal of Theoretical Biology, 558(??):??, February 7, 2023. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322003241.

Carrera-Pineyro:2020:CAV

[CPHL^{+}20] Daniel Carrera-Pineyro, Harley Hanes, Adam Litzler, Andrea McCormack, Josean Velazquez-Molina, Anuj Mubayi, Karen Ríos-Soto, and Christopher Kribs. Cost analysis of vaccination in tick-mouse transmission of Lyme disease. Journal of Theoretical Biology, 494(??):??, June 7, 2020. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519320301004.

Coletti:2021:MEI

[CPM21] Roberta Coletti, Andrea Pugliese, and Luca Marchetti. Modeling the effect of immunotherapies on human castrationresistant prostate cancer. Journal of Theoretical Biology, 509(??):??, January 21, 2021. CODEN JTBIAP.

ISSN 0022-5193 (print), 1095-8541 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0022519320303556.

Chekroun:2022:TEO

[CPMF^{+}22] Abdennasser Chekroun, Laurent Pujo-Menjouet, Steve Falcoz, Kamyine Tsuen, Kevin Yueh-Hsun Yang, and Jean-Philippe Berteau. Theoretical evidence of osteoblast selfinhibition after activation of the genetic regulatory network controlling mineralization. Journal of Theoretical Biology, 537(??):??, March 21, 2022. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519322000030.

Couto:2020:GRP

[CPS20] Marta C. Couto, Jorge M. Pacheco, and Francisco C. Santos. Governance of risky public goods under graduated punishment. Journal of Theoretical Biology, 505(??):??, November 21, 2020. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519320302782.

Cope:2020:IRT

[CR20] Robert C. Cope and Joshua V. Ross. Identification of the relative timing of infectiousness and symptom onset for outbreak control. Journal of Theoretical Biology, 486(??):??, February 7, 2020. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519319304485.

Capello:2022:SHA

[CRDD22] Manuela Capello, Jonathan Rault, Jean-Louis Deneubourg, and Laurent Dagorn. Schooling in habitats with aggregative sites: the case of tropical tuna and floating objects. Journal of Theoretical Biology, 547(??):??, August 21, 2022. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322001618.

Creswell:2023:BNM

[CRG^{+}23] Richard Creswell, Martin Robinson, David Gavaghan, Kris V. Parag, Chon Lok Lei, and Ben Lambert. A Bayesian nonparametric method for detecting rapid changes in disease transmission. Journal of Theoretical Biology, 558(??):??, February 7, 2023. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519322003423.

Capo-Rangel:2020:MPC

[CRGGSC20] G. Capo-Rangel, L. Gerardo-Giorda, E. Somersalo, and D. Calvetti. Metabolism plays a central role in the cortical spreading depression: Evidence from a mathematical model. Journal of Theoretical Biology, 486(??):??, February 7, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251931930462X.

Cruz-Rosas:2020:MSK

[CRRRP^{+}20] Hugo I. Cruz-Rosas, Francisco Riquelme, Alejandra Ramírez-Padrón, Thomas Buhse, Germinal Cocho, and Pedro Miramontes. Molecular shape as a key source of prebiotic information. Journal of Theoretical Biology, 499(??):??, August 21, 2020. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519320301715.

Choi:2020:OSV

[CS20] Wongyeong Choi and Eunha Shim. Optimal strategies for vaccination and social distancing in a gametheoretic epidemiologic model. Journal of Theoretical Biology, 505(??):??, November 21, 2020. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519320302770.

Choi:2021:OSS

[CS21] Wongyeong Choi and Eunha Shim. Optimal strategies for social distancing and testing to control COVID-19. Journal of Theoretical Biology, 512(??):??, March 7, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519320304239.

Czaran:2022:WSH

[CS22] Tamás Czárán and István Scheuring. Weak selection helps cheap but harms expensive cooperation in spatial threshold dilemmas. Journal of Theoretical Biology, 536(??):??, March 7, 2022. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251932100415X.

Curran-Sebastian:2025:PEP

[CSDH^{+}25] Jacob Curran-Sebastian, Louise Dyson, Edward M. Hill, Ian Hall, Lorenzo Pellis, and Thomas House. Probability of extinction and peak time for multi-type epidemics with application to COVID-19 variants of concern. Journal of Theoretical Biology, 608(??):??, July 7, 2025. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519325001018.

Cavallo:2020:DDO

[CSF20] James C. Cavallo, Steffen Scholpp, and Mark B. Flegg. Delay-driven oscillations via Axin2 feedback in the Wnt/ β-catenin signalling pathway. Journal of Theoretical Biology, 507(??):??, December 21, 2020. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519320303131.

Cueno:2025:NMD

[CSH^{+}25] Marni E. Cueno, Noa Shintaku, Noe Hayasaki, Yuna Migita, and Kenichi Imai. Network modeling of the different SARS-CoV-2 spike protein infection points within the human hematopoietic network. Journal of Theoretical Biology, 609(??):??, July 21, 2025. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519325001055.

Cao:2020:EPG

[CSHH20] Xiulei Cao, Zilong Song, Tzyy-Leng Horng, and Huaxiong Huang. Electric potential generation of electrocytes: Modelling, analysis, and computation. Journal of Theoretical Biology, 487(??):??, February 21, 2020. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931930476X.

Colombi:2021:HID

[CSP21] Annachiara Colombi, Marco Scianna, and Luigi Preziosi. A hybrid integro-differential model for the early development of the zebrafish posterior lateral line. Journal of Theoretical Biology, 514(??):??, April 7, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519320304331.

Colnaghi:2025:PAD

[CSVB25] Marco Colnaghi, Fernando P. Santos, Paul A. M. Van Lange, and Daniel Balliet. Power asymmetry destabilizes reciprocal cooperation in social dilemmas. Journal of Theoretical Biology, 606(??):??, June 7, 2025. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519325000724.

Chaplain:2024:PDK

[CSW24] Mark Chaplain, Akira Sasaki, and Joshua Weitz. Professor Denise Kirschner. Journal of Theoretical Biology, 595(??):??, December 7, 2024. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0022519324002613.

Crowley:2022:RCW

[CT22] Philip H. Crowley and Cédric Tentelier. Resolving conflict over within-pair mating rate in external fertilizers. Journal of Theoretical Biology, 532(??):??, January 7, 2022. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321003453.

Coton:2022:EEL

[CTL^{+}22] Charlotte Coton, Grégoire Talbot, Maud Le Louarn, Christine Dillmann, and Dominique de Vienne. Evolution of enzyme levels in metabolic pathways: a theoretical approach. Part 1. Journal of Theoretical Biology, 538(??):??, April 7, 2022. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519322000133.

Chen:2023:APS

[CWF23] Alex Chen, Timothy Wessler, and M. Gregory Forest. Antibody protection from SARS-CoV-2 respiratory tract exposure and infection. Journal of Theoretical Biology, 557(??):??, January 21, 2023. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519322003253.

Cho:2020:SDD

[CWL20] Heyrim Cho, Zuping Wang, and Doron Levy. Study of dose-dependent combination immunotherapy using engineered T cells and IL-2 in cervical cancer. Journal of Theoretical Biology, 505(??):??, November 21, 2020. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519320302587.

Clifton:2021:TCV

[CWR21] Sara M. Clifton, Rachel J. Whitaker, and Zoi Rapti. Temperate and chronic virus competition leads to low lysogen frequency. Journal of Theoretical Biology, 523(??):??, August 21, 2021. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519321001326.

Cueno:2023:SPS

[CWT^{+}23] Marni E. Cueno, Kanta Wada, Arisa Tsuji, Kouta Ishikawa, and Kenichi Imai. Structural patterns of SARS-CoV-2 variants of concern (alpha, beta, gamma, delta) spike protein are influenced by variant-specific amino acid mutations: a computational study with implications on viral evolution. Journal of Theoretical Biology, 558(??):??, February 7, 2023. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322003678.

Chen:2022:EDZ

[CWW22] Fang Chen, Te Wu, and Long Wang. Evolutionary dynamics of zero-determinant strategies in repeated multiplayer games. Journal of Theoretical Biology, 549(??):??, September 21, 2022. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519322002077.

Cai:2021:MEC

[CZN^{+}21] Yongli Cai, Shi Zhao, Yun Niu, Zhihang Peng, Kai Wang, Daihai He, and Weiming Wang. Modelling the effects of the contaminated environments on tuberculosis in Jiangsu, China. Journal of Theoretical Biology, 508(??):??, January 7, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519320303088.

DAgata:2020:CDS

[D’A20a] Antonio D’Agata. Corrigendum to ‘Dynamics of spatially dispersed population under ideal free distribution’ [Journal of Theoretical Biology 486 (2020) 110071]. Journal of Theoretical Biology, 487(??):??, February 21, 2020. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319304850. See [D’A20b].

DAgata:2020:DSD

[D’A20b] Antonio D’Agata. Dynamics of spatially dispersed population under ideal free distribution. Journal of Theoretical Biology, 486(??):??, February 7, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S0022519319304400. See corrigendum [D’A20a].

Dangerfield:2023:GMM

[DAB^{+}23] Ciara E. Dangerfield, I. David Abrahams, Chris Budd, Matt Butchers, Michael E. Cates, Alan R. Champneys, Christine S. M. Currie, Jessica Enright, Julia R. Gog, Alain Goriely, T. Déirdre Hollingsworth, Rebecca B. Hoyle, INI Professional Services, Valerie Isham, Joanna Jordan, Maha H. Kaouri, Kostas Kavoussanakis, Jane Leeks, Philip K. Maini, Christie Marr, Clare Merritt, Denis Mollison, Surajit Ray, Robin N. Thompson, Alexandra Wakefield, and Dawn Wasley. Getting the most out of maths: How to coordinate mathematical modelling research to support a pandemic, lessons learnt from three initiatives that were part of the COVID-19 response in the UK. Journal of Theoretical Biology, 557(??):??, January 21, 2023. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251932200323X.

Dalton:2022:HEB

[Dal22] Dave Dalton. How evolutionary biology can explain why human and a few marine mammal females are the only ones that are menopausal. Journal of Theoretical Biology, 543(??):??, June 21, 2022. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519322001217.

Prado:2020:NMD

[dAPdBMTS20] Carlos Henrique Britto de Assis Prado, Dilma Maria de Brito Melo Trovão, and João Paulo Souza. A network model for determining decomposition, topology, and properties of the woody crown. Journal of Theoretical Biology, 499(??):??, August 21, 2020. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519320301739.

Das:2022:PSB

[DB22a] Soutrick Das and Debashis Barik. Pulsatile signaling of bistable switches reveal the distinct nature of pulse processing by mutual activation and mutual inhibition loop. Journal of Theoretical Biology, 540(??):??, May 7, 2022. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251932200073X.

Diekman:2022:BLC

[DB22b] Casey O. Diekman and Amitabha Bose. Beyond the limits of circadian entrainment: Non-24-h sleep–wake disorder, shift work, and social jet lag. Journal of Theoretical Biology, 545(??):??, July 21, 2022. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519322001461.

Deb:2023:ETP

[DBD23] Smita Deb, Subhendu Bhandary, and Partha Sharathi Dutta. Evading tipping points in socio-mutualistic networks via structure mediated optimal strategy. Journal of Theoretical Biology, 567(??):??, June 21, 2023. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519323000905.

Delamonica:2023:CBM

[DBS23] Brenda Delamonica, Gábor Balázsi, and Michael Shub. Cusp bifurcation in a metastatic regulatory network. Journal of Theoretical Biology, 575(??):??, November 7, 2023. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519323002278.

Daegling:2021:ESS

[DBTI21] David J. Daegling, Henna D. Bhramdat, and Viviana Toro-Ibacache. Efficacy of shear strain gradients as an osteogenic stimulus. Journal of Theoretical Biology, 524(??):??, September 7, 2021. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519321001521.

Duran:2022:MIR

[DBW22] Cecilia Duran, Manuel Barcenas, and Qixuan Wang. Modeling of ionizing radiation induced hair follicle regenerative dynamics. Journal of Theoretical Biology, 555(??):??, December 21, 2022. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519322002740.

dePillis:2023:MMW

[dCC^{+}23] Lisette dePillis, Rebecca Caffrey, Ge Chen, Mark D. Dela, Leif Eldevik, Joseph McConnell, Shahrokh Shabahang, and Stephen A. Varvel. A mathematical model of the within-host kinetics of SARS-CoV-2 neutralizing antibodies following COVID-19 vaccination. Journal of Theoretical Biology, 556(??):??, January 7, 2023. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519322002715.

Docken:2023:RDE

[DCL23] Steffen S. Docken, Colleen E. Clancy, and Timothy J. Lewis. Rate-dependent effects of state-specific sodium channel blockers in cardiac tissue: Insights from idealized models. Journal of Theoretical Biology, 573(??):??, September 21, 2023. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519323001923.

Detmer:2022:FCD

[DCP^{+}22] A. Raine Detmer, Ross Cunning, Ferdinand Pfab, Alexandra L. Brown, Adrian C. Stier, Roger M. Nisbet, and Holly V. Moeller. Fertilization by coral-dwelling fish promotes coral growth but can exacerbate bleaching response. Journal of Theoretical Biology, 541(??):??, May 21, 2022. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322000856.

Diabate:2020:PET

[DCS20] Modibo Diabaté, Loren Coquille, and Adeline Samson. Parameter estimation and treatment optimization in a stochastic model for immunotherapy of cancer. Journal of Theoretical Biology, 502(??):??, October 7, 2020. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519320302149.

delValleRafo:2021:DDM

[dDA21] María del Valle Rafo, Juan Pablo Di Mauro, and Juan Pablo Aparicio. Disease dynamics and mean field models for clustered networks. Journal of Theoretical Biology, 526(??):??, October 7, 2021. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519320304094.

Djidjou-Demasse:2022:UDP

[DDDMT22] Ramsès Djidjou-Demasse, Arnaud Ducrot, Nicole Mideo, and Gaëtan Texier. Understanding dynamics of Plasmodium falciparum gametocytes production: Insightsl from an agestructured model. Journal of Theoretical Biology, 539(??):??, April 21, 2022. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519322000546.

Dupin:2023:PCA

[DEM^{+}23] Isabelle Dupin, Edmée Eyraud, Élise Maurat, Jean-Marc Sac-Épée, and Pierre Vallois. Probabilistic cellular automata modelling of intercellular interactions in airways: complex pattern formation in patients with chronic obstructive pulmonary disease. Journal of Theoretical Biology, 564(??):??, May 7, 2023. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519323000449.

Descary:2022:EBR

[DF22] Marie-Hélène Descary and Sorana Froda. Estimating the basic reproduction number from noisy daily data. Journal of Theoretical Biology, 549(??):??, September 21, 2022. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322002089.

Ladron-de-Guevara:2022:PPM

[dGSNK22] Antonio Ladrón de Guevara, Jessica K. Shang, Maiken Nedergaard, and Douglas H. Kelley. Perivascular pumping in the mouse brain: Improved boundary conditions reconcile theory, simulation, and experiment. Journal of Theoretical Biology, 542(??):??, June 7, 2022. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322001011.

Doro:2022:FTQ

[DH22] Stephen Doro and Matthew A. Herman. On the Fourier transform of a quantitative trait: Implications for compressive sensing. Journal of Theoretical Biology, 540(??):??, May 7, 2022. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519321004057.

Dale:2025:AMC

[DHC25] Renee Dale, Hongyu He, and Yingqing Chen. Absorbing Markov chain model of PrEP drug adherence to estimate adherence decay rate and probability distribution in clinical trials. Journal of Theoretical Biology, 604(??):??, May 7, 2025. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519325000529.

Duchen:2020:LMM

[DHLS20] Pablo Duchen, Sophie Hautphenne, Laurent Lehmann, and Nicolas Salamin. Linking micro and macroevolution in the presence of migration. Journal of Theoretical Biology, 486(??):??, February 7, 2020. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519319304564.

Didelot:2025:API

[DHR25] Xavier Didelot, David Helekal, and Ian Roberts. Ancestral process for infectious disease outbreaks with superspreading. Journal of Theoretical Biology, 607(??):??, June 21, 2025. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251932500075X.

DiGiulio:2020:CAE

[Di 20] Massimo Di Giulio. Common ancestry of eukaryotes and Asgardarchaeota: Three, two or more cellular domains of life? Journal of Theoretical Biology, 486(??):??, February 7, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519319304527.

DiGiulio:2021:OGC

[Di 21] Massimo Di Giulio. The origin of the genetic code and origin of ideas. Journal of Theoretical Biology, 516(??):??, May 7, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519321000370.

dOnofrio:2024:MPW

[dIMM24] Alberto d’Onofrio, Mimmo Iannelli, Gabriela Marinoschi, and Piero Manfredi. Multiple pandemic waves vs multi-period/ multi-phasic epidemics: Global shape of the COVID-19 pandemic. Journal of Theoretical Biology, 593(??):??, October 7, 2024. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519324001656.

Duan:2022:HMM

[DJ22] Moran Duan and Zhen Jin. The heterogeneous mixing model of COVID-19 with interventions. Journal of Theoretical Biology, 553(??):??, November 21, 2022. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322002508.

deJong:2020:MFH

[dJHP20] T. G. de Jong, J. Hulshof, and G. Prokert. Modelling fungal hypha tip growth via viscous sheet approximation. Journal of Theoretical Biology, 492(??):??, May 7, 2020. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251932030045X.

deKemmeter:2022:SSH

[dKCA22] Jean-François de Kemmeter, Timoteo Carletti, and Malbor Asllani. Self-segregation in heterogeneous metapopulation landscapes. Journal of Theoretical Biology, 554(??):??, December 7, 2022. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519322002624.

Dufek:2024:EDD

[DKSP24] Grant Dufek, Guy Katriel, Sagi Snir, and Matteo Pellegrini. Exponential dynamics of DNA methylation with age. Journal of Theoretical Biology, 579(??):??, February 21, 2024. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519323002941.

deLeon:2021:PBM

[dLFO21] Marco A. Avila Ponce de León, Bryan Félix, and Hans G. Othmer. A phosphoinositide-based model of actin waves in frustrated phagocytosis. Journal of Theoretical Biology, 527(??):??, October 21, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0022519321001867.

Donovan:2020:PPS

[DLN20] Graham M. Donovan, David Langton, and Peter B. Noble. Phenotype- and patient-specific modelling in asthma: Bronchial thermoplasty and uncertainty quantification. Journal of Theoretical Biology, 501(??):??, September 21, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519320301922.

Donovan:2024:ESE

[DLSA24] Graham M. Donovan, Congping Lin, Imogen Sparkes, and Peter Ashwin. Emergence and stability of endoplasmic reticulum network streaming in plant cells. Journal of Theoretical Biology, 595(??):??, December 7, 2024. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S002251932400239X.

Diouf:2021:CCA

[DMAC21] Abdoulaye Diouf, Houda Mokrani, Evans Afenya, and Baba Issa Camara. Computation of the conditions for antiangiogenesis and gene therapy synergistic effects: Sensitivity analysis and robustness of target solutions. Journal of Theoretical Biology, 528(??):??, November 7, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321002691.

Dague:2024:TQT

[DMPIG24] Philippe Dague, Laurent Muller, Loïc Paulevé, and Marc Irigoin-Guichandut. Towards a qualitative theory of the interruption of eating behavior change. Journal of Theoretical Biology, 581(??):??, March 21, 2024. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519324000122.

Quintela:2022:TAS

[dMQHRGA^{+}22] Bárbara de Melo Quintela, Silvia Hervas-Raluy, Jose Manuel Garcia-Aznar, Dawn Walker, Kenneth Y. Wertheim, and Marco Viceconti. A theoretical analysis of the scale separation in a model to predict solid tumour growth. Journal of Theoretical Biology, 547(??):??, August 21, 2022. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322001710.

Denes:2021:GAC

[DMR21] Attila Dénes, Sadegh Marzban, and Gergely Röst. Global analysis of a cancer model with drug resistance due to Lamarckian induction and microvesicle transfer. Journal of Theoretical Biology, 527(??):??, October 21, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321002319.

Denysova:2022:CAO

[DN22] Olga Denysova and Johannes M. Nitsche. Conclusions about osmotically inactive volume and osmotic fragility from a detailed erythrocyte model. Journal of Theoretical Biology, 539(??):??, April 21, 2022. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519321004021.

Gaetano:2024:SLM

[DNK^{+}24] Andrea De Gaetano, Ilona Nagy, Daniel Kiss, Valery G. Romanovski, and Thomas A. Hardy. A simplified longitudinal model for the development of Type 2 diabetes mellitus. Journal of Theoretical Biology, 587(??):??, June 21, 2024. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519324001036.

Domogo:2021:PSP

[DO21] Andrei A. Domogo and Johnny T. Ottesen. Patient-specific parameter estimation: Coupling a heart model and experimental data. Journal of Theoretical Biology, 526(??):??, October 7, 2021. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519321002101.

Dari:2025:URC

[DOF25] Sonia Dari, Reuben D. O’dea, and Nabil T. Fadai. Understanding the regulation of chronic wounds by tissue inhibitors of matrix metalloproteinases through mathematical modelling. Journal of Theoretical Biology, 604(??):??, May 7, 2025. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519325000499.

Donovan:2023:PTB

[Don23] Graham M. Donovan. Pillars of theoretical biology: Airway stability and heterogeneity in the constricted lung. Journal of Theoretical Biology, 575(??):??, November 7, 2023. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519323002254.

Deeley:2022:PIP

[DP22] Bradly Deeley and Natalia Petrovskaya. Propagation of invasive plant species in the presence of a road. Journal of Theoretical Biology, 548(??):??, September 7, 2022. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322001941.

Diaz-Pachon:2021:SCC

[DPR21] Daniel Andrés Díaz-Pachón and J. Sunil Rao. A simple correction for COVID-19 sampling bias. Journal of Theoretical Biology, 512(??):??, March 7, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519320304112.

Deyranlou:2023:EDA

[DRK23] Amin Deyranlou, Alistair Revell, and Amir Keshmiri. Exergy destruction in atrial fibrillation and a new ‘Exergy Age Index’. Journal of Theoretical Biology, 575(??):??, November 7, 2023. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519323002205.

Domogo:2023:MMM

[DRO23] Andrei A. Domogo, Peter Reinstrup, and Johnny T. Ottesen. Mechanistic-mathematical modeling of intracranial pressure (ICP) profiles over a single heart cycle. The fundament of the ICP curve form. Journal of Theoretical Biology, 564(??): ??, May 7, 2023. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519323000474.

DeGaetano:2021:MCG

[DSB^{+}21] Andrea De Gaetano, Sasikarn Sakulrang, Alessandro Borri, Dario Pitocco, Surattana Sungnul, and Elvin J. Moore. Modeling continuous glucose monitoring with fractional differential equations subject to shocks. Journal of Theoretical Biology, 526(??):??, October 7, 2021. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519321001983.

Dela:2022:MMG

[DSdP22] An Dela, Blerta Shtylla, and Lisette de Pillis. Multi-method global sensitivity analysis of mathematical models. Journal of Theoretical Biology, 546(??):??, August 7, 2022. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322001576.

Dodd:2023:MMM

[DSM^{+}23] Josie Dodd, Peter K. Sweby, Sean Mayes, Erik H. Murchie, Asha S. Karunaratne, Festo Massawe, and Marcus J. Tindall. A multiscale mathematical model describing the growth and development of bambara groundnut. Journal of Theoretical Biology, 560(??):??, March 7, 2023. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322003642.

Deng:2024:LIP

[DTLS24] Jie Deng, Washington Taylor, Simon A. Levin, and Serguei Saavedra. On the limits to invasion prediction using coexistence outcomes. Journal of Theoretical Biology, 577(??):??, January 21, 2024. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519323002710.

Deng:2021:JIM

[DTS21] Jiawei Deng, Sanyi Tang, and Hongying Shu. Joint impacts of media, vaccination and treatment on an epidemic Filippov model with application to COVID-19. Journal of Theoretical Biology, 523(??):??, August 21, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251932100120X.

Rafo:2020:SEN

[dVRA20] María del Valle Rafo and Juan Pablo Aparicio. Simple epidemic network model for highly heterogeneous populations. Journal of Theoretical Biology, 486(??):??, February 7, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319304254.

Donovan:2024:ATS

[DWNW24] Graham M. Donovan, Carolyn J. Wang, Peter B. Noble, and Kimberley C. W. Wang. Adipose tissue in the small airways: How much is enough to drive functional changes? Journal of Theoretical Biology, 588(??):??, July 7, 2024. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519324001164.

Dokukina:2021:CDI

[DYS^{+}21] Irina V. Dokukina, Mikhail V. Yamashev, Ekaterina A. Samarina, Oksana M. Tilinova, and Eugene A. Grachev. Calciumdependent insulin resistance in hepatocytes: mathematical model. Journal of Theoretical Biology, 522(??):??, August 7, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519321001065.

Eftimie:2021:MII

[EB21] Raluca Eftimie and Charlotte Barelle. Mathematical investigation of innate immune responses to lung cancer: the role of macrophages with mixed phenotypes. Journal of Theoretical Biology, 524(??):??, September 7, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321001612.

ElKhalifi:2024:SEI

[EB24] Mohamed El Khalifi and Tom Britton. SIRS epidemics with individual heterogeneity of immunity waning. Journal of Theoretical Biology, 587(??):??, June 21, 2024. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519324000961.

Ecoffet:2021:NBD

[EBA21] Paul Ecoffet, Nicolas Bredeche, and Jean-Baptiste André. Nothing better to do? Environment quality and the evolution of cooperation by partner choice. Journal of Theoretical Biology, 527(??):??, October 21, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321002241.

Ewing:2021:PBS

[EBK21] David A. Ewing, Vivian Blok, and Helen Kettle. A processbased, stage-structured model of potato cyst nematode population dynamics: Effects of temperature and resistance. Journal of Theoretical Biology, 522(??):??, August 7, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321001235.

Etienne:2020:CEB

[ECBR20] Thibault A. Etienne, Muriel Cocaign-Bousquet, and Delphine Ropers. Competitive effects in bacterial mRNA decay. Journal of Theoretical Biology, 504(??):??, November 7, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519320301880.

Escosio:2023:MBS

[ECH^{+}23] Rey Audie S. Escosio, Olive R. Cawiding, Bryan S. Hernandez, Renier G. Mendoza, Victoria May P. Mendoza, Rhudaina Z. Mohammad, Carlene P. C. Pilar-Arceo, Pamela Kim N. Salonga, Fatima Lois E. Suarez, Polly W. Sy, Thomas Herald M. Vergara, and Aurelio A. de los Reyes. A model-based strategy for the COVID-19 vaccine roll-out in the Philippines. Journal of Theoretical Biology, 573(??):??, September 21, 2023. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519323001935.

ElHajj:2023:EPI

[EE23] W. El Hajj and N. El Khatib. Effect of permeability on the initiation of Atherosclerosis modeled as an inflammatory process. Journal of Theoretical Biology, 564(??):??, May 7, 2023. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519323000577.

Erkan:2025:CNI

[EE25] Yasemin Erkan and Erdem Erkan. Channel noise induced stochastic effect of Hodgkin–Huxley neurons in a real classification task. Journal of Theoretical Biology, 599(??):??, February 21, 2025. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519324003138.

Esmaeili:2021:DDR

[EHA^{+}21] Shadisadat Esmaeili, Alan Hastings, Karen Abbott, Jonathan Machta, and Vahini Reddy Nareddy. Density dependent Resource Budget Model for alternate bearing. Journal of Theoretical Biology, 509(??):??, January 21, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519320303532.

Engelhard:2021:MMG

[EHP^{+}21] Cesar A. Gonzalez Engelhard, Allison P. Hodgkins, Essence E. Pearl, Paul K. Spears, Jan Rychtář, and Dewey Taylor. A mathematical model of Guinea worm disease in Chad with fish as intermediate transport hosts. Journal of Theoretical Biology, 521(??):??, July 21, 2021. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519321001053.

Ei:2021:ENK

[EIK^{+}21] Shin-Ichiro Ei, Hiroshi Ishii, Shigeru Kondo, Takashi Miura, and Yoshitaro Tanaka. Effective nonlocal kernels on reaction– diffusion networks. Journal of Theoretical Biology, 509(??):??, January 21, 2021. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519320303519.

Estavoyer:2025:MMF

[ELM25] Maxime Estavoyer, Thomas Lepoutre, and Marie Manceau. Mathematical modeling of the feather follicle morphogenetic wave in birds. Journal of Theoretical Biology, 612(??):??, September 7, 2025. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0022519325001663.

Esteve:2021:FDS

[ELRG21] Jorge Esteve, Matheo López, Carlos-Guillermo Ramírez, and Iván Gómez. Fluid dynamic simulation suggests hopping locomotion in the Ordovician trilobite Placoparia. Journal of Theoretical Biology, 531(??):??, December 21, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321003350.

Eksin:2021:RON

[ENMW21] Ceyhun Eksin, Martial Ndeffo-Mbah, and Joshua S. Weitz. Reacting to outbreaks at neighboring localities. Journal of Theoretical Biology, 520(??):??, July 7, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321000540.

Ellis:2020:CSD

[EP20] John R. Ellis and Natalia B. Petrovskaya. A computational study of density-dependent individual movement and the formation of population clusters in two-dimensional spatial domains. Journal of Theoretical Biology, 505(??):??, November 21, 2020. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519320302769.

Emami:2020:CPA

[EPF20] Neda Emami, Parvin Samadi Pakchin, and Reza Ferdousi. Computational predictive approaches for interaction and structure of aptamers. Journal of Theoretical Biology, 497(??):??, July 21, 2020. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519320301235.

Eigentler:2020:SSO

[ES20] L. Eigentler and J. A. Sherratt. Spatial self-organisation enables species coexistence in a model for savanna ecosystems. Journal of Theoretical Biology, 487(??):??, February 21, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519319304916.

Eigentler:2024:DLS

[ES24] Lukas Eigentler and Mattia Sensi. Delayed loss of stability of periodic travelling waves: Insights from the analysis of essential spectra. Journal of Theoretical Biology, 595(??):??, December 7, 2024. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0022519324002303.

Eilertsen:2024:UET

[ESW24] Justin Eilertsen, Santiago Schnell, and Sebastian Walcher. The unreasonable effectiveness of the total quasi-steady state approximation, and its limitations. Journal of Theoretical Biology, 583(??):??, April 21, 2024. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519324000511.

Eguchi:2024:DBF

[ETTT24] Go Eguchi, Tsutomu Takagi, Shinsuke Torisawa, and Kohsei Takehara. Drafting behaviors in fish induced by a local pressure drop around a hydrofoil model. Journal of Theoretical Biology, 588(??):??, July 7, 2024. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519324001024.

Eastman:2020:EPP

[EWK20] Brydon Eastman, Dominik Wodarz, and Mohammad Kohandel. The effects of phenotypic plasticity on the fixation probability of mutant cancer stem cells. Journal of Theoretical Biology, 503(??):??, October 21, 2020. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519320302393.

Ekanayake:2023:EPM

[EWS23] E. M. E. K. Ekanayake, J. Angus Webb, and Michael J. Stewardson. The effect of population model choice and network topologies on extinction time patterns in dendritic ecological networks. Journal of Theoretical Biology, 556(??):??, January 7, 2023. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519322003058.

Fischer:2023:TGT

[FB23] Matthias M. Fischer and Nils Blüthgen. On tumoural growth and treatment under cellular dedifferentiation. Journal of Theoretical Biology, 557(??):??, January 21, 2023. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322003186.

Fischer:2024:MTG

[FB24] Matthias M. Fischer and Nils Blüthgen. On minimising tumoural growth under treatment resistance. Journal of Theoretical Biology, 579(??):??, February 21, 2024. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519323003132.

Farahbakhsh:2021:BRD

[FBA21] Isaiah Farahbakhsh, Chris T. Bauch, and Madhur Anand. Best response dynamics improve sustainability and equity outcomes in common-pool resources problems, compared to imitation dynamics. Journal of Theoretical Biology, 509(??):??, January 21, 2021. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519320303313.

Fotinos:2023:EDT

[FBC23] J. Fotinós, L. Barberis, and C. A. Condat. Effects of a differentiating therapy on cancer-stem-cell-driven tumors. Journal of Theoretical Biology, 572(??):??, September 7, 2023. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519323001601.

Fraboul:2023:ASC

[FBD23] Jules Fraboul, Giulio Biroli, and Silvia De Monte. Artificial selection of communities drives the emergence of structured interactions. Journal of Theoretical Biology, 571(??):??, August 21, 2023. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519323001546.

Ferdousi:2020:NWW

[FBJR20] M. Ferdousi, T. Babaie-Janvier, and P. A. Robinson. Nonlinear wave-wave interactions in the brain. Journal of Theoretical Biology, 500(??):??, September 7, 2020. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519320301636.

Fic:2025:CMS

[FBMG25] Małgorzata Fic, Frank Bastian, Jacek Miȩkisz, and Chaitanya S. Gokhale. Compartment model of strategy-dependent time delays in replicator dynamics. Journal of Theoretical Biology, 601(??):??, March 21, 2025. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0022519325000104.

Fiandaca:2022:PSM

[FBSD22] Giada Fiandaca, Sara Bernardi, Marco Scianna, and Marcello Edoardo Delitala. A phenotype-structured model to reproduce the avascular growth of a tumor and its interaction with the surrounding environment. Journal of Theoretical Biology, 535(??):??, February 21, 2022. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321004008.

Fouedji:2024:MLP

[FET^{+}24] Chenceline Fouedji, Armand Sylvin Etémé, Conrad Bertrand Tabi, Henri Paul Ekobena Fouda, and Timoléon Crépin Kofané. Multisolitons-like patterns in a one-dimensional MARCKS protein cyclic model. Journal of Theoretical Biology, 579(??):??, February 21, 2024. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519323002990.

Flores:2021:SBP

[FFAV21] Lucas S. Flores, Heitor C. M. Fernandes, Marco A. Amaral, and Mendeli H. Vainstein. Symbiotic behaviour in the public goods game with altruistic punishment. Journal of Theoretical Biology, 524(??):??, September 7, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321001594.

Farina:2025:HDA

[FFGR25] Angiolo Farina, Antonio Fasano, Massimo Grimaldi, and Fabio Rosso. Heat delivery accompanying pulsed field ablation. Journal of Theoretical Biology, 610(??):??, August 7, 2025. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519325001110.

Flores-Ferrer:2021:EGD

[FFNG^{+}21] Alheli Flores-Ferrer, Anne Nguyen, Sylvain Glémin, Jean-Marc Deragon, Olivier Panaud, and Sébastien Gourbière. The ecology of the genome and the dynamics of the biological dark matter. Journal of Theoretical Biology, 518(??):??, June 7, 2021. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519321000631.

Farina:2022:MVA

[FFR22] Angiolo Farina, Antonio Fasano, and Fabio Rosso. Modeling of vasomotion in arterioles. Journal of Theoretical Biology, 544(??):??, July 7, 2022. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519322001229.

Farina:2023:TMF

[FFR23] Angiolo Farina, Antonio Fasano, and Fabio Rosso. A theoretical model for the Fåhræus effect in medium-large microvessels. Journal of Theoretical Biology, 558(??):??, February 7, 2023. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519322003460.

Fan:2021:VMC

[FG21] Ruili Fan and Stefan A. H. Geritz. Virulence management: Closing the feedback loop between healthcare interventions and virulence evolution. Journal of Theoretical Biology, 531(??):??, December 21, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321003192.

Fan:2022:MOH

[FG22] Ruili Fan and Stefan A. H. Geritz. Modelling and optimising healthcare interventions in a model with explicit within- and between-host dynamics. Journal of Theoretical Biology, 554(??):??, December 7, 2022. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322002673.

Fan:2023:EPC

[FG23] Ruili Fan and Stefan A. H. Geritz. Evolution of pathogens with cross-immunity in response to healthcare interventions. Journal of Theoretical Biology, 572(??):??, September 7, 2023. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519323001728.

Fotso:2024:MLM

[FGA^{+}24] Christian Tayou Fotso, Simon Girel, Fabienne Anjuère, Véronique M. Braud, Florence Hubert, and Thierry Goudon. A mixture-like model for tumor-immune system interactions. Journal of Theoretical Biology, 581(??):??, March 21, 2024. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519324000195.

Fung:2024:LTE

[FGC24] Tak Fung, Jonah Goh, and Ryan A. Chisholm. Longterm effects of non-pharmaceutical interventions on total disease burden in parsimonious epidemiological models. Journal of Theoretical Biology, 587(??):??, June 21, 2024. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519324000985.

Febvre:2024:TPE

[FGES24] Camille Febvre, Colin Goldblatt, and Rana El-Sabaawi. Thermal performance of ecosystems: Modeling how physiological responses to temperature scale up in communities. Journal of Theoretical Biology, 585(??):??, May 21, 2024. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519324000730.

Farmer:2023:MMC

[FH23] Adam Farmer and Paul J. Harris. A mathematical model of cell movement and clustering due to chemotaxis. Journal of Theoretical Biology, 575(??):??, November 7, 2023. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519323002436.

Flaig:2020:CEP

[FHM20] Julien Flaig, Nicolas Houy, and Philippe Michel. Cost effectiveness and policy announcement: the case of measles mandatory vaccination. Journal of Theoretical Biology, 485(??):??, January 21, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0022519319303984.

Fischer-Holzhausen:2022:HRO

[FHR22] Sophie Fischer-Holzhausen and Susanna Röblitz. Hormonal regulation of ovarian follicle growth in humans: Model-based exploration of cycle variability and parameter sensitivities. Journal of Theoretical Biology, 547(??):??, August 21, 2022. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322001485.

Fujita:2020:SRR

[FHTK20] Hironori Fujita, Mika Hayashi-Tsugane, and Masayoshi Kawaguchi. Spatial regulation of resource allocation in response to nutritional availability. Journal of Theoretical Biology, 486(??):??, February 7, 2020. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519319304473.

FitzGerald:2021:PMR

[FK21] Cody E. FitzGerald and James P. Keener. The PafBCmediated response sensitizes a bistable DNA damage response in Mycobacteria. Journal of Theoretical Biology, 508(??):??, January 7, 2021. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519320303179.

Falconer:2025:TLT

[FLO25] Alistair D. Falconer, Zhiyong Li, and Dietmar B. Oelz. Twolane track geometry facilitates coordination of collective cell migration. Journal of Theoretical Biology, 606(??):??, June 7, 2025. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519325000712.

Frank:2021:BSA

[FLR^{+}21] Anna S. Frank, Kamila Larripa, Hwayeon Ryu, Ryan G. Snodgrass, and Susanna Röblitz. Bifurcation and sensitivity analysis reveal key drivers of multistability in a model of macrophage polarization. Journal of Theoretical Biology, 509(??):??, January 21, 2021. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519320303660.

Flandoli:2023:MMC

[FLR23] Franco Flandoli, Marta Leocata, and Cristiano Ricci. The mathematical modeling of cancer growth and angiogenesis by an individual based interacting system. Journal of Theoretical Biology, 562(??):??, April 7, 2023. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519323000280.

Frank:2023:MPT

[FLRR23] Anna-Simone Josefine Frank, Kamila Larripa, Hwayeon Ryu, and Susanna Röblitz. Macrophage phenotype transitions in a stochastic gene-regulatory network model. Journal of Theoretical Biology, 575(??):??, November 7, 2023. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251932300231X.

Fahimi:2023:SMP

[FMO23] Peyman Fahimi, Chérif F. Matta, and Jordan G. Okie. Are size and mitochondrial power of cells inter-determined? Journal of Theoretical Biology, 572(??):??, September 7, 2023. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519323001625.

Foroutannia:2021:CMF

[FNGJ21] Ali Foroutannia, Fahimeh Nazarimehr, Mahdieh Ghasemi, and Sajad Jafari. Chaos in memory function of sleep: a nonlinear dynamical analysis in thalamocortical study. Journal of Theoretical Biology, 528(??):??, November 7, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321002563.

Forsdyke:2024:GCC

[For24] Donald R. Forsdyke. Genomic compliance with Chargaff’s second parity rule may have originated non-adaptively, but stem-loops now function adaptively. Journal of Theoretical Biology, 595(??):??, December 7, 2024. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0022519324002285.

Fornalski:2025:SCI

[For25] Krzysztof Wojciech Fornalski. The short comment on the individual response to ionizing radiation. Journal of Theoretical Biology, 604(??):??, May 7, 2025. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251932500058X.

Fagundes:2021:RMF

[FOY21] Tadeu M. Fagundes, Juan C. Ordonez, and Neda Yaghoobian. The role of mound functions and local environment in the diversity of termite mound structures. Journal of Theoretical Biology, 527(??):??, October 21, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321002423.

Fernandez-Peralta:2021:SMC

[FPGC21] R. Fernández-Peralta and A. Gómez-Corral. A structured Markov chain model to investigate the effects of pre-exposure vaccines in tuberculosis control. Journal of Theoretical Biology, 509(??):??, January 21, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519320303453.

Flores-Perez:2025:TDE

[FPGOCP^{+}25] Anahí Flores-Pérez, Marcos A. González-Olvera, Gustavo Chávez-Peña, Ana G. Gallardo-Hernández, and Lizeth Torres. Time-delay enhanced SIR model for COVID-19 waves in Mexico: Parameter estimation using evolutionary algorithms. Journal of Theoretical Biology, 613(??):??, October 7, 2025. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251932500195X.

Franco:2024:NIC

[FPS24] Daniel Franco, Juan Perán, and Juan Segura. New insights into the combined effect of dispersal and local dynamics in a two-patch population model. Journal of Theoretical Biology, 595(??):??, December 7, 2024. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0022519324002273.

Fang:2020:SMP

[FPX20] Yinhai Fang, Matjaz Perc, and Haiyan Xu. The Singaporean model in public goods dilemmas with benevolent leaders and bribery. Journal of Theoretical Biology, 501(??):??, September 21, 2020. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519320302009.

Frascoli:2022:DRD

[FRMK22] Federico Frascoli, Izanne Roos, Charles B. Malpas, and Tomas Kalincik. The dynamics of relapses during treatment switch in relapsing-remitting multiple sclerosis. Journal of Theoretical Biology, 541(??):??, May 21, 2022. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322000893.

Filar:2022:SRI

[FS22] Jerzy Filar and Sabrina Streipert. Square root identities for harvested Beverton–Holt models. Journal of Theoretical Biology, 549(??):??, September 21, 2022. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322001977.

Fontanari:2024:SPD

[FS24] José F. Fontanari and Mauro Santos. Solving the prisoner’s dilemma trap in Hamilton’s model of temporarily formed random groups. Journal of Theoretical Biology, 595(??):??, December 7, 2024. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0022519324002315.

Franssen:2021:NAC

[FSC21] Linnea C. Franssen, Nikolaos Sfakianakis, and Mark A. J. Chaplain. A novel 3D atomistic-continuum cancer invasion model: In silico simulations of an in vitro organotypic invasion assay. Journal of Theoretical Biology, 522(??):??, August 7, 2021. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519321000990.

Forouzannia:2021:IRM

[FSK21] Farinaz Forouzannia, Vahid Shahrezaei, and Mohammad Kohandel. The impact of random microenvironmental fluctuations on tumor control probability. Journal of Theoretical Biology, 509(??):??, January 21, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519320303490.

Fornari:2020:SEN

[FSKG20] Sveva Fornari, Amelie Schäfer, Ellen Kuhl, and Alain Goriely. Spatially-extended nucleation–aggregation–fragmentation models for the dynamics of prion-like neurodegenerative proteinspreading in the brain and its connectome. Journal of Theoretical Biology, 486(??):??, February 7, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519319304710. See comment [GMMR22].

Frank:2021:PDR

[FSKG21] A. Frank, S. Subbey, M. Kobras, and H. Gjøsæter. Population dynamic regulators in an empirical predator–prey system. Journal of Theoretical Biology, 527(??):??, October 21, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519321002332.

Farzanfar:2020:PCV

[FSNS20] Arezoo Farzanfar, Farzaneh Shayegh, Behzad Nazari, and Saeid Sadri. Physiological constraints of visual pathway lead to more efficient coding of information in retina. Journal of Theoretical Biology, 506(??):??, December 7, 2020. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519320302733.

Fazli:2020:PBM

[FVB20] Mehran Fazli, Theodore Vo, and Richard Bertram. Phantom bursting may underlie electrical bursting in single pancreatic β-cells. Journal of Theoretical Biology, 501(??):??, September 21, 2020. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519320302010.

Fung:2020:PDE

[FVC20] Tak Fung, Sonali Verma, and Ryan A. Chisholm. Probability distributions of extinction times, species richness, and immigration and extinction rates in neutral ecological models. Journal of Theoretical Biology, 485(??):??, January 21, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519319304205.

Fletcher:2022:NCM

[FZE22] Alvaro Fletcher, Ruonan Zhao, and German Enciso. Noncooperative mechanism for bounded and ultrasensitive chromatin remodeling. Journal of Theoretical Biology, 534(??):??, February 7, 2022. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519321003659.

Garzon-Alvarado:2024:PIN

[GADDVG^{+}24] Diego A. Garzón-Alvarado, Carlos A. Duque-Daza, Juan Jairo Vaca-González, Abdelkader Boucetta, Dorian L. Linero, Gregory de Boer, Raj Das, and Salah Ramtani. Part II: a new perspective for modeling the bone remodeling process: Biology, mechanics, and pathologies. Journal of Theoretical Biology, 593(??):??, October 7, 2024. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519324001784.

Gardner:2023:WDH

[Gar23] Andy Gardner. W. D. Hamilton and the golden sex ratio. Journal of Theoretical Biology, 573(??):??, September 21, 2023. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519323001960.

Gardner:2024:GAE

[Gar24] Andy Gardner. A geometric approach to the evolution of altruism. Journal of Theoretical Biology, 576(??):??, January 7, 2024. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519323002503.

Gavish:2025:REP

[Gav25] Nir Gavish. Revisiting the exclusion principle in epidemiology at the limit of a large competitive advantage. Journal of Theoretical Biology, 600(??):??, March 7, 2025. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0022519325000116.

Gabriel:2023:TID

[GB23] Jean-Pierre Gabriel and Louis-Félix Bersier. Theoretical investigation of the dependence of population parameters on initial population size. Journal of Theoretical Biology, 568(??):??, July 7, 2023. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519323000887.

Garcia:2020:CII

[GBF20] Victor Garcia, Sebastian Bonhoeffer, and Feng Fu. Cancerinduced immunosuppression can enable effectiveness of immunotherapy through bistability generation: a mathematical and computational examination. Journal of Theoretical Biology, 492(??):??, May 7, 2020. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519320300412.

Giral-Barajas:2023:PAS

[GBHNHVL23] José Giral-Barajas, Carlos Ignacio Herrera-Nolasco, Marco Arieli Herrera-Valdez, and Sergio I. López. A probabilistic approach for the study of epidemiological dynamics of infectious diseases: Basic model and properties. Journal of Theoretical Biology, 572(??):??, September 7, 2023. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251932300173X.

Ghosh:2021:RAC

[GBRM21] Aditi Ghosh, Isaac Berger, Christopher H. Remien, and Anuj Mubayi. The role of alcohol consumption on acetaminophen induced liver injury: Implications from a mathematical model. Journal of Theoretical Biology, 519(??):??, June 21, 2021. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519320304148.

Gray:2020:IAT

[GC20] Catheryn W. Gray and Adelle C. F. Coster. From insulin to Akt: Time delays and dominant processes. Journal of Theoretical Biology, 507(??):??, December 21, 2020. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251932030309X.

Guo:2020:DSI

[GCA^{+}20] Xiaohui Guo, Jun Chen, Asma Azizi, Jennifer Fewell, and Yun Kang. Dynamics of social interactions, in the flow of information and disease spreading in social insects colonies: Effects of environmental events and spatial heterogeneity. Journal of Theoretical Biology, 492(??):??, May 7, 2020. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519320300461.

Gawthrop:2020:PPM

[GCC20] Peter J. Gawthrop, Peter Cudmore, and Edmund J. Crampin. Physically-plausible modelling of biomolecular systems: a simplified, energy-based model of the mitochondrial electron transport chain. Journal of Theoretical Biology, 493(??):??, May 21, 2020. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519320300783.

Garay:2023:SSA

[GCM23] József Garay, Villő Csiszár, and Tamás F. Móri. Subsistence of sib altruism in different mating systems and Haldane’s arithmetic. Journal of Theoretical Biology, 557(??):??, January 21, 2023. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519322003216.

Garzon:2021:DPP

[GCNZ^{+}21] Doménica N. Garzón, Yair Castillo, M. Gabriela Navas-Zuloaga, Leah Darwin, Abigail Hardin, Nora Culik, Anji Yang, Carlos Castillo-Garsow, Karen Ríos-Soto, Leon Arriola, and Aditi Ghosh. Dynamics of prion proliferation under combined treatment of pharmacological chaperones and interferons. Journal of Theoretical Biology, 527(??):??, October 21, 2021. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519321002162.

Garay:2020:WOF

[GCX^{+}20] József Garay, Ross Cressman, Fei Xu, Mark Broom, Villő Csiszár, and Tamás F. Móri. When optimal foragers meet in a game theoretical conflict: a model of kleptoparasitism. Journal of Theoretical Biology, 502(??):??, October 7, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519320301612.

Ghosh:2025:DCO

[GDC^{+}25] Adrita Ghosh, Parthasakha Das, Tanujit Chakraborty, Pritha Das, and Dibakar Ghosh. Developing cholera outbreak forecasting through qualitative dynamics: Insights into Malawi case study. Journal of Theoretical Biology, 605(??):??, May 21, 2025. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519325000633.

Gandolfi:2021:AVC

[GDd^{+}21] A. Gandolfi, S. De Franciscis, A. d’Onofrio, A. Fasano, and C. Sinisgalli. Angiogenesis and vessel co-option in a mathematical model of diffusive tumor growth: the role of chemotaxis. Journal of Theoretical Biology, 512(??):??, March 7, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519320303817.

Gunnarsson:2020:URP

[GDLF20] Einar Bjarki Gunnarsson, Subhajyoti De, Kevin Leder, and Jasmine Foo. Understanding the role of phenotypic switching in cancer drug resistance. Journal of Theoretical Biology, 490(??):??, April 7, 2020. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519320300187.

Guo:2023:HID

[GDQR23] Ting Guo, Qi Deng, Zhipeng Qiu, and Libin Rong. HIV infection dynamics and viral rebound: Modeling results from humanized mice. Journal of Theoretical Biology, 567(??):??, June 21, 2023. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519323000863.

Gomez:2022:SES

[GEHL^{+}22] Angela Maria Gomez, J. F. Escobar-Huertas, D. L. Linero, F. P. Cardenas, and D. A. Garzón-Alvarado. Simulation of the electrical stimulation of the rat brain using sleep frequencies: a finite element modeling approach. Journal of Theoretical Biology, 542(??):??, June 7, 2022. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322000911.

Goodin:2023:EIA

[GF23] Dylan A. Goodin and Hermann B. Frieboes. Evaluation of innate and adaptive immune system interactions in the tumor microenvironment via a 3D continuum model. Journal of Theoretical Biology, 559(??):??, February 21, 2023. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322003745.

Garakani:2025:EHR

[GFAP^{+}25] Spalding Garakani, Luis Flores, Guillermo Alvarez-Pardo, Jan Rychtár, and Dewey Taylor. The effect of heterogeneity of relative vaccine costs on the mean population vaccination rate with mpox as an example. Journal of Theoretical Biology, 602–603(??):??, April 7, 2025. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519325000281.

Gomes:2022:IVS

[GFC^{+}22] M. Gabriela M. Gomes, Marcelo U. Ferreira, Rodrigo M. Corder, Jessica G. King, Caetano Souto-Maior, Carlos Penha-Gonçalves, Guilherme Gonçalves, Maria Chikina, Wesley Pegden, and Ricardo Aguas. Individual variation in susceptibility or exposure to SARS-CoV-2 lowers the herd immunity threshold. Journal of Theoretical Biology, 540(??):??, May 7, 2022. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519322000613.

Ganyani:2020:IGG

[GFH20] Tapiwa Ganyani, Christel Faes, and Niel Hens. Inference of the generalized-growth model via maximum likelihood estimation: a reflection on the impact of overdispersion. Journal of Theoretical Biology, 484(??):??, January 7, 2020. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319303996.

Gunnarsson:2023:SIR

[GFL23] Einar Bjarki Gunnarsson, Jasmine Foo, and Kevin Leder. Statistical inference of the rates of cell proliferation and phenotypic switching in cancer. Journal of Theoretical Biology, 568(??):??, July 7, 2023. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519323000930.

Ghasemi:2023:PCS

[GFN23] Mahdieh Ghasemi, Ali Foroutannia, and Fatemeh Nikdelfaz. A PID controller for synchronization between master-slave neurons in fractional-order of neocortical network model. Journal of Theoretical Biology, 556(??):??, January 7, 2023. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322003022.

Glasser:2023:ASS

[GFV^{+}23] John W. Glasser, Zhilan Feng, MyVan Vo, Jefferson N. Jones, and Kristie E. N. Clarke. Analysis of serological surveys of antibodies to SARS-CoV-2 in the United States to estimate parameters needed for transmission modeling and to evaluate and improve the accuracy of predictions. Journal of Theoretical Biology, 556(??):??, January 7, 2023. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322002879.

Gutierrez:2023:IVI

[GG23] Maria A. Gutierrez and Julia R. Gog. The importance of vaccinated individuals to population-level evolution of pathogens. Journal of Theoretical Biology, 567(??):??, June 21, 2023. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519323000899.

Goy:2021:AVT

[GGG21] Natascha Goy, Simone M. Glaser, and Christoph Grüter. The adaptive value of tandem communication in ants: Insights from an agent-based model. Journal of Theoretical Biology, 526(??):??, October 7, 2021. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519321001843.

Gantseva:2024:KAP

[GGS^{+}24] A. R. Gantseva, E. R. Gantseva, A. N. Sveshnikova, M. A. Panteleev, and T. A. Kovalenko. Kinetic analysis of prothrombinase assembly and substrate delivery mechanisms. Journal of Theoretical Biology, 594(??):??, November 7, 2024. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519324002108.

Gevertz:2021:NCE

[GGSTS21] Jana L. Gevertz, James M. Greene, Cynthia H. Sanchez-Tapia, and Eduardo D. Sontag. A novel COVID-19 epidemiological model with explicit susceptible and asymptomatic isolation compartments reveals unexpected consequences of timing social distancing. Journal of Theoretical Biology, 510(??):??, February 7, 2021. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519320303945.

Ghaderi:2020:SCM

[GHA^{+}20] Susan Ghaderi, Hulda S. Haraldsdóttir, Masoud Ahookhosh, Sylvain Arreckx, and Ronan M. T. Fleming. Structural conserved moiety splitting of a stoichiometric matrix. Journal of Theoretical Biology, 499(??):??, August 21, 2020. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519320301314.

Gasior:2022:MIC

[GHB22] Kelsey Gasior, Marlene Hauck, and Sudin Bhattacharya. Modeling the influence of cell-cell contact and TGF-β signaling on the epithelial mesenchymal transition in MCF7 breast carcinoma cells. Journal of Theoretical Biology, 546(??):??, August 7, 2022. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519322001588.

Grunnill:2022:CYA

[GHF22] Martin Grunnill, Ian Hall, and Thomas Finnie. Check your assumptions: Further scrutiny of basic model frameworks of antimicrobial resistance. Journal of Theoretical Biology, 554(??):??, December 7, 2022. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519322002685.

Gunther:2018:BMS

[GHS18] Michael Günther, Daniel F. B. Haeufle, and Syn Schmitt. The basic mechanical structure of the skeletal muscle machinery: One model for linking microscopic and macroscopic scales. Journal of Theoretical Biology, 456(??):137–167, November 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519318303473. See [GHS20].

Gunther:2020:CBM

[GHS20] Michael Günther, Daniel F. B. Haeufle, and Syn Schmitt. Corrigendum to “The basic mechanical structure of the skeletal muscle machinery: One model for linking microscopic and macroscopic scales” [Journal of Theoretical Biology 456 (2018) 137–167]. Journal of Theoretical Biology, 488(??):??, March 7, 2020. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519319305132. See [GHS18].

Gutierrez-Jara:2020:SOP

[GJCLMQC20] J. P. Gutiérrez-Jara, F. D. Córdova-Lepe, M. T. Muñoz-Quezada, and G. Chowell. Susceptibility to organophosphates pesticides and the development of infectious-contagious respiratory diseases. Journal of Theoretical Biology, 488(??):??, March 7, 2020. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519319305028.

Ghasemi:2020:SNO

[GJDS20] Maryam Ghasemi, Benjamin Jenkins, Andrew C. Doxey, and Sivabal Sivaloganathan. A study of nitric oxide dynamics in a growing biofilm using a density dependent reaction–diffusion model. Journal of Theoretical Biology, 485(??):??, January 21, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519319304229.

Gibo:2020:TSR

[GK20] Shingo Gibo and Gen Kurosawa. Theoretical study on the regulation of circadian rhythms by RNA methylation. Journal of Theoretical Biology, 490(??):??, April 7, 2020. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319305090.

Giri:2021:IMB

[GK21] Amitava Giri and Sandip Kar. Incoherent modulation of bistable dynamics orchestrates the Mushroom and Isola bifurcations. Journal of Theoretical Biology, 530(??):??, December 7, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519321003015.

Giverso:2023:COU

[GLLP23] Chiara Giverso, Nadia Loy, Giulio Lucci, and Luigi Preziosi. Cell orientation under stretch: a review of experimental findings and mathematical modelling. Journal of Theoretical Biology, 572(??):??, September 7, 2023. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519323001613.

Gao:2022:TSM

[GMMR22] Shasha Gao, Maia Martcheva, Hongyu Miao, and Libin Rong. A two-sex model of human papillomavirus infection: Vaccination strategies and a case study. Journal of Theoretical Biology, 536(??):??, March 7, 2022. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322000042.

Giacobbi:2024:MMA

[GMR^{+}24] Anne-Sophie Giacobbi, Leo Meyer, Magali Ribot, Romain Yvinec, Hedi Soula, and Chloe Audebert. Mathematical modeling of adipocyte size distributions: Identifiability and parameter estimation from rat data. Journal of Theoretical Biology, 581(??):??, March 21, 2024. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519324000286.

Guerrero-Morin:2020:CDB

[GMS20] José G. Guerrero-Morín and Moisés Santillán. Crosstalk dynamics between the circadian clock and the mTORC1 pathway. Journal of Theoretical Biology, 501(??):??, September 21, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519320302150.

Gupta:2020:PDR

[GMV^{+}20] Ankit Gupta, Marc Manceau, Timothy Vaughan, Mustafa Khammash, and Tanja Stadler. The probability distribution of the reconstructed phylogenetic tree with occurrence data. Journal of Theoretical Biology, 488(??):??, March 7, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519319304849.

Germano:2021:MMC

[GO21] Domenic P. J. Germano and James M. Osborne. A mathematical model of cell fate selection on a dynamic tissue. Journal of Theoretical Biology, 514(??):??, April 7, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519320303908.

Goyal:2021:MRN

[Goy21] Ashish Goyal. Modeling reveals no direct role of the extent of HBV DNA integrations on the outcome of infection. Journal of Theoretical Biology, 526(??):??, October 7, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321002125.

Greenman:2020:CGR

[GPDW20] Chris D. Greenman, Luca Penso-Dolfin, and Taoyang Wu. The complexity of genome rearrangement combinatorics under the infinite sites model. Journal of Theoretical Biology, 501(??):??, September 21, 2020. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519320301909.

Golnaraghi:2023:OFS

[GQG23] Farnaz Golnaraghi, David A. Quint, and Ajay Gopinathan. Optimal foraging strategies for mutually avoiding competitors. Journal of Theoretical Biology, 570(??):??, August 7, 2023. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519323001339.

Guo:2021:MHM

[GQK^{+}21] Ting Guo, Zhipeng Qiu, Kosaku Kitagawa, Shingo Iwami, and Libin Rong. Modeling HIV multiple infection. Journal of Theoretical Biology, 509(??):??, January 21, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251932030357X.

Greulich:2025:CSD

[Gre25] Philip Greulich. Cooperative SIR dynamics as a model for spontaneous blood clot initiation. Journal of Theoretical Biology, 598(??):??, February 7, 2025. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519324002765.

Ghosh:2024:EED

[GRPG24] Subrata Ghosh, Sourav Roy, Matjaz Perc, and Dibakar Ghosh. The eco-evolutionary dynamics of two strategic species: From the predator-prey to the innocent-spreader rumor model. Journal of Theoretical Biology, 595(??):??, December 7, 2024. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0022519324002406.

Gunther:2021:RNF

[GRW^{+}21] Michael Günther, Robert Rockenfeller, Tom Weihmann, Daniel F. B. Haeufle, Thomas Götz, and Syn Schmitt. Rules of nature’s Formula Run: Muscle mechanics during late stance is the key to explaining maximum running speed. Journal of Theoretical Biology, 523(??):??, August 21, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321001363.

Goddard:2023:ERP

[GSC23] Jerome Goddard, Ratnasingham Shivaji, and James T. Cronin. Ecological release and patch geometry can cause nonlinear density-area relationships. Journal of Theoretical Biology, 557(??):??, January 21, 2023. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519322003162.

Gholampour:2022:ESB

[GSMG22] Maryam Gholampour, Ali Khaki Sedigh, Mohammad Ghassem Mahjani, and Abdorasoul Ghasemi. Eigenvalue sensitivity-based analysis for evaluation of biological network stability versus disturbances. Journal of Theoretical Biology, 533(??):??, January 21, 2022. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251932100360X.

Ghosh:2023:COW

[GSR23] Salil Ghosh, Shubhankar Saha, and Priti Kumar Roy. Critical observation of WHO recommended multidrug therapy on the disease leprosy through mathematical study. Journal of Theoretical Biology, 567(??):??, June 21, 2023. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519323000929.

Gao:2023:MSM

[GSW^{+}23] Shasha Gao, Mingwang Shen, Xueying Wang, Jin Wang, Maia Martcheva, and Libin Rong. A multi-strain model with asymptomatic transmission: Application to COVID-19 in the US. Journal of Theoretical Biology, 565(??):??, May 21, 2023. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519323000644.

Ganguly:2020:MCM

[GT20] Krishnendu Ganguly and Surendra Kumar Trigun. Mapping connectome in mammalian brain: a novel approach by bioengineering neuro-glia specific vectors. Journal of Theoretical Biology, 496(??):??, July 7, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519320300990.

Galimzyanov:2022:DMP

[GT22] A. V. Galimzyanov and R. N. Tchuraev. Dynamic mechanism of phase variation in bacteria based on multistable gene regulatory networks. Journal of Theoretical Biology, 549(??):??, September 21, 2022. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519322002107.

Giaretta:2020:SMH

[GTE20] Alberto Giaretta, Gianna Maria Toffolo, and Timothy C. Elston. Stochastic modeling of human papillomavirus early promoter gene regulation. Journal of Theoretical Biology, 486(??):??, February 7, 2020. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519319304266.

Gallo:2021:MUN

[GTGB21] Anthony J. Gallo, Hayden Tronnolone, J. Edward F. Green, and Benjamin J. Binder. Modelling uniaxial nonuniform yeast colony growth: Comparing an agent-based model and continuum approximations. Journal of Theoretical Biology, 523(??):??, August 21, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519321001375.

Goerlitz:2020:NRB

[GtHH20] Holger R. Goerlitz, Hannah M. ter Hofstede, and Marc W. Holderied. Neural representation of bat predation risk and evasive flight in moths: a modelling approach. Journal of Theoretical Biology, 486(??):??, February 7, 2020. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319304515.

Glover:2020:VHM

[GW20] Andrew Glover and Andrew White. A vector–host model to assess the impact of superinfection exclusion on vaccination strategies using dengue and yellow fever as case studies. Journal of Theoretical Biology, 484(??):??, January 7, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519319303844.

Gumbel:2021:MLC

[GW21] M. Gumbel and P. Wiedemann. Motif lengths of circular codes in coding sequences. Journal of Theoretical Biology, 523(??):??, August 21, 2021. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519321001302.

Gao:2022:DDM

[GWZ^{+}22] Jia-Xing Gao, Zhen-Yi Wang, Michael Q. Zhang, Min-Ping Qian, and Da-Quan Jiang. A data-driven method to learn a jump diffusion process from aggregate biological gene expression data. Journal of Theoretical Biology, 532(??):??, January 7, 2022. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519321003428.

Guo:2023:ACD

[GYY^{+}23] Yuchen Guo, Ling Yong, Qingyu Yao, Mengyi Han, Junsheng Xue, Weizhe Jian, and Tianyan Zhou. Application of a count data model to evaluate the anti-metastatic efficacy of QAP14 in 4T1 breast cancer allografts. Journal of Theoretical Biology, 557(??):??, January 21, 2023. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322003149.

Georgescu:2024:GSC

[GZ24] Paul Georgescu and Hong Zhang. Global stability of coexistence equilibria for n-species models of facultative mutualism. Journal of Theoretical Biology, 595(??):??, December 7, 2024. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0022519324002467.

Giuraniuc:2022:MMP

[GZGH22] Claudiu V. Giuraniuc, Shabana Zain, Shahmama Ghafoor, and Stefan Hoppler. A mathematical modelling portrait of Wnt signalling in early vertebrate embryogenesis. Journal of Theoretical Biology, 551–552(??):??, ???? 2022. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251932200234X.

Grubelnik:2020:MED

[GZM^{+}20] Vladimir Grubelnik, Jan Zmazek, Rene Markovic, Marko Gosak, and Marko Marhl. Modelling of energy-driven switch for glucagon and insulin secretion. Journal of Theoretical Biology, 493(??):??, May 21, 2020. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519320300680.

Hernandez-Aristizabal:2024:BSM

[HAGADDM24] David Hernandez-Aristizabal, Diego-Alexander Garzon-Alvarado, Carlos-Alberto Duque-Daza, and Anotida Madzvamuse. A bulk-surface mechanobiochemical modelling approach for single cell migration in two-space dimensions. Journal of Theoretical Biology, 595(??):??, December 7, 2024. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0022519324002510.

Haji:2020:CAT

[Haj20] Amir Hossein Haji. Cells adapt their shapes with their experienced stiffness: a geometrical approach to differentiation. Journal of Theoretical Biology, 486(??):??, February 7, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519319304552.

Hanin:2023:CSM

[Han23] Leonid Hanin. The circulation stage of the metastatic cascade: a mathematical description and its clinical implications. Journal of Theoretical Biology, 572(??):??, September 7, 2023. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519323001790.

Halkola:2022:TMM

[HAP22] Anni S. Halkola, Tero Aittokallio, and Kalle Parvinen. Tumor microenvironment as a metapopulation model: the effects of angiogenesis, emigration and treatment modalities. Journal of Theoretical Biology, 545(??):??, July 21, 2022. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251932200145X.

Hossjer:2021:WTU

[HBG21] Ola Hössjer, Günter Bechly, and Ann Gauger. On the waiting time until coordinated mutations get fixed in regulatory sequences. Journal of Theoretical Biology, 524(??):??, September 7, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519321000795. See cautionary

note [TRKB25].

Heitzman-Breen:2021:MDU

[HBGV^{+}21] Nora Heitzman-Breen, Jacob Golden, Ana Vazquez, Sarah C. Kuchinsky, Nisha K. Duggal, and Stanca M. Ciupe. Modeling the dynamics of Usutu virus infection in birds. Journal of Theoretical Biology, 531(??):??, December 21, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321003155.

Heitzman-Breen:2025:RNV

[HBGW^{+}25] Nora Heitzman-Breen, Jacob Golden, Yuyao Wang, James Weger-Lucarelli, and Stanca M. Ciupe. The role of nutrition in virus dynamics in vector-borne viral infections. Journal of Theoretical Biology, 611(??):??, August 21, 2025. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251932500150X.

Hahn:2020:GPF

[HBK^{+}20] Artur Hahn, Julia Bode, Thomas Krüwel, Thomas Kampf, Lukas R. Buschle, Volker J. F. Sturm, Ke Zhang, Björn Tews, Heinz-Peter Schlemmer, Sabine Heiland, Martin Bendszus, Christian H. Ziener, Michael O. Breckwoldt, and Felix T. Kurz. Gibbs point field model quantifies disorder in microvasculature of U87-glioblastoma. Journal of Theoretical Biology, 494(??):??, June 7, 2020. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519320300850.

Herron-Bedoya:2021:PBM

[HBWLLPA21] Alejandro Herron-Bedoya, Manuela Walteros-Leon, Laura Lema-Perez, and Hernan Alvarez. Phenomenologicalbased model of glucose transport from liver to abdominal subcutaneous adipose tissue. Journal of Theoretical Biology, 530(??):??, December 7, 2021. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519321003027.

Hansen:2024:PDG

[HC24] Matheus Hansen and Fabio A. C. C. Chalub. Population dynamics and games of variable size. Journal of Theoretical Biology, 589(??):??, July 21, 2024. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519324001231.

Hengenius:2021:ONR

[HCC^{+}21] James B. Hengenius, Erin G. Connor, John P. Crimaldi, Nathaniel N. Urban, and G. Bard Ermentrout. Olfactory navigation in the real world: Simple local search strategies for turbulent environments. Journal of Theoretical Biology, 516(??):??, May 7, 2021. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519321000291.

Hernandez:2023:MMW

[HCJT23] Julio A. Hernández, Silvia Chifflet, Cristian Justet, and Alicia Torriglia. A mathematical model of wound healing in bovine corneal endothelium. Journal of Theoretical Biology, 559(??):??, February 21, 2023. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519322003654.

Hillmann:2020:AIH

[HCR20] Andreas Hillmann, Martin Crane, and Heather J. Ruskin. Assessing the impact of HIV treatment interruptions using stochastic cellular Automata. Journal of Theoretical Biology, 502(??):??, October 7, 2020. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519320302319.

Hegarty-Cremer:2021:MCG

[HCSAB21] Solene G. D. Hegarty-Cremer, Matthew J. Simpson, Thomas L. Andersen, and Pascal R. Buenzli. Modelling cell guidance and curvature control in evolving biological tissues. Journal of Theoretical Biology, 520(??):??, July 7, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519321000801.

Holt:2022:CMD

[HD22] Alan G. Holt and Adrian M. Davies. A comparison of mtDNA deletion mutant proliferation mechanisms. Journal of Theoretical Biology, 551–552(??):??, ???? 2022. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322002387.

Holt:2024:LTE

[HD24] Alan G. Holt and Adrian M. Davies. The long term effects of uncoupling interventions as a therapy for dementia in humans. Journal of Theoretical Biology, 587(??):??, June 21, 2024. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519324001061.

Hudnall:2025:WDT

[HD25] Kevin Hudnall and Raissa M. D’Souza. What does the tree of life look like as it grows? Evolution and the multifractality of time. Journal of Theoretical Biology, 607(??):??, June 21, 2025. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519325000876.

Hayakawa:2020:BRC

[HDM20] Tomohiro Hayakawa, Shigeto Dobata, and Fumitoshi Matsuno. Behavioral responses to colony-level properties affect disturbance resistance of red harvester ant colonies. Journal of Theoretical Biology, 492(??):??, May 7, 2020. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519320300424.

Henry:2020:HME

[Hen20] John M. Henry. A hybrid model for the effects of treatment and demography on malaria superinfection. Journal of Theoretical Biology, 491(??):??, April 21, 2020. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519320300497.

Haun:2023:ECR

[HFD23] Asher Haun, Baylor Fain, and Hana M. Dobrovolny. Effect of cellular regeneration and viral transmission mode on viral spread. Journal of Theoretical Biology, 558(??):??, February 7, 2023. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519322003617.

Halder:2022:UNC

[HGCC22] Suvankar Halder, Sumana Ghosh, Joydev Chattopadhyay, and Samrat Chatterjee. Understanding noise in cell signalling in the prospect of drug-targets. Journal of Theoretical Biology, 555(??):??, December 21, 2022. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322002892.

Harris:2025:IAE

[HGD^{+}25] Jeremy D. Harris, Esther Gallmeier, Jonathan Dushoff, Stephen J. Beckett, and Joshua S. Weitz. Infections are not alike: the effects of covariation between individual susceptibility and transmissibility on epidemic dynamics. Journal of Theoretical Biology, 613(??):??, October 7, 2025. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519325001778.

Huber:2023:SBP

[HGF23] Holly A. Huber, Senta K. Georgia, and Stacey D. Finley. Systematic Bayesian posterior analysis guided by Kullback– Leibler divergence facilitates hypothesis formation. Journal of Theoretical Biology, 558(??):??, February 7, 2023. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322003320.

Hu:2025:PRM

[HGZM25] Bing Hu, Yaqi Guo, JinDong Zhao, and Xunfu Ma. Possible regulatory mechanisms of typical and atypical absence seizures through an equivalent projection from the subthalamic nucleus to the cortex: Evidence in a computational model. Journal of Theoretical Biology, 602–603(??):??, April 7, 2025. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519325000256.

Hedh:2020:MPM

[HH20] Linus Hedh and Anders Hedenström. The migration pattern of a monogamous shorebird challenges existing hypotheses explaining the evolution of differential migration. Journal of Theoretical Biology, 487(??):??, February 21, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319304801.

Huang:2021:MCF

[HH21] Yan-San Huang and Fu-Lan Hsu. Mechanism of cracking failure in curved stems due to transverse stress under a bending moment. Journal of Theoretical Biology, 509(??):??, January 21, 2021. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519320303714.

Hayashi:2024:VRO

[HHI24] Rena Hayashi, Akane Hara, and Yoh Iwasa. Viral rebound occurrence immediately after drug discontinuation involving neither drug resistance nor latent reservoir. Journal of Theoretical Biology, 582(??):??, April 7, 2024. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519324000481.

He:2025:SIN

[HHT25] Sha He, Mengqi He, and Sanyi Tang. Statistical inference and neural network training based on stochastic difference model for air pollution and associated disease transmission. Journal of Theoretical Biology, 596(??):??, January 7, 2025. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519324002728.

Hara:2020:ADI

[HI20] Akane Hara and Yoh Iwasa. Autoimmune diseases initiated by pathogen infection: Mathematical modeling. Journal of Theoretical Biology, 498(??):??, August 7, 2020. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251932030151X.

Hishida:2021:SDG

[HI21] Shintaro Hishida and Yoh Iwasa. Spatial distribution of gut microbes along the intestinal duct. Journal of Theoretical Biology, 523(??):??, August 21, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321001478.

Hadjicharalambous:2022:CAA

[HIA^{+}22] Myrianthi Hadjicharalambous, Eleftherios Ioannou, Nicolas Aristokleous, Kristaq Gazeli, Charalambos Anastassiou, and Vasileios Vavourakis. Combined anti-angiogenic and cytotoxic treatment of a solid tumour: in silico investigation of a xenograft animal model’s digital twin. Journal of Theoretical Biology, 553(??):??, November 21, 2022. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322002399.

Hayashi:2022:ESE

[HII22] Rena Hayashi, Shingo Iwami, and Yoh Iwasa. Escaping stochastic extinction of mutant virus: Temporal pattern of emergence of drug resistance within a host. Journal of Theoretical Biology, 537(??):??, March 21, 2022. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322000273.

Hill:2023:MEI

[Hil23] Edward M. Hill. Modelling the epidemiological implications for SARS-CoV-2 of Christmas household bubbles in England. Journal of Theoretical Biology, 557(??):??, January 21, 2023. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519322003228.

Hajij:2020:GBA

[HJKS20] Mustafa Hajij, Natasa Jonoska, Denys Kukushkin, and Masahico Saito. Graph based analysis for gene segment organization in a scrambled genome. Journal of Theoretical Biology, 494(??):??, June 7, 2020. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519320300709.

Hamley:2021:PEA

[HK21] Jonathan I. D. Hamley and Jacob C. Koella. Parasite evolution in an age-structured population. Journal of Theoretical Biology, 527(??):??, October 21, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321001545.

Harkess:2025:HTS

[HKBK25] Miranda Harkess, Sudha Kumari, Trilochan Bagarti, and Niraj Kumar. HIV transactivation: Stochastic modeling for studying the effects of BET inhibitors on the modulation of P-TEFb levels. Journal of Theoretical Biology, 599(??):??, February 21, 2025. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519324002960.

Hawi:2025:ONP

[HKL25] Georgio Hawi, Peter S. Kim, and Peter P. Lee. Optimisation of neoadjuvant pembrolizumab therapy for locally advanced MSI-H/dMMR colorectal cancer using data-driven delay integro-differential equations. Journal of Theoretical Biology, 613(??):??, October 7, 2025. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519325001973.

Herath:2020:CIT

[HL20] Samantha Herath and Daniel Lobo. Cross-inhibition of Turing patterns explains the self-organized regulatory mechanism of planarian fission. Journal of Theoretical Biology, 485(??):??, January 21, 2020. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519319304114.

Holden:2021:PPD

[HL21] Matthew H. Holden and Jakeb Lockyer. Poacher-population dynamics when legal trade of naturally deceased organisms funds anti-poaching enforcement. Journal of Theoretical Biology, 517(??):??, May 21, 2021. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519321000400.

Hunter:2024:FRN

[HL24] Iain Hunter and Raz Leib. A framework for relating natural movement to length and quality of life in human and nonhuman animals. Journal of Theoretical Biology, 576(??):??, January 7, 2024. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519323002461.

Huang:2025:NIE

[HLS^{+}25] Ting Huang, Hengmin Lv, Yiting Shu, Jian Luo, Linxuan Yu, Bing Chen, Xin Sun, Xilin Hou, Xiong You, and Tonghua Zhang. Noise-induced entrainment of the circadian clock by thermoperiods in tomato: a computational approach. Journal of Theoretical Biology, 598(??):??, February 7, 2025. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519324002844.

Hu:2024:ICR

[HLW24] Qing Hu, Min Luo, and Ruiqi Wang. Identifying critical regulatory interactions in cell fate decision and transition by systematic perturbation analysis. Journal of Theoretical Biology, 577(??):??, January 21, 2024. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519323002709.

Hurford:2023:PMR

[HMLO^{+}23] Amy Hurford, Maria M. Martignoni, J. Concepción Loredo-Osti, Francis Anokye, Julien Arino, Bilal Saleh Husain, Brian Gaas, and James Watmough. Pandemic modelling for regions implementing an elimination strategy. Journal of Theoretical Biology, 561(??):??, March 21, 2023. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322003691.

Hasegawa:2021:OPS

[HMS21] Takuya M. Hasegawa, Takashi T. Makino, and Satoki Sakai. Optimal pollen stickiness to pollinators for maximizing paternal fitness: Increased number of recipient flowers or increased pollen deposition on recipient flowers? Journal of Theoretical Biology, 524(??):??, September 7, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321001533.

Hoetzlein:2024:FMO

[Hoe24] Rama Carl Hoetzlein. Flock2: a model for orientation-based social flocking. Journal of Theoretical Biology, 593(??):??, October 7, 2024. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519324001644.

Hong:2020:MGN

[Hon20] Ze Hong. Modelling the on-going natural selection of educational attainment in contemporary societies. Journal of Theoretical Biology, 493(??):??, May 21, 2020. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519320300655.

Hall:2020:DSA

[HPD20] Cameron L. Hall, Mason A. Porter, and Marian S. Dawkins. Dominance, sharing, and assessment in an iterated Hawk– Dove game. Journal of Theoretical Biology, 493(??):??, May 21, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519319304709.

Hay:2024:MMW

[HPG^{+}24] Quintessa Hay, Eunice Pak, Luke Gardner, Anna Shaw, Liza M. Roger, Nastassja A. Lewinski, Rebecca A. Segal, and Angela M. Reynolds. A mathematical model for wound healing in the reef-building coral Pocillopora damicornis. Journal of Theoretical Biology, 593(??):??, October 7, 2024. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519324001814.

Hinch:2025:ECA

[HPGF25] Robert Hinch, Jasmina Panovska-Griffiths, and Christophe Fraser. Estimating the contribution of age-structure to the COVID-19 epidemic in England. Journal of Theoretical Biology, 611(??):??, August 21, 2025. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519325001432.

Halkola:2020:MKC

[HPK^{+}20] Anni S. Halkola, Kalle Parvinen, Henna Kasanen, Satu Mustjoki, and Tero Aittokallio. Modelling of killer T-cell and cancer cell subpopulation dynamics under immuno- and chemotherapies. Journal of Theoretical Biology, 488(??):??, March 7, 2020. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519319305053.

Halimubieke:2021:HDB

[HPSA21] Naerhulan Halimubieke, Alistair Pirrie, Tamás Székely, and Ben Ashby. How do biases in sex ratio and disease characteristics affect the spread of sexually transmitted infections? Journal of Theoretical Biology, 527(??):??, October 21, 2021. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519321002514. See corri-

gendum [HPSA23].

Halimubieke:2023:CHD

[HPSA23] Naerhulan Halimubieke, Alistair Pirrie, Tamás Székely, and Ben Ashby. Corrigendum to “How do biases in sex ratio and disease characteristics affect the spread of sexually transmitted infections?” [J. Theor. Biol. 527 (2021) 110832]. Journal of Theoretical Biology, 571(??):??, August 21, 2023. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251932300142X. See [HPSA21].

Helps:2020:DOI

[HPWvdB20] Joseph Christopher Helps, Neil D. Paveley, Sacha White, and Frank van den Bosch. Determinants of optimal insecticide resistance management strategies. Journal of Theoretical Biology, 503(??):??, October 21, 2020. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519320302381.

Henningsen:2024:URD

[HRD^{+}24] Maj Beldring Henningsen, Mossa Merhi Reimert, Matt Denwood, Maya Katrin Gussmann, Carsten Thure Kirkeby, and Søren Saxmose Nielsen. Using registry data to identify individual dairy cows with abnormal patterns in routinely recorded somatic cell counts. Journal of Theoretical Biology, 579(??):??, February 21, 2024. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519323003156.

Hassanpoor:2020:IER

[HS20] Hossein Hassanpoor and Maryam Saidi. An investigation into the effective role of astrocyte in the hippocampus pattern separation process: a computational modeling study. Journal of Theoretical Biology, 487(??):??, February 21, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319304837.

Harkos:2024:NSE

[HS24] Constantinos Harkos and Triantafyllos Stylianopoulos. Investigating the synergistic effects of immunotherapy and normalization treatment in modulating tumor microenvironment and enhancing treatment efficacy. Journal of Theoretical Biology, 583(??):??, April 21, 2024. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519324000493.

Hilker:2020:SSI

[HSAH20] F. M. Hilker, T. A. Sun, L. J. S. Allen, and F. M. Hamelin. Separate seasons of infection and reproduction can lead to multi-year population cycles. Journal of Theoretical Biology, 489(??):??, March 21, 2020. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251932030014X.

Huang:2024:DCT

[HSL24] Rongsheng Huang, Qiaojun Situ, and Jinzhi Lei. Dynamics of cell-type transition mediated by epigenetic modifications. Journal of Theoretical Biology, 577(??):??, January 21, 2024. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519323002618.

Holmes:2022:ASS

[HTD22] Alex Holmes, Mike Tildesley, and Louise Dyson. Approximating steady state distributions for household structured epidemic models. Journal of Theoretical Biology, 534(??):??, February 7, 2022. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519321003945.

Hardwick:2021:SSH

[HVC^{+}21] Robert J. Hardwick, Carolin Vegvari, Benjamin Collyer, James E. Truscott, and Roy M. Anderson. Spatial scales in human movement between reservoirs of infection. Journal of Theoretical Biology, 524(??):??, September 7, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251932100148X.

Hardwick:2020:BST

[HVTA20] Robert J. Hardwick, Carolin Vegvari, James E. Truscott, and Roy M. Anderson. The ‘breakpoint’ of soil-transmitted helminths with infected human migration. Journal of Theoretical Biology, 486(??):??, February 7, 2020. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931930445X.

Huo:2020:SSN

[HXK^{+}20] Yanhao Huo, Lihui Xin, Chuanze Kang, Minghui Wang, Qin Ma, and Bin Yu. SGL-SVM: a novel method for tumor classification via support vector machine with sparse group Lasso. Journal of Theoretical Biology, 486(??):??, February 7, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519319304679.

Hu:2021:WBW

[HXS21] Di Hu, Cai-Hua Xiong, and Ronglei Sun. Working out the bipedal walking expenditure of energy based on foot morphology of different hominid genera: Implications for foot evolution. Journal of Theoretical Biology, 519(??):??, June 21, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519321000680.

Hu:2022:BHB

[HXZ^{+}22] Bing Hu, Minbo Xu, Luyao Zhu, Jiahui Lin, Zhizhi Wang, Dingjiang Wang, and Dongmei Zhang. A bidirectional Hopf bifurcation analysis of Parkinson’s oscillation in a simplified basal ganglia model. Journal of Theoretical Biology, 536(??): ??, March 7, 2022. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519321003994.

Hayashi:2022:PFE

[HYM^{+}22] Tatsuya Hayashi, Fumitaka Yura, Jun Mada, Hiroki Kurihara, and Tetsuji Tokihiro. Pattern formation of elliptic particles by two-body interactions: a model for dynamics of endothelial cells in angiogenesis. Journal of Theoretical Biology, 555(??):??, December 21, 2022. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322002910.

He:2021:HRK

[HZM^{+}21] Qiao-Qiao He, Xiu-Deng Zheng, Ruth Mace, Yi Tao, and Ting Ji. Hamilton’s rule and kin competition in a finite kin population. Journal of Theoretical Biology, 529(??):??, November 21, 2021. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519321002812.

He:2021:HHB

[HZS^{+}21] Zike He, Wei Zhang, Yu Sun, Chuchu Li, Jianing Wu, and Zhigang Wu. How honey bees dip nectar: Dynamic spacing of tongue hairs facilitates to collect nectar of various viscosities. Journal of Theoretical Biology, 512(??):??, March 7, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519320303933.

Itsko:2021:NTC

[IB21] Mark Itsko and Yuval Ben Abu. Novel type of coagulation equation and its application to DNA repeat expansion process. Journal of Theoretical Biology, 511(??):??, February 21, 2021. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519320304100.

Ispolatov:2023:EEP

[IDD23] Yaroslav Ispolatov, Carlos Doebeli, and Michael Doebeli. On the evolutionary emergence of predation. Journal of Theoretical Biology, 572(??):??, September 7, 2023. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519323001753.

Imizcoz:2024:ORA

[IDMG24] J. Innerarity Imizcoz, W. Djema, F. Mairet, and J.-L. Gouzé. Optimal resource allocation in micro-organisms under periodic nutrient fluctuations. Journal of Theoretical Biology, 595(??):??, December 7, 2024. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0022519324002388.

Imirzian:2021:ABM

[IH21] Natalie Imirzian and David P. Hughes. An agent-based model shows zombie ants exhibit search behavior. Journal of Theoretical Biology, 526(??):??, October 7, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321002083.

Iwasa:2023:WIE

[IH23] Yoh Iwasa and Rena Hayashi. Waves of infection emerging from coupled social and epidemiological dynamics. Journal of Theoretical Biology, 558(??):??, February 7, 2023. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322003575.

Islier:2021:AIS

[IHG21] Zeynep Gökçe Islier, Wolfgang Hörmann, and Refik Güllü. Assessing intervention strategies for non-homogeneous populations using a closed form formula for R0. Journal of Theoretical Biology, 511(??):??, February 21, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519320304161.

Irons:2020:SBV

[IHO^{+}20] Linda Irons, Huang Huang, Markus R. Owen, Reuben D. O’Dea, Gerald A. Meininger, and Bindi S. Brook. Switching behaviour in vascular smooth muscle cell-matrix adhesion during oscillatory loading. Journal of Theoretical Biology, 502(??):??, October 7, 2020. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519320302423.

Iwasa:2021:VVH

[IHO21] Yoh Iwasa, Akane Hara, and Shihomi Ozone. Virulence of a virus: How it depends on growth rate, effectors, memory cells, and immune escape. Journal of Theoretical Biology, 530(??):??, December 7, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321002940.

Iwasa:2025:OSS

[IHS25] Yoh Iwasa, Rena Hayashi, and Akiko Satake. Optimal seasonal schedule for producing biogenic volatile organic compounds for tree defense. Journal of Theoretical Biology, 596(??):??, January 7, 2025. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519324002716.

Izquierdo:2025:PNS

[IIH25] Segismundo S. Izquierdo, Luis R. Izquierdo, and Christoph Hauert. Positive and negative selective assortment. Journal of Theoretical Biology, 608(??):??, July 7, 2025. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519325000955.

Ikegawa:2021:SMF

[IIHH21] Yusuke Ikegawa, Koichi Ito, Chihiro Himuro, and Atsushi Honma. Sterile males and females can synergistically suppress wild pests targeted by sterile insect technique. Journal of Theoretical Biology, 530(??):??, December 7, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321002976.

Ibarra-Junquera:2022:ETM

[IJROTE22] Vrani Ibarra-Junquera, Diego Radillo-Ochoa, and César A. Terrero-Escalante. Evolutionary timeline of a modeled cell. Journal of Theoretical Biology, 551–552(??):??, ???? 2022. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322002302.

Ichinose:2021:DRS

[IKYW21] Natsuhiro Ichinose, Takeshi Kawashima, Tetsushi Yada, and Hiroshi Wada. Dynamical robustness and its structural dependence in biological networks. Journal of Theoretical Biology, 526(??):??, October 7, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519321002277.

Ippolito:2020:SST

[ILGI20] Stephen Ippolito, Jose Laborde, Tim Gottwald, and Michael S. Irey. Studying the spatial temporal spread of the citrus tristeza virus through ODEs and Bernoulli trials. Journal of Theoretical Biology, 497(??):??, July 21, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251932030134X.

Ionel:2024:PPA

[IMM24] Ciuperca S. Ionel, Moncef Mahjoub, and Tine Léon Matar. A β-protein polymerization in Alzheimer disease: Optimal control for nucleation parameter estimation. Journal of Theoretical Biology, 589(??):??, July 21, 2024. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251932400095X.

Iosilevskii:2021:FOS

[Ios21] Gil Iosilevskii. A foraging ocean sunfish and the ‘nearest neighbor’ problem. Journal of Theoretical Biology, 520(??): ??, July 7, 2021. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519321001016.

Ito:2020:EBD

[IS20] Hiroshi C. Ito and Akira Sasaki. Evolutionary branching in distorted trait spaces. Journal of Theoretical Biology, 489(??):??, March 21, 2020. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519320300084.

Idumah:2023:SDM

[ISC23] Gideon Idumah, Erkki Somersalo, and Daniela Calvetti. A spatially distributed model of brain metabolism highlights the role of diffusion in brain energy metabolism. Journal of Theoretical Biology, 572(??):??, September 7, 2023. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519323001649.

Ishii:2023:TAV

[Ish23] Yuichiro Ishii. Topological analysis of vertebrate phylogenetic trees utilizing Horton’s first law. Journal of Theoretical Biology, 562(??):??, April 7, 2023. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519323000152.

Iwasa:2022:EMN

[IY22] Yoh Iwasa and Sachi Yamaguchi. Evolution of male nuptial gift and female remating: a quantitative genetic model. Journal of Theoretical Biology, 533(??):??, January 21, 2022.

CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321003581.

Islam:2025:FAP

[IYL^{+}25] Rafiqul Islam, Junyuan Yang, Miao Li, Sri Prakash Mokshagundam, Lu Cai, and Jiaxu Li. Fatty acids and physical activity are critical for β-cell mass and insulin sensitivity: Pathways to T2D and prevention. Journal of Theoretical Biology, 611(??):??, August 21, 2025. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519325001547.

Iwasa:2022:EGL

[IYY22] Yoh Iwasa, Yoichi Yusa, and Sachi Yamaguchi. Evolutionary game of life-cycle types in marine benthic invertebrates: Feeding larvae versus nonfeeding larvae versus direct development. Journal of Theoretical Biology, 537(??):??, March 21, 2022. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519322000170.

Jiang:2023:IVM

[JA23] Jianfei Jiang and Christof M. Aegerter. An integrated vertex model of the mesoderm invagination during the embryonic development of Drosophila. Journal of Theoretical Biology, 572(??):??, September 7, 2023. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519323001789.

Jnawali:2022:SIP

[JAB22] Kamal Jnawali, Madhur Anand, and Chris T. Bauch. Stochasticity-induced persistence in coupled social-ecological systems. Journal of Theoretical Biology, 542(??):??, June 7, 2022. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519322000868.

Jackson:2025:KTC

[Jac25] Tristen M. Jackson. Kinetics, thresholds, and a comparison of mechanisms underlying systemic infection by Listeria monocytogenes. Journal of Theoretical Biology, 599(??):??, February 21, 2025. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519324002947.

James:2021:DTM

[Jam21] William H. James. Does testosterone in males facilitate evolution by natural selection? Journal of Theoretical Biology, 529(??):??, November 21, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321002836.

Ji:2025:HAS

[JAW25] Juping Ji, Shohel Ahmed, and Hao Wang. A hybrid approach to study and forecast climate-sensitive norovirus infections in the USA. Journal of Theoretical Biology, 598(??):??, February 7, 2025. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519324002923.

Jiao:2020:PPR

[JCC20] Yuhang Jiao, Tong Chen, and Qiao Chen. Probabilistic punishment and reward under rule of trust-based decisionmaking in continuous public goods game. Journal of Theoretical Biology, 486(??):??, February 7, 2020. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319304722.

Jordan-Cooley:2011:BDE

[JCLS^{+}11] William C. Jordan-Cooley, Romuald N. Lipcius, Leah B. Shaw, Jian Shen, and Junping Shi. Bistability in a differential equation model of oyster reef height and sediment accumulation. Journal of Theoretical Biology, 289(??):1–11, November 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519311004085. See correction

[JCLS^{+}21].

Jordan-Cooley:2021:CBD

[JCLS^{+}21] William C. Jordan-Cooley, Romuald N. Lipcius, Leah B. Shaw, Jian Shen, and Junping Shi. Corrigendum to “Bistability in a differential equation model of oyster reef height and sediment accumulation” [J. Theor. Biol. 289 (2011) 1– 11]. Journal of Theoretical Biology, 525(??):??, September 21, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519321001570. See [JCLS^{+}11].

Jessie:2021:RSD

[JD21] Benjamin Jessie and Hana M. Dobrovolny. The role of syncytia during viral infections. Journal of Theoretical Biology, 525(??):??, September 21, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321001715.

Jenner:2020:EOV

[JFCK20] Adrianne L. Jenner, Federico Frascoli, Adelle C. F. Coster, and Peter S. Kim. Enhancing oncolytic virotherapy: Observations from a Voronoi cell-based model. Journal of Theoretical Biology, 485(??):??, January 21, 2020. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319304217.

Juayerk-Herrera:2021:DMS

[JHFMP^{+}21] Kenia L. Juayerk-Herrera, Gerardo J. Félix-Martínez, Arturo Picones, José L. Del-Río-Correa, and J. Rafael Godínez-Fernández. Deterministic modeling of singlechannel and whole-cell currents. Journal of Theoretical Biology, 508(??):??, January 7, 2021. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519320303143.

Joshi:2020:PAF

[JJR20] Shripad Joshi, Kung-Ming Jan, and David Rumschitzki. Pre-atherosclerotic flow and oncotically active solute transport across the arterial endothelium. Journal of Theoretical Biology, 499(??):??, August 21, 2020. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519320301302.

Jung:2023:MSE

[JKH^{+}23] Seungpil Jung, Jong-Hoon Kim, Seung-Sik Hwang, Junyoung Choi, and Woojoo Lee. Modified susceptible-exposed-infectious-recovered model for assessing the effectiveness of non-pharmaceutical interventions during the COVID-19 pandemic in Seoul. Journal of Theoretical Biology, 557(??):??, January 21, 2023. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519322003204.

Jonas:2023:MAI

[JKSD23] Daniel Jonas, Michael Kirby, Alan R. Schenkel, and Gerhard Dangelmayr. Modeling of adaptive immunity uncovers disease tolerance mechanisms. Journal of Theoretical Biology, 568(??):??, July 7, 2023. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519323000942.

Joslyn:2022:VHM

[JLK22] Louis R. Joslyn, Jennifer J. Linderman, and Denise E. Kirschner. A virtual host model of Mycobacterium tuberculosis infection identifies early immune events as predictive of infection outcomes. Journal of Theoretical Biology, 539(??):??, April 21, 2022. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519322000406.

Jacobs:2020:RBP

[JMD20] Bas Jacobs, Jaap Molenaar, and Eva E. Deinum. Robust banded protoxylem pattern formation through microtubulebased directional ROP diffusion restriction. Journal of Theoretical Biology, 502(??):??, October 7, 2020. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251932030206X.

Jegatheesan:2024:MMC

[JME24] Thulasi Jegatheesan, Arun S. Moorthy, and Hermann J. Eberl. A mathematical model of competition between fiber and mucin degraders in the gut provides a possible explanation for mucus thinning. Journal of Theoretical Biology, 587(??):??, June 21, 2024. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251932400105X.

Jing:2023:VHP

[JMWX23] Shuanglin Jing, Russell Milne, Hao Wang, and Ling Xue. Vaccine hesitancy promotes emergence of new SARS-CoV-2 variants. Journal of Theoretical Biology, 570(??):??, August 7, 2023. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519323001182.

Jacquet:2023:RMC

[JS23] Pierre Jacquet and Angélique Stéphanou. A reduced model of cell metabolism to revisit the glycolysis-OXPHOS relationship in the deregulated tumor microenvironment. Journal of Theoretical Biology, 562(??):??, April 7, 2023. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519323000309.

Jones:2021:AIF

[JSCW21] Eric Jones, Jiming Sheng, Jean Carlson, and Shenshen Wang. Aging-induced fragility of the immune system. Journal of Theoretical Biology, 510(??):??, February 7, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519320303283.

Jones:2025:BTS

[JTB25] Michael L. Jones, Ronald E. Thresher, and Nicholas J. Bax. Biased tertiary sex ratios enhance the efficacy of sex-ratio distorting genetic techniques to control invasive species. Journal of Theoretical Biology, 608(??):??, July 7, 2025. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519325001031.

Johnson:2021:DMO

[JWKA21] Joseph D. Johnson, Nathan L. White, Alain Kangabire, and Daniel M. Abrams. A dynamical model for the origin of anisogamy. Journal of Theoretical Biology, 521(??):??, July 21, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519321000916.

Kania:2021:HIT

[Kan21] Adrian Kania. Harnessing the information theory and chaos game representation for pattern searching among essential and non-essential genes in bacteria. Journal of Theoretical Biology, 531(??):??, December 21, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321003362.

Kaushik:2023:EBS

[Kau23] Sachin Kaushik. Effect of beneficial sweeps and background selection on genetic diversity in changing environments. Journal of Theoretical Biology, 562(??):??, April 7, 2023. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519323000279.

Kushchenko:2020:EHT

[KB20] Yulia K. Kushchenko and Aleksey V. Belyaev. Effects of hydrophobicity, tethering and size on flow-induced activation of von Willebrand factor multimers. Journal of Theoretical Biology, 485(??):??, January 21, 2020. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319304199.

King:2022:CRD

[KB22] Clarence C. King and S. P. Beckman. Coupled reactiondiffusion transport into a core-shell geometry. Journal of Theoretical Biology, 546(??):??, August 7, 2022. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322001369.

Korosec:2023:MCS

[KBD^{+}23] Chapin S. Korosec, Matthew I. Betti, David W. Dick, Hsu Kiang Ooi, Iain R. Moyles, Lindi M. Wahl, and Jane M. Heffernan. Multiple cohort study of hospitalized SARS-CoV-2 in-host infection dynamics: Parameter estimates, identifiability, sensitivity and the eclipse phase profile. Journal of Theoretical Biology, 564(??):??, May 7, 2023. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519323000450.

King:2019:MRD

[KBS^{+}19] Clarence C. King, Amelia Ann Brown, Irmak Sargin, K. M. Bratlie, and S. P. Beckman. Modeling of reaction–diffusion transport into a core-shell geometry. Journal of Theoretical Biology, 460(??):204–208, January 7, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318304624. See [KBS^{+}20].

King:2020:CMR

[KBS^{+}20] Clarence C. King, Amelia Ann Brown, Irmak Sargin, K. M. Bratlie, and S. P. Beckman. Corrigendum to: Modeling of reaction–diffusion transport into a core-shell geometry. Journal of Theoretical Biology, 507(??):??, December 21, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519320302940. See [KBS^{+}19].

KhudaBukhsh:2023:PCC

[KBW^{+}23] Wasiur R. KhudaBukhsh, Caleb Deen Bastian, Matthew Wascher, Colin Klaus, Saumya Yashmohini Sahai, Mark H. Weir, Eben Kenah, Elisabeth Root, Joseph H. Tien, and Grzegorz A. Rempała. Projecting COVID-19 cases and hospital burden in Ohio. Journal of Theoretical Biology, 561(??):??, March 21, 2023. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519322003952.

Krivan:2022:AHD

[KC22] Vlastimil Krivan and Ross Cressman. The asymmetric Hawk– Dove game with costs measured as time lost. Journal of Theoretical Biology, 547(??):??, August 21, 2022. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322001606.

Krivan:2024:IFD

[KC24] Vlastimil Krivan and Ross Cressman. The Ideal Free Distribution with travel costs. Journal of Theoretical Biology, 579(??):??, February 21, 2024. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519323003144.

Kristensen:2025:MSG

[KCO25] Nadiah P. Kristensen, Ryan A. Chisholm, and Hisashi Ohtsuki. Many-strategy games in groups with relatives and the evolution of coordinated cooperation. Journal of Theoretical Biology, 605(??):??, May 21, 2025. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519325000554.

Kirschner:2020:D

[KCS20] Denise Kirschner, Mark Chaplain, and Akira Sasaki. Disclaimer. Journal of Theoretical Biology, 506(??):??, December 7, 2020. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519320303118.

Kettle:2025:PBL

[KCW^{+}25] Helen Kettle, Duncan J. Coston, Sacha White, Tom W. Pope, Joe M. Roberts, and David Ewing. A processbased life cycle model of the cabbage stem flea beetle on winter oilseed rape: Effects of temperature. Journal of Theoretical Biology, 612(??):??, September 7, 2025. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519325001511.

Keeling:2023:PRI

[Kee23] Matt J. Keeling. Patterns of reported infection and reinfection of SARS-CoV-2 in England. Journal of Theoretical Biology, 556(??):??, January 7, 2023. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519322002909.

Koyama:2020:IEE

[KF20] Hiroshi Koyama and Toshihiko Fujimori. Isotropic expansion of external environment induces tissue elongation and collective cell alignment. Journal of Theoretical Biology, 496(??): ??, July 7, 2020. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251932030103X.

Kabir:2022:NBA

[KG22] M. Humayun Kabir and M. Osman Gani. Numerical bifurcation analysis and pattern formation in a minimal reaction–diffusion model for vegetation. Journal of Theoretical Biology, 536(??):??, March 7, 2022. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519321004173.

Karst:2023:MTT

[KG23] Nathaniel J. Karst and John B. Geddes. Modeling transit time distributions in microvascular networks. Journal of Theoretical Biology, 572(??):??, September 7, 2023. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519323001819.

Kahramanogullari:2022:SDC

[KGP^{+}22] Ozan Kahramanogullari, Bruna Giordano, Jonathan Perrin, Daniel Vielzeuf, and Lorenzo Bramanti. Stochastic diffusion characterises early colony formation in Mediterranean coral Corallium rubrum. Journal of Theoretical Biology, 553(??):??, November 21, 2022. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519322002405.

Kazarnikov:2020:SAP

[KH20] Alexey Kazarnikov and Heikki Haario. Statistical approach for parameter identification by Turing patterns. Journal of Theoretical Biology, 501(??):??, September 21, 2020. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519320301740.

Kushal:2024:EMS

[KH24] Appilineni Kushal and Alan Hastings. Effect of migrations on synchrony in host-parasitoid system. Journal of Theoretical Biology, 590(??):??, August 7, 2024. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251932400136X.

Koyama:2021:AGR

[KHAK21] Kento Koyama, Satoko Hiura, Hiroki Abe, and Shige Koseki. Application of growth rate from kinetic model to calculate stochastic growth of a bacteria population at low contamination level. Journal of Theoretical Biology, 525(??):??, September 21, 2021. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519321001806.

Kaye:2022:DCM

[KHB^{+}22] A. R. Kaye, W. S. Hart, J. Bromiley, S. Iwami, and R. N. Thompson. A direct comparison of methods for assessing the threat from emerging infectious diseases in seasonally varying environments. Journal of Theoretical Biology, 548(??):??, September 7, 2022. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251932200193X.

Kaitala:2020:CEI

[KHBS20] Veijo Kaitala, Teppo Hiltunen, Lutz Becks, and Thomas Scheuerl. Co-evolution as an important component explaining microbial predator–prey interaction. Journal of Theoretical Biology, 486(??):??, February 7, 2020. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319304643.

Kotsubo:2025:ADF

[KHHI25] Yuna Kotsubo, Akane Hara, Rena Hayashi, and Yoh Iwasa. Age-dependence of food allergy due to decreased supply of naïve T cells. Journal of Theoretical Biology, 602–603(??): ??, April 7, 2025. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519325000268.

Kim:2021:RAT

[KHST21] Bo-Moon Kim, Junnosuke Horita, Jun-Ichirou Suzuki, and Yuuya Tachiki. Resource allocation in tragedy of the commons game in plants for belowground competition. Journal of Theoretical Biology, 529(??):??, November 21, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321002770.

Knight:2022:MMC

[KHW^{+}22] Martin A. Knight, Michael R. Hutchings, Piran C. L. White, Ross S. Davidson, and Glenn Marion. A mechanistic model captures livestock trading, disease dynamics, and compensatory behaviour in response to control measures. Journal of Theoretical Biology, 539(??):??, April 21, 2022. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322000571.

Kuwabara:2020:TOS

[KI20] Tomohiko Kuwabara and Kensuke Igarashi. Thermotogales origin scenario of eukaryogenesis. Journal of Theoretical Biology, 492(??):??, May 7, 2020. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519320300473.

Kanayama:2021:WDS

[KI21] Yuki Kanayama and Yoh Iwasa. Why did sauropod dinosaurs grow so big? — A possible answer from the life history theory. Journal of Theoretical Biology, 508(??):??, January 7, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519320303404.

Kadelka:2022:EHA

[KIM^{+}22] Claus Kadelka, Md Rafiul Islam, Audrey McCombs, Jake Alston, and Noah Morton. Ethnic homophily affects vaccine prioritization strategies. Journal of Theoretical Biology, 555(??):??, December 21, 2022. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519322002867.

Koyama:2025:CPB

[KIO^{+}25] Hiroshi Koyama, Atsushi M. Ito, Hisashi Okumura, Tetsuhisa Otani, Kazuyuki Nakamura, and Toshihiko Fujimori. Cell position-based evaluation of mechanical features of cells in multicellular systems. Journal of Theoretical Biology, 604(??): ??, May 7, 2025. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519325000360.

Kisdi:2024:OVS

[Kis24] Eva Kisdi. Optimal vaccination strategies for imperfect vaccines and variable host susceptibility. Journal of Theoretical Biology, 594(??):??, November 7, 2024. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519324001838.

Kopfer:2020:MMR

[KJM20] Kai H. Kopfer, Willi Jäger, and Franziska Matthäus. A mechanochemical model for rho GTPase mediated cell polarization. Journal of Theoretical Biology, 504(??):??, November 7, 2020. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519320302411.

Kauffman:2020:TCE

[KJV20] Stuart A. Kauffman, Dávid Péter Jelenfi, and Gábor Vattay. Theory of chemical evolution of molecule compositions in the universe, in the Miller–Urey experiment and the mass distribution of interstellar and intergalactic molecules. Journal of Theoretical Biology, 486(??):??, February 7, 2020. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319304667.

Khosravi:2022:NRT

[KK22a] Ali Khosravi and Gholamreza Karimi. A new real-time analog circuit of Ca^{2+} Li–Rinzel astrocyte model based on analytical method. Journal of Theoretical Biology, 547(??):??, August 21, 2022. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251932200162X.

Kuznetsov:2022:BUU

[KK22b] Ivan A. Kuznetsov and Andrey V. Kuznetsov. Bidirectional, unlike unidirectional transport, allows transporting axonal cargos against their concentration gradient. Journal of Theoretical Biology, 546(??):??, August 7, 2022. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251932200159X.

Kuznetsov:2023:OAR

[KK23] Maxim Kuznetsov and Andrey Kolobov. Optimization of antitumor radiotherapy fractionation via mathematical modeling with account of 4 R’s of radiobiology. Journal of Theoretical Biology, 558(??):??, February 7, 2023. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322003629.

Kurusu:2020:QAE

[KKK^{+}20] Tatsuya Kurusu, Kwang Su Kim, Yoshiki Koizumi, Shinji Nakaoka, Keisuke Ejima, Naoko Misawa, Yoshio Koyanagi, Kei Sato, and Shingo Iwami. Quantifying the antiviral effect of APOBEC3 on HIV-1 infection in humanized mouse model. Journal of Theoretical Biology, 498(??):??, August 7, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519320301508.

Kumbhari:2020:OAC

[KKL20] Adarsh Kumbhari, Peter S. Kim, and Peter P. Lee. Optimisation of anti-cancer peptide vaccines to preferentially elicit high-avidity T cells. Journal of Theoretical Biology, 486(??):??, February 7, 2020. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519319304369.

Khan:2023:MES

[KKS23] Raheel Khan, D. Kulasiri, and S. Samarasinghe. A multifarious exploration of synaptic tagging and capture hypothesis in synaptic plasticity: Development of an integrated mathematical model and computational experiments. Journal of Theoretical Biology, 556(??):??, January 7, 2023. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322003174.

Kayser:2021:SDE

[KL21a] Kirk Kayser and Adam Lampert. Survivor’s dilemma: the evolution of cooperation in volatile environments. Journal of Theoretical Biology, 516(??):??, May 7, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321000254.

Kroumi:2021:EVP

[KL21b] Dhaker Kroumi and Sabin Lessard. The effect of variability in payoffs on average abundance in two-player linear games under symmetric mutation. Journal of Theoretical Biology, 513(??):??, March 21, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519320304240.

Kubatko:2024:IST

[KLC24] Laura Kubatko, Alexander Leonard, and Julia Chifman. Identifiability of speciation times under the multispecies coalescent. Journal of Theoretical Biology, 595(??):??, December 7, 2024. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0022519324002121.

Kinnunen:2022:CMI

[KLLL22] Patrick C. Kinnunen, Gary D. Luker, Kathryn E. Luker, and Jennifer J. Linderman. Computational modeling implicates protein scaffolding in p38 regulation of Akt. Journal of Theoretical Biology, 555(??):??, December 21, 2022. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322002855.

Krieg:2020:TPM

[KM20] Michael Krieg and Kamran Mohseni. Transient pressure modeling in jetting animals. Journal of Theoretical Biology, 494(??):??, June 7, 2020. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519320300928.

Kribs:2022:CBO

[KM22] Christopher M. Kribs and Crystal Mackey. Competition between obligate and facultative scavengers and infection: vulture-jackal-anthrax dynamics in etosha national park. Journal of Theoretical Biology, 537(??):??, March 21, 2022. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251932100401X.

Kye:2021:SBF

[KMA^{+}21] Geunho Kye, Jonathan Machta, Karen C. Abbott, Alan Hastings, William Huffmyer, Fang Ji, Andrew M. Liebhold, and Julie C. Blackwood. Sharp boundary formation and invasion between spatially adjacent periodical cicada broods. Journal of Theoretical Biology, 515(??):??, April 21, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321000229.

Kaji:2023:EPC

[KMI23] Hotaka Kaji, Fumito Mori, and Hiroshi Ito. Enhanced precision of circadian rhythm by output system. Journal of Theoretical Biology, 574(??):??, October 7, 2023. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519323002187.

Khudiakova:2021:TLL

[KNK21] Ksenia A. Khudiakova, Tatiana Yu. Neretina, and Alexey S. Kondrashov. Two linked loci under mutation-selection balance and Muller’s ratchet. Journal of Theoretical Biology, 524(??):??, September 7, 2021. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251932100151X.

Kubo:2025:MAS

[KNY25] Hidaka Kubo, Shinji Nakaoka, and Ryo Yamaguchi. Modeling alternative strategies of male emergence timing. Journal of Theoretical Biology, 612(??):??, September 7, 2025. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519325001808.

Koch:2020:HOS

[Koc20] Daniel Koch. Homo-oligomerisation in signal transduction: Dynamics, homeostasis, ultrasensitivity, bistability. Journal of Theoretical Biology, 499(??):??, August 21, 2020. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519320301600.

Kumar:2021:CTA

[KP21] Abhimanyu Kumar and Om Prakash Pandey. Causal theory on acceleration of seed germination in the vicinity of high voltage direct current transmission line. Journal of Theoretical Biology, 531(??):??, December 21, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321003180.

Kemp:2021:MCD

[KPA^{+}21] Françoise Kemp, Daniele Proverbio, Atte Aalto, Laurent Mombaerts, Aymeric Fouquier d’Hérouël, Andreas Husch, Christophe Ley, Jorge Gonçalves, Alexander Skupin, and Stefano Magni. Modelling COVID-19 dynamics and potential for herd immunity by vaccination in Austria, Luxembourg and Sweden. Journal of Theoretical Biology, 530(??):??, December 7, 2021. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519321002939.

Kassis:2024:DSM

[KPH^{+}24] George Kassis, Mukta G. Palshikar, Shannon P. Hilchey, Martin S. Zand, and Juilee Thakar. Discrete-state models identify pathway specific B cell states across diseases and infections at single-cell resolution. Journal of Theoretical Biology, 583(??): ??, April 21, 2024. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251932400050X.

Kim:2020:CTL

[KPP20] Seyoung Kim, Cheolhoon Park, and Chanhun Park. Change in telescoping leg strategy with varying walking speed to modulate force advantage. Journal of Theoretical Biology, 496(??): ??, July 7, 2020. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519320301041.

Krantz:2020:TEG

[KPR20] Steven G. Krantz, Peter Polyakov, and Arni S. R. Srinivasa Rao. True epidemic growth construction through harmonic analysis. Journal of Theoretical Biology, 494(??):??, June 7, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519320300989.

Karimi:2022:TMS

[KR22] Gholamreza Karimi and Soraya Rastegar. A TSTDP memristive synapse based on a comprehensive mathematical model of memory-TFT threshold voltage shift. Journal of Theoretical Biology, 544(??):??, July 7, 2022. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322001175.

Kumaran:2024:SAO

[KRG24] Shyni Unni Kumaran, Lavanya Rajagopal, and Manavaalan Gunasekaran. Sensitivity assessment of optimal control strategies and cost-effectiveness analysis of a novel Candida auris environmental transmission model in intensive care facilities. Journal of Theoretical Biology, 595(??):??, December 7, 2024. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0022519324002169.

Kroumi:2025:RSE

[Kro25] Dhaker Kroumi. The role of selective extinction in promoting cooperation in a many demes model. Journal of Theoretical Biology, 613(??):??, October 7, 2025. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251932500178X.

Krishnan:2024:CSH

[KRSG24] Nandakishor Krishnan, Lajos Rózsa, András Szilágyi, and József Garay. Coevolutionary stability of host-symbiont systems with mixed-mode transmission. Journal of Theoretical Biology, 576(??):??, January 7, 2024. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519323002175.

Ke:2021:ERN

[KRSSH21] Ruian Ke, Ethan Romero-Severson, Steven Sanche, and Nick Hengartner. Estimating the reproductive number R_{0} of SARS-CoV-2 in the United States and eight European countries and implications for vaccination. Journal of Theoretical Biology, 517(??):??, May 21, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321000436.

Kohnke:2020:TIF

[KSSM20] Merlin C. Köhnke, Ivo Siekmann, Hiromi Seno, and Horst Malchow. A type IV functional response with different shapes in a predator-prey model. Journal of Theoretical Biology, 505(??):??, November 21, 2020. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519320302745.

Kabir:2020:CEA

[KT20a] K. M. Ariful Kabir and Jun Tanimoto. Cost-efficiency analysis of voluntary vaccination against n-serovar diseases using antibody-dependent enhancement: a game approach. Journal of Theoretical Biology, 503(??):??, October 21, 2020. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519320302344.

Kabir:2020:RAL

[KT20b] K. M. Ariful Kabir and Jun Tanimoto. The role of advanced and late provisions in a co-evolutionary epidemic game model for assessing the social triple-dilemma aspect. Journal of Theoretical Biology, 503(??):??, October 21, 2020. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251932030254X.

Kabir:2021:AIS

[KT21] K. M. Ariful Kabir and Jun Tanimoto. Analysis of individual strategies for artificial and natural immunity with imperfectness and durability of protection. Journal of Theoretical Biology, 509(??):??, January 21, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519320303866.

Konstorum:2020:SBF

[KTP^{+}20] Anna Konstorum, Lia Tesfay, Bibbin T. Paul, Frank M. Torti, Reinhard C. Laubenbacher, and Suzy V. Torti. Systems biology of ferroptosis: a modeling approach. Journal of Theoretical Biology, 493(??):??, May 21, 2020. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519320300771.

Kitaguchi:2022:CSH

[KTU22] Yuta Kitaguchi, Hajime Tei, and Koichiro Uriu. Cell size homeostasis under the circadian regulation of cell division in cyanobacteria. Journal of Theoretical Biology, 553(??):??, November 21, 2022. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251932200251X.

Kurokawa:2021:EGS

[Kur21] Shun Kurokawa. Effect of the group size on the evolution of cooperation when an exit option is present. Journal of Theoretical Biology, 521(??):??, July 21, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321001004.

Kurokawa:2023:ESP

[Kur23] Shun Kurokawa. Evolution of spite in an n-player game with an opting-out option. Journal of Theoretical Biology, 572(??):??, September 7, 2023. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519323001820.

Kuznetsov:2024:GRS

[Kuz24a] Andrey V. Kuznetsov. The growth rate of senile plaques is determined by the competition between the rate of deposition of free Aβ aggregates into plaques and the autocatalytic production of free Aβ aggregates. Journal of Theoretical Biology, 593(??):??, October 7, 2024. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251932400184X.

Kuznetsov:2024:LBR

[Kuz24b] Andrey V. Kuznetsov. Lewy body radius growth: the hypothesis of the cube root of time dependency. Journal of Theoretical Biology, 581(??):??, March 21, 2024. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519324000158.

Kuznetsov:2024:NMS

[Kuz24c] Andrey V. Kuznetsov. Numerical modeling of senile plaque development under conditions of limited diffusivity of amyloid-β monomers. Journal of Theoretical Biology, 587(??): ??, June 21, 2024. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519324001048.

Khalili:2023:SIR

[KV23] Pariya Khalili and Ramin Vatankhah. Studying the importance of regulatory T cells in chemoimmunotherapy mathematical modeling and proposing new approaches for developing a mathematical dynamic of cancer. Journal of Theoretical Biology, 563(??):??, April 21, 2023. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519323000334.

Kandiah:2025:CDS

[KvLBD25] Joel Kandiah, Edwin van Leeuwen, Paul J. Birrell, and Daniela De Angelis. Contact data and SARS-CoV-2: Retrospective analysis of the estimated impact of the first UK lockdown. Journal of Theoretical Biology, 610(??):??, August 7, 2025. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519325001249.

Kumakura:2023:DGC

[KYHN23] Daiki Kumakura, Ryo Yamaguchi, Akane Hara, and Shinji Nakaoka. Disentangling the growth curve of microbial culture. Journal of Theoretical Biology, 573(??):??, September 21, 2023. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519323001947.

Kobayashi:2021:OCR

[KYI21] Ryota Kobayashi, Sachi Yamaguchi, and Yoh Iwasa. Optimal control of root nodulation — Prediction of life history theory of a mutualistic system. Journal of Theoretical Biology, 510(??):??, February 7, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519320303994.

Li:2024:RLH

[LAP24] Wentian Li, Yannis Almirantis, and Astero Provata. Rangelimited Heaps’ law for functional DNA words in the human genome. Journal of Theoretical Biology, 592(??):??, September 7, 2024. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519324001620.

Liu:2020:RDE

[LB20] Lin Liu and Alexander Bockmayr. Regulatory dynamic enzyme-cost flux balance analysis: a unifying framework for constraint-based modeling. Journal of Theoretical Biology, 501(??):??, September 21, 2020. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519320301727.

Le:2025:RSN

[LB25] Quang Anh Le and Seung Ki Baek. Response of social norms to individual differences in error-proneness. Journal of Theoretical Biology, 612(??):??, September 7, 2025. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519325001651.

Liao:2024:CAP

[LBF24] Kang-Ling Liao, Xue-Feng Bai, and Avner Friedman. IL-27 in combination with anti-PD-1 can be anti-cancer or procancer. Journal of Theoretical Biology, 579(??):??, February 21, 2024. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519323003016.

Liao:2021:MME

[LCMW21] Kang-Ling Liao, Wei-Chen Chang, Jeffrey M. Marcus, and Jenn-Nan Wang. Mathematical modeling of the eyespots in butterfly wings. Journal of Theoretical Biology, 531(??):??, December 21, 2021. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519321003179.

Lu:2023:RBD

[LCYD23] Zhengshu Lu, Xu Chen, Jing Yang, and Yanrui Ding. RSC-based differential model with correlation removal for improving multi-omics clustering. Journal of Theoretical Biology, 556(??):??, January 7, 2023. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519322003198.

Li:2022:TSR

[LDaH22] Cancan Li, Qi Dai, and Ping an He. A time series representation of protein sequences for similarity comparison. Journal of Theoretical Biology, 538(??):??, April 7, 2022. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322000376.

Lappo:2023:CAC

[LDF23] Egor Lappo, Kaleda K. Denton, and Marcus W. Feldman. Conformity and anti-conformity in a finite population. Journal of Theoretical Biology, 562(??):??, April 7, 2023. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519323000255.

Li:2024:APE

[LDL^{+}24] Shulan Li, Chunpeng Du, Xingxu Li, Chen Shen, and Lei Shi. Antisocial peer exclusion does not eliminate the effectiveness of prosocial peer exclusion in structured populations. Journal of Theoretical Biology, 576(??):??, January 7, 2024. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251932300262X.

Lehtinen:2021:EEC

[Leh21] Sami O. Lehtinen. Ecological and evolutionary consequences of predator-prey role reversal: Allee effect and catastrophic predator extinction. Journal of Theoretical Biology, 510(??):??, February 7, 2021. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519320303970.

Lehmann:2022:HRE

[Leh22] Laurent Lehmann. Hamilton’s rule, the evolution of behavior rules and the wizardry of control theory. Journal of Theoretical Biology, 555(??):??, December 21, 2022. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322002739.

Lei:2020:GMF

[Lei20] Jinzhi Lei. A general mathematical framework for understanding the behavior of heterogeneous stem cell regeneration. Journal of Theoretical Biology, 492(??):??, May 7, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519320300515. See corrigendum [Lei24].

Lei:2024:CGM

[Lei24] Jinzhi Lei. Corrigendum to “A general mathematical framework for understanding the behavior of heterogeneous stem cell regeneration” [J. Theoret. Biol. 492 (2020) 110196]. Journal of Theoretical Biology, 593(??):??, October 7, 2024. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519324001802. See [Lei20].

Lazebnik:2025:STM

[LF25] Teddy Lazebnik and Avner Friedman. Spatio-temporal model of combining ADT and chemotherapy with senolytic treatment in metastatic prostate cancer. Journal of Theoretical Biology, 602–603(??):??, April 7, 2025. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519325000359.

Lin:2021:MDM

[LFJ21] Hsin-Yi Lin, William F. Fagan, and Pierre-Emmanuel Jabin. Memory-driven movement model for periodic migrations. Journal of Theoretical Biology, 508(??):??, January 7, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519320303416.

Li:2023:WSS

[LFT^{+}23] Cong Li, Tianjiao Feng, Yi Tao, Xiudeng Zheng, and Jiajia Wu. Weak selection and stochastic evolutionary stability in a stochastic replicator dynamics. Journal of Theoretical Biology, 570(??):??, August 7, 2023. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519323001200.

Liu:2024:MST

[LFY^{+}24] Jiashuai Liu, Zhenmin Fan, Xia Ye, Yingying Zhang, Mingyuan Liu, and Xiaoyan Deng. Modelling of the in-stent thrombus formation by dissipative particle dynamics. Journal of Theoretical Biology, 582(??):??, April 7, 2024. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519324000390.

Li:2024:MMF

[LFZ^{+}24] Xiong Li, Xuan Feng, Juan Zhou, Yuchao Luo, Xiao Chen, Jiapeng Zhao, Haowen Chen, Guoming Xiong, and Guoliang Luo. A multi-modal feature fusion method based on deep learning for predicting immunotherapy response. Journal of Theoretical Biology, 586(??):??, June 7, 2024. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519324000973.

Lindwall:2023:BIA

[LG23] Gustav Lindwall and Philip Gerlee. Bayesian inference on the Allee effect in cancer cell line populations using time-lapse microscopy images. Journal of Theoretical Biology, 574(??):??, October 7, 2023. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519323002217.

Li:2023:MMN

[LGBG23] K. Li, A. J. Gallo, B. J. Binder, and J. E. F. Green. A mathematical model for nutrient-limited uniaxial growth of a compressible tissue. Journal of Theoretical Biology, 575(??):??, November 7, 2023. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251932300228X.

Lindsly:2021:UMB

[LGSR21] Stephen Lindsly, Maya Gupta, Cooper Stansbury, and Indika Rajapakse. Understanding memory B cell selection. Journal of Theoretical Biology, 531(??):??, December 21, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321003246.

Lambert:2021:MCM

[LGT21] Ben Lambert, David J. Gavaghan, and Simon J. Tavener. A Monte Carlo method to estimate cell population heterogeneity from cell snapshot data. Journal of Theoretical Biology, 511(??):??, February 21, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519320303969.

Loos:2020:RCH

[LH20] Carolin Loos and Jan Hasenauer. Robust calibration of hierarchical population models for heterogeneous cell populations. Journal of Theoretical Biology, 488(??):??, March 7, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519319304874.

Leech:2022:MMA

[LHG^{+}22] V. Leech, J. W. Hazel, J. C. Gatlin, A. E. Lindsay, and A. Manhart. Mathematical modeling accurately predicts the dynamics and scaling of nuclear growth in discrete cytoplasmic volumes. Journal of Theoretical Biology, 533(??):??, January 21, 2022. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519321003556.

Lion:2018:PES

[Lio18] Sébastien Lion. From the Price equation to the selection gradient in class-structured populations: a quasi-equilibrium route. Journal of Theoretical Biology, 447(??):178–189, June 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519318301474. See corrigen-

dum [Lio24].

Lion:2024:CPE

[Lio24] Sébastien Lion. Corrigendum to “From the Price equation to the selection gradient in class-structured populations: a quasi-equilibrium route” [J. Theor. Biol. 447 (2018) 178– 189]. Journal of Theoretical Biology, 591(??):??, August 21, 2024. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519324001577. See [Lio18].

Leng:2020:IPA

[LK20] Trystan Leng and Matt J. Keeling. Improving pairwise approximations for network models with susceptibleinfected-susceptible dynamics. Journal of Theoretical Biology, 500(??):??, September 7, 2020. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519320301831.

Lopez:2025:CON

[LK25a] Samuel Lopez and Natalia L. Komarova. Corrigendum to “An optimal network that promotes the spread of an advantageous variant in an SIR epidemic”. [J. Theor. Biol. 605 (2025) 112095]. Journal of Theoretical Biology, 613(??):??, October 7, 2025. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519325001833. See [LK25b].

Lopez:2025:ONP

[LK25b] Samuel Lopez and Natalia L. Komarova. An optimal network that promotes the spread of an advantageous variant in an SIR epidemic. Journal of Theoretical Biology, 605(??):??, May 21, 2025. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251932500061X. See corri-

gendum [LK25a].

Lechthaler:2020:TPL

[LKP^{+}20] Silvia Lechthaler, Natasa Kiorapostolou, Andrea Pitacco, Tommaso Anfodillo, and Giai Petit. The total path length hydraulic resistance according to known anatomical patterns: What is the shape of the root-to-leaf tension gradient along the plant longitudinal axis? Journal of Theoretical Biology, 502(??):??, October 7, 2020. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519320302241.

Li:2020:EMI

[LL20] Wenjie Li and Yao Li. Entropy, mutual information, and systematic measures of structured spiking neural networks. Journal of Theoretical Biology, 501(??):??, September 21, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251932030165X.

Li:2021:EOS

[LL21] Cong Li and Sabin Lessard. The effect of the optingout strategy on conditions for selection to favor the evolution of cooperation in a finite population. Journal of Theoretical Biology, 510(??):??, February 7, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519320303982.

Lenstra:2025:CFA

[LLJ25] Reijer Lenstra, Antonia Claudio Lenstra, and Marco V. José. Crick’s “Frozen accident theory” of the genetic code explored by Ising models. Journal of Theoretical Biology, 609(??):??, July 21, 2025. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519325001122.

Leschiera:2022:MMS

[LLS^{+}22] Emma Leschiera, Tommaso Lorenzi, Shensi Shen, Luis Almeida, and Chloe Audebert. A mathematical model to study the impact of intra-tumour heterogeneity on anti-tumour CD8^{+} T cell immune response. Journal of Theoretical Biology, 538(??):??, April 7, 2022. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322000261.

Liu:2022:IGL

[LLZ^{+}22] Liyue Liu, Xiangyu Liu, Juan Zhou, Weihua Wu, and Xiong Li. An improved graph Laplacian regularization method for identifying biomarkers of Alzheimer’s disease. Journal of Theoretical Biology, 543(??):??, June 21, 2022. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322001199.

Liu:2025:SRF

[LLZS25] Xiaoting Liu, Maoxing Liu, Donghua Zhao, and Yongzheng Sun. Stability and reactivity of food webs with trophic interactions. Journal of Theoretical Biology, 613(??):??, October 7, 2025. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519325001821.

Lopez:2021:MRA

[LMA21] Alejandra Rojas López, Pablo Monzón, and Luis Acerenza. A model for the regulation of apoptosis intrinsic pathway: the potential role of the transcriptional regulator E2F in the point of no return. Journal of Theoretical Biology, 525(??):??, September 21, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321001879.

Lee:2022:SOS

[LMB22] Sanghun Lee, Yohsuke Murase, and Seung Ki Baek. A second-order stability analysis for the continuous model of indirect reciprocity. Journal of Theoretical Biology, 548(??):??, September 7, 2022. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0022519322002004.

Liu:2020:NSE

[LMC^{+}20] Wei Liu, Liangyu Ma, Ling Chen, Bolun Chen, Byeungwoo Jeon, and Jipeng Qiang. A novel scheme for essential protein discovery based on multi-source biological information. Journal of Theoretical Biology, 504(??):??, November 7, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519320302691.

Li:2021:MWH

[LMC21] Ke Li, James M. McCaw, and Pengxing Cao. Modelling within-host macrophage dynamics in influenza virus infection. Journal of Theoretical Biology, 508(??):??, January 7, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519320303477.

Le:2022:DHM

[LMG22] Thi Minh Thao Le, Sten Madec, and Erida Gjini. Disentangling how multiple traits drive 2 strain frequencies in SIS dynamics with coinfection. Journal of Theoretical Biology, 538(??):??, April 7, 2022. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251932200039X.

Lubenia:2024:CAK

[LML24] Patrick Vincent N. Lubenia, Eduardo R. Mendoza, and Angelyn R. Lao. Comparative analysis of kinetic realizations of insulin signaling. Journal of Theoretical Biology, 577(??):??, January 21, 2024. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519323002692.

Li:2020:BSM

[LMLP20] Yuan Li, Ian C. McPhee, Michael P. H. Lau, and Leonard F. Pease. Biomimetic surgical mesh to replace fascia with tunable force-displacement. Journal of Theoretical Biology, 486(??):??, February 7, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519319304278.

Levesque:2021:MCP

[LMS21] Jérôme Levesque, David W. Maybury, and R. H. A. David Shaw. A model of COVID-19 propagation based on a gamma subordinated negative binomial branching process. Journal of Theoretical Biology, 512(??):??, March 7, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251932030391X.

Liu:2021:PNR

[LMW21] Z. Liu, P. Magal, and G. Webb. Predicting the number of reported and unreported cases for the COVID-19 epidemics in China, South Korea, Italy, France, Germany and United Kingdom. Journal of Theoretical Biology, 509(??):??, January 21, 2021. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519320303568.

Le:2023:SMD

[LNT23] Dieu Huong Le, Trong Hieu Nguyen, and Fugo Takasu. Spatial metapopulation dynamics with local and global colonization. Journal of Theoretical Biology, 572(??):??, September 7, 2023. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519323001765.

Laplanche:2025:EFW

[LPA25] Christophe Laplanche, Benjamin Pey, and Robin Aguilée. Emergence of food webs with a multi-trophic hierarchical structure driven by nonlinear trait-matching. Journal of Theoretical Biology, 605(??):??, May 21, 2025. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519325000578.

Lema-Perez:2021:PBS

[LPBMA21] Laura Lema-Perez, Carlos E. Builes-Montaño, and Hernan Alvarez. A phenomenological-based semi-physical model of the kidneys and its role in glucose metabolism. Journal of Theoretical Biology, 508(??):??, January 7, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519320303441.

Lie-Panis:2023:RSE

[LPD23] Julien Lie-Panis and Jean-Louis Dessalles. Runaway signals: Exaggerated displays of commitment may result from second-order signaling. Journal of Theoretical Biology, 572(??):??, September 7, 2023. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519323001832.

Landi:2025:OSD

[LPLM25] Alberto Landi, Giulio Pisaneschi, Marco Laurino, and Piero Manfredi. Optimal social distancing in pandemic preparedness and lessons from COVID-19: Intervention intensity and infective travelers. Journal of Theoretical Biology, 604(??): ??, May 7, 2025. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519325000384.

Lauro:2025:LCC

[LPP^{+}25] Francesco Di Lauro, William J. M. Probert, Michael Pickles, Anne Cori, Robert Hinch, Luca Ferretti, Jasmina Panovska-Griffiths, Lucie Abeler-Dörner, Rory Dunbar, Peter Bock, Deborah J. Donnell, Helen Ayles, Sarah Fidler, Richard Hayes, and Christophe Fraser. Large connected components in sexual networks and their role in HIV transmission in Sub-Saharan Africa: a model-based analysis of HPTN 071(PopART) data. Journal of Theoretical Biology, 613(??):??, October 7, 2025. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519325001845.

Legaria-Pena:2023:EEF

[LPSMCP23] Juan Uriel Legaria-Peña, Félix Sánchez-Morales, and Yuriria Cortés-Poza. Evaluation of entropy and fractal dimension as biomarkers for tumor growth and treatment response using cellular automata. Journal of Theoretical Biology, 564(??): ??, May 7, 2023. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519323000589.

Lee:2021:DGT

[LR21] Tae Kyu Lee and Isaac I. Rosen. Development of generalized time-dependent TCP model and the investigation of the effect of repopulation and weekend breaks in fractionated external beam therapy. Journal of Theoretical Biology, 512(??):??, March 7, 2021. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519320304203.

Larripa:2025:MMM

[LR25] Kamila Larripa and Anca Rădulescu. A mathematical model of microglia glucose metabolism and lactylation with positive feedback. Journal of Theoretical Biology, 602–603(??): ??, April 7, 2025. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519325000153.

Lacy:2022:CDD

[LRAP22] Thomas C. Lacy, Peter A. Robinson, Kevin M. Aquino, and James C. Pang. Cortical depth-dependent modeling of visual hemodynamic responses. Journal of Theoretical Biology, 535(??):??, February 21, 2022. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321003982.

Lovell-Read:2022:ELO

[LRST22] Francesca A. Lovell-Read, Silvia Shen, and Robin N. Thompson. Estimating local outbreak risks and the effects of non-pharmaceutical interventions in age-structured populations: SARS-CoV-2 as a case study. Journal of Theoretical Biology, 535(??):??, February 21, 2022. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321004033.

Lasri:2021:IMS

[LS21] Ayoub Lasri and Marc Sturrock. The influence of methylation status on a stochastic model of MGMT dynamics in glioblastoma: Phenotypic selection can occur with and without a downshift in promoter methylation status. Journal of Theoretical Biology, 521(??):??, July 21, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321000849.

Lotter:2025:PEW

[LS25a] Clara R. Lotter and Jonathan A. Sherratt. Pulmonary epithelial wound healing and the immune system. Mathematical modeling and bifurcation analysis of a bistable system. Journal of Theoretical Biology, 596(??):??, January 7, 2025. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519324002534.

Luo:2025:DCD

[LS25b] Lanjiang Luo and Jianwei Shen. Decoding cancer dynamics: the role of miRNA in E2F/Myc networks and bifurcation points. Journal of Theoretical Biology, 613(??):??, October 7, 2025. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519325001857.

Lachance:2020:LSS

[LSW20] Joseph Lachance, Corinne N. Simonti, and Joshua S. Weitz. Large sample spaces do not imply biological systems are ‘finetuned’. Journal of Theoretical Biology, 507(??):??, December 21, 2020. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251932030312X.

Lange:2024:MCA

[LSWVB24] Steffen Lange, Jannik Schmied, Paul Willam, and Anja Voss-Böhme. Minimal cellular automaton model with heterogeneous cell sizes predicts epithelial colony growth. Journal of Theoretical Biology, 592(??):??, September 7, 2024. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519324001668.

Lehotzky:2021:CAM

[LSZ21] Dávid Lehotzky, Rifat Sipahi, and Günther K. H. Zupanc. Cellular automata modeling suggests symmetric stem-cell division, cell death, and cell drift as key mechanisms driving adult spinal cord growth in teleost fish. Journal of Theoretical Biology, 509(??):??, January 21, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519320303295.

Loo:2022:RPI

[LT22] Sara L. Loo and Mark M. Tanaka. The role of a programmatic immune response on the evolution of pathogen traits. Journal of Theoretical Biology, 534(??):??, February 7, 2022. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321003829.

Legried:2023:IIP

[LT23] Brandon Legried and Jonathan Terhorst. Identifiability and inference of phylogenetic birth-death models. Journal of Theoretical Biology, 568(??):??, July 7, 2023. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519323001169.

Liu:2025:LHD

[LTB25] Deze Liu, Mohammad Tuqan, and Daniel Burbano. Learning to hunt: a data-driven stochastic feedback control model of predator–prey interactions. Journal of Theoretical Biology, 599(??):??, February 21, 2025. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519324003060.

Leyshon:2021:DPD

[LTS^{+}21] Thomas Leyshon, Elisa Tonello, David Schnoerr, Heike Siebert, and Michael P. H. Stumpf. The design principles of discrete Turing patterning systems. Journal of Theoretical Biology, 531(??):??, December 21, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321003209.

Lutscher:2020:RCE

[LW20] Frithjof Lutscher and Xiaoying Wang. Reactivity of communities at equilibrium and periodic orbits. Journal of Theoretical Biology, 493(??):??, May 21, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519320300953.

Li:2023:EDT

[LWM23] Ao Li, Jianhong Wu, and Seyed M. Moghadas. Epidemic dynamics with time-varying transmission risk reveal the role of disease stage-dependent infectiousness. Journal of Theoretical Biology, 573(??):??, September 21, 2023. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519323001911.

Li:2020:IAE

[LWXM20] Meili Li, Manting Wang, Shuyang Xue, and Junling Ma. The influence of awareness on epidemic spreading on random networks. Journal of Theoretical Biology, 486(??):??, February 7, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251931930459X.

Li:2021:LDT

[LX21] Tangjuan Li and Yanni Xiao. Linking the disease transmission to information dissemination dynamics: an insight from a multi-scale model study. Journal of Theoretical Biology, 526(??):??, October 7, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519321002150.

Lee:2024:EBC

[LYE^{+}24] Lloyd Lee, Ryan Yoast, Scott Emrich, Mohamed Trebak, Vivien Kirk, and James Sneyd. Emergence of broad cytosolic Ca^{2+} oscillations in the absence of CRAC channels: a model for CRAC-mediated negative feedback on PLC and Ca^{2+} oscillations through PKC. Journal of Theoretical Biology, 581(??):??, March 21, 2024. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519324000213.

Lee:2020:PVR

[LYYN20] Joung Hun Lee, Ryo Yamaguchi, Hiroyuki Yokomizo, and Mayuko Nakamaru. Preservation of the value of rice paddy fields: Investigating how to prevent farmers from abandoning the fields by means of evolutionary game theory. Journal of Theoretical Biology, 495(??):??, June 21, 2020. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519320301028.

Lv:2023:HUE

[LZ23] Yanzhao Lv and Zupeng Zhou. Humeri under external load: Mechanical implications of differing bone curvature in American otter and honey badger. Journal of Theoretical Biology, 558(??):??, February 7, 2023. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322003496.

Li:2025:OCV

[LZL^{+}25] Li Li, Na Zheng, Chen Liu, Zhen Wang, and Zhen Jin. Optimal control of vaccination for an epidemic model with standard incidence rate. Journal of Theoretical Biology, 598(??):??, February 7, 2025. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519324002789.

Lu:2023:EMR

[LZW23] Ruijin Lu, Hongxiao Zhu, and Xiaowei Wu. Estimating mutation rates in a Markov branching process using approximate Bayesian computation. Journal of Theoretical Biology, 565(??):??, May 21, 2023. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519323000632.

Li:2020:SAD

[LZZ20] Chentong Li, Yingying Zhang, and Yicang Zhou. Spatially antiviral dynamics determines HCV in vivo replication and evolution. Journal of Theoretical Biology, 503(??):??, October 21, 2020. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519320302332.

Li:2023:ABV

[LZZ^{+}23] Zhaowan Li, Jianguo Zhao, Yuhao Zhou, Lina Tian, Qihuai Liu, Huaiping Zhu, and Guanghu Zhu. Adaptive behaviors and vaccination on curbing COVID-19 transmission: Modeling simulations in eight countries. Journal of Theoretical Biology, 559(??):??, February 21, 2023. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322003708.

Montaseri:2020:MMF

[MAHMH20] Ghazal Montaseri, Juan Carlos López Alfonso, Haralampos Hatzikirou, and Michael Meyer-Hermann. A minimal modeling framework of radiation and immune system synergy to assist radiotherapy planning. Journal of Theoretical Biology, 486(??):??, February 7, 2020. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319304680.

Martinelli:2024:ROI

[Mar24] Agostino Martinelli. Revisiting the observability and identifiability properties of a popular HIV model. Journal of Theoretical Biology, 584(??):??, May 7, 2024. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519324000614.

Mairet:2021:PDM

[MB21] Francis Mairet and Térence Bayen. The promise of dawn: Microalgae photoacclimation as an optimal control problem of resource allocation. Journal of Theoretical Biology, 515(??): ??, April 21, 2021. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519321000199.

Mukund:2022:TSP

[MB22] Adi Mukund and Lacramioara Bintu. Temporal signaling, population control, and information processing through chromatin-mediated gene regulation. Journal of Theoretical Biology, 535(??):??, February 21, 2022. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519321003970.

Moyse:2024:CNF

[MB24] Lisa Blum Moyse and Hugues Berry. A coupled neural field model for the standard consolidation theory. Journal of Theoretical Biology, 588(??):??, July 7, 2024. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519324000997.

Mayer-Bacon:2021:BCU

[MBF21] Christopher Mayer-Bacon and Stephen J. Freeland. A broader context for understanding amino acid alphabet optimality. Journal of Theoretical Biology, 520(??):??, July 7, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519321000837.

Mosleh:2024:EID

[MBHG^{+}24] Rahele Mosleh, Mortaza Baky-Haskuee, Abbas Ghasemi, Martin Grunnill, Julien Arino, Mohammadali Tofighi, Edward W. Thommes, and Jianhong Wu. Evaluating infectious disease outbreak potential and mitigation effectiveness on cruise ships. Journal of Theoretical Biology, 592(??):??, September 7, 2024. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519324001590.

Michael:2025:RTA

[MBM^{+}25] Christian T. Michael, Maral Budak, Pauline Maiello, Kara Kracinovsky, Mark Rodgers, Jaime Tomko, Philana Ling Lin, JoAnne Flynn, Jennifer J. Linderman, and Denise Kirschner. Rankings of tuberculosis antibiotic treatment regimens are sensitive to spatial scale, detection limit, and initial host bacterial burden. Journal of Theoretical Biology, 611(??):??, August 21, 2025. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519325001420.

Marenduzzo:2025:OSM

[MBMA25] Davide Marenduzzo, Aidan T. Brown, Craig W. Miller, and Graeme J. Ackland. Oscillation in the SIRS model. Journal of Theoretical Biology, 611(??):??, August 21, 2025. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519325001353.

MacQueen:2022:MGF

[MBT22] Sarah A. MacQueen, W. John Braun, and Rebecca C. Tyson. Memory-guided foraging and landscape design interact to determine ecosystem services. Journal of Theoretical Biology, 534(??):??, February 7, 2022. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519321003775.

Mazzolini:2020:GSE

[MC20a] Andrea Mazzolini and Antonio Celani. Generosity, selfishness and exploitation as optimal greedy strategies for resource sharing. Journal of Theoretical Biology, 485(??):??, January 21, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519319304102.

Mukhopadhyay:2020:POC

[MC20b] Archan Mukhopadhyay and Sagar Chakraborty. Periodic orbit can be evolutionarily stable: Case study of discrete replicator dynamics. Journal of Theoretical Biology, 497(??):??, July 21, 2020. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519320301430.

Munoz:2022:PMS

[MC22] Estefanía Muñoz and Jorge Carneiro. Plant-microbe symbiosis widens the habitability range of the Daisyworld. Journal of Theoretical Biology, 554(??):??, December 7, 2022. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322002661.

Melfo:2020:IBN

[MCL20] Alejandra Melfo, Ragan M. Callaway, and Luis D. Llambí. Interactions between nurse plants and parasitic beneficiaries: a theoretical approach to indirect facilitation. Journal of Theoretical Biology, 494(??):??, June 7, 2020. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251932030093X.

Meyer:2021:DAS

[MCL21] Emily E. Meyer, Colleen E. Clancy, and Timothy J. Lewis. Dynamics of adrenergic signaling in cardiac myocytes and implications for pharmacological treatment. Journal of Theoretical Biology, 519(??):??, June 21, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321000412.

Meehan:2020:PSI

[MCM20] Michael T. Meehan, Robert C. Cope, and Emma S. McBryde. On the probability of strain invasion in endemic settings: Accounting for individual heterogeneity and control in multistrain dynamics. Journal of Theoretical Biology, 487(??):??, February 21, 2020. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519319304783.

McGrane-Corrigan:2024:ISG

[MCMdAM24] Blake McGrane-Corrigan, Oliver Mason, and Rafael de Andrade Moral. Inferring stochastic group interactions within structured populations via coupled autoregression. Journal of Theoretical Biology, 584(??):??, May 7, 2024. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519324000742.

Miller:2021:MPC

[MCO21] Claire Miller, Edmund Crampin, and James M. Osborne. Maintaining the proliferative cell niche in multicellular models of epithelia. Journal of Theoretical Biology, 527(??):??, October 21, 2021. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519321002265.

Montefusco:2020:HAC

[MCP20] Francesco Montefusco, Giuliana Cortese, and Morten G. Pedersen. Heterogeneous alpha-cell population modeling of glucose-induced inhibition of electrical activity. Journal of Theoretical Biology, 485(??):??, January 21, 2020. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319304059.

Manjoo-Docrat:2022:SSM

[MD22] Raeesa Manjoo-Docrat. A spatio-stochastic model for the spread of infectious diseases. Journal of Theoretical Biology, 533(??):??, January 21, 2022. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519321003623.

Meaney:2023:DLC

[MDCK23] Cameron Meaney, Sunit Das, Errol Colak, and Mohammad Kohandel. Deep learning characterization of brain tumours with diffusion weighted imaging. Journal of Theoretical Biology, 557(??):??, January 21, 2023. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322003332.

Moyles:2021:QSU

[MDF21] I. R. Moyles, J. G. Donohue, and A. C. Fowler. Quasisteady uptake and bacterial community assembly in a mathematical model of soil-phosphorus mobility. Journal of Theoretical Biology, 509(??):??, January 21, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519320303854.

Meca:2024:MTE

[MDG24] Esteban Meca, Concepción M. Díez, and Brandon S. Gaut. Modeling transposable elements dynamics during polyploidization in plants. Journal of Theoretical Biology, 579(??):??, February 21, 2024. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519323002989.

McNicol:2025:TMF

[MDS25] Gordon R. McNicol, Matthew J. Dalby, and Peter S. Stewart. A theoretical model for focal adhesion and cytoskeleton formation in non-motile cells. Journal of Theoretical Biology, 596(??):??, January 7, 2025. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519324002509.

Miele:2021:RST

[MEA21] Leonardo Miele, R. M. L. Evans, and Sandro Azaele. Redundancy-selection trade-off in phenotype-structured populations. Journal of Theoretical Biology, 531(??):??, December 21, 2021. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519321003039.

Mohsin:2024:STV

[MEBNZ24] Nuverah Mohsin, Heiko Enderling, Renee Brady-Nicholls, and Mohammad U. Zahid. Simulating tumor volume dynamics in response to radiotherapy: Implications of model selection. Journal of Theoretical Biology, 576(??):??, January 7, 2024. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519323002539.

Massard:2022:MSE

[MEPS22] Mathilde Massard, Raluca Eftimie, Antoine Perasso, and Bruno Saussereau. A multi-strain epidemic model for COVID-19 with infected and asymptomatic cases: Application to French data. Journal of Theoretical Biology, 545(??): ??, July 21, 2022. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519322001151.

Moise:2021:MMM

[MF21] Nicolae Moise and Avner Friedman. A mathematical model of the multiple sclerosis plaque. Journal of Theoretical Biology, 512(??):??, March 7, 2021. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519320303878.

Moise:2022:MMI

[MF22] Nicolae Moise and Avner Friedman. A mathematical model of immunomodulatory treatment in myocardial infarction. Journal of Theoretical Biology, 544(??):??, July 7, 2022. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322001205.

Mangrum:2024:MHA

[MF24] Diamond S. Mangrum and Stacey D. Finley. Modeling the heterogeneous apoptotic response of caspase-mediated signaling in tumor cells. Journal of Theoretical Biology, 590(??):??, August 7, 2024. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519324001383.

Marquez-Florez:2023:CMA

[MFGACST23] K. Márquez-Flórez, D. A. Garzón-Alvarado, C. Carda, and M. Sancho-Tello. Computational model of articular cartilage regeneration induced by scaffold implantation in vivo. Journal of Theoretical Biology, 561(??):??, March 21, 2023. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322003848.

Milne:2021:CPA

[MG21] Russell Milne and Frederic Guichard. Coupled phaseamplitude dynamics in heterogeneous metacommunities. Journal of Theoretical Biology, 523(??):??, August 21, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321000989.

Mesterton-Gibbons:2021:EET

[MGCH21] Mike Mesterton-Gibbons, Antonino Cusumano, and Ian C. W. Hardy. Escaping the evolutionary trap: Can sizerelated contest advantage compensate for juvenile mortality disadvantage when parasitoids develop in unnatural invasive hosts? Journal of Theoretical Biology, 527(??):??, October 21, 2021. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251932100240X.

Martinez-Garcia:2020:HRR

[MGFS^{+}20] Ricardo Martinez-Garcia, Christen H. Fleming, Ralf Seppelt, William F. Fagan, and Justin M. Calabrese. How range residency and long-range perception change encounter rates. Journal of Theoretical Biology, 498(??):??, August 7, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519320301223.

Manceau:2021:PDA

[MGVS21] Marc Manceau, Ankit Gupta, Timothy Vaughan, and Tanja Stadler. The probability distribution of the ancestral population size conditioned on the reconstructed phylogenetic tree with occurrence data. Journal of Theoretical Biology, 509(??):??, January 21, 2021. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519320302551.

Martignoni:2021:CIA

[MGZ^{+}21] Maria M. Martignoni, Jimmy Garnier, Xinlu Zhang, Daniel Rosa, Vasilis Kokkoris, Rebecca C. Tyson, and Miranda M. Hart. Co-inoculation with arbuscular mycorrhizal fungi differing in carbon sink strength induces a synergistic effect in plant growth. Journal of Theoretical Biology, 531(??):??, December 21, 2021. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519321002782.

Mallela:2022:OMS

[MH22a] Abhishek Mallela and Alan Hastings. Optimal management of stochastic invasion in a metapopulation with Allee effects. Journal of Theoretical Biology, 549(??):??, September 21, 2022. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519322002193.

McAlister:2022:ADM

[MH22b] John S. McAlister and Ian M. Hamilton. An adaptive dynamic model of a vigilance game among group foragers. Journal of Theoretical Biology, 538(??):??, April 7, 2022. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322000285.

Matsuo:2023:MCC

[MHI23] Kosei Matsuo, Rena Hayashi, and Yoh Iwasa. Multiple colonies of cancer involved in mutual suppression with the immune system. Journal of Theoretical Biology, 572(??):??, September 7, 2023. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519323001741.

Monter:2023:DRS

[MHPB23] Samuel Monter, Veit-Lorenz Heuthe, Emanuele Panizon, and Clemens Bechinger. Dynamics and risk sharing in groups of selfish individuals. Journal of Theoretical Biology, 562(??): ??, April 7, 2023. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519323000292.

Minucci:2021:MMV

[MHV^{+}21] Sarah Minucci, Rebecca L. Heise, Michael S. Valentine, Franck J. Kamga Gninzeko, and Angela M. Reynolds. Mathematical modeling of ventilator-induced lung inflammation. Journal of Theoretical Biology, 526(??):??, October 7, 2021.

CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321001600.

Melkikh:2020:AIT

[MI20] A. V. Melkikh and K. D. Izmozherova. Active ion transport as the basis for water movement in plants. Journal of Theoretical Biology, 500(??):??, September 7, 2020. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519320301879.

Mimura:2023:CCB

[MI23] Tomohiro Mimura and Yasuhiro Inoue. Cell-center-based model for simulating three-dimensional monolayer tissue deformation. Journal of Theoretical Biology, 571(??):??, August 21, 2023. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519323001571.

Mardanpour:2021:IWF

[MIP^{+}21] Pezhman Mardanpour, Ehsan Izadpanahi, Shanae Powell, Siavash Rastkar, and Adrian Bejan. Inflected wings in flight: Uniform flow of stresses makes strong and light wings for stable flight. Journal of Theoretical Biology, 508(??):??, January 7, 2021. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519320303076.

Mitarai:2020:HPP

[Mit20] Namiko Mitarai. How pirate phage interferes with helper phage: Comparison of the two distinct strategies. Journal of Theoretical Biology, 486(??):??, February 7, 2020. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319304655.

Mackey:2021:MAR

[MK21] Crystal Mackey and Christopher Kribs. Modeling anthrax– rabies interactions in zebra–jackal cycles. Journal of Theoretical Biology, 511(??):??, February 21, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519320304082.

Miki:2024:MVP

[MK24] Takeshi Miki and Po-Ju Ke. Macroscale vertical powerlaw distribution of bacteria in dark oceans can emerge from microscale bacteria-particle interactions. Journal of Theoretical Biology, 595(??):??, December 7, 2024. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0022519324002418.

Morris:2021:SHM

[ML21] Patricia Morris and Frithjof Lutscher. A seasonal hybrid model for the evolution of flowering onset. Journal of Theoretical Biology, 523(??):??, August 21, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321000904.

Martin:2024:ECS

[ML24] Éloi Martin and Sabin Lessard. Evolution of cooperation in social dilemmas with assortment in finite populations. Journal of Theoretical Biology, 592(??):??, September 7, 2024. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519324001759.

Mafessoni:2021:FIC

[MLG21] Fabrizio Mafessoni, Michael Lachmann, and Chaitanya S. Gokhale. On the fitness of informative cues in complex environments. Journal of Theoretical Biology, 527(??):??, October 21, 2021. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519321002381.

MacKay:2021:DDL

[MLK21] Laurent MacKay, Etienne Lehman, and Anmar Khadra. Deciphering the dynamics of lamellipodium in a fish keratocytes model. Journal of Theoretical Biology, 512(??):??, March 7, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519320303891.

Meng:2022:EDM

[MM22] Lingqi Meng and Naoki Masuda. Epidemic dynamics on metapopulation networks with node2vec mobility. Journal of Theoretical Biology, 534(??):??, February 7, 2022. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321003805.

Maciel:2021:ESC

[MMG21] Gabriel Andreguetto Maciel and Ricardo Martinez-Garcia. Enhanced species coexistence in Lotka–Volterra competition models due to nonlocal interactions. Journal of Theoretical Biology, 530(??):??, December 7, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321002915.

Mamiya:2021:CEZ

[MMI21] Azumi Mamiya, Daiki Miyagawa, and Genki Ichinose. Conditions for the existence of zero-determinant strategies under observation errors in repeated games. Journal of Theoretical Biology, 526(??):??, October 7, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321002290.

Miyagawa:2022:APA

[MMI22] Daiki Miyagawa, Azumi Mamiya, and Genki Ichinose. Adapting paths against zero-determinant strategies in repeated prisoner’s dilemma games. Journal of Theoretical Biology, 549(??):??, September 21, 2022. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519322002090.

McCauley:2024:BMF

[MMK^{+}24] Mark E. McCauley, Peter McCauley, Leonid V. Kalachev, Samantha M. Riedy, Siobhan Banks, Adrian J. Ecker, David F. Dinges, and Hans P. A. Van Dongen. Biomathematical modeling of fatigue due to sleep inertia. Journal of Theoretical Biology, 590(??):??, August 7, 2024. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519324001322.

Morikawa:2023:URD

[MMS^{+}23] Kentaro Morikawa, Shinichi Morita, Kazuki Sakura, Akiteru Maeno, Hiroki Gotoh, Teruyuki Niimi, and Yasuhiro Inoue. Unveiling the role of differential growth in 3D morphogenesis: an inference method to analyze area expansion rate distribution in biological systems. Journal of Theoretical Biology, 575(??):??, November 7, 2023. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519323002473.

Morita:2023:EIH

[MN23] Satoru Morita and Kuninori Nakagawa. Evaluating the impact of human flow on the spread of infectious diseases. Journal of Theoretical Biology, 558(??):??, February 7, 2023. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322003587.

Ma:2022:MCI

[MP22] Linda Ma and Mainak Patel. Mechanism of carbacholinduced 40 Hz gamma oscillations and the effects of NMDA activation on oscillatory dynamics in a model of the CA3 subfield of the hippocampus. Journal of Theoretical Biology, 548(??):??, September 7, 2022. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519322001989.

Mahmoodi:2023:NIP

[MPA23] Mohammad Mahmoodi, Ahmadreza Pishevar, and Farzaneh Azargoshasbi. Numerical investigation of the pharmacokinetics and pharmacodynamics of the chemotherapeutic drug in avascular and vascular stages of a brain tumor. Journal of Theoretical Biology, 575(??):??, November 7, 2023. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519323002308.

Morsky:2024:CRU

[MPA24] Bryce Morsky, Joshua B. Plotkin, and Erol Akçay. Convergence of reputations under indirect reciprocity. Journal of Theoretical Biology, 595(??):??, December 7, 2024. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0022519324002327.

Mitchinson:2024:SMT

[MPC^{+}24] A. J. Mitchinson, M. Pogson, G. Czanner, D. Conway, R. R. Wilkinson, M. F. Murphy, I. Siekmann, and S. D. Webb. A stochastic model for topographically influenced cell migration. Journal of Theoretical Biology, 581(??):??, March 21, 2024. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519324000262.

Malige:2022:MML

[MPH^{+}22] Franck Malige, Julie Patris, Maxime Hauray, Pascale Giraudet, and Hervé Glotin. Mathematical models of long term evolution of blue whale song types’ frequencies. Journal of Theoretical Biology, 548(??):??, September 7, 2022. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322001825.

Marrero-Ponce:2020:LBG

[MPTCT^{+}20] Yovani Marrero-Ponce, Julio E. Teran, Ernesto Contreras-Torres, César R. García-Jacas, Yunierkis Perez-Castillo, Nestor Cubillan, Facundo Peréz-Giménez, and José R. Valdés-Martini. LEGO-based generalized set of two linear algebraic 3D bio-macro-molecular descriptors: Theory and validation by QSARs. Journal of Theoretical Biology, 485(??):??, January 21, 2020. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519319304084.

Marcou:2024:EEP

[MRK24] Thomas Marcou, Tomás A. Revilla, and Vlastimil Krivan. Evolutionary emergence of plant and pollinator polymorphisms in consumer-resource mutualisms. Journal of Theoretical Biology, 594(??):??, November 7, 2024. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519324001954.

Mirchi:2020:EDT

[MS20] Pedram Mirchi and M. Soltani. Estimation of drug and tumor properties using novel hybrid meta-heuristic methods. Journal of Theoretical Biology, 488(??):??, March 7, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319304904.

Murgel:2025:ESS

[MS25] Felipe A. Murgel and Max O. Souza. Evolutionarily stable strategies in population games: an invader’s perspective. Journal of Theoretical Biology, 607(??):??, June 21, 2025. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519325000748.

Mandal:2024:NEP

[MSB24] Arindam Mandal, Nazmul Sk, and Saswati Biswas. Nutrient enrichment and phytoplankton toxicity influence a diversity of complex dynamics in a fear-induced plankton-fish model. Journal of Theoretical Biology, 578(??):??, February 7, 2024. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519323002953.

Mehra:2022:HDP

[MSK^{+}22] Somya Mehra, Eva Stadler, David Khoury, James M. Mc-Caw, and Jennifer A. Flegg. Hypnozoite dynamics for Plasmodium vivax malaria: the epidemiological effects of radical cure. Journal of Theoretical Biology, 537(??):??, March 21, 2022. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519322000121.

Matsufuji:2023:OSI

[MSL23] Taiga Matsufuji and Sungrim Seirin-Lee. The optimal strategy of incompatible insect technique (IIT) using Wolbachia and the application to malaria control. Journal of Theoretical Biology, 569(??):??, July 21, 2023. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519323001157.

Magalie:2023:ODS

[MSLW23] Andreea Magalie, Daniel A. Schwartz, Jay T. Lennon, and Joshua S. Weitz. Optimal dormancy strategies in fluctuating environments given delays in phenotypic switching. Journal of Theoretical Biology, 561(??):??, March 21, 2023. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519323000097.

McHuron:2023:IPA

[MSM23] Elizabeth A. McHuron, Jeremy T. Sterling, and Marc Mangel. The influence of prey availability on behavioral decisions and reproductive success of a central-place forager during lactation. Journal of Theoretical Biology, 560(??):??, March 7, 2023. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519322003836.

Marick:2024:HDP

[MTB24] Sounov Marick, Fugo Takasu, and Nandadulal Bairagi. How do productivity gradient and diffusion shape patterns in a plant-herbivore grazing system? Journal of Theoretical Biology, 590(??):??, August 7, 2024. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519324001371.

Marcus:2025:EVI

[MTB25] Stav Marcus, Ari M. Turner, and Guy Bunin. Extinctions as a vestige of instability: the geometry of stability and feasibility. Journal of Theoretical Biology, 608(??):??, July 7, 2025. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519325001079.

Mei:2022:EGD

[MTLZ22] Jie Mei, Yi Tao, Cong Li, and Xiu-Deng Zheng. Evolutionary game dynamics with non-uniform interaction rates in finite population. Journal of Theoretical Biology, 540(??):??, May 7, 2022. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519322000844.

Magero:2024:RBM

[MUMS24] Eileen Joan Magero, Koichi Unami, Osama Mohawesh, and Marie Sato. Resource budget model with Duffing oscillator for dynamics of synchronized biennial-bearing olives in the Levant. Journal of Theoretical Biology, 595(??):??, December 7, 2024. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0022519324002583.

Murase:2025:CPS

[Mur25] Yohsuke Murase. Costly punishment sustains indirect reciprocity under low defection detectability. Journal of Theoretical Biology, 600(??):??, March 7, 2025. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0022519325000098.

Mitchell:2020:PHB

[MW20] Evan Mitchell and Geoff Wild. Prophylactic host behaviour discourages pathogen exploitation. Journal of Theoretical Biology, 503(??):??, October 21, 2020. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519320302435.

Mullon:2021:CDG

[MWO21] Charles Mullon, Joe Yuichiro Wakano, and Hisashi Ohtsuki. Coevolutionary dynamics of genetic traits and their longterm extended effects under non-random interactions. Journal of Theoretical Biology, 525(??):??, September 21, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321001727.

Murray:2020:NCD

[MY20] Ryan Murray and Glenn Young. Neutral competition in a deterministically changing environment: Revisiting continuum approaches. Journal of Theoretical Biology, 486(??):??, February 7, 2020. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519319304734.

Musa:2020:MML

[MZG^{+}20] Salihu S. Musa, Shi Zhao, Daozhou Gao, Qianying Lin, Gerardo Chowell, and Daihai He. Mechanistic modelling of the large-scale Lassa fever epidemics in Nigeria from 2016 to 2019. Journal of Theoretical Biology, 493(??):??, May 21, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519320300643.

Meng:2020:NAT

[MZLJ20] Jingbo Meng, Wensheng Zhu, Canhui Li, and Kyongson Jon. A novel association test for rare variants based on algebraic statistics. Journal of Theoretical Biology, 493(??):??, May 21, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519320300837.

Nahalkova:2020:MMA

[Nah20] Jarmila Nahálková. The molecular mechanisms associated with PIN7, a protein-protein interaction network of seven pleiotropic proteins. Journal of Theoretical Biology, 487(??):??, February 21, 2020. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251931930493X.

Nakamasu:2025:PSL

[Nak25] Akiko M. Nakamasu. Peripheral straightness leads to shape diversification during formations of entire leaves. Journal of Theoretical Biology, 597(??):??, January 21, 2025. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519324002753.

Nah:2020:IIV

[NAR^{+}20] Kyeongah Nah, Mahnaz Alavinejad, Ashrafur Rahman, Jane M. Heffernan, and Jianhong Wu. Impact of influenza vaccine-modified infectivity on attack rate, case fatality ratio and mortality. Journal of Theoretical Biology, 492(??):

??, May 7, 2020. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519320300400.

Nitschke:2023:EBS

[NBBR23] Matthew C. Nitschke, Andrew J. Black, Pierrick Bourrat, and Paul B. Rainey. The effect of bottleneck size on evolution in nested Darwinian populations. Journal of Theoretical Biology, 561(??):??, March 21, 2023. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519323000103.

Nakajima:2020:CSD

[NCL20] Rafael Takahiro Nakajima, Pedro Rafael Costa, and Ney Lemke. Cooperative and sequence-dependent model for RNAP dynamics: Application to ribosomal gene transcription. Journal of Theoretical Biology, 488(??):??, March 7, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251931930503X.

Nicolis:2022:EIA

[ND22] Stamatios C. Nicolis and Jean-Louis Deneubourg. The effect of idiosyncrasy on aggregation in group-living organisms. Journal of Theoretical Biology, 542(??):??, June 7, 2022. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519322001187.

Noffel:2024:MBE

[ND24] Zakarya Noffel and Hana M. Dobrovolny. Modeling the bystander effect during viral coinfection. Journal of Theoretical Biology, 594(??):??, November 7, 2024. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519324002133.

Nisal:2020:PMV

[NDB20] Apoorva Nisal, Urmila Diwekar, and Vibha Bhalerao. Personalized medicine for in vitro fertilization procedure using modeling and optimal control. Journal of Theoretical Biology, 487(??):??, February 21, 2020. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519319304746.

Newman:2020:CDW

[New20] Stuart A. Newman. Cell differentiation: What have we learned in 50 years? Journal of Theoretical Biology, 485(??):??, January 21, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0022519319304011.

Ngonghala:2021:HCS

[NGKB21] Calistus N. Ngonghala, Palak Goel, Daniel Kutor, and Samit Bhattacharyya. Human choice to self-isolate in the face of the COVID-19 pandemic: a game dynamic modelling approach. Journal of Theoretical Biology, 521(??):??, July 21, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519321001144.

Ngonghala:2022:AII

[Ngo22] Calistus N. Ngonghala. Assessing the impact of insecticidetreated nets in the face of insecticide resistance on malaria control. Journal of Theoretical Biology, 555(??):??, December 21, 2022. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519322002727.

Nemati:2023:CPE

[NKE23] Hossein Nemati, Kamran Kaveh, and Mohammad Reza Ejtehadi. Counterintuitive properties of evolutionary measures: a stochastic process study in cyclic population structures with periodic environments. Journal of Theoretical Biology, 564(??):??, May 7, 2023. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519323000322.

Noel:2023:PII

[NM23] Frédérique Noël and Benjamin Mauroy. Propagation of an idealized infection in an airway tree, consequences of the inflammation on the oxygen transfer to blood. Journal of Theoretical Biology, 561(??):??, March 21, 2023. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519323000012.

Naim:2021:SMI

[NMC^{+}21] Wan Naimah Wan Ab Naim, Mohd Jamil Mohamed Mokhtarudin, Bee Ting Chan, Einly Lim, Azam Ahmad Bakir, and Nik Abdullah Nik Mohamed. The study of myocardial ischemia-reperfusion treatment through computational modelling. Journal of Theoretical Biology, 509(??):??, January 21, 2021. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519320303829.

Ni:2020:MRG

[NNG20] Junjun Ni, Charles Wang Wai Ng, and Yufeng Gao. Modelling root growth and soil suction due to plant competition. Journal of Theoretical Biology, 484(??):??, January 7, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519319303893.

Nguyen-Ngoc:2024:CBT

[NNNHA24] Doanh Nguyen-Ngoc, Tri Nguyen-Huu, and Pierre Auger. Connectivity between two fishing sites can lead to an emergence phenomenon related to Maximum Sustainable Yield. Journal of Theoretical Biology, 595(??):??, December 7, 2024. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S002251932400198X.

Nasab:2023:MSM

[NNOBS23] Mohammad Salar Ghasemi Nasab, Hanieh Niroomand-Oscuii, Hossein Bazmara, and Majid Soltani. Multi-scale model of lumen formation via inverse membrane blebbing mechanism during sprouting angiogenesis process. Journal of Theoretical Biology, 556(??):??, January 7, 2023. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322003034.

Nishihara:2024:MPT

[NO24] Shin Nishihara and Toru Ohira. The mechanism of pattern transitions between formation and dispersion. Journal of Theoretical Biology, 581(??):??, March 21, 2024. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519324000171.

Nakano:2025:CEC

[NO25] Raiki Nakano and Hisashi Ohtsuki. Conditions for the establishment of creole languages from an evolutionary game theoretic perspective. Journal of Theoretical Biology, 605(??):??, May 21, 2025. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519325000566.

Nourani:2020:AEB

[NR20] Esmaeil Nourani and Vahideh Reshadat. Association extraction from biomedical literature based on representation and transfer learning. Journal of Theoretical Biology, 488(??):??, March 7, 2020. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519319304813.

Nugent:2022:ERP

[NSD22] Andrew Nugent, Emma Southall, and Louise Dyson. Exploring the role of the potential surface in the behaviour of early warning signals. Journal of Theoretical Biology, 554(??):??, December 7, 2022. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519322002600.

Nava-Sedeno:2025:OMI

[NSMH25] Josué Manik Nava-Sedeño, Abraham Martínez, and Haralampos Hatzikirou. On the onset of multicellular invasive behavior in hierarchical lineage: the role of inhibitory feedback and local fluctuations. Journal of Theoretical Biology, 613(??):??, October 7, 2025. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519325001869.

Nicol:2025:ANE

[NSZ^{+}25] Matthew Nicol, Julian D. J. Sng, Yanshan Zhu, Sissy Therese Sonnleitner, Kirsty R. Short, and Meagan Carney. An analytical, numerical and experimental study of in-vitro SARS-CoV-2 evolution in Vero B4 cells. Journal of Theoretical Biology, 598(??):??, February 7, 2025. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519324002856.

Nishiyama:2022:MEL

[NTTI22] Takara Nishiyama, Toru Takada, Hiroaki Takeuchi, and Shingo Iwami. Maternal embryonic leucine zipper kinase (MELK) optimally regulates the HIV-1 uncoating process. Journal of Theoretical Biology, 545(??):??, July 21, 2022. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322001503.

Nguyen:2022:TGE

[NV22] Vu A. T. Nguyen and Dervis Can Vural. Theoretical guidelines for editing ecological communities. Journal of Theoretical Biology, 534(??):??, February 7, 2022. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321003647.

Nematollahi:2020:NAC

[NVS20] Mohammad Hassan Nematollahi, Ramin Vatankhah, and Mojtaba Sharifi. Nonlinear adaptive control of tuberculosis with consideration of the risk of endogenous reactivation and exogenous reinfection. Journal of Theoretical Biology, 486(??):??, February 7, 2020. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519319304503.

Northrup:2024:EDH

[NWP^{+}24] Graham R. Northrup, Andy White, Steven R. Parratt, Carly Rozins, Anna-Liisa Laine, and Mike Boots. The evolutionary dynamics of hyperparasites. Journal of Theoretical Biology, 582(??):??, April 7, 2024. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519324000225.

Nikolov:2020:RCG

[NWVN20] Svetoslav Nikolov, Olaf Wolkenhauer, Julio Vera, and Momchil Nenov. The role of cooperativity in a p53– miR34 dynamical mathematical model. Journal of Theoretical Biology, 495(??):??, June 21, 2020. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519320301077.

Osojnik:2020:ICI

[OGD^{+}20] Ana Osojnik, Eamonn A. Gaffney, Michael Davies, James W. T. Yates, and Helen M. Byrne. Identifying and characterising the impact of excitability in a mathematical model of tumour-immune interactions. Journal of Theoretical Biology, 501(??):??, September 21, 2020. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519320301053.

Oliver:2025:UQC

[OGM^{+}25] Mary Chriselda Antony Oliver, Matthew Graham, Ioanna Manolopoulou, Graham F. Medley, Lorenzo Pellis, Koen B. Pouwels, Matthew Thorpe, and T. Deirdre Hollingsworth. Uncertainty quantification in cost-effectiveness analysis for stochastic-based infectious disease models: Insights from surveillance on lymphatic filariasis. Journal of Theoretical Biology, 611(??):??, August 21, 2025. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519325001638.

Omori:2024:HMA

[OIK^{+}24] Ryosuke Omori, Koichi Ito, Shunsuke Kanemitsu, Ryusuke Kimura, and Yoh Iwasa. Human movement avoidance decisions during Coronavirus disease 2019 in Japan. Journal of Theoretical Biology, 585(??):??, May 21, 2024. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519324000766.

Oda:2021:QAE

[OKF^{+}21] Takafumi Oda, Kwang Su Kim, Yasuhisa Fujita, Yusuke Ito, Tomoyuki Miura, and Shingo Iwami. Quantifying antiviral effects against simian/human immunodeficiency virus induced by host immune response. Journal of Theoretical Biology, 509(??):??, January 21, 2021. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519320303489.

OConnell:2021:CFS

[OKT^{+}21] Delaney O’Connell, Catherine E. Kehl, Brian K. Taylor, Joseph Piacenza, Susan Piacenza, and Kenneth John Faller II. A computational framework for studying energetics and resource management in sea turtle migration and autonomous systems. Journal of Theoretical Biology, 527(??):??, October 21, 2021. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519321002344.

Omta:2024:MTD

[OLF^{+}24] Anne Willem Omta, Justin D. Liefer, Zoe V. Finkel, Andrew J. Irwin, Daniel Sher, and Michael J. Follows. A model of time-dependent macromolecular and elemental composition of phytoplankton. Journal of Theoretical Biology, 592(??):??, September 7, 2024. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251932400167X.

Olejarz:2024:GDE

[ON24] Jason W. Olejarz and Martin A. Nowak. Gene drives for the extinction of wild metapopulations. Journal of Theoretical Biology, 577(??):??, January 21, 2024. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519323002515.

Onifade:2024:DGL

[ORT24] Akindele Akano Onifade, Jan Rychtár, and Dewey Taylor. A dynamic game of lymphatic filariasis prevention by voluntary use of insecticide treated nets. Journal of Theoretical Biology, 585(??):??, May 21, 2024. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519324000778.

Ohtsuki:2020:CDD

[ORW^{+}20] Hisashi Ohtsuki, Claus Rueffler, Joe Yuichiro Wakano, Kalle Parvinen, and Laurent Lehmann. The components of directional and disruptive selection in heterogeneous group-structured populations. Journal of Theoretical Biology, 507(??):??, December 21, 2020. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519320303040.

Otero:2021:TMP

[OSAR21] Marcelo Otero, Silvina N. Sarno, Sofía L. Acebedo, and Javier A. Ramírez. Tracing molecular properties throughout evolution: a chemoinformatic approach. Journal of Theoretical Biology, 515(??):??, April 21, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321000230.

Osborne:2024:ANM

[Osb24] James M. Osborne. An adaptive numerical method for multicellular simulations of tissue development and maintenance. Journal of Theoretical Biology, 594(??):??, November 7, 2024. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519324002078.

Oster:2023:LMJ

[Ost23] Andrew M. Oster. A laminar model for the joint development of ocular dominance columns and CO blobs in the primary visual cortex. Journal of Theoretical Biology, 572(??):??, September 7, 2023. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519323001856.

Ouyang:2024:SON

[OTCE24] Bowei Ouyang, Aaron C. True, John P. Crimaldi, and Bard Ermentrout. Simple olfactory navigation in air and water. Journal of Theoretical Biology, 595(??):??, December 7, 2024. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0022519324002261.

Omta:2020:QNT

[OTI^{+}20] A. W. Omta, D. Talmy, K. Inomura, A. J. Irwin, Z. V. Finkel, D. Sher, J. D. Liefer, and M. J. Follows. Quantifying nutrient throughput and DOM production by algae in continuous culture. Journal of Theoretical Biology, 494(??):??, June 7, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519320300692.

Ohtsuki:2022:ISD

[OU22] Hisashi Ohtsuki and Shumpei Ujiyama. Impact of social dominance on the evolution of individual learning. Journal of Theoretical Biology, 535(??):??, February 21, 2022. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321004069.

OCallaghan:2025:SIT

[OV25] Alan O’Callaghan and Catalina A. Vallejos. Scalable inference of transcriptional variability with BASiCS. Journal of Theoretical Biology, 611(??):??, August 21, 2025. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519325001237.

Okabe:2021:TSM

[OY21] Takuya Okabe and Jin Yoshimura. Two-step mechanism of spiral phyllotaxis. Journal of Theoretical Biology, 508(??):??, January 7, 2021. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519320303398.

Ohsaki:2025:SAP

[OYT25] Haruna Ohsaki, Akira Yamawo, and Yuuya Tachiki. The spatial aggregation of phytophagous insects driven by the evolution of preference for plant chemicals. Journal of Theoretical Biology, 605(??):??, May 21, 2025. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519325000608.

Pattni:2023:EED

[PABS23] Karan Pattni, Wajid Ali, Mark Broom, and Kieran J. Sharkey. Eco-evolutionary dynamics in finite networkstructured populations with migration. Journal of Theoretical Biology, 572(??):??, September 7, 2023. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519323001844.

Pagano:2025:SPA

[Pag25] Ugo Pagano. Selfish parents, altruistic children and helpers’ infertility. Journal of Theoretical Biology, 611(??):??, August 21, 2025. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519325001560.

Plaugher:2022:UPI

[PAM22] Daniel Plaugher, Boris Aguilar, and David Murrugarra. Uncovering potential interventions for pancreatic cancer patients via mathematical modeling. Journal of Theoretical Biology, 548(??):??, September 7, 2022. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322001953.

Pandula:2024:IRA

[PAM24] Neel Pandula, Erol Akçay, and Bryce Morsky. Indirect reciprocity with abductive reasoning. Journal of Theoretical Biology, 580(??):??, March 7, 2024. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519323003120.

Park:2022:ECB

[Par22] Peter S. Park. The evolution of cognitive biases in human learning. Journal of Theoretical Biology, 541(??):??, May 21, 2022. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519322000297.

Pedersen:2021:MMH

[PASO21] Rasmus Kristoffer Pedersen, Morten Andersen, Thomas Stiehl, and Johnny T. Ottesen. Mathematical modelling of the hematopoietic stem cell-niche system: Clonal dominance based on stem cell fitness. Journal of Theoretical Biology, 518(??):??, June 7, 2021. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519321000424.

Patsatzis:2022:AAA

[Pat22] Dimitrios G. Patsatzis. Algorithmic asymptotic analysis: Extending the arsenal of cancer immunology modeling. Journal of Theoretical Biology, 534(??):??, February 7, 2022. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321003957.

Pinti:2021:CAB

[PAV21] Jérôme Pinti, Ken H. Andersen, and André W. Visser. Coadaptive behavior of interacting populations in a habitat selection game significantly impacts ecosystem functions. Journal of Theoretical Biology, 523(??):??, August 21, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321000850.

Phillips:2021:NSM

[PB21] Brendon Phillips and Chris T. Bauch. Network structural metrics as early warning signals of widespread vaccine refusal in social-epidemiological networks. Journal of Theoretical Biology, 531(??):??, December 21, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321003003.

Picoche:2022:SBC

[PB22] Coralie Picoche and Frédéric Barraquand. Seed banks can help to maintain the diversity of interacting phytoplankton species. Journal of Theoretical Biology, 538(??):??, April 7, 2022. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519322000182.

Piltz:2020:HSD

[PBB20] Sofia H. Piltz, Cecilia G. Diniz Behn, and Victoria Booth. Habitual sleep duration affects recovery from acute sleep deprivation: a modeling study. Journal of Theoretical Biology, 504(??):??, November 7, 2020. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519320302563.

Pitcher:2020:MEE

[PBD^{+}20] Michael J. Pitcher, Ruth Bowness, Simon Dobson, Raluca Eftimie, and Stephen H. Gillespie. Modelling the effects of environmental heterogeneity within the lung on the tuberculosis life-cycle. Journal of Theoretical Biology, 506(??):??, December 7, 2020. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519320302368.

Papa:2023:TGC

[PBP23] Federico Papa, Alessandro Borri, and Pasquale Palumbo. Tumour growth control: analysis of alternative approaches. Journal of Theoretical Biology, 562(??):??, April 7, 2023. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519323000164.

Pan:2021:RVS

[PCLL21] Jichao Pan, Yan Cai, Mengchen Liu, and Zhiyong Li. Role of vascular smooth muscle cell phenotypic switching in plaque progression: a hybrid modeling study. Journal of Theoretical Biology, 526(??):??, October 7, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321002137.

Palmer:2021:AAE

[PCR21] Ryan A. Palmer, Isaac V. Chenchiah, and Daniel Robert. Analysis of aerodynamic and electrostatic sensing in mechanoreceptor arthropod hairs. Journal of Theoretical Biology, 530(??):??, December 7, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0022519321002903.

Palmer:2023:PET

[PCR23] Ryan A. Palmer, Isaac V. Chenchiah, and Daniel Robert. Passive electrolocation in terrestrial arthropods: Theoretical modelling of location detection. Journal of Theoretical Biology, 558(??):??, February 7, 2023. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519322003484.

Padmanabhan:2023:MHI

[PD23] Pranesh Padmanabhan and Narendra M. Dixit. Modelling how increased Cathepsin B/L and decreased TMPRSS2 usage for cell entry by the SARS-CoV-2 Omicron variant may affect the efficacy and synergy of TMPRSS2 and Cathepsin B/L inhibitors. Journal of Theoretical Biology, 572(??):??, September 7, 2023. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519323001650.

Polavarapu:2024:ETS

[PDD24] Nishnath Polavarapu, Madison Doty, and Hana M. Dobrovolny. Exploring the treatment of SARS-CoV-2 with modified vesicular stomatitis virus. Journal of Theoretical Biology, 595(??):??, December 7, 2024. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0022519324002443.

Portet:2022:INR

[PEMLD22] Stéphanie Portet, Sandrine Etienne-Manneville, Cécile Leduc, and J. C. Dallon. Impact of noise on the regulation of intracellular transport of intermediate filaments. Journal of Theoretical Biology, 547(??):??, August 21, 2022. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519322001813.

Pfeuty:2022:MTB

[Pfe22] Benjamin Pfeuty. Multistability and transitions between spatiotemporal patterns through versatile Notch–Hes signaling. Journal of Theoretical Biology, 539(??):??, April 21, 2022. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322000583.

Pozzi:2022:CTA

[PGC^{+}22] G. Pozzi, B. Grammatica, L. Chaabane, M. Catucci, A. Mondino, P. Zunino, and P. Ciarletta. T cell therapy against cancer: a predictive diffuse-interface mathematical model informed by pre-clinical studies. Journal of Theoretical Biology, 547(??):??, August 21, 2022. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322001709.

Ponce:2020:QAE

[PGML20] Luis F. Ponce, Karina García-Martínez, and Kalet León. Quantitative assessment of extracellular IL-1 regulation. Journal of Theoretical Biology, 487(??):??, February 21, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319304825.

Padilla:2023:MRL

[PGP23] Dustin G. Padilla, Xiaoqian Gong, and Charles Perrings. Modeling the role of land conversion on the spread of an epizootic disease. Journal of Theoretical Biology, 557(??):??, January 21, 2023. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519322003150.

Pei:2023:DBM

[PHL^{+}23] Yongzhen Pei, Siqi Han, Changguo Li, Jinzhi Lei, and Fengxi Wen. Data-based modeling of breast cancer and optimal therapy. Journal of Theoretical Biology, 573(??):??, September 21, 2023. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251932300190X.

Park:2023:VGH

[PHN^{+}23] Hye Jin Park, Christian Hilbe, Martin A. Nowak, Beom Jun Kim, and Hyeong-Chai Jeong. Vacancies in growing habitats promote the evolution of cooperation. Journal of Theoretical Biology, 575(??):??, November 7, 2023. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519323002266.

Pierson:2021:UGR

[Pie21] Kawika Pierson. Using geometry to rank evenness measures: Towards a deeper understanding of divergence. Journal of Theoretical Biology, 531(??):??, December 21, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321003210.

Piskovsky:2025:EPP

[Pis25] Vit Piskovsky. Evolution of predators and prey kills Turing patterns. Journal of Theoretical Biology, 607(??):??, June 21, 2025. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519325000736.

Pramanik:2021:MAR

[PJB21] D. Pramanik, M. K. Jolly, and R. Bhat. Matrix adhesion and remodeling diversifies modes of cancer invasion across spatial scales. Journal of Theoretical Biology, 524(??):??, September 7, 2021. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519321001557.

Pugh:2025:SPR

[PJPH25] Kira Pugh, Rhys D. O. Jones, Gibin Powathil, and Sara Hamis. Simulations probe the role of space in the interplay between drug-sensitive and drug-resistant cancer cells. Journal of Theoretical Biology, 602–603(??):??, April 7, 2025. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519325000141.

Pandey:2023:MIC

[PJVR23] Rakesh Pandey, Amit Jangid, Ramalakshmi Gayathri Vinjamuri, and Ramakrishna Ramaswamy. Modelling of indirect cell-cell interaction networks mediated by IFNγ / IL-4 cytokine involved in atopic dermatitis. Journal of Theoretical Biology, 556(??):??, January 7, 2023. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251932200282X.

Plazen:2023:EDM

[PK23] Lucie Plazen and Anmar Khadra. Excitable dynamics in a molecularly-explicit model of cell motility: Mixed-mode oscillations and beyond. Journal of Theoretical Biology, 564(??): ??, May 7, 2023. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519323000462.

Posner:2020:CMM

[PL20] Russell Posner and Reinhard Laubenbacher. The contribution of microRNA-mediated regulation to short- and long-term gene expression predictability. Journal of Theoretical Biology, 486(??):??, February 7, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519319304242.

Parisey:2022:OSM

[PLAC22] Nicolas Parisey, Melen Leclerc, and Katarzyna Adamczyk-Chauvat. Optimal spatial monitoring of populations described by reaction–diffusion models. Journal of Theoretical Biology, 534(??):??, February 7, 2022. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519321003969.

Peyhardi:2024:PSD

[PLM24] Jean Peyhardi, Fabien Laroche, and Frédéric Mortier. Pólyasplitting distributions as stationary solutions of multivariate birth-death processes under extended neutral theory. Journal of Theoretical Biology, 582(??):??, April 7, 2024. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519324000365.

Pontarp:2024:SED

[PLR24] Mikael Pontarp, Per Lundberg, and Jörgen Ripa. The succession of ecological divergence and reproductive isolation in adaptive radiations. Journal of Theoretical Biology, 587(??): ??, June 21, 2024. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519324001000.

Pattanaik:2024:ELE

[PLSS24a] Deepak K. Pattanaik, Vasudevan Lakshminarayanan, Nachieketa K. Sharma, and Amir Prasad Sahu. Erratum to “Leading edge of the a-wave of the electroretinogram and sodium iodateinduced age-related macular degeneration: a Model” [J. Theor. Biol. 592 (2024) 01-17/111879]. Journal of Theoretical Biology, 593(??):??, October 7, 2024. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519324001978. See [PLSS24b].

Pattanaik:2024:LEW

[PLSS24b] Deepak K. Pattanaik, Vasudevan Lakshminarayanan, Nachieketa K. Sharma, and Amir Prasad Sahu. Leading edge of the awave of the electroretinogram and sodium iodate-induced age-related macular degeneration: a model. Journal of Theoretical Biology, 592(??):??, September 7, 2024. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519324001632. See erratum [PLSS24a].

Paquin-Lefebvre:2020:PFC

[PLXD^{+}20] Frédéric Paquin-Lefebvre, Bin Xu, Kelsey L. DiPietro, Alan E. Lindsay, and Alexandra Jilkine. Pattern formation in a coupled membrane-bulk reaction–diffusion model for intracellular polarization and oscillations. Journal of Theoretical Biology, 497(??):??, July 21, 2020. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519320300977.

Paredes:2023:LEA

[PMCLC^{+}23] Omar Paredes, Elsa Magaña-Cuevas, Sorpresa Lora-Castro, Juan D. Serna-Grilló, D. Y. De la Mora, Diana Ríos Patiño, Rebeca Romo-Vázquez, and J. Alejandro Morales. A letter to the editor about the article “The modularity codes”. Journal of Theoretical Biology, 559(??):??, February 21, 2023. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322003605.

Parvinen:2021:WED

[PMD21] Kalle Parvinen, Johan A. J. Metz, and Ulf Dieckmann. WITHDRAWN: Environmental dimensionality determines species coexistence. Journal of Theoretical Biology, 526 (??):??, October 7, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519320301351.

Pizarro:2020:DCW

[PO20] Gonzalo Pizarro and J. Fernando Olivera. The dynamics of Ca ^{2+} within the sarcoplasmic reticulum of frog skeletal muscle. A simulation study. Journal of Theoretical Biology, 504(??):??, November 7, 2020. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519320302265.

Pattni:2021:EGT

[POS21] Karan Pattni, Christopher E. Overton, and Kieran J. Sharkey. Evolutionary graph theory derived from ecoevolutionary dynamics. Journal of Theoretical Biology, 519 (??):??, June 21, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519321000709.

Parvinen:2023:EDU

[POW23] Kalle Parvinen, Hisashi Ohtsuki, and Joe Yuichiro Wakano. Evolution of dispersal under spatio-temporal heterogeneity. Journal of Theoretical Biology, 574(??):??, October 7, 2023. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519323002096.

Pozzobon:2020:MNW

[PP20] Victor Pozzobon and Patrick Perré. Multiscale numerical workflow describing microalgae motion and light pattern incidence towards population growth in a photobioreactor. Journal of Theoretical Biology, 498(??):??, August 7, 2020. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251932030148X. See correction [PP24].

Perret:2022:IRL

[PP22] Cedric Perret and Simon T. Powers. An investigation of the role of leadership in consensus decision-making. Journal of Theoretical Biology, 543(??):??, June 21, 2022. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322000923.

Pozzobon:2024:CMN

[PP24] Victor Pozzobon and Patrick Perré. Corrigendum to “Multiscale numerical workflow describing microalgae motion and light pattern incidence towards population growth in a photobioreactor” [J. Theoret. Biol. 498 (2020) 110293]. Journal of Theoretical Biology, 581(??):??, March 21, 2024. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519324000237. See [PP20].

Pak:2024:MFE

[PPFB24] Thomas F. Pak, Joe Pitt-Francis, and Ruth E. Baker. A mathematical framework for the emergence of winners and losers in cell competition. Journal of Theoretical Biology, 577(??):??, January 21, 2024. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519323002631.

Peng:2020:EFA

[PPGB20] Zhen Peng, Alex M. Plum, Praful Gagrani, and David A. Baum. An ecological framework for the analysis of prebiotic chemical reaction networks. Journal of Theoretical Biology, 507(??):??, December 21, 2020. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519320303064.

Panday:2021:DSS

[PPS^{+}21] Pijush Panday, Nikhil Pal, Sudip Samanta, Piotr Tryjanowski, and Joydev Chattopadhyay. Dynamics of a stage-structured predator–prey model: cost and benefit of fear-induced group defense. Journal of Theoretical Biology, 528(??):??, November 7, 2021. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519321002654.

Patel:2021:OEW

[PR21] Mainak Patel and Aaditya Rangan. Olfactory encoding within the insect antennal lobe: the emergence and role of higher order temporal correlations in the dynamics of antennal lobe spiking activity. Journal of Theoretical Biology, 522(??):??, August 7, 2021. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519321001223.

Parra-Rojas:2020:QMF

[PRFSHV20] César Parra-Rojas, Boris Fürtig, Harald Schwalbe, and Esteban A. Hernandez-Vargas. Quantitative modeling of the function of kinetically driven transcriptional riboswitches. Journal of Theoretical Biology, 506(??):??, December 7, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519320302617.

Perez:2020:MGD

[PRHS20] Pablo Pérez, Alfonso Ruiz-Herrera, and Ana María San Luis. Management guidelines in disturbance-prone populations: the importance of the intervention time. Journal of Theoretical Biology, 486(??):??, February 7, 2020. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319304448.

Payandeh:2020:DEA

[PRJ^{+}20] Zahra Payandeh, Mohammad Reza Rahbar, Abolfazl Jahangiri, Zahra Sadat Hashemi, Alireza Zakeri, Moslem Jafarisani, Mohammad Javad Rasaee, and Saeed Khalili. Design of an engineered ACE2 as a novel therapeutics against COVID-19. Journal of Theoretical Biology, 505(??):??, November 21, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0022519320302800.

Proverbio:2024:CHE

[Pro24] Daniele Proverbio. Chemotaxis in heterogeneous environments: a multi-agent model of decentralized gathering past obstacles. Journal of Theoretical Biology, 586(??):??, June 7, 2024. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519324001012.

Palacios:2020:MBH

[PRP^{+}20] Ensor Rafael Palacios, Adeel Razi, Thomas Parr, Michael Kirchhoff, and Karl Friston. On Markov blankets and hierarchical self-organisation. Journal of Theoretical Biology, 486(??):??, February 7, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0022519319304588.

Prusinkiewicz:2025:PTB

[Pru25] Przemyslaw Prusinkiewicz. Pillars of theoretical biology: ‘Mathematical models for cellular interaction in development, I and II’. Journal of Theoretical Biology, 609(??):??, July 21, 2025. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519325001080.

Pande:2022:HTE

[PS22a] Jayant Pande and Nadav M. Shnerb. How temporal environmental stochasticity affects species richness: Destabilization, neutralization and the storage effect. Journal of Theoretical Biology, 539(??):??, April 21, 2022. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322000510.

Pouryahya:2022:PSU

[PS22b] Fatemeh Pouryahya and David Sankoff. Peripheral structures in unlabelled trees and the accumulation of subgenomes in the evolution of polyploids. Journal of Theoretical Biology, 532(??):??, January 7, 2022. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251932100343X.

Piephoff:2024:BTM

[PTK^{+}24] Faye Piephoff, Brian K. Taylor, Catherine E. Kehl, Bruno Mota, and Cynthia M. Harley. Biomechanics of transduction by mechanosensory cilia for prey detection in aquatic organisms. Journal of Theoretical Biology, 583(??):??, April 21, 2024. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519324000638.

Preziosi:2021:CTG

[PTZ21] Luigi Preziosi, Giuseppe Toscani, and Mattia Zanella. Control of tumor growth distributions through kinetic methods. Journal of Theoretical Biology, 514(??):??, April 7, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321000011.

Padovano:2024:DDR

[PV24] Federica Padovano and Chiara Villa. The development of drug resistance in metastatic tumours under chemotherapy: an evolutionary perspective. Journal of Theoretical Biology, 595(??):??, December 7, 2024. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S002251932400242X.

Pearson:2023:MIV

[PWC^{+}23] Jason Pearson, Timothy Wessler, Alex Chen, Richard C. Boucher, Ronit Freeman, Samuel K. Lai, Raymond Pickles, and M. Gregory Forest. Modeling identifies variability in SARS-CoV-2 uptake and eclipse phase by infected cells as principal drivers of extreme variability in nasal viral load in the 48 h post infection. Journal of Theoretical Biology, 565(??):??, May 21, 2023. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519323000668.

Petrovskaya:2020:WSB

[PZ20] Natalia Petrovskaya and Wenxin Zhang. When seeing is not believing: Comparative study of various spatial distributions of invasive species. Journal of Theoretical Biology, 488(??):??, March 7, 2020. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519319305107.

Poormohammadi:2020:NMC

[PZG20] Hadi Poormohammadi, Mohsen Sardari Zarchi, and Hossein Ghaneai. NCHB: a method for constructing rooted phylogenetic networks from rooted triplets based on height function and binarization. Journal of Theoretical Biology, 489(??):??, March 21, 2020. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519319305144.

Paryad-Zanjani:2023:OHR

[PZJPZ23] Sasan Paryad-Zanjani, Aditya Jagarapu, Michael J. Piovoso, and Ryan Zurakowski. Ongoing HIV replication in lymph node sanctuary sites in treated individuals contributes to the total latent HIV at a very slow rate. Journal of Theoretical Biology, 575(??):??, November 7, 2023. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519323002485.

Pan:2022:RCD

[PZXC22] Yufeng Pan, Hudong Zhang, Yan Xie, and Yuan Chai. Role of coupling distances in a coupled thalamocortical network in the treatment of epilepsy. Journal of Theoretical Biology, 550(??):??, October 7, 2022. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519322002041.

Qu:2023:BMG

[QHS23] Hongchun Qu, Aviad Heifetz, and Merav Seifan. Batesian mimicry or general food deception? An evolutionary game between plants for pollinator services. Journal of Theoretical Biology, 575(??):??, November 7, 2023. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519323002060.

Quevreux:2021:PAF

[QPL21] Pierre Quévreux, Rémi Pigeault, and Michel Loreau. Predator avoidance and foraging for food shape synchrony and response to perturbations in trophic metacommunities. Journal of Theoretical Biology, 528(??):??, November 7, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321002551.

Quillien:2020:ECU

[Qui20] Tadeg Quillien. Evolution of conditional and unconditional commitment. Journal of Theoretical Biology, 492(??):??, May 7, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251932030059X.

Qin:2025:IER

[QVE25] Tianjian Qin, Luis Valente, and Rampal S. Etienne. Impact of evolutionary relatedness on species diversification and tree shape. Journal of Theoretical Biology, 598(??):??, February 7, 2025. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519324002777.

Qian:2024:TVR

[QWZK24] Guangjing Qian, Dan Wu, Lai Zhang, and Susanne Kortsch. Temperature variability regulates the interactive effects of warming and pharmaceutical on aquatic ecosystem dynamics. Journal of Theoretical Biology, 595(??):??, December 7, 2024. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0022519324002339.

Qi:2023:MOM

[QYY^{+}23] Hong Qi, Yu-Song Yin, Zhi-Yong Yin, Xiang Li, and Jian-Wei Shuai. Mitochondrial outer membrane permeabilization and inner membrane permeabilization in regulating apoptosis and inflammation. Journal of Theoretical Biology, 571(??):??, August 21, 2023. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519323001558.

Ravoni:2021:LTB

[Rav21] Alessandro Ravoni. Long-term behaviours of Autocatalytic Sets. Journal of Theoretical Biology, 529(??):??, November 21, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519321002794.

Rahimabadi:2022:CAS

[RB22] Arsalan Rahimabadi and Habib Benali. Comment on the article “Spatially-extended nucleation–aggregation–fragmentation models for the dynamics of prion-like neurodegenerative protein-spreading in the brain and its connectome 486 (2020) 110102”. Journal of Theoretical Biology, 534(??):??, February 7, 2022. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519321003854. See [FSKG20].

Roy:2020:ERT

[RBD^{+}20] Suvam Roy, Niraja V. Bapat, Julien Derr, Sudha Rajamani, and Supratim Sengupta. Emergence of ribozyme and tRNAlike structures from mineral-rich muddy pools on prebiotic earth. Journal of Theoretical Biology, 506(??):??, December 7, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519320303015.

Rostami:2023:ILM

[RBG^{+}23] Minghao W. Rostami, Brittany E. Bannish, Kelsey Gasior, Rebecca L. Pinals, Calina Copos, and Adriana T. Dawes. Inferring local molecular dynamics from the global actin network structure: a case study of 2D synthetic branching actin networks. Journal of Theoretical Biology, 575(??):??, November 7, 2023. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519323002102.

Rudrauf:2020:MIE

[RBW20] David Rudrauf, Daniel Bennequin, and Kenneth Williford. The Moon illusion explained by the projective consciousness model. Journal of Theoretical Biology, 507(??):??, December 21, 2020. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519320303106.

Roques:2024:PPE

[RBZP24] L. Roques, N. Boutillon, P. Zamberletti, and J. Papaïx. Polymorphic population expansion velocity in a heterogeneous environment. Journal of Theoretical Biology, 595(??):??, December 7, 2024. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0022519324002170.

Ranjit:2025:BPF

[RCM^{+}25] Buddhadev Ranjit, Arnab Chattopadhyay, Arindam Mandal, Santosh Biswas, and Joydev Chattopadhyay. Beyond predation: Fish–coral interactions can tip the scales of coral disease. Journal of Theoretical Biology, 599(??):??, February 21, 2025. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519324003163.

Reyne:2025:MET

[RDDSA25] Bastien Reyné, Ramsès Djidjou-Demasse, Mircea T. Sofonea, and Samuel Alizon. Mutant emergence timing and population immunisation status impact epidemiological dynamics. Journal of Theoretical Biology, 608(??):??, July 7, 2025. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519325001067.

Robinson:2025:UAM

[RE25] P. A. Robinson and Rawan Khalil El Zghir. Unification of alpha, mu, and tau rhythms and their beta-band harmonics via eigenmodes: spectral peaks, topography, and reactivity. Journal of Theoretical Biology, 608(??):??, July 7, 2025. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251932500102X.

Renardy:2020:PSW

[REK20] Marissa Renardy, Marisa Eisenberg, and Denise Kirschner. Predicting the second wave of COVID-19 in Washtenaw County, MI. Journal of Theoretical Biology, 507(??):??, December 21, 2020. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519320303167.

Revilla:2025:PPI

[Rev25] Tomás A. Revilla. Plant-pollinator interaction model with separate pollen and nectar dynamics. Journal of Theoretical Biology, 606(??):??, June 7, 2025. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519325000621.

Risau-Gusman:2021:CDP

[RG21] Sebastián Risau-Gusman. Color discrimination properties arising from optimal decoding in the early stages of visual systems. Journal of Theoretical Biology, 526(??):??, October 7, 2021. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519321001958.

Reimert:2025:CLD

[RGBD25] Mossa Merhi Reimert, Maya Katrin Gussmann, Anette Ella Boklund, and Matt Denwood. Choice of landscape discretisation method affects the inferred rate of spread in wildlife disease spread models. Journal of Theoretical Biology, 596(??):??, January 7, 2025. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519324002480.

Ruiz-Herrera:2023:RST

[RH23] Alfonso Ruiz-Herrera. The role of the spatial topology in trophic metacommunities: Species with reduced mobility and total population size. Journal of Theoretical Biology, 566(??): ??, June 7, 2023. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519323000759.

Rubin:2023:ADN

[RID23] Ilan N. Rubin, Yaroslav Ispolatov, and Michael Doebeli. Adaptive diversification and niche packing on rugged fitness landscapes. Journal of Theoretical Biology, 562(??):??, April 7, 2023. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519323000176.

Ranjan:2022:HRS

[RK22a] Ravi Ranjan and Christopher A. Klausmeier. How the resource supply distribution structures competitive communities. Journal of Theoretical Biology, 538(??):??, April 7, 2022. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322000522.

Richards:2022:SPP

[RK22b] A. Richards and L. Kubatko. Site pattern probabilities under the multispecies coalescent and a relaxed molecular clock: Theory and applications. Journal of Theoretical Biology, 542(??):??, June 7, 2022. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519322000765.

Rottschafer:2025:AAC

[RKCvB25] Vivi Rottschäfer, Willem G. N. Kuppers, Jiao Chen, and Ed van Bavel. Arterial arcades and collaterals regress under hemodynamics-based diameter adaptation: a computational and mathematical analysis. Journal of Theoretical Biology, 607(??):??, June 21, 2025. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519325000773.

Raguin:2020:SMC

[RKP20] Adélaïde Raguin, Norbert Kern, and Andrea Parmeggiani. Stochastic modelling of collective motor protein transport through a crossing of microtubules. Journal of Theoretical Biology, 505(??):??, November 21, 2020. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519320302253.

Rostami:2025:DSD

[RLB25] Ati Rostami, Brodie A. J. Lawson, and Kevin Burrage. An in-depth study of the dynamics of Thornley’s mathematical model in plant biology with a view to an improved model. Journal of Theoretical Biology, 604(??):??, May 7, 2025. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519325000372.

Rui:2025:KCV

[RLL25] Yu Rui, Shu Liu, and Suyu Liu. Kir4.1 channel and voltagegated calcium channel of astrocyte account for the transition dynamics of seizures. Journal of Theoretical Biology, 604(??): ??, May 7, 2025. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519325000487.

Ren:2021:RFA

[RLW^{+}21] Hua-Rong Ren, Ming-Tao Li, You-Ming Wang, Zhen Jin, and Juan Zhang. The risk factor assessment of the spread of foot-and-mouth disease in mainland China. Journal of Theoretical Biology, 512(??):??, March 7, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519320304136.

Rose:2021:HSD

[RMG^{+}21] Christopher Rose, Andrew J. Medford, C. Franklin Goldsmith, Tejs Vegge, Joshua S. Weitz, and Andrew A. Peterson. Heterogeneity in susceptibility dictates the order of epidemic models. Journal of Theoretical Biology, 528(??):??, November 7, 2021. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519321002587.

Rouf:2025:PER

[RMS^{+}25] Samiha Rouf, Casey Moore, Debabrata Saha, Dan Nguyen, MaryLena Bleile, Robert Timmerman, Hao Peng, and Steve Jiang. PULSAR Effect: Revealing potential synergies in combined radiation therapy and immunotherapy via differential equations. Journal of Theoretical Biology, 596(??):??, January 7, 2025. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519324002595.

Ravaee:2024:IPD

[RMSS24] Hamid Ravaee, Mohammad Hossein Manshaei, Mehran Safayani, and Javad Salimi Sartakhti. Intelligent phenotypedetection and gene expression profile generation with generative adversarial networks. Journal of Theoretical Biology, 577(??):??, January 21, 2024. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519323002333.

Reckell:2021:MSP

[RNP^{+}21] Trevor Reckell, Kyle Nguyen, Tin Phan, Sharon Crook, Eric J. Kostelich, and Yang Kuang. Modeling the synergistic properties of drugs in hormonal treatment for prostate cancer. Journal of Theoretical Biology, 514(??):??, April 7, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519320304252.

Rajagopal:2025:PSA

[RNP^{+}25] Karthikeyan Rajagopal, Nafise Naseri, Fatemeh Parastesh, Farnaz Ghassemi, and Sajad Jafari. Phase synchronization analysis of EEG functional connectivity in Parkinson’s disease. Journal of Theoretical Biology, 598(??):??, February 7, 2025. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519324002820.

Roberts:2022:MMR

[Rob22] Paul A. Roberts. Mathematical models of retinitis pigmentosa: the trophic factor hypothesis. Journal of Theoretical Biology, 534(??):??, February 7, 2022. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251932100357X.

Rosenberg:2021:DML

[Ros21] Eric Rosenberg. On deriving Murray’s law from constrained minimization of flow resistance. Journal of Theoretical Biology, 512(??):??, March 7, 2021. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519320304185.

Rosenberg:2022:TFL

[Ros22] Eric Rosenberg. Three or four levels of hierarchy minimize hydraulic power in leaves with pinnate dendritic venation. Journal of Theoretical Biology, 539(??):??, April 21, 2022. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322000595.

Royce:2021:ANM

[Roy21] Katherine Royce. Application of a novel mathematical model to identify intermediate hosts of SARS-CoV-2. Journal of Theoretical Biology, 526(??):??, October 7, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321001831.

Renton:2021:CSE

[RP21a] Jessie Renton and Karen M. Page. Cooperative success in epithelial public goods games. Journal of Theoretical Biology, 528(??):??, November 7, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321002575.

Rosandic:2021:NIS

[RP21b] Marija Rosandić and Vladimir Paar. The novel Ideal Symmetry Genetic Code table — common purine–pyrimidine symmetry net for all RNA and DNA species. Journal of Theoretical Biology, 524(??):??, September 7, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321001703.

Rodriguez:2024:APU

[RR24] Christopher W. Rodriguez and Peter W. Reddien. Agelessness is possible under the disposable soma theory but system complexity makes it unlikely. Journal of Theoretical Biology, 595(??):??, December 7, 2024. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0022519324002431.

Randall:2021:GSA

[RRAO21] E. Benjamin Randall, Nicholas Z. Randolph, Alen Alexanderian, and Mette S. Olufsen. Global sensitivity analysis informed model reduction and selection applied to a Valsalva maneuver model. Journal of Theoretical Biology, 526(??):??, October 7, 2021. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519321001818.

Ruiz-Ramirez:2020:MPL

[RRZDF20] Javier Ruiz-Ramírez, Arturas Ziemys, Prashant Dogra, and Mauro Ferrari. A modeling platform for the lymphatic system. Journal of Theoretical Biology, 493(??):??, May 21, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519320300485.

Radunskaya:2024:KRI

[RS24] Ami Radunskaya and Joshua Sack. Kill rates by immune cells: Ratio-dependent, or mass action? Journal of Theoretical Biology, 582(??):??, April 7, 2024. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519324000298.

Rudrauf:2022:MSP

[RSPB^{+}22] D. Rudrauf, G. Sergeant-Perthuis, O. Belli, Y. Tisserand, and G. Di Marzo Serugendo. Modeling the subjective perspective of consciousness and its role in the control of behaviours. Journal of Theoretical Biology, 534(??):??, February 7, 2022. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519321003763.

Rahnsch:2024:NBU

[RT24a] Beatrix Rahnsch and Leila Taghizadeh. Network-based uncertainty quantification for mathematical models in epidemiology. Journal of Theoretical Biology, 577(??):??, January 21, 2024. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519323002680.

Richards:2024:DIS

[RT24b] Kelvin J. Richards and Axel Timmermann. Dispersal and interbreeding as survival strategies for species exposed to environment change. Journal of Theoretical Biology, 585(??): ??, May 21, 2024. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251932400078X.

Reda:2020:AIG

[RW20] Clémence Réda and Bartek Wilczyński. Automated inference of gene regulatory networks using explicit regulatory modules. Journal of Theoretical Biology, 486(??):??, February 7, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519319304606.

Rybnikov:2021:FDP

[RWHK21] Sviatoslav Rybnikov, Daniel B. Weissman, Sariel Hübner, and Abraham B. Korol. Fitness dependence preserves selection for recombination across diverse mixed mating strategies. Journal of Theoretical Biology, 528(??):??, November 7, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251932100268X.

Rubio:2022:MME

[RY22] Felipe Alves Rubio and Hyun Mo Yang. A mathematical model to evaluate the role of memory B and T cells in heterologous secondary dengue infection. Journal of Theoretical Biology, 534(??):??, February 7, 2022. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321003817.

Ramirez-Zuniga:2022:DDM

[RZRS^{+}22] Ivan Ramirez-Zuniga, Jonathan. E. Rubin, David Swigon, Heinz Redl, and Gilles Clermont. A data-driven model of the role of energy in sepsis. Journal of Theoretical Biology, 533(??):??, January 21, 2022. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519321003672.

Smith:2025:ECW

[SA25] Cameron A. Smith and Ben Ashby. Efficient coupling of within-and between-host infectious disease dynamics. Journal of Theoretical Biology, 602–603(??):??, April 7, 2025. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251932500027X.

Sachs:2022:PGC

[Sac22] Gottfried Sachs. Powered-gliding/climbing flight. Journal of Theoretical Biology, 547(??):??, August 21, 2022. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322001448.

Sole:2021:ECD

[SAG21] Ricard Solé and Guim Aguadé-Gorgorió. The ecology of cancer differentiation therapy. Journal of Theoretical Biology, 511(??):??, February 21, 2021. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519320304070.

Sakai:2020:MHI

[Sak20] Satoki Sakai. Maintenance of high inbreeding depression in selfing populations: Two-stage effect of early- and late-acting mutations. Journal of Theoretical Biology, 502(??):??, October 7, 2020. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519320301624.

Salamatin:2021:LMA

[SAK21] Artur A. Salamatin, Peter H. Adler, and Konstantin G. Kornev. Lepidopteran mouthpart architecture suggests a new mechanism of fluid uptake by insects with long proboscises. Journal of Theoretical Biology, 510(??):??, February 7, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519320303805.

Sakai:2022:EMM

[Sak22a] Satoki Sakai. Evolution of mixed mating: Effects of among-parent variation in early-acting inbreeding depression coupled with overproduction of ovules. Journal of Theoretical Biology, 550(??):??, October 7, 2022. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322002363.

Sakai:2022:WDM

[Sak22b] Satoki Sakai. Why are deleterious mutations maintained in selfing populations? An analysis of the effects of earlyand late-acting mutations by a two-locus two-allele model. Journal of Theoretical Biology, 533(??):??, January 21, 2022. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321003751.

Salvador:2024:PTB

[Sal24] Armindo Salvador. Pillars of theoretical biology: “Biochemical systems analysis, I, II and III”. Journal of Theoretical Biology, 576(??):??, January 7, 2024. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519323002527.

Saldana:2025:VSP

[Sal25] Fernando Saldaña. Vaccination strategies in a pair formation model for human papillomavirus infection: an optimal control approach. Journal of Theoretical Biology, 597(??):??, January 21, 2025. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519324002790.

Santoni:2021:ICC

[San21] Daniele Santoni. The impact of codon choice on translation process in Saccharomyces cerevisiae: folding class, protein function and secondary structure. Journal of Theoretical Biology, 526(??):??, October 7, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321002253.

Simpson:2022:REP

[SBB^{+}22] Matthew J. Simpson, Ruth E. Baker, Pascal R. Buenzli, Ruanui Nicholson, and Oliver J. Maclaren. Reliable and efficient parameter estimation using approximate continuum limit descriptions of stochastic models. Journal of Theoretical Biology, 549(??):??, September 21, 2022. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322001990.

Schweinoch:2020:MME

[SBC^{+}20] Darius Schweinoch, Pia Bachmann, Diana Clausznitzer, Marco Binder, and Lars Kaderali. Mechanistic modeling explains the dsRNA length-dependent activation of the RIG-I mediated immune response. Journal of Theoretical Biology, 500(??):??, September 7, 2020. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519320301910.

Sun:2021:MMK

[SBJR21] Shan Sun, Mark Broom, Michal Johanis, and Jan Rychtár. A mathematical model of kin selection in floral displays. Journal of Theoretical Biology, 509(??):??, January 21, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519320303258.

Sharp:2020:DCT

[SBM^{+}20] Jesse A. Sharp, Alexander P. Browning, Tarunendu Mapder, Christopher M. Baker, Kevin Burrage, and Matthew J. Simpson. Designing combination therapies using multiple optimal controls. Journal of Theoretical Biology, 497(??):??, July 21, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519320301326.

Soltanzadeh:2023:LBS

[SBSB23] Milad Soltanzadeh, Solenna Blanchard, Jean-Paul Soucy, and Habib Benali. Lactate’s behavioral switch in the brain: an in-silico model. Journal of Theoretical Biology, 575(??):??, November 7, 2023. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251932300245X.

Simpson:2022:PIM

[SBW^{+}22a] Matthew J. Simpson, Alexander P. Browning, David J. Warne, Oliver J. Maclaren, and Ruth E. Baker. Parameter identifiability and model selection for sigmoid population growth models. Journal of Theoretical Biology, 535(??):??, February 21, 2022. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519321004185.

Sund:2022:UMH

[SBW^{+}22b] Derrick T. Sund, Andrew F. Brouwer, Heather M. Walline, Thomas E. Carey, Rafael Meza, Trachette Jackson, and Marisa C. Eisenberg. Understanding the mechanisms of HPVrelated carcinogenesis: Implications for cell cycle dynamics. Journal of Theoretical Biology, 551–552(??):??, ???? 2022. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322002326.

Serodio:2021:POT

[SC21] João Serôdio and Douglas A. Campbell. Photoinhibition in optically thick samples: Effects of light attenuation on chlorophyll fluorescence-based parameters. Journal of Theoretical Biology, 513(??):??, March 21, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321000023.

Sherlock:2023:OVT

[SC23a] Brock D. Sherlock and Adelle C. F. Coster. Oncolytic virus treatment of human breast cancer cells: Modelling therapy efficacy. Journal of Theoretical Biology, 560(??):??, March 7, 2023. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251932200385X.

Sibly:2023:AFS

[SC23b] Richard M. Sibly and Robert N. Curnow. Allele frequencies and selection coefficients in locally adapted populations. Journal of Theoretical Biology, 565(??):??, May 21, 2023. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519323000590.

Skedros:2023:ESR

[SCD^{+}23] John G. Skedros, John T. Cronin, Michael R. Dayton, Roy D. Bloebaum, and Kent N. Bachus. Exploration of the synergistic role of cortical thickness asymmetry (“Trabecular Eccentricity” concept) in reducing fracture risk in the human femoral neck and a control bone (Artiodactyl Calcaneus). Journal of Theoretical Biology, 567(??):??, June 21, 2023. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519323000917.

Schulman:2017:BRA

[Sch17] Lawrence S. Schulman. Bacterial resistance to antibodies: a model evolutionary study. Journal of Theoretical Biology, 417(??):61–67, March 21, 2017. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S002251931730022X. See corrigendum [Sch23].

Schlomann:2018:SMD

[Sch18] Brandon H. Schlomann. Stationary moments, diffusion limits, and extinction times for logistic growth with random catastrophes. Journal of Theoretical Biology, 454(??):154–163, October 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318302996. See corri-

gendum [Sch21a].

Schlomann:2021:CSM

[Sch21a] Brandon H. Schlomann. Corrigendum to “Stationary moments, diffusion limits, and extinction times for logistic growth with random catastrophes” [J. Theor. Biol. 454 (2018) 154–163]. Journal of Theoretical Biology, 514(??):??, April 7, 2021. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519321000175. See [Sch18].

Schutz:2021:MMT

[Sch21b] G. M. Schütz. Molecular motor traffic with a slow binding site. Journal of Theoretical Biology, 518(??):??, June 7, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519321000667.

Schulman:2023:CBR

[Sch23] Lawrence S. Schulman. Corrigendum to “Bacterial resistance to antibodies: a model evolutionary study” [J. Theor. Biol. 417 (2017) 61–67]. Journal of Theoretical Biology, 565(??):??, May 21, 2023. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519323000607. See [Sch17].

Su:2024:PBI

[SCH24] Min Su, Xiaowei Chen, and Cang Hui. Persistent biological invasions alter ecological network topology, impacting disease transmission during community assembly. Journal of Theoretical Biology, 595(??):??, December 7, 2024. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0022519324002352.

Song:2020:EDE

[SCHH20] Zilong Song, Xiulei Cao, Tzyy-Leng Horng, and Huaxiong Huang. Electric discharge of electrocytes: Modelling, analysis and simulation. Journal of Theoretical Biology, 498(??):??, August 7, 2020. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519320301491.

Scott:2024:CPK

[Sco24] Thomas W. Scott. Crozier’s paradox and kin recognition: Insights from simplified models. Journal of Theoretical Biology, 581(??):??, March 21, 2024. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251932400016X.

Schmid:2020:ASI

[SD20] Harald Schmid and Hana M. Dobrovolny. An approximate solution of the interferon-dependent viral kinetics model of influenza. Journal of Theoretical Biology, 498(??):??, August 7, 2020. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519320301211.

Storch:2022:TEW

[SD22] Laura S. Storch and Sarah L. Day. Topological early warning signals: Quantifying varying routes to extinction in a spatially distributed population model. Journal of Theoretical Biology, 554(??):??, December 7, 2022. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251932200265X.

Sutradhar:2025:EHB

[SD25] Rupchand Sutradhar and D. C. Dalal. Exploration of hepatitis B virus infection dynamics through an intracellular model. Journal of Theoretical Biology, 611(??):??, August 21, 2025. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519325001559.

Salles:2024:ECT

[SdAM24] Luis Francisco Prado Salles, Marcus Aloizio Martinez de Aguiar, and Flavia Maria Darcie Marquitti. Evolution of cooperation in a two-species system with a common resource pool. Journal of Theoretical Biology, 577(??):??, January 21, 2024. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519323002679.

Schroter:2023:WEP

[SdB23] Juliane Schröter and Rob J. de Boer. What explains the poor contraction of the viral load during paediatric HIV infection? Journal of Theoretical Biology, 570(??):??, August 7, 2023. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519323001170.

Siewe:2023:CTI

[SF23] Nourridine Siewe and Avner Friedman. Cancer therapy with immune checkpoint inhibitor and CSF-1 blockade: a mathematical model. Journal of Theoretical Biology, 556(??):??, January 7, 2023. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519322002880.

Selvitella:2024:ASS

[SF24a] Alessandro Maria Selvitella and Kathleen Lois Foster. An approximate solution of the SLIP model under the regime of linear angular dynamics during stance and the stability of symmetric periodic running gaits. Journal of Theoretical Biology, 595(??):??, December 7, 2024. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0022519324002194.

Siewe:2024:MMO

[SF24b] Nourridine Siewe and Avner Friedman. A mathematical model of obesity-induced type 2 diabetes and efficacy of anti-diabetic weight reducing drug. Journal of Theoretical Biology, 581(??):??, March 21, 2024. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519324000377.

Sasanami:2023:PCI

[SFK^{+}23] Misaki Sasanami, Marie Fujimoto, Taishi Kayano, Katsuma Hayashi, and Hiroshi Nishiura. Projecting the COVID-19 immune landscape in Japan in the presence of waning immunity and booster vaccination. Journal of Theoretical Biology, 559(??):??, February 21, 2023. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322003757.

Song:2020:DNP

[SFYL20] Da Song, Meng Fan, Shihan Yan, and Meng Liu. Dynamics of a nutrient-phytoplankton model with random phytoplankton mortality. Journal of Theoretical Biology, 488(??): ??, March 7, 2020. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519319304886.

Shen:2024:OAC

[SGJ^{+}24] Anglu Shen, Shufei Gao, Jie Jiang, Qingjing Hu, Hao Wang, and Sanling Yuan. Oscillations of algal cell quota: Considering two-stage phosphate uptake kinetics. Journal of Theoretical Biology, 581(??):??, March 21, 2024. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519324000201.

Seurat:2025:DOI

[SGM^{+}25] Jérémy Seurat, Krista R. Gerbino, Justin R. Meyer, Joshua M. Borin, and Joshua S. Weitz. Design, optimization, and inference of biphasic decay of infectious virus particles. Journal of Theoretical Biology, 600(??):??, March 7, 2025. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0022519325000086.

Sun:2021:CBB

[SH21] T. Anthony Sun and Frank M. Hilker. Comparison between best-response dynamics and replicator dynamics in a social-ecological model of lake eutrophication. Journal of Theoretical Biology, 509(??):??, January 21, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519320303465.

Sorensen:2020:DMS

[SHH20] Anna Irene Vedel Sørensen, Julie Elvekjær Hansen, and Tariq Halasa. A dynamic model for spread of livestock-associated methicillin-resistant Staphylococcus aureus on a pig farm, incorporating bacterial load and human exposure through air. Journal of Theoretical Biology, 505(??):??, November 21, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519320302575.

Steel:2020:DBD

[SHK20] Mike Steel, Wim Hordijk, and Stuart A. Kauffman. Dynamics of a birth-death process based on combinatorial innovation. Journal of Theoretical Biology, 491(??):??, April 21, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519320300436.

Sinha:2021:SAO

[SHY21] Nirvik Sinha, C. J. Heckman, and Yuan Yang. Slowly activating outward membrane currents generate input-output sub-harmonic cross frequency coupling in neurons. Journal of Theoretical Biology, 509(??):??, January 21, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519320303647.

Simpson:2025:DIM

[Sim25] Matthew J. Simpson. Data-informed model reduction for inference and prediction from non-identifiable models. Journal of Theoretical Biology, 611(??):??, August 21, 2025. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519325001213.

Singh:2021:ACP

[Sin21] Abhyudai Singh. Attack by a common parasitoid stabilizes population dynamics of multi-host communities. Journal of Theoretical Biology, 531(??):??, December 21, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321003167.

Silva:2020:MES

[SJNRS20] Telma Silva, Willi Jäger, Maria Neuss-Radu, and Adélia Sequeira. Modeling of the early stage of atherosclerosis with emphasis on the regulation of the endothelial permeability. Journal of Theoretical Biology, 496(??):??, July 7, 2020. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519320300849.

Sathishkumar:2025:EBB

[SJS^{+}25] M. Sathishkumar, Maya Joby, Srimanta Santra, Yong-Ki Ma, and S. Marshal Anthoni. Event-based biological pest control: an LMI approach. Journal of Theoretical Biology, 596(??):??, January 7, 2025. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519324002601.

Samanta:2021:UOG

[SK21] Tagari Samanta and Sandip Kar. Unraveling the origin of glucose mediated disparate proliferation dynamics of cancer stem cells. Journal of Theoretical Biology, 526(??):??, October 7, 2021. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251932100196X.

Sudersan:2023:MDF

[SK23] Pranav Sudersan and Michael Kappl. Mechanisms of detachment in fibrillar adhesive systems. Journal of Theoretical Biology, 557(??):??, January 21, 2023. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251932200306X.

Schirmer:2022:EMR

[SKNvRL22] Saskia Schirmer, Fränzi Korner-Nievergelt, Jan A. C. von Rönn, and Volkmar Liebscher. Estimation in the multinomial reencounter model — where do migrating animals go and how do they survive in their destination area? Journal of Theoretical Biology, 543(??):??, June 21, 2022. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322001060.

Schirmer:2023:ESC

[SKNvRL23] Saskia Schirmer, Fränzi Korner-Nievergelt, Jan A. C. von Rönn, and Volkmar Liebscher. Estimating survival in continuous space from mark-dead-recovery data — Towards a continuous version of the multinomial dead recovery model. Journal of Theoretical Biology, 574(??):??, October 7, 2023. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519323002229.

Safdari:2021:PLH

[SKS21] Hadiseh Safdari, Ata Kalirad, and Mehdi Sadeghi. Populationlevel heterogeneity as a reflection of mixed strategy: a computational perspective on the Crabtree effect. Journal of Theoretical Biology, 531(??):??, December 21, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519321003313.

Senneff:2021:EEP

[SL21] Sageanne Senneff and Madeleine M. Lowery. Effects of extracellular potassium on calcium handling and force generation in a model of excitation-contraction coupling in skeletal muscle. Journal of Theoretical Biology, 519(??):??, June 21, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519321000783.

Shtossel:2022:SBM

[SL22] Oshrit Shtossel and Yoram Louzoun. Sampling bias minimization in disease frequency estimates. Journal of Theoretical Biology, 534(??):??, February 7, 2022. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519321003921.

Stadt:2023:MMC

[SL23] Melissa M. Stadt and Anita T. Layton. Mathematical modeling of calcium homeostasis in female rats: an analysis of sex differences and maternal adaptations. Journal of Theoretical Biology, 572(??):??, September 7, 2023. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519323001807.

Siddharth:2024:PPO

[SL24] Thiru Siddharth and Nathan E. Lewis. Predicting pathways for old and new metabolites through clustering. Journal of Theoretical Biology, 578(??):??, February 7, 2024. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519323002813.

Sim:2022:DRL

[SLCC22] Hong Jhun Sim, Weng Ngai Lam, Ryan A. Chisholm, and Kwek Yan Chong. Downstream resource leakage a necessary condition for the stress-gradient hypothesis in processing chain commensalisms. Journal of Theoretical Biology, 538(??):??, April 7, 2022. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519322000418.

Sequeira:2020:TEB

[SLML20] João Sequeira, Jorge Louçã, António M. Mendes, and Pedro G. Lind. Transition from endemic behavior to eradication of malaria due to combined drug therapies: an agent-model approach. Journal of Theoretical Biology, 484(??):??, January 7, 2020. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251931930400X.

Schiavo:2022:TVT

[SLTD22] Filippo Schiavo, Emely Kjellsson Lindblom, and Iuliana Toma-Dasu. Towards the virtual tumor for optimizing radiotherapy treatments of hypoxic tumors: a novel model of heterogeneous tissue vasculature and oxygenation. Journal of Theoretical Biology, 547(??):??, August 21, 2022. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322001734.

Smerlak:2021:QSE

[Sme21] Matteo Smerlak. Quasi-species evolution maximizes genotypic reproductive value (not fitness or flatness). Journal of Theoretical Biology, 522(??):??, August 7, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321001211.

Sego:2022:MMS

[SMEG22] T. J. Sego, Ericka D. Mochan, G. Bard Ermentrout, and James A. Glazier. A multiscale multicellular spatiotemporal model of local influenza infection and immune response. Journal of Theoretical Biology, 532(??):??, January 7, 2022. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321003374.

Sandhu:2020:ERS

[SMJ20] Simran Kaur Sandhu, Andrew Morozov, and Lourdes Juan. Exploring the role of spatial and stoichiometric heterogeneity in the top-down control in eutrophic planktonic ecosystems. Journal of Theoretical Biology, 499(??):??, August 21, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519320301661.

Sriram:2023:ISP

[SMM^{+}23] Sridevi Sriram, Simin Mirzaei, Mahtab Mehrabbeik, Karthikeyan Rajagopal, Mehdi Rostami, and Sajad Jafari. The influence of synaptic pathways on the synchronization patterns of regularly structured mChialvo map network. Journal of Theoretical Biology, 572(??):??, September 7, 2023. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519323001881.

Simpson:2024:MCD

[SMM24] Matthew J. Simpson, Ryan J. Murphy, and Oliver J. Maclaren. Modelling count data with partial differential equation models in biology. Journal of Theoretical Biology, 580(??):??, March 7, 2024. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519324000134.

Singh:2020:MCN

[SMS20] Shakti Nath Singh, Md. Zubbair Malik, and R. K. Brojen Singh. Molecular crosstalk: Notch can manipulate Hes1 and miR-9 behavior. Journal of Theoretical Biology, 504(??):??, November 7, 2020. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519320302599.

Sinha:2020:MMA

[SNK20] Arvind Kumar Sinha, Nishant Namdev, and Awanish Kumar. A mathematical model of adiponectin resistance. Journal of Theoretical Biology, 494(??):??, June 7, 2020. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519320301016.

Southall:2023:PGM

[SOGK^{+}23] E. Southall, Z. Ogi-Gittins, A. R. Kaye, W. S. Hart, F. A. Lovell-Read, and R. N. Thompson. A practical guide to mathematical methods for estimating infectious disease outbreak risks. Journal of Theoretical Biology, 562(??):??, April 7, 2023. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519323000139.

Sharma:2022:BNI

[SOKZ22] Akshay Sharma, Louise Olsen-Kettle, and Tonghua Zhang. Bistability and noise-induced transient behaviour of steady states in a cancer network with the regulation of microRNA. Journal of Theoretical Biology, 554(??):??, December 7, 2022. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322002533.

Steel:2020:ESA

[SP20] Harrison Steel and Antonis Papachristodoulou. The effect of spatiotemporal antibiotic inhomogeneities on the evolution of resistance. Journal of Theoretical Biology, 486(??):??, February 7, 2020. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519319304461.

Sachak-Patwa:2023:TCL

[SPLS^{+}23] Rahil Sachak-Patwa, Erin I. Lafferty, Claude J. Schmit, Robin N. Thompson, and Helen M. Byrne. A targetcell limited model can reproduce influenza infection dynamics in hosts with differing immune responses. Journal of Theoretical Biology, 567(??):??, June 21, 2023. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519323000875.

Sakref:2024:EEA

[SR24] Yann Sakref and Olivier Rivoire. On the exclusion of exponential autocatalysts by sub-exponential autocatalysts. Journal of Theoretical Biology, 579(??):??, February 21, 2024. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519323003119.

Sanchez:2024:TGR

[SRB^{+}24] Juan Felipe Sánchez, Salah Ramtani, Abdelkader Boucetta, Marco Antonio Velasco, Juan Jairo Vaca-González, Carlos A. Duque-Daza, and Diego A. Garzón-Alvarado. Tumor growth for remodeling process: a 2D approach. Journal of Theoretical Biology, 585(??):??, May 21, 2024. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519324000626.

Senapati:2021:IIS

[SRDC21] Abhishek Senapati, Sourav Rana, Tamalendu Das, and Joydev Chattopadhyay. Impact of intervention on the spread of COVID-19 in India: a model based study. Journal of Theoretical Biology, 523(??):??, August 21, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321001338.

Steinmetz:2021:CAD

[SS21] Bnaya Steinmetz and Nadav M. Shnerb. Competition with abundance-dependent fitness and the dynamics of heterogeneous populations in fluctuating environment. Journal of Theoretical Biology, 531(??):??, December 21, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251932100299X.

Stepien:2022:SMD

[SS22] Tracy L. Stepien and Timothy W. Secomb. Spreading mechanics and differentiation of astrocytes during retinal development. Journal of Theoretical Biology, 549(??):??, September 21, 2022. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519322002065.

Sagitov:2023:CUM

[SS23] Serik Sagitov and Anders Ståhlberg. Counting unique molecular identifiers in sequencing using a multi-type branching process with immigration. Journal of Theoretical Biology, 558(??):??, February 7, 2023. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519322003563.

Sau:2020:ESA

[SSB20] Anurag Sau, Bapi Saha, and Sabyasachi Bhattacharya. An extended stochastic Allee model with harvesting and the risk of extinction of the herring population. Journal of Theoretical Biology, 503(??):??, October 21, 2020. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519320302307.

Souza:2020:CGA

[SSBG20] Paulo Victor Santos Souza, Rafael Silva, Chris Bauch, and Daniel Girardi. Cooperation in a generalized age-structured spatial game. Journal of Theoretical Biology, 484(??):??, January 7, 2020. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519319303455.

Sturman:2025:CPI

[SSK^{+}25] Fabian Sturman, Ben Swallow, Cliff Kerr, Robyn M. Stuart, and Jasmina Panovska-Griffiths. Can pruning improve agent-based models’ calibration? An application to HPVsim. Journal of Theoretical Biology, 611(??):??, August 21, 2025. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519325000967.

Sveshnikova:2024:KRC

[SSKP24] Anastasia N. Sveshnikova, Alexey M. Shibeko, Tatiana A. Kovalenko, and Mikhail A. Panteleev. Kinetics and regulation of coagulation factor X activation by intrinsic tenase on phospholipid membranes. Journal of Theoretical Biology, 582(??): ??, April 7, 2024. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519324000389.

Stadler:2018:SAS

[SSKS18] Tanja Stadler, Stavroula Skylaki, Konstantinos D. Kokkaliaris, and Timm Schroeder. On the statistical analysis of single cell lineage trees. Journal of Theoretical Biology, 439(??):160–165, February 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519317305301. See erra-

tum [SSKS20].

Stadler:2020:ESA

[SSKS20] Tanja Stadler, Stavroula Skylaki, Konstantinos D. Kokkaliaris, and Timm Schroeder. Erratum to “On the statistical analysis of single cell lineage trees” [J. Theor. Biol. 439 (2018) 160– 165]. Journal of Theoretical Biology, 506(??):??, December 7, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519320302794. See [SSKS18].

Sangregorio-Soto:2025:PBA

[SSLMGC25] Viyils Sangregorio-Soto, Edgar Yesid Mayorga Lancheros, Gianfranco Mazzanti, and Claudia L. Garzón-Castro. PD-based ADRC using time-varying gains: an application to microalgal-based bioprocess. Journal of Theoretical Biology, 604(??):??, May 7, 2025. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519325000402.

Spencer:2022:DVR

[SSM^{+}22] Julie A. Spencer, Deborah P. Shutt, S. Kane Moser, Hannah Clegg, Helen J. Wearing, Harshini Mukundan, and Carrie A. Manore. Distinguishing viruses responsible for influenza-like illness. Journal of Theoretical Biology, 545(??):??, July 21, 2022. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519322001436.

Susko:2021:CUW

[SSR21] Edward Susko, Mike Steel, and Andrew J. Roger. Conditions under which distributions of edge length ratios on phylogenetic trees can be used to order evolutionary events. Journal of Theoretical Biology, 526(??):??, October 7, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321002071.

Schonfeld:2021:APU

[SSS21] Ethan Schonfeld, Elan Schonfeld, and Dan Schonfeld. Autocatalytic-protection for an unknown locus CRISPR-Cas countermeasure for undesired mutagenic chain reactions. Journal of Theoretical Biology, 528(??):??, November 7, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321002502.

Shen:2022:ISI

[SSX^{+}22] Mingwang Shen, Xiang Sun, Yanni Xiao, Yuanbao Liu, Congyue Wang, Zhiguo Wang, Libin Rong, and Zhihang Peng. The impact of supplementary immunization activities on measles transmission dynamics and implications for measles elimination goals: a mathematical modelling study. Journal of Theoretical Biology, 551–552(??):??, ???? 2022. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322002375.

Shuttleworth:2020:MDH

[ST20] Robyn Shuttleworth and Dumitru Trucu. Multiscale dynamics of a heterotypic cancer cell population within a fibrous extracellular matrix. Journal of Theoretical Biology, 486(??):??, February 7, 2020. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519319304096.

Sauve:2020:ESS

[STB20] Alix M. C. Sauve, Rachel A. Taylor, and Frédéric Barraquand. The effect of seasonal strength and abruptness on predator–prey dynamics. Journal of Theoretical Biology, 491(??):??, April 21, 2020. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251932030031X.

Sanchez-Taltavull:2021:MSO

[STCSC^{+}21] Daniel Sánchez-Taltavull, Violeta Castelo-Szekely, Daniel Candinas, Edgar Roldán, and Guido Beldi. Modelling strategies to organize healthcare workforce during pandemics: Application to COVID-19. Journal of Theoretical Biology, 523(??):??, August 21, 2021. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519321001405.

Steel:2021:MAC

[Ste21a] Mike Steel. Modelling aspects of consciousness: a topological perspective. Journal of Theoretical Biology, 523(??):??, August 21, 2021. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519321001351.

Stenseke:2021:PHS

[Ste21b] Jakob Stenseke. Persistent homology and the shape of evolutionary games. Journal of Theoretical Biology, 531(??):??, December 21, 2021. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519321003222.

Suarez:2020:GAM

[SUA^{+}20] Gonzalo P. Suarez, Oyita Udiani, Brian F. Allan, Candice Price, Sadie J. Ryan, Eric Lofgren, Alin Coman, Chris M. Stone, Lazaros K. Gallos, and Nina H. Fefferman. A generic arboviral model framework for exploring trade-offs between vector control and environmental concerns. Journal of Theoretical Biology, 490(??):??, April 7, 2020. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519320300175.

Suarez:2021:ABS

[SVHM^{+}21] Alejandro Suarez, Pedro A. Valdes-Hernandez, Arash Moshkforoush, Nikolaos Tsoukias, and Jorge Riera. Arterial blood stealing as a mechanism of negative BOLD response: From the steady-flow with nonlinear phase separation to a windkessel-based model. Journal of Theoretical Biology, 529(??):??, November 21, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321002757.

Sanchez-Vargas:2024:FMS

[SVVPPR^{+}24] J. Sánchez-Vargas, F. J. Valdés-Parada, L. Peraza-Reyes, D. Lasseux, and M. A. Trujillo-Roldán. Flow modeling and structural characterization in fungal pellets. Journal of Theoretical Biology, 590(??):??, August 7, 2024. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519324001346.

Shan:2022:SRA

[SW22] Xu Shan and Bin Wu. Social relationship adjustments within the same sex promote marital bliss. Journal of Theoretical Biology, 550(??):??, October 7, 2022. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322002223.

Sweatman:2020:MMD

[Swe20] Catherine Z. W. Hassell Sweatman. Mathematical model of diabetes and lipid metabolism linked to diet, leptin sensitivity, insulin sensitivity and VLDLTG clearance predicts paths to health and type II diabetes. Journal of Theoretical Biology, 486(??):??, February 7, 2020. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319304060.

Shi:2020:TMF

[SWK^{+}20] Lianhui Shi, Jianing Wu, Harald W. Krenn, Yunqiang Yang, and Shaoze Yan. Temporal model of fluid-feeding mechanisms in a long proboscid orchid bee compared to the short proboscid honey bee. Journal of Theoretical Biology, 484(??):??, January 7, 2020. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251931930387X.

Spinelli:2025:IEP

[SWVW25] Kamryn Spinelli, Chaozhen Wei, Luis Vidali, and Min Wu. Inferring exocytosis profiles from cell shapes using a dualconfiguration model of walled cell tip growth. Journal of Theoretical Biology, 612(??):??, September 7, 2025. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519325001626.

Sun:2020:WLL

[SXJ20] Xiaodan Sun, Yanni Xiao, and Xiangting Ji. When to lift the lockdown in Hubei province during COVID-19 epidemic? An insight from a patch model and multiple source data. Journal of Theoretical Biology, 507(??):??, December 21, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519320303246.

Shen:2020:LTP

[SXL^{+}20] Mingwang Shen, Ruyi Xia, Zheng Luo, Hongmei Zeng, Wenqiang Wei, Guihua Zhuang, Lei Zhang, and Wanqing Chen. The long-term population impact of endoscopic screening programmes on disease burdens of gastric cancer in China: a mathematical modelling study. Journal of Theoretical Biology, 484(??):??, January 7, 2020. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319303467.

Song:2024:RMD

[SY24] Yan Song and Qian Yang. Revisiting the modularitydisease transmission link: uncovering the importance of intramodular structure. Journal of Theoretical Biology, 583(??):??, April 21, 2024. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519324000535.

Song:2025:AEA

[SY25] Xinlin Song and Feifei Yang. Analysis of electrical activities in a functional neuron with dual memristors. Journal of Theoretical Biology, 599(??):??, February 21, 2025. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519324003199.

Su:2020:HIP

[SYH20] Min Su, Yuanqi Yang, and Cang Hui. How intraguild predation affects the host diversity–disease relationship in a multihost community. Journal of Theoretical Biology, 490(??):

??, April 7, 2020. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519320300308.

Sekiya:2020:BMO

[SYI20] Tokuharu Sekiya, Sachi Yamaguchi, and Yoh Iwasa. Bovine mastitis and optimal disease management: Dynamic programming analysis. Journal of Theoretical Biology, 498 (??):??, August 7, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519320301478.

Sun:2025:PRI

[SZR^{+}25] Xiaodan Sun, Weike Zhou, Yuhua Ruan, Guanghua Lan, Qiuying Zhu, and Yanni Xiao. Perceived risk induced multiscale model: Coupled within-host and between-host dynamics and behavioral dynamics. Journal of Theoretical Biology, 599(??):??, February 21, 2025. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519324002832.

Takashina:2024:EDE

[Tak24] Nao Takashina. Effects of delay and error in the feedback structure of ecological management. Journal of Theoretical Biology, 595(??):??, December 7, 2024. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S002251932400211X.

Tatematsu:2023:RBI

[TAP^{+}23] Daiki Tatematsu, Marwa Akao, Hyeongki Park, Shingo Iwami, Keisuke Ejima, and Shoya Iwanami. Relationship between the inclusion/exclusion criteria and sample size in randomized controlled trials for SARS-CoV-2 entry inhibitors. Journal of Theoretical Biology, 561(??):??, March 21, 2023. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322003940.

Tatsukawa:2021:FTD

[TATT21] Yuichi Tatsukawa, Md. Rajib Arefin, Masaki Tanaka, and Jun Tanimoto. Free ticket, discount ticket or intermediate of the best of two worlds — which subsidy policy is socially optimal to suppress the disease spreading? Journal of Theoretical Biology, 520(??):??, July 7, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321001041.

Taufer:2020:RLS

[Täu20] Matthias Täufer. Rapid, large-scale, and effective detection of COVID-19 via non-adaptive testing. Journal of Theoretical Biology, 506(??):??, December 7, 2020. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519320303052.

Taylor:2023:MFT

[Tay23] David Taylor. Mechanical fatigue in trees mitigated by annual growth: a theoretical model. Journal of Theoretical Biology, 569(??):??, July 21, 2023. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519323001194.

Toth:2023:RIS

[TBNS23] Zoltán Tóth, Roland Bartók, Zsófia Nagy, and Viktor R. Szappanos. The relative importance of social information use for population abundance in group-living and non-grouping prey. Journal of Theoretical Biology, 575(??):??, November 7, 2023. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519323002230.

Terman:2021:MMC

[TCH21] David Terman, Liam Chen, and Yousef Hannawi. Mathematical modeling of cerebral capillary blood flow heterogeneity and its effect on brain tissue oxygen levels. Journal of Theoretical Biology, 527(??):??, October 21, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321002368.

Thenard:2021:CPE

[TCMA21] Thomas Thenard, Anita Catapano, Michel Mesnard, and Rachele Allena. A Cellular Potts energy-based approach to analyse the influence of the surface topography on single cell motility. Journal of Theoretical Biology, 509(??):??, January 21, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519320303428.

Thakuri:2023:CRA

[TDRC23] Bikash Thakuri, Jayanta Kumar Das, Amit Kumar Roy, and Amit Chakraborty. Circulating renin-angiotensin systems mediated feedback controls over the mean-arterial pressure. Journal of Theoretical Biology, 572(??):??, September 7, 2023. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519323001868.

Terman:2023:MRH

[Ter23] David Terman. Modeling the response of homogeneous and heterogeneous cerebral capillary networks to local changes in vessel diameters. Journal of Theoretical Biology, 568(??):??, July 7, 2023. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519323001054.

Teboh-Ewungkem:2021:MAI

[TEWN21] Miranda I. Teboh-Ewungkem, Woldegebriel Assefa Woldegerima, and Gideon A. Ngwa. Mathematical assessment of the impact of human-antibodies on sporogony during the within-mosquito dynamics of Plasmodium falciparum parasites. Journal of Theoretical Biology, 515(??):??, April 21, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519320304173.

Tverskoi:2022:EGS

[TG22] Denis Tverskoi and Sergey Gavrilets. The evolution of germsoma specialization under different genetic and environmental effects. Journal of Theoretical Biology, 534(??):??, February 7, 2022. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519321003842.

Tellioglu:2022:MEW

[TGC22] Nefel Tellioglu, Nicholas Geard, and Rebecca H. Chisholm. Modelling the effect of within-host dynamics on the diversity of a multi-strain pathogen. Journal of Theoretical Biology, 548(??):??, September 7, 2022. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519322001837.

Thorvaldsen:2020:USM

[TH20] Steinar Thorvaldsen and Ola Hössjer. Using statistical methods to model the fine-tuning of molecular machines and systems. Journal of Theoretical Biology, 501(??):??, September 21, 2020. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519320302071.

Tupper:2021:RCS

[TH21] Andrew S. Tupper and Paul G. Higgs. Rolling-circle and strand-displacement mechanisms for non-enzymatic RNA replication at the time of the origin of life. Journal of Theoretical Biology, 527(??):??, October 21, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321002411.

Terman:2025:MES

[TH25] David Terman and Yousef Hannawi. Modeling the effects of superoxide production on the cerebral blood flow. Journal of Theoretical Biology, 611(??):??, August 21, 2025. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519325001614.

Thakar:2024:PBB

[Tha24]  Juilee Thakar.  Pillars of biology: Boolean model-

ing of gene-regulatory networks. Journal of Theoretical Biology, 578(??):??, February 7, 2024. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519323002795.

Tomimoto:2025:BAA

[TIS25] Sou Tomimoto, Yoh Iwasa, and Akiko Satake. Branching architecture affects genetic diversity within an individual tree. Journal of Theoretical Biology, 605(??):??, May 21, 2025. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519325000591.

Tuschhoff:2025:HCB

[TK25] Beth M. Tuschhoff and David A. Kennedy. Heterogeneity in and correlation between host transmissibility and susceptibility can greatly impact epidemic dynamics. Journal of Theoretical Biology, 611(??):??, August 21, 2025. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519325001523.

Tanwer:2020:IPT

[TKB^{+}20] Pooja Tanwer, Sree Rohit Raj Kolora, Anshu Babbar, Daman Saluja, and Uma Chaudhry. Identification of potential therapeutic targets in Neisseria gonorrhoeae by an in-silico approach. Journal of Theoretical Biology, 490(??):??, April 7, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251932030028X.

Tuckman:2021:DSI

[TKR^{+}21] Harrison Tuckman, Jungmin Kim, Aaditya Rangan, Hong Lei, and Mainak Patel. Dynamics of sensory integration of olfactory and mechanical stimuli within the response patterns of moth antennal lobe neurons. Journal of Theoretical Biology, 509(??):??, January 21, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519320303659.

Tsur:2020:RPM

[TKRA20] N. Tsur, Y. Kogan, M. Rehm, and Z. Agur. Response of patients with melanoma to immune checkpoint blockade — insights gleaned from analysis of a new mathematical mechanistic model. Journal of Theoretical Biology, 485(??):??, January 21, 2020. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519319304023.

Tang:2023:LNG

[TMC^{+}23] Shan Tang, Shanjun Mao, Yangyang Chen, Falong Tan, Lihua Duan, Cong Pian, and Xiangxiang Zeng. LRBmat: a novel gut microbial interaction and individual heterogeneity inference method for colorectal cancer. Journal of Theoretical Biology, 571(??):??, August 21, 2023. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519323001340.

Tan:2020:MBC

[TML20] Pei Yen Tan, Marcos, and Yu Liu. Modelling bacterial chemotaxis for indirectly binding attractants. Journal of Theoretical Biology, 487(??):??, February 21, 2020. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319304898.

Tay:2022:IPR

[TMTK22] Chai Jian Tay, Mohd Hafiz Mohd, Su Yean Teh, and Hock Lye Koh. Internal phosphorus recycling promotes rich and complex dynamics in an algae-phosphorus model: Implications for eutrophication management. Journal of Theoretical Biology, 532(??):??, January 7, 2022. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321003325.

Tetteh:2021:NME

[TNHV21] Josephine N. A. Tetteh, Van Kinh Nguyen, and Esteban A. Hernandez-Vargas. Network models to evaluate vaccine strategies towards herd immunity in COVID-19. Journal of Theoretical Biology, 531(??):??, December 21, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321003131.

Tosato:2021:HCS

[TNW21] Marco Tosato, Kyeongah Nah, and Jianhong Wu. Are host control strategies effective to eradicate tick-borne diseases (TBD)? Journal of Theoretical Biology, 508(??):??, January 7, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519320303386.

Tamre:2024:SDS

[TP24] Erik Tamre and Chris Parsons. Selection by differential survival among marine animals in the Phanerozoic. Journal of Theoretical Biology, 590(??):??, August 7, 2024. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519324001309.

Traulsen:2025:MET

[TRKB25] Arne Traulsen, Mikkel Nif Rasmussen, Joachim Krug, and Andreas Beyer. On the misuse of evolutionary theory to bolster the ‘scientific’ case for intelligent design: a cautionary note. Journal of Theoretical Biology, 596(??):??, January 7, 2025. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519324002704. See [HBG21].

Tomimoto:2023:MSM

[TS23] Sou Tomimoto and Akiko Satake. Modelling somatic mutation accumulation and expansion in a long-lived tree with hierarchical modular architecture. Journal of Theoretical Biology, 565(??):??, May 21, 2023. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519323000619.

Tsugawa:2020:SSS

[Tsu20] Satoru Tsugawa. Suppression of soft spots and excited modes in the shape deformation model with spatio-temporal growth noise. Journal of Theoretical Biology, 486(??):??, February 7, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519319304618.

Tsuruyama:2021:NTB

[Tsu21] Tatsuaki Tsuruyama. Nonlinear thermodynamics of biological signal transduction for predicting conservation of entropy production rate. Journal of Theoretical Biology, 523(??):??, August 21, 2021. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519321001387.

Tanaka:2020:SVH

[TT20] Masaki Tanaka and Jun Tanimoto. Is subsidizing vaccination with hub agent priority policy really meaningful to suppress disease spreading? Journal of Theoretical Biology, 486(??):??, February 7, 2020. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251931930428X.

Tomizuka:2024:EED

[TT24] Haruto Tomizuka and Yuuya Tachiki. The eco-evolutionary dynamics of Batesian mimicry. Journal of Theoretical Biology, 577(??):??, January 21, 2024. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519323002801.

Teytsa:2021:CMP

[TTBL21] Hyacinthe M. Ndongmo Teytsa, Berge Tsanou, Samuel Bowong, and Jean Lubuma. Coupling the modeling of phage-bacteria interaction and cholera epidemiological model with and without optimal control. Journal of Theoretical Biology, 512(??):??, March 7, 2021. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519320303921.

Tu:2023:ECT

[TWT23] Han Tu, Xia Wang, and Sanyi Tang. Exploring COVID-19 transmission patterns and key factors during epidemics caused by three major strains in Asia. Journal of Theoretical Biology, 557(??):??, January 21, 2023. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519322003277.

Uchiyama:2022:OCC

[UIY22] Yuka Uchiyama, Yoh Iwasa, and Sachi Yamaguchi. Optimal composition of chloride cells for osmoregulation in a randomly fluctuating environment. Journal of Theoretical Biology, 537(??):??, March 21, 2022. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519322000145.

Uhl:2023:MED

[UV23] Peter M. Uhl and Naveen K. Vaidya. Modeling the effects of drugs of abuse on within-host dynamics of two HIV species. Journal of Theoretical Biology, 562(??):??, April 7, 2023. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519323000310.

Ventura:2022:CEM

[VA22] Rafael Ventura and Erol Akcay. A cognitive-evolutionary model for the evolution of teaching. Journal of Theoretical Biology, 533(??):??, January 21, 2022. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519321003520.

VanderMude:2020:SED

[Van20] Antony Van der Mude. Structure encoding in DNA. Journal of Theoretical Biology, 492(??):??, May 7, 2020. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519320300606.

Venkateswaran:2021:SCE

[VB21] Vignesh Venkateswaran and Renee M. Borges. Staying in the club: Exploring criteria governing metacommunity membership for obligate symbionts under host-symbiont feedback. Journal of Theoretical Biology, 510(??):??, February 7, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519320303672.

vanBenthem:2025:QEI

[vBBB^{+}25] Koen J. van Benthem, Rishabh Bagawade, Chantal Blüml, Peter Nabutanyi, Frans M. Thon, and Meike J. Wittmann. Quantifying the effects of intraspecific trait variation and interspecific trait correlations on interacting populations — a nonlinear averaging approach. Journal of Theoretical Biology, 609(??):??, July 21, 2025. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519325001006.

vanBenthem:2020:DDM

[vBW20] Koen J. van Benthem and Meike J. Wittmann. Density dependence on multiple spatial scales maintains spatial variation in both abundance and traits. Journal of Theoretical Biology, 491(??):??, April 21, 2020. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319305119.

Vajdi:2024:ADR

[VCS24] Aram Vajdi, Lee W. Cohnstaedt, and Caterina M. Scoglio. Assessing dengue risk globally using non-Markovian models. Journal of Theoretical Biology, 591(??):??, August 21, 2024. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519324001462.

vandenBosch:2023:VVS

[vdBMB^{+}23] Frank van den Bosch, Neil McRoberts, Yoann Bourhis, Stephen Parnell, and Kirsty L. Hassall. The value of volunteer surveillance for the early detection of biological invaders. Journal of Theoretical Biology, 560(??):??, March 7, 2023. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519322003769.

vanderVegt:2022:MMA

[vdVPW^{+}22] Solveig A. van der Vegt, Liudmila Polonchuk, Ken Wang, Sarah L. Waters, and Ruth E. Baker. Mathematical modelling of autoimmune myocarditis and the effects of immune checkpoint inhibitors. Journal of Theoretical Biology, 537(??):??, March 21, 2022. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519321004227.

Vazquez:2025:GBM

[VG25] Alexei Vazquez and Tomás Gedeon. Geometrically balanced model of cell growth. Journal of Theoretical Biology, 604(??): ??, May 7, 2025. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519325000517.

Villa:2022:NNP

[VGC22] Chiara Villa, Alf Gerisch, and Mark A. J. Chaplain. A novel nonlocal partial differential equation model of endothelial progenitor cell cluster formation during the early stages of vasculogenesis. Journal of Theoretical Biology, 534(??):??, February 7, 2022. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519321003830.

Vilches:2021:AED

[VGTY21] Thomas N. Vilches, Carlos Graeff-Teixeira, and Hyun Mo Yang. Assessing the effects of diagnostic sensitivity on schistosomiasis dynamics. Journal of Theoretical Biology, 523(??):??, August 21, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0022519321001491.

Vafadar:2020:IDH

[VKS^{+}20] Safar Vafadar, Kaveh Kavousi, Hadiseh Safdari, Ata Kalirad, and Mehdi Sadeghi. From indication to decision: a hierarchical approach to model the chemotactic behavior of Escherichia coli. Journal of Theoretical Biology, 495(??):??, June 21, 2020. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519320301089.

Valega-Mackenzie:2023:OCA

[VMRSL23] Wencel Valega-Mackenzie, Karen Ríos-Soto, and Suzanne Lenhart. Optimal control applied to Zika virus epidemics in Colombia and Puerto Rico. Journal of Theoretical Biology, 575(??):??, November 7, 2023. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519323002448.

Volpert:2024:MMR

[Vol24] V. Volpert. Mathematical model of repressive response to collective action and protest waves. Journal of Theoretical Biology, 595(??):??, December 7, 2024. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0022519324002558.

Vissio:2022:IIH

[VP22] Gabriele Vissio and Antonello Provenzale. On–off intermittency and irruptions in host–parasitoid dynamics. Journal of Theoretical Biology, 546(??):??, August 7, 2022. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322001722.

Vundavilli:2021:NMI

[VTDM21] Haswanth Vundavilli, Lokesh P. Tripathi, Aniruddha Datta, and Kenji Mizuguchi. Network modeling and inference of peroxisome proliferator-activated receptor pathway in high fat diet-linked obesity. Journal of Theoretical Biology, 519(??): ??, June 21, 2021. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519321000692.

Vibishan:2024:RBM

[VVD24] B. Vibishan, Harshavardhan B. V., and Sutirth Dey. A resource-based mechanistic framework for castrationresistant prostate cancer (CRPC). Journal of Theoretical Biology, 587(??):??, June 21, 2024. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519324000870.

Wedgwood:2022:MDC

[WA22] Kyle C. A. Wedgwood and Peter Ashwin. Morphogendirected cell fate boundaries: slow passage through bifurcation and the role of folded saddles. Journal of Theoretical Biology, 549(??):??, September 21, 2022. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322002181.

Wadhwa:2022:SFN

[Wad22] Dinkar Wadhwa. The series of a four-node motif and coherent type-1 feed-forward loop can provide sensitive detection over arbitrary range of input signal, thereby explain Weber’s law in higher-order sensory processes, and compute logarithm. Journal of Theoretical Biology, 543(??):??, June 21, 2022. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519322001059.

Wallace:2020:STC

[Wal20] Rodrick Wallace. Signal transduction in cognitive systems: Origin and dynamics of the inverted-U/U dose-response relations. Journal of Theoretical Biology, 504(??):??, November 7, 2020. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519320302320.

Walch:2022:MPI

[Wal22] Jean-Paul Walch. Modeling phyllotaxis: From the inhibition potential to the real plant. Journal of Theoretical Biology, 553(??):??, November 21, 2022. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322002521.

Wattis:2022:APG

[WBKR22] Jonathan A. D. Wattis, Sian M. Bray, Panagiota Kyratzi, and Cyril Rauch. Analysis of phenotype-genotype associations using genomic informational field theory (GIFT). Journal of Theoretical Biology, 548(??):??, September 7, 2022. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322001965.

Warne:2020:PGP

[WBS20] David J. Warne, Ruth E. Baker, and Matthew J. Simpson. A practical guide to pseudo-marginal methods for computational inference in systems biology. Journal of Theoretical Biology, 496(??):??, July 7, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519320301107.

Williamson:2021:CLL

[WCD21] Carly Williamson, Helen M. Chamberlin, and Adriana T. Dawes. Coordination of local and long range signaling modulates developmental patterning. Journal of Theoretical Biology, 517(??):??, May 21, 2021. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519321000187.

Wang:2025:IAC

[WCL^{+}25] Yuman Wang, Shuli Chen, Zhaolian Lu, Yu Liu, Jie Hu, and Da Zhou. Inferring absolute cell numbers from relative proportion in stochastic models with cell plasticity. Journal of Theoretical Biology, 608(??):??, July 7, 2025. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519325000992.

Waites:2021:RBE

[WCM^{+}21] W. Waites, M. Cavaliere, D. Manheim, J. Panovska-Griffiths, and V. Danos. Rule-based epidemic models. Journal of Theoretical Biology, 530(??):??, December 7, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321002708.

Wifvat:2021:RRM

[WCWL21] Kathryn Wifvat, Erika T. Camacho, Stephen Wirkus, and Thierry Léveillard. The role of RdCVFL in a mathematical model of photoreceptor interactions. Journal of Theoretical Biology, 520(??):??, July 7, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321000643.

Wang:2025:QIM

[WCZL25] Xu Wang, Xi Chen, Jinhui Zhu, and Sheng Li. Quantifying the impact of metronomic chemotherapy chemo-switch regimen and the sequencing of chemotherapy and radiotherapy on pancreatic ductal adenocarcinoma treatment. Journal of Theoretical Biology, 599(??):??, February 21, 2025.

CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519324003187.

Willard:2023:THC

[WDA^{+}23] Helena F. Willard, Eva S. Deutekom, Denis Allemand, Sylvie Tambutté, and Jaap A. Kaandorp. Testing hypotheses on the calcification in scleractinian corals using a spatio-temporal model that shows a high degree of robustness. Journal of Theoretical Biology, 561(??):??, March 21, 2023. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322003733.

Whitman:2020:STD

[WDH^{+}20] John Whitman, Aleya Dhanji, Fernand Hayot, Stuart C. Sealfon, and Ciriyam Jayaprakash. Spatio-temporal dynamics of Host–Virus competition: a model study of influenza A. Journal of Theoretical Biology, 484(??):??, January 7, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319303960.

Wickman:2020:HGP

[WDHB20] Jonas Wickman, Ulf Dieckmann, Cang Hui, and Åke Brännström. How geographic productivity patterns affect food-web evolution. Journal of Theoretical Biology, 506(??):??, December 7, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0022519320302290.

Wardah:2020:PPP

[WDT^{+}20] Wafaa Wardah, Abdollah Dehzangi, Ghazaleh Taherzadeh, Mahmood A. Rashid, M. G. M. Khan, Tatsuhiko Tsunoda, and Alok Sharma. Predicting protein-peptide binding sites with a deep convolutional neural network. Journal of Theoretical Biology, 496(??):??, July 7, 2020. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519320301338.

Wang:2024:ISS

[WFL^{+}24] Jian Wang, Zhenmin Fan, Jiashuai Liu, Kailei Liu, Chao-Jun Yan, Xia Ye, and Xiaoyan Deng. Influence of stent strut and its associated injury on thrombus formation: a dissipative particle dynamics study. Journal of Theoretical Biology, 595(??):??, December 7, 2024. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0022519324002145.

Wild:2023:KSM

[WFT23] Geoff Wild, Vonica J. Flear, and Graham J. Thompson. A kin-selection model of fairness in heterogeneous populations. Journal of Theoretical Biology, 565(??):??, May 21, 2023. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519323000656.

Waldeland:2021:SII

[WGRE21] Jahn Otto Waldeland, Jon-Vidar Gaustad, Einar K. Rofstad, and Steinar Evje. In silico investigations of intratumoral heterogeneous interstitial fluid pressure. Journal of Theoretical Biology, 526(??):??, October 7, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251932100206X.

Wang:2021:GPS

[WH21] Xiaoliang Wang and Andrew Harrison. A general principle for spontaneous genetic symmetry breaking and pattern formation within cell populations. Journal of Theoretical Biology, 526(??):??, October 7, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519321002289.

Woolley:2022:ADD

[WHH22] Thomas E. Woolley, William Hill, and Catherine Hogan. Accounting for dimensional differences in stochastic domain invasion with applications to precancerous cell removal. Journal of Theoretical Biology, 541(??):??, May 21, 2022. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322000224.

Whitehead:2020:CSG

[Whi20] Hal Whitehead. Cultural specialization and genetic diversity: Killer whales and beyond. Journal of Theoretical Biology, 490(??):??, April 7, 2020. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519320300205.

Whitmeyer:2021:SIS

[Whi21] Mark Whitmeyer. Strategic inattention in the Sir Philip Sidney Game. Journal of Theoretical Biology, 509(??):??, January 21, 2021. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519320303684.

Whittaker:2023:UES

[WHRH^{+}23] Dominic G. Whittaker, Alejandra D. Herrera-Reyes, Maurice Hendrix, Markus R. Owen, Leah R. Band, Gary R. Mirams, Kirsty J. Bolton, and Simon P. Preston. Uncertainty and error in SARS-CoV-2 epidemiological parameters inferred from population-level epidemic models. Journal of Theoretical Biology, 558(??):??, February 7, 2023. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322003289.

Wild:2023:PBG

[Wil23] Geoff Wild. Pillars of biology: ‘the genetical evolution of social behaviour, I and II’. Journal of Theoretical Biology, 562(??):??, April 7, 2023. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519323000267.

Wickman:2025:ETS

[WK25] Jonas Wickman and Christopher A. Klausmeier. The effects of trade-off shape and dimensionality on eco-evolutionary dynamics in resource competition. Journal of Theoretical Biology, 604(??):??, May 7, 2025. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519325000530.

Wang:2020:SAM

[WLJ^{+}20] Guan Wang, Hamid Laga, Jinyuan Jia, Stanley J. Miklavcic, and Anuj Srivastava. Statistical analysis and modeling of the geometry and topology of plant roots. Journal of Theoretical Biology, 486(??):??, February 7, 2020. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319304771.

Wei:2019:VDE

[WLW19] Yuting Wei, Yaosen Lin, and Bin Wu. Vaccination dilemma on an evolving social network. Journal of Theoretical Biology, 483(??):??, December 21, 2019. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S0022519319303212. See corrigendum [WLW20].

Wei:2020:CVD

[WLW20] Yuting Wei, Yaosen Lin, and Bin Wu. Corrigendum to “Vaccination dilemma on an evolving social network” [Journal of Theoretical Biology 483 (2019) 109978]. Journal of Theoretical Biology, 490(??):??, April 7, 2020. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519320300102. See [WLW19].

West:2020:FPR

[WM20] Robert West and Mauro Mobilia. Fixation properties of rock–paper–scissors games in fluctuating populations. Journal of Theoretical Biology, 491(??):??, April 21, 2020. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319305041.

Wang:2021:PSM

[WM21a] Jia-Zeng Wang and Ying-Tao Meng. The power spectrum of the membrane voltage driven by a single nACh receptor. Journal of Theoretical Biology, 509(??):??, January 21, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519320303830.

Wang:2021:VDP

[WM21b] Jia-Zeng Wang and Ying-Tao Meng. The voltage-depolarization performance vs the free energy cost of a single nACh receptor. Journal of Theoretical Biology, 531(??):??, December 21, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519321003234.

Whipple:2024:MCC

[WMHV24] Benjamin Whipple, Tanya A. Miura, and Esteban A. Hernandez-Vargas. Modeling the CD8+ T cell immune response to influenza infection in adult and aged mice. Journal of Theoretical Biology, 593(??):??, October 7, 2024. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519324001826.

Williams:2023:CSD

[WMO23] Thomas Williams, James M. McCaw, and James M. Osborne. Choice of spatial discretisation influences the progression of viral infection within multicellular tissues. Journal of Theoretical Biology, 573(??):??, September 21, 2023. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519323001893.

Week:2021:WNA

[WNHK21] Bob Week, Scott L. Nuismer, Luke J. Harmon, and Stephen M. Krone. A white noise approach to evolutionary ecology. Journal of Theoretical Biology, 521(??):??, July 21, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519321000825.

Worden:2020:CCF

[Wor20] Lee Worden. Conservation of community functional structure across changes in composition in consumer-resource models. Journal of Theoretical Biology, 493(??):??, May 21, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519320300941.

Wang:2021:MMH

[WP21] Jiayu Wang and Stephen J. Payne. Mathematical modelling of haemorrhagic transformation after ischaemic stroke. Journal of Theoretical Biology, 531(??):??, December 21, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321003398.

Weathered:2022:RBB

[WPEP22] Catherine Weathered, Kelly Pennington, Patricio Escalante, and Elsje Pienaar. The role of biofilms, bacterial phenotypes, and innate immune response in Mycobacterium avium colonization to infection. Journal of Theoretical Biology, 534(??):??, February 7, 2022. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519321003684.

White:2022:IDL

[WRB22] C. White, V. Rottschäfer, and L. J. Bridge. Insights into the dynamics of ligand-induced dimerisation via mathematical modelling and analysis. Journal of Theoretical Biology, 538(??):??, April 7, 2022. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519321004161.

Wang:2022:MEP

[WS22] Jia-Zeng Wang and Zhuo-Jun Shen. Mechanisms of the endplate potential noise and its implication for the neuromuscular junction anatomy. Journal of Theoretical Biology, 540(??):??, May 7, 2022. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251932200087X.

Watanabe:2024:SAE

[WS24] Yu Watanabe and Satoki Sakai. Simulation analyses of the evolution of intra-inflorescence flowering patterns assuming selection on anthesis interval among individual flowers. Journal of Theoretical Biology, 594(??):??, November 7, 2024.

CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519324002157.

Wilasang:2022:CEH

[WSC^{+}22] Chaiwat Wilasang, Pikkanet Suttirat, Sudarat Chadsuthi, Anuwat Wiratsudakul, and Charin Modchang. Competitive evolution of H1N1 and H3N2 influenza viruses in the United States: a mathematical modeling study. Journal of Theoretical Biology, 555(??):??, December 21, 2022. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322002831.

Williams:2023:AMM

[WSEK23] Katherine S. Williams, Timothy W. Secomb, and Ardith W. El-Kareh. An autonomous mathematical model for the mammalian cell cycle. Journal of Theoretical Biology, 569(??):??, July 21, 2023. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519323001297.

Wakamoto:2025:OTT

[WSL25] Tamaki Wakamoto and Sungrim Seirin-Lee. Optimal targeted therapy for multiple cancers based on contrastive Notch signaling networks. Journal of Theoretical Biology, 610(??):??, August 7, 2025. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519325001092.

Wang:2021:ESD

[WSW21] Guocheng Wang, Qi Su, and Long Wang. Evolution of state-dependent strategies in stochastic games. Journal of Theoretical Biology, 527(??):??, October 21, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251932100237X.

Wei:2021:DRD

[WSY21] Yingying Wei, Baojun Song, and Sanling Yuan. Dynamics of a ratio-dependent population model for Green Sea Turtle with age structure. Journal of Theoretical Biology, 516(??): ??, May 7, 2021. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519321000369.

Wu:2024:SIS

[Wu24] Min Wu. Simulating irregular symmetry breaking in gut cross sections using a novel energy-optimization approach in growth-elasticity. Journal of Theoretical Biology, 595(??):??, December 7, 2024. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S002251932400256X.

Whitfield:2023:MIR

[WUH23] Carl A. Whitfield, University of Manchester COVID-19 Modelling Group, and Ian Hall. Modelling the impact of repeat asymptomatic testing policies for staff on SARS-CoV-2 transmission potential. Journal of Theoretical Biology, 557(??):??, January 21, 2023. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519322003265.

Weiss:2021:EFR

[WvdDL^{+}21] Lora D. Weiss, P. van den Driessche, John S. Lowengrub, Dominik Wodarz, and Natalia L. Komarova. Effect of feedback regulation on stem cell fractions in tissues and tumors: Understanding chemoresistance in cancer. Journal of Theoretical Biology, 509(??):??, January 21, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519320303544.

Wang:2022:ISG

[WW22] Yue Wang and Zikun Wang. Inference on the structure of gene regulatory networks. Journal of Theoretical Biology, 539(??): ??, April 21, 2022. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519322000534.

Wang:2024:SMM

[WW24] Tianxu Wang and Hao Wang. Stoichiometric microplastics models in natural and laboratory environments. Journal of Theoretical Biology, 580(??):??, March 7, 2024. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519324000146.

Walker:2024:MMB

[WWR^{+}24] N. C. Walker, S. M. White, S. A. Ruiz, D. McKay Fletcher, M. Saponari, and T. Roose. A mathematical model of biofilm growth and spread within plant xylem: Case study of Xylella fastidiosa in olive trees. Journal of Theoretical Biology, 581(??):??, March 21, 2024. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519324000183.

Wang:2025:DCD

[WXS25] Jintao Wang, Yanni Xiao, and Pengfei Song. Discovering the climate dependent disease transmission mechanism through learning-explaining framework. Journal of Theoretical Biology, 601(??):??, March 21, 2025. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S002251932500013X.

Wang:2021:ITT

[WZKM21] Yue Wang, Boyu Zhang, Jérémie Kropp, and Nadya Morozova. Inference on tissue transplantation experiments. Journal of Theoretical Biology, 520(??):??, July 7, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321000679.

Wang:2023:CPG

[WZP^{+}23] Yue Wang, Joseph X. Zhou, Edoardo Pedrini, Irit Rubin, May Khalil, Roberto Taramelli, Hong Qian, and Sui Huang. Cell population growth kinetics in the presence of stochastic heterogeneity of cell phenotype. Journal of Theoretical Biology, 575(??):??, November 7, 2023. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519323002424.

Xu:2023:SWH

[XCFH23] Jingsi Xu, Jonathan Carruthers, Thomas Finnie, and Ian Hall. Simplified within-host and dose-response models of SARS-CoV-2. Journal of Theoretical Biology, 565(??):??, May 21, 2023. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519323000437.

Xu:2020:MMM

[XCH^{+}20] Bo Xu, Jun Cai, Daihai He, Gerardo Chowell, and Bing Xu. Mechanistic modelling of multiple waves in an influenza epidemic or pandemic. Journal of Theoretical Biology, 486(??):??, February 7, 2020. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519319304394.

Xia:2020:ERN

[XDT^{+}20] Fan Xia, Feng Deng, Hui Tian, Wei He, Yanni Xiao, and Xiaodan Sun. Estimation of the reproduction number and identification of periodicity for HFMD infections in northwest China. Journal of Theoretical Biology, 484(??):??, January 7, 2020. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519319303972.

Xu:2021:SDI

[XDZ^{+}21] Nuo Xu, Gustav W. Delius, Lai Zhang, Uffe H. Thygesen, and Ken H. Andersen. Spatial drivers of instability in marine sizespectrum ecosystems. Journal of Theoretical Biology, 517(??): ??, May 21, 2021. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519321000539.

Xie:2020:NTT

[Xie20] Ping Xie. Non-tight and tight chemomechanical couplings of biomolecular motors under hindering loads. Journal of Theoretical Biology, 490(??):??, April 7, 2020. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519320300291.

Xie:2021:DKM

[Xie21] Ping Xie. Dynamics of kinesin motor proteins under longitudinal and sideways loads. Journal of Theoretical Biology, 530(??):??, December 7, 2021. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S0022519321002988. See corrigendum [Xie22].

Xie:2022:CDK

[Xie22] Ping Xie. Corrigendum to “Dynamics of kinesin motor proteins under longitudinal and sideways loads” [J. Theor. Biol. 530 (2021) 110879]. Journal of Theoretical Biology, 544(??):??, July 7, 2022. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519322001357. See [Xie21].

Xie:2023:MMI

[Xie23] Ping Xie. Molecular mechanism of interaction between kinesin motors affecting their residence times on microtubule lattice and end. Journal of Theoretical Biology, 571(??):??, August 21, 2023. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519323001534.

Xie:2024:ESM

[Xie24] Ping Xie. Effect of small molecular crowders on dynamics of kinesin molecular motors. Journal of Theoretical Biology, 578(??):??, February 7, 2024. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519323002825.

Xie:2025:MTR

[Xie25] Ping Xie. A model of tubulin removal and exchange caused by kinesin motor walking on microtubule lattices. Journal of Theoretical Biology, 605(??):??, May 21, 2025. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519325000542.

Xu:2020:GPD

[XKV20] Yifang Xu, Andrew L. Krause, and Robert A. Van Gorder. Generalist predator dynamics under Kolmogorov versus non-Kolmogorov models. Journal of Theoretical Biology, 486(??):??, February 7, 2020. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519319304291.

Xu:2022:SSC

[XML22] Jingjing Xu, Evelyn H. Merrill, and Mark A. Lewis. Spreading speed of chronic wasting disease across deer groups with overlapping home ranges. Journal of Theoretical Biology, 547(??):??, August 21, 2022. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519322001333.

Xiao:2024:CCA

[XSH^{+}24] Chun Xiao, Yishui Sun, Huaxiong Huang, Xingye Yue, Zilong Song, Tim David, and Shixin Xu. Cellular communication among smooth muscle cells: the role of membrane potential via connexins. Journal of Theoretical Biology, 576(??):??, January 7, 2024. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519323002242.

Xu:2024:GSS

[XWS24] Jin Xu, H. Steven Wiley, and Herbert M. Sauro. Generating synthetic signaling networks for in silico modeling studies. Journal of Theoretical Biology, 593(??):??, October 7, 2024. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519324001851.

Xie:2020:WHM

[XYK^{+}20] Xiao-Ting Xie, Alexander Yitbarek, Salah Uddin Khan, Shayan Sharif, Zvonimir Poljak, and Amy L. Greer. A withinhost mathematical model of H9N2 avian influenza infection and type-I interferon response pathways in chickens. Journal of Theoretical Biology, 499(??):??, August 21, 2020. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519320301752.

Xie:2024:MAR

[XZW^{+}24] Qingxin Xie, Han Zhang, Maoxiang Wang, Kexun Yan, Fenglan Hu, and Meng Xu. A model about regulation on three division modes of stem cell. Journal of Theoretical Biology, 581(??):??, March 21, 2024. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519324000274.

Xue:2023:TMD

[XZZ^{+}23] Ling Xue, Hongyu Zhang, Xiaoming Zheng, Wei Sun, and Jinzhi Lei. Treatment of melanoma with dendritic cell vaccines and immune checkpoint inhibitors: a mathematical modeling study. Journal of Theoretical Biology, 568(??):??, July 7, 2023. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519323000851.

Yamauchi:2024:ELD

[Yam24] Atsushi Yamauchi. Evolution of labor division in reproduction and multiple group tasks. Journal of Theoretical Biology, 593(??):??, October 7, 2024. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519324001942.

Yang:2021:MET

[YAN21] Yichi Yang, Yusuke Asai, and Hiroshi Nishiura. A method for estimating the transmissibility of influenza using serial cross-sectional seroepidemiological data. Journal of Theoretical Biology, 511(??):??, February 21, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519320304215.

Yadipour:2023:OFE

[YBF23] Mehdi Yadipour, Md Arif Billah, and Imraan A. Faruque. Optic flow enrichment via Drosophila head and retina motions to support inflight position regulation. Journal of Theoretical Biology, 562(??):??, April 7, 2023. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519323000127.

Yoon:2024:OSC

[YCKK24] Changyeon Yoon, Jaemoo Choi, Hee-Dae Kwon, and Myungjoo Kang. Optimal STI controls for HIV patients based on an efficient deep Q learning method. Journal of Theoretical Biology, 594(??):??, November 7, 2024. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519324001991.

Yin:2020:EMN

[YdKdJ^{+}20] Shenglai Yin, Henrik J. de Knegt, Mart C. M. de Jong, Yali Si, Herbert H. T. Prins, Zheng Y. X. Huang, and Willem F. de Boer. Effects of migration network configuration and migration synchrony on infection prevalence in geese. Journal of Theoretical Biology, 502(??):??, October 7, 2020. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519320301703.

Yang:2024:IEM

[YDS24] Junyuan Yang, Xinyi Duan, and Guiquan Sun. An immunoepidemiological model with non-exponentially distributed disease stage on complex networks. Journal of Theoretical Biology, 595(??):??, December 7, 2024. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0022519324002492.

Younes:2021:MMA

[YE21] G. Abi Younes and N. El Khatib. Mathematical modeling of atherogenesis: Atheroprotective role of HDL. Journal of Theoretical Biology, 529(??):??, November 21, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321002745.

Yin:2025:ASS

[YFF25] Yuan Yin, Jennifer A. Flegg, and Mark B. Flegg. Accurate stochastic simulation algorithm for multiscale models of infectious diseases. Journal of Theoretical Biology, 612(??):??, September 7, 2025. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519325001602.

Yu:2020:TCF

[YFZ^{+}20] Jie-Ru Yu, Tian-Jiao Feng, Xiu-Deng Zheng, Da-Hua Chen, and Yi Tao. Transitions in the cell-fate induction induced by colored noise associated with the inductive stimulus. Journal of Theoretical Biology, 484(??):??, January 7, 2020. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319303881.

Young:2025:UCG

[YG25] Stephanie Young and Jérôme Gilles. Use of 3D chaos game representation to quantify DNA sequence similarity with applications for hierarchical clustering. Journal of Theoretical Biology, 596(??):??, January 7, 2025. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519324002571.

Yan:2024:UQR

[YHT24] Dingding Yan, Mengqi He, and Sanyi Tang. Uncertainty quantification for the random HIV dynamical model driven by drug adherence. Journal of Theoretical Biology, 592(??):??, September 7, 2024. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519324001796.

Yamaguchi:2021:EGA

[YI21] Sachi Yamaguchi and Yoh Iwasa. Evolutionary game in an androdioecious population: Coupling of outcrossing and male production. Journal of Theoretical Biology, 513(??):??, March 21, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519321000163.

Yin:2021:LPC

[Yin21] Changchuan Yin. Latent periodicity-2 in coronavirus SARS-CoV-2 genome: Evolutionary implications. Journal of Theoretical Biology, 515(??):??, April 21, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321000266.

Yule:2020:IAI

[YJBF20] Kelsey M. Yule, Christopher A. Johnson, Judith L. Bronstein, and Régis Ferrière. Interactions among interactions: the dynamical consequences of antagonism between mutualists. Journal of Theoretical Biology, 501(??):??, September 21, 2020. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519320301892.

Yao:2024:NMF

[YJZ^{+}24] Zhihao Yao, Suoqin Jin, Fuling Zhou, Junbai Wang, Kai Wang, and Xiufen Zou. A novel multiscale framework for delineating cancer evolution from subclonal compositions. Journal of Theoretical Biology, 582(??):??, April 7, 2024. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519324000249.

Yao:2025:IAP

[YL25] Lingxing Yao and Yizeng Li. Implementation of actin polymerization and depolymerization in a two-dimensional cell migration model and its implications on mammalian cell morphology and velocity. Journal of Theoretical Biology, 596(??):??, January 7, 2025. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519324002625.

Yuan:2022:QMP

[YLT22] Baoyin Yuan, Rui Liu, and Sanyi Tang. A quantitative method to project the probability of the end of an epidemic: Application to the COVID-19 outbreak in Wuhan, 2020. Journal of Theoretical Biology, 545(??):??, July 21, 2022. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519322001473.

Yuan:2021:SST

[YLZ^{+}21] Ye Yuan, Jian Liu, Peng Zhao, Hong Huo, and Tao Fang. Spike signal transmission between modules and the predictability of spike activity in modular neuronal networks. Journal of Theoretical Biology, 526(??):??, October 7, 2021.

CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321002307.

Yu:2023:NID

[YM23] Xingwang Yu and Yuanlin Ma. Noise-induced dynamics in a single species model with Allee effect driven by correlated colored noises. Journal of Theoretical Biology, 573(??):??, September 21, 2023. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519323002072.

Yang:2024:JJC

[YMR24] Feifei Yang, Jun Ma, and Guodong Ren. A Josephson junction-coupled neuron with double capacitive membranes. Journal of Theoretical Biology, 578(??):??, February 7, 2024. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519323002837.

Yan:2024:BPI

[YOIS24] Xin Yan, Goshi Ogita, Shuji Ishihara, and Kaoru Sugimura. Bayesian parameter inference for epithelial mechanics. Journal of Theoretical Biology, 595(??):??, December 7, 2024. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0022519324002455.

Ye:2020:QPU

[YRMWT20] Yanan Ye, Alvaro Ruiz-Martinez, Peng Wang, and Daniel M. Tartakovsky. Quantification of predictive uncertainty in models of FtsZ ring assembly in Escherichia coli. Journal of Theoretical Biology, 484(??):??, January 7, 2020. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319303765.

Yadav:2025:EIR

[YS25] K. S. Yadav and Gopinath Sadhu. Effect of inosine on recurrence of tumor after radiation therapy: a mathematical investigation. Journal of Theoretical Biology, 609(??):??, July 21, 2025. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519325001043.

Yadav:2023:CTP

[YSK23] Saumya Yadav, Aritra Sen, and Ambarish Kunwar. Cargo transport properties are enhanced by cylindrical microtubule geometry and elliptical contact zone on cargo surface. Journal of Theoretical Biology, 565(??):??, May 21, 2023. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519323000620.

Yin:2023:HTC

[YSP^{+}23] Fulian Yin, Yuwei She, Yanyan Pan, Xinyi Tang, Haotong Hou, and Jianhong Wu. Hot-topics cross-propagation and opinion-transfer dynamics in the Chinese Sina-microblog social media: a modeling study. Journal of Theoretical Biology, 566(??):??, June 7, 2023. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519323000760.

Yang:2021:CHB

[YST21] Lei Yang, Weigang Sun, and Marc Turcotte. Coexistence of Hopf-born rotation and heteroclinic cycling in a time-delayed three-gene auto-regulated and mutuallyrepressed core genetic regulation network. Journal of Theoretical Biology, 527(??):??, October 21, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321002320.

Yang:2024:ABE

[YTG24] Qin Yang, Yi Tang, and Dehua Gao. Agent-based evolutionary game dynamics uncover the dual role of resource heterogeneity in the evolution of cooperation. Journal of Theoretical Biology, 595(??):??, December 7, 2024. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0022519324002376.

Yang:2021:UMM

[YTH^{+}21] Boya Yang, Xi Tang, Michael J. Haller, Desmond A. Schatz, and Libin Rong. A unified mathematical model of thyroid hormone regulation and implication for personalized treatment of thyroid disorders. Journal of Theoretical Biology, 528(??):??, November 7, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321002721.

Yokoyama:2021:CFF

[YTW21] Naoto Yokoyama, Naoki Takeishi, and Shigeo Wada. Cerebrospinal fluid flow driven by arterial pulsations in axisymmetric perivascular spaces: Analogy with Taylor’s swimming sheet. Journal of Theoretical Biology, 523(??):??, August 21, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519321001314.

Yan:2020:IMR

[YTY^{+}20] Qinling Yan, Yingling Tang, Dingding Yan, Jiaying Wang, Linqian Yang, Xinpei Yang, and Sanyi Tang. Impact of media reports on the early spread of COVID-19 epidemic. Journal of Theoretical Biology, 502(??):??, October 7, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251932030240X.

Yan:2023:CMA

[YWQX23] Kexun Yan, Maoxiang Wang, Zhipeng Qiu, and Meng Xu. A cell model about symmetric and asymmetric stem cell division. Journal of Theoretical Biology, 560(??):??, March 7, 2023. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251932200371X.

Yu:2025:TEV

[YWTC25] Zhuojun Yu, Yangyang Wang, Peter J. Thomas, and Hillel J. Chiel. Tradeoffs in the energetic value of neuromodulation in a closed-loop neuromechanical system. Journal of Theoretical Biology, 604(??):??, May 7, 2025. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519325000165.

Yang:2025:ERC

[YWWF25] Liu Yang, Qi Wang, Youming Wang, and Meng Fan. Establishment of risk classification index for animal brucellosis by stochastic models. Journal of Theoretical Biology, 611(??):??, August 21, 2025. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519325001535.

Yu:2021:GPM

[YXW21] Chong Yu, Hong Xu, and Jin Wang. A global and physical mechanism of gastric cancer formation and progression. Journal of Theoretical Biology, 520(??):??, July 7, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321000655.

Yin:2021:IRC

[YY21] Changchuan Yin and Stephen S.-T. Yau. Inverted repeats in coronavirus SARS-CoV-2 genome manifest the evolution events. Journal of Theoretical Biology, 530(??):??, December 7, 2021. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519321003040.

Yamaguchi:2021:ELC

[YYI21] Sachi Yamaguchi, Yoichi Yusa, and Yoh Iwasa. Evolution of life cycle dimorphism: an example of sacoglossan sea slugs. Journal of Theoretical Biology, 525(??):??, September 21, 2021. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251932100182X.

Yin:2020:SRP

[YZB20] Zhanyuan Yin and Leif Zinn-Björkman. Simulating rolling paths and reorientation behavior of ball-rolling dung beetles. Journal of Theoretical Biology, 486(??):??, February 7, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519319304758.

Yan:2020:CAE

[YZSS20] Luyao Yan, Honghui Zhang, Zhongkui Sun, and Zhuan Shen. Control analysis of electrical stimulation for epilepsy waveforms in a thalamocortical network. Journal of Theoretical Biology, 504(??):??, November 7, 2020. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519320302460.

Zhang:2022:NMO

[ZB22] Yi Zhang and Daniel Boley. Nonlinear multi-objective flux balance analysis of the Warburg Effect. Journal of Theoretical Biology, 550(??):??, October 7, 2022. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322002211.

Zhou:2022:EVC

[ZBT22] Weike Zhou, Yao Bai, and Sanyi Tang. The effectiveness of various control strategies: an insight from a comparison modelling study. Journal of Theoretical Biology, 549(??):??, September 21, 2022. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251932200203X.

Zhdanova:2021:GPU

[ZF21] Oksana L. Zhdanova and Efim Ya. Frisman. Genetic polymorphism under cyclical selection in long-lived species: the complex effect of age structure and maternal selection. Journal of Theoretical Biology, 512(??):??, March 7, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519320304197.

Zhang:2021:SSM

[ZG21] Ruqiang Zhang and Didier Gonze. Stochastic simulation of a model for circadian rhythms in plants. Journal of Theoretical Biology, 527(??):??, October 21, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321002095.

Zhang:2020:USU

[ZL20] Xinyu Zhang and Alan J. Levy. On the uniform stress/ uniform stretch states of prestressed arteries. Journal of Theoretical Biology, 486(??):??, February 7, 2020. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319304692.

Zhao:2020:MPE

[ZLYL20] Yu Zhao, Rong Liang, Yiying Yang, and Songyi Lin. The mechanism of pulsed electric field (PEF) targeting location on the spatial conformation of pine nut peptide. Journal of Theoretical Biology, 492(??):??, May 7, 2020. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519320300503.

Zhong:2025:HGG

[ZLZ^{+}25] Fulian Zhong, Yujian Liu, Jianquan Zhong, Ling He, Zhonglan Tang, and Jing Zhang. Hybrid of glioma growth model and deformable image registration for longitudinal brain MRIs synthesis. Journal of Theoretical Biology, 610(??):??, August 7, 2025. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519325001134.

Zeaiter:2020:PHT

[ZM20] Zeaiter Zeaiter and Mary R. Myerscough. Poor hive thermoregulation produces an Allee effect and leads to colony collapse. Journal of Theoretical Biology, 503(??):??, October 21, 2020. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519320302162.

Zobolas:2022:BFM

[ZMKF22] John Zobolas, Pedro T. Monteiro, Martin Kuiper, and Åsmund Flobak. Boolean function metrics can assist modelers to check and choose logical rules. Journal of Theoretical Biology, 538(??):??, April 7, 2022. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519322000236.

Zhang:2023:DVS

[ZPP23] Peng Zhang, Sean D. Peterson, and Maurizio Porfiri. Dipoleand vortex sheet-based models of fish swimming. Journal of Theoretical Biology, 556(??):??, January 7, 2023. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322003046.

Zhang:2024:DMC

[ZQHN24] Tong Zhang, Jiaying Qiao, Katsuma Hayashi, and Hiroshi Nishiura. Decomposing mechanisms of COVID-19 mortality in empirical datasets: a modeling study. Journal of Theoretical Biology, 584(??):??, May 7, 2024. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519324000523.

Zhang:2021:MOC

[ZRL^{+}21] Xianghong Zhang, Xinmiao Rong, Juan Li, Meng Fan, Youming Wang, Xiangdong Sun, Baoxu Huang, and Huaiping Zhu. Modeling the outbreak and control of African swine fever virus in large-scale pig farms. Journal of Theoretical Biology, 526(??):??, October 7, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321002174.

Zhao:2023:MTA

[ZSL23] Na Zhao, Jian Song, and Shenquan Liu. Multi-timescale analysis of midbrain dopamine neuronal firing activities. Journal of Theoretical Biology, 556(??):??, January 7, 2023. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322003010.

Zeltser:2022:MMM

[ZSN22] Gregory Zeltser, Ilya M. Sukhanov, and Alexey J. Nevorotin. MMM — the molecular model of memory. Journal of Theoretical Biology, 549(??):??, September 21, 2022. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251932200217X.

Zaytseva:2022:PFM

[ZSS^{+}22] Sofya Zaytseva, Leah B. Shaw, Junping Shi, Matthew L. Kirwan, and Romuald N. Lipcius. Pattern formation in marsh ecosystems modeled through the interaction of marsh vegetation, mussels and sediment. Journal of Theoretical Biology, 543(??):??, June 21, 2022. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251932200100X.

Zhao:2022:NPI

[ZWC^{+}22] Shi Zhao, Kai Wang, Marc K. C. Chong, Salihu S. Musa, Mu He, Lefei Han, Daihai He, and Maggie H. Wang. The non-pharmaceutical interventions may affect the advantage in transmission of mutated variants during epidemics: a conceptual model for COVID-19. Journal of Theoretical Biology, 542(??):??, June 7, 2022. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322001035.

Zhao:2024:IST

[ZWT24] Wuqiong Zhao, Xia Wang, and Biao Tang. The impacts of spatial-temporal heterogeneity of human-to-human contacts on the extinction probability of infectious disease from branching process model. Journal of Theoretical Biology, 579(??):??, February 21, 2024. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519323003004.

Zhong:2022:SPR

[ZWW22] Haonan Zhong, Kaifa Wang, and Wendi Wang. Spatiotemporal pattern recognition and dynamical analysis of COVID-19 in Shanghai, China. Journal of Theoretical Biology, 554(??):??, December 7, 2022. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519322002703.

Zhang:2023:EDU

[ZWWW23] Feipeng Zhang, Xiaofeng Wang, Te Wu, and Long Wang. Evolutionary dynamics under partner preferences. Journal of Theoretical Biology, 557(??):??, January 21, 2023. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322003319.

Zhou:2022:ECF

[ZX22] Xiao Zhou and BingKan Xue. Effect of compositional fluctuation on the survival of bet-hedging species. Journal of Theoretical Biology, 553(??):??, November 21, 2022. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322002612.

Zhang:2023:TRA

[ZXLH23] Kai Zhang, Ling Xue, Xuezhi Li, and Daihai He. Transmission risk assessment of HIV/AIDS epidemic resulting from sexual transmission in China, 2013–2017. Journal of Theoretical Biology, 575(??):??, November 7, 2023. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519323002321.

Zhao:2025:WLC

[ZYF25] Xin Zhao, Shiyang Yu, and Meng Fan. White lies for coral reefs: Dynamics of two-patch coral reefs model with asymmetric dispersal. Journal of Theoretical Biology, 601(??):??, March 21, 2025. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0022519325000128.

Zhou:2021:VTG

[ZYLW21] Jian Zhou, Kailiang Yu, Guanghui Lin, and Zhiheng Wang. Variance in tree growth rates provides a key link for completing the theory of forest size structure formation. Journal of Theoretical Biology, 529(??):??, November 21, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321002769.

Zhang:2022:MEA

[ZZ22] Rui Zhang and Yunhu Zhang. Modeling the effects of air pollutants and meteorological factors on scarlet fever in five provinces, Northwest China, 2013–2018. Journal of Theoretical Biology, 544(??):??, July 7, 2022. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519322001321.

Zhang:2021:OCS

[ZZB^{+}21] Bo Zhang, Lu Zhai, Jason Bintz, Suzanne M. Lenhart, Wencel Valega-Mackenzie, and J. David Van Dyken. The optimal controlling strategy on a dispersing population in a two-patch system: Experimental and theoretical perspectives. Journal of Theoretical Biology, 528(??):??, November 7, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251932100254X.

Zhang:2021:CME

[ZZGG21] Manyu Zhang, Alice Zhao, Chun Guo, and Lingzhong Guo. A combined modelling and experimental study of heat shock factor SUMOylation in response to heat shock. Journal of Theoretical Biology, 530(??):??, December 7, 2021. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321002964.

Zhao:2021:SSI

[ZZT^{+}21] Shi Zhao, Yu Zhao, Biao Tang, Daozhou Gao, Zihao Guo, Marc K. C. Chong, Salihu S. Musa, Yongli Cai, Weiming Wang, Daihai He, and Maggie H. Wang. Shrinkage in serial intervals across transmission generations of COVID-19. Journal of Theoretical Biology, 529(??):??, November 21, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519321002800.